### v1

# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

Abstract
In recent years, spatio-temporal graph neural networks (GNNs) have attracted considerable interest in the field of time series analysis, due to their ability to capture dependencies among variables and across time points. The objective of the presented systematic literature review is hence to provide a comprehensive overview of the various modeling approaches and application domains of GNNs for time series classification and forecasting. A database search was conducted, and over 150 journal papers were selected for a detailed examination of the current state-of-the-art in the field. This examination is intended to offer to the reader a comprehensive collection of proposed models, links to related source code, available datasets, benchmark models, and fitting results. All this information is hoped to assist researchers in future studies. To the best of our knowledge, this is the first systematic literature review presenting a detailed comparison of the results of current spatio-temporal GNN models in different domains. In addition, in its final part this review discusses current limitations and challenges in the application of spatio-temporal GNNs, such as comparability, reproducibility, explainability, poor information capacity, and scalability.
1 Introduction
In recent years, graph neural networks (GNNs) have emerged as a powerful class of artificial neural network models for processing data that can be represented as graphs. GNNs are in fact particularly well-suited for a wide range of practical and engineering applications where data naturally lend themselves to be represented as graph structures, such as applications to transportation networks, image analysis, and natural language processing. The intuitive concept behind this approach is that nodes in a graph can be put in correspondence to objects or concepts, whereas edges can be assumed to represent their mutual relationships. Data for these quantities can then be processed at once by a GNN model. GNNs can be used for three different classes of problems: at graph level, at edge level, and at node level. In a graph-level problem, the goal is to predict a property of a graph based on its entire structure, rather than on individual nodes or edges. For example, a molecule from a sample can be represented as a graph, and global properties of this molecule can be inferred from data coming from its entire structure in relation to the sample. In an edge-level problem, the goal is to predict the presence or absence of edges between pairs of nodes. As an example, this task can be used in recommendation systems to predict potential connections between users and items on the basis of past interactions. In node-level problems, the goal is to predict the identity or role of each node within a graph, in order to solve either a classification or a regression task. Noticeably, regression tasks can be ”autoregressions”, that is, they can include time.
This review examines then the specific use of GNNs on time series related tasks, with a particular focus on time series classification and forecasting. The idea behind the use of GNNs for time series problems is that GNNs can be made capable of capturing complex relationships, both inter-variable (connections between different variables within a multivariate series) but also inter-temporal (dependencies between different points in time) at once. This results in a spatio-temporal GNN approach, where the spatial dimension is related to the multivariate framework, and the temporal dimension is related to the temporal nature of the data. Namely, spatio-temporal GNNs are a class of GNN models designed to handle data in both spatial and temporal dimensions at once. In this case, in a spatio-temporal graph, data in each node or edge can evolve over time, and the challenge is here to model the related dynamic interactions over time.
Notwithstanding all this potential interest, to the best of our knowledge there is a lack of systematic literature reviews (SLRs) on spatio-temporal GNN models for time series applications. Most of the existing surveys are not SLRs, they are not recent, or only focus on a limited number of application domains. Individually, recent reviews focus for example on algorithmic features [cit_rev1], [cit_rev2], [cit_rev3], [survey_Lucheroni], GNN characterization [cit_rev8], specific fields [cit_rev4], and distributed training [cit_rev5], [cit_rev6]. Only a SLR on spatio-temporal GNNs was indeed recently published by Zeghina et al. [survey_spatiotemporal], but due to the restrictive search query, it covers only 52 publications.
The intention of the SLR presented here is to provide a broad and comprehensive overview of the applications of spatio-temporal GNNs for time series classification and forecasting in different areas. This SLR also aims to assess if the current popularity of spatio-temporal GNN models is indeed due to their effectiveness, and to evaluate their efficacy and accuracy in different fields. Additionally, it collects results and benchmarks to assist researchers in their work by synthesizing the available literature on GNN applications. To the best of our knowledge, this is the first review to present comprehensive tables with all the results from the various models and benchmarks proposed in a highly fragmented literature. It is thus hoped that this collected and distilled knowledge can become a valuable resource and reference for researchers.
In this review two sets of questions will be posed. The answers to the first set want to provide a general overview of the publications in the field, while the answers to the second set focus on specific aspects of the proposed GNN models.
The general overview questions (GQs) are:
  1. GQ1)
    Trend: What is the temporal trend of publications on GNN models? Is there a growing interest in GNNs?
  2. GQ2)
    Fields: In which fields are GNN models most commonly applied?
  3. GQ3)
    Journals: In which journals are the papers on GNNs published? Are these journals chosen because of the specific domain of application or not?
  4. GQ4)
    Research groups: Which are the most active research groups on GNNs?
  5. GQ5)
    Tools: What tools (programming languages, libraries, frameworks) are used to implement the GNN models?
  6. GQ6)
    Fundings: Did the authors of the papers receive public or private funding for their research?
The specific questions (SQs) are:
  1. SQ1)
    Applications: Which are the most studied applications of spatio-temporal GNNs? Are there differences in approaches and results across different application areas?
  2. SQ2)
    Graph construction: Is the graph structure predetermined? If not, how do researchers define it?
  3. SQ3)
    Taxonomy: Among the various taxonomy classes, which are the most common? Are there some recurring mechanisms?
  4. SQ4)
    Benchmark models: Which are the most common benchmark models? Are they classical ML models or also GNN models?
  5. SQ5)
    Benchmark datasets: Which are the most common benchmark datasets?
  6. SQ6)
    Modeling paradigms: What is the most common paradigm? Is it modeling complex interacting systems, or modeling a system with multiple interacting quantities? Specifically, do graphs typically condense the relationships among multiple entities, or do they describe different aspects of the same entity? In other words, is the graph structure homogeneous or heterogeneous?
  7. SQ7)
    Metrics: Which are the most common metrics used to assess the accuracy of a given model?
The answers to both the general and the specific questions will be integrated into the following discussion in order to provide a comprehensive overview of the present state-of-the-art of spatio-temporal GNN models. A general overview of such GNN models will be assembled by highlighting similarities and differences between the approaches across different fields. However, in case, for a more detailed analysis of a specific discipline or other aspects of GNN models, the reader is invited to consult the specialized reviews mentioned before ([cit_rev1], [cit_rev2], [cit_rev3], [survey_Lucheroni], [cit_rev8], [cit_rev4], [cit_rev5], [cit_rev6]).
The paper is organized as follows. After this Introduction, Sec. 2 illustrates the methodology used to collect the papers. Sec. 3 provides a general overview of the selected papers. Sec. 4 introduces some fundamental definitions and notions at the basis of spatio-temporal GNN models, and discusses their taxonomy and the determination of the graph structure. Sec. 5 is the core of this review. It explores the diverse domains of application of the spatio-temporal GNN models proposed in the selected papers for the identified domains. A discussion of the findings and the answers to the research questions mentioned above are provided in Sec. 6. Finally, limits, challenges and future research directions are presented in Sec. 7. Sec. 8 concludes the review. Appendix  A provides a list of all journal papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (e.g., classification or forecasting).
2 Methodology for the systematic literature review
To conduct this SLR, four primary databases were consulted: Scopus, IEEE Xplore, Web of Science, and ACM. These databases were chosen for their comprehensive coverage of academic and scientific literature, ensuring a broad and inclusive search across multiple disciplines.
Prior to commencing the literature search, some exclusion criteria were defined. First, papers from the current year 2024 were excluded because it has not yet ended, hence the review only covers up to year 2023. Second, only journal articles were considered, in order to guarantee the inclusion of ”established” sources, whereas both conference papers and book chapters were excluded. The exclusion of conference papers is also made in order to keep the number of papers manageable. However, it is important to notice that relevant conference papers which provide useful benchmarks are anyway referenced throughout the text. Third, only articles published in English were included because of facility of fruition and comparison. Review articles are also not included, although they may be mentioned in the test.
An advanced search query submitted to the databases was designed to capture the breadth of the topic while maintaining specificity. The query used is (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting). The first two elements of query aim to restrict the pool of GNN models to the time series domain, and the second part focuses the search on classification and forecasting applications.
The selection process of the papers followed the multi-stage approach described below:
  1. 1.
    Database search: the search query was submitted to each of the four databases, and all records were imported into a csv file.
  2. 2.
    Duplicate removal: duplicate records were identified and removed.
  3. 3.
    Title and abstract screening: the remaining records were screened by reading their titles and abstracts. Studies that at first sight did not meet the criteria were discarded.
  4. 4.
    Full-text screening: the full texts of the remaining papers were analyzed one by one, which led to a further assessment of eligibility.
  5. 5.
    Data extraction: relevant data were extracted from the included studies, and analyzed in order to better discuss the approach and present the most common datasets and models for each identified field.
A total of 473 records were identified, 188 from Scopus, 31 from IEEE, 104 from Web of Science, and 150 from ACM. A scheme of this selection process is sketched in Fig. 1. First, 142 duplicates were removed, then the titles and abstracts of the remaining 331 records were manually evaluated to select the pertinent papers. Among these, 170 were discarded because they were considered outside the scope of this review, and 5 were discarded because they were not in English. Hence, 156 records were included in the review.
Figure 1: PRISMA flowchart to summarize the identification of the studies and the selection process, adapted from Ref. [PRISMA_statement].
Next, Sec. 3 provides an overview of the selected publications.
3 Overview of the publications
This section provides a comprehensive overview of the 156 selected journal articles. After their selection, the papers were grouped according to their domain of application, as better explained in the thematic analysis of Sec. 5. In the following, some bibliographic information about the reviewed articles is presented.
Fig. 2 illustrates the number of publications over time across different groups. A cumulative positive trend can be observed over time, reflecting an increasing interest in spatio-temporal GNN models across various fields, since 2020. The three most common groups that have seen the greatest increase in publications are ”Environment”, ”Generic”, and ”Mobility”. For more details, see also the group numerosity reported in Tab. 1.
Figure 2: Number of journal publications over time across the different groups.
Table 1: Pivot table illustrating the number of publications by year and group.
Field / Year 2020 2021 2022 2023 Total
Energy 1 6 6 13
Environment 2 4 21 27
Finance 3 6 9
Generic 1 4 19 24
Health 1 2 5 8
Mobility 2 4 15 18 39
Other 2 4 14 20
Predictive monitoring 2 3 11 16
Total 3 12 41 100 156
An inspection of the publication sources indicates that 95 different journals have published articles on GNN models for time series forecasting or classification. Tab. 2 lists the 15 journals which contain the highest number of published papers.
Table 2: Top 15 journals with the highest number of published paper.
Journal Number of papers
IEEE Access 7
ACM Transactions on Knowledge Discovery from Data 6
Applied Intelligence 6
IEEE Transactions on Instrumentation and Measurement 6
ACM Transactions on Intelligent Systems and Technology 5
Expert Systems with Applications 5
Sensors 4
Applied Energy 3
Applied Soft Computing 3
Energy 3
Engineering Applications of Artificial Intelligence 3
IEEE Transactions on Intelligent Transportation Systems 3
Information Sciences 3
International Journal of Data Science and Analytics 3
Knowledge-Based Systems 3
The diversity of the journals indicates a vast range of potential applications for spatio-temporal GNN models in different fields. This suggests that researchers often select journals based on the specific domain of application, allowing them to reach targeted audiences and contribute within their respective fields. Such diversity also underlines the interdisciplinary nature of GNNs research, as it spans multiple domains.
Figure 3: Pie chart of the distribution of corresponding authors in different countries.
In order to examine the degree of collaboration among researchers, a co-authorship network was generated by using the VOSviewer software [VOSviewer]. Fig. 4 illustrates authors as nodes and collaborations as connections, revealing some research clusters. Two of these clusters are enlarged and shown in Fig. 4.
(a) Visualization of the complete collaboration network.
(b) Zoom in on two different clusters.
Figure 4: Network of collaboration between researchers, created using VOSviewer software.
In the figure, the color of the nodes is related to the average publication year of an author, so that the evolution of research groups over time can be evaluated. Two key observations can be made. First, the number of clusters suggests that there are many distinct research groups, indicating a high degree of fragmentation within the community. Second, the color of the nodes indicates that some groups were active only at the beginning and have since disappeared, while others have recently formed, as indicated by their yellow color. Some authors have various collaborations over time, as also shown in Fig. 4.
Fig. 3 shows a pie chart of the distribution of corresponding authors among different countries. This visualization provides an overview of the global distribution of research activities. More than the 70% of the authors are affiliated with institutions in China.
4 Graph neural networks
This section introduces some fundamental definitions and notions at the basis of the study of time series with spatio-temporal GNNs.
4.1 Definitions and notations
Definition 1 (Time series).
A time series is a sequence of data points indexed in time order (that here is assumed to be regularly sampled, i.e., equispaced). An equispaced univariate time series of length $T$ is a sequence of scalar observations collected over time, denoted as $\vec{x}_{t}\in\mathbb{R}^{T}$. An equispaced multivariate time series of length $T$ is a sequence of $D$-dimensional vector observations collected over time, denoted as $\vec{x}_{t}\in\mathbb{R}^{D\times T}$.
Definition 2 (Graph).
A graph $\mathcal{G}$ is a pair of finite sets $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ is a set of $n$ nodes (also called vertices) and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. In an undirected graph each edge is an unordered pair of nodes $\{v_{i},v_{j}\}$, while in a directed graph the edges have an orientation and correspond to ordered pairs $(v_{i},v_{j})$. When an edge $(v_{i},v_{j})$ or $\{v_{i},v_{j}\}$ exists, the nodes $v_{i}$ and $v_{j}$ are called adjacent.
Definition 3 (Spatio-temporal graph).
A spatio-temporal graph is a 4-tuple $\mathcal{G}=(\mathcal{V}_{t},\mathcal{E}_{t},\mathcal{X}_{t},\mathcal{T})$, where $\mathcal{T}=\{t_{1},\dots,t_{T}\}$ is a set of $T$ timestamps $t$, $\mathcal{V}_{t}=\{v_{1}(t),\dots,v_{n_{t}}(t)\}$ is a set of $n_{t}$ nodes representing spatial entities at time $t$, $\mathcal{E}_{t}=\{e_{1}(t),\dots,e_{m_{t}}(t)\}$ is a set of $m_{t}$ edges representing the relationships between nodes at time $t$, $\mathcal{X}_{t}=\{x_{v_{1}}(t),\dots,x_{v_{n_{t}}}(t)\}\cup\{r_{e_{1}}(t),\dots,r_{e_{m_{t}}}(t)\}$ is a set of attributes associated with nodes and edges ($x_{v_{i}}$ and $r_{e_{i}}$ are respectively the attributes of node $v_{i}$ and edge $e_{i}$, and can be either scalars of vectors) at time $t$. Notice that the sets of nodes and edges can change over time, as well as the set of attributes $\mathcal{X}_{t}$.
Definition 4 (Adjacency matrix).
The adjacency matrix $\mathcal{A}$ of a graph $\mathcal{G}$ with $|\mathcal{V}|=n$ nodes is a $n\times n$ square matrix, with $\mathcal{A}_{ij}$ specifying the number of connections from node $v_{i}$ to node $v_{j}$ for $i,j=1,\dots,n$. Here it is assumed that a graph cannot have more than one edge between any two nodes, so $\mathcal{A}_{ij}$ can only be 0 or 1 depending on whether there is a connection or not.
Definition 5 (Degree matrix).
The degree matrix $\mathcal{D}$ of a graph with $|\mathcal{V}|=n$ nodes is a diagonal matrix whose entries are given by the degree of each node, i.e., the number of edges attached to the node. In formulas, its diagonal elements are given by
$\mathcal{D}_{ii}=\sum_{j}\mathcal{A}_{ij}\>.$ (1)
4.2 Overview of graph neural networks
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, a GNN model (not necessarily changing in time, i.e., spatio-temporal) aims to generate an embedding (i.e., a vector of real entries) for each node in the graph. An overall graph embedding consists of representing a graph into a possibly different-dimensional space while preserving its structural information.
The notion of GNN was initially proposed by Gori et al. in 2005 in Ref. [gori_aggiunto] and by Scarselli et al. in 2009 in Ref. [Scarselli2009] as an architecture to implement a function that maps a graph $\mathcal{G}$ with $n$ nodes into an $m$-dimensional Euclidean space. In Ref. [Scarselli2009] the target node’s embedding was learned by first propagating forward at fixed weights neighbor information within the graph in an iterative manner, until a stable fixed point is reached (whose existence is guaranteed under certain assumptions by Banach’s fixed point theorem), and then propagating errors on targets back across the graph. The calculation of node embeddings is in general quite complex, and over time many other methods have been developed. These methods are based on different architectures and sophisticated mechanisms. Based on these aggregating mechanisms, a taxonomy of GNN models with different classes has been defined. In this review, we adopt a taxonomy for the discussed spatio-temporal GNNs similar to the one proposed by Chen et al. in Ref. [survey_Lucheroni]. More details will be given in Subsec. 4.3.
4.3 Spatio-temporal GNNs and their taxonomy
The majority of research on spatio-temporal GNNs is focused on multivariate time series, as they can be naturally abstracted into spatio-temporal graphs, as depicted in Fig. 5. In this context, nodes correspond to different variables at a given time step, and edges represent relationships between variables.
Figure 5: Representation of a spatio-temporal graph. Here, each node has a dynamic feature, given by the points of the time series.
This spatio-temporal modeling assumes that the feature information of a node depends on both its own historical values and on the historical data from its neighboring nodes [survey_spatiotemporal]. Once the the graph structure is given, information is propagated within it until the output is obtained, as better explained in the following.
When it comes to the development of a GNN model, there are two primary approaches: one that handles spatial and temporal substructures in separate blocks, and another that integrates and processes these two substructures together [survey_spatiotemporal]. This is implemented by means of modules that are purely spatial, purely temporal, or hybrid combinations of the two. These modules are then organized into a series of blocks to create the final spatio-temporal GNN model.
In the taxonomy of spatio-temporal GNNs used in this review, the focus is exclusively on the description of the spatial module that propagates the information within the nodes at fixed time $t$ (not explicitly included in the following equations for simplicity), which is in the reviewed literature a fundamental aspect that differentiates one GNN model from another. Based on the mechanism at the basis of the information propagation in the graph at a given time step, spatio-temporal GNNs can be further classified as recurrent GNNs, convolutional GNNs, and attentional GNNs, as explained in the following. These classes are not mutually exclusive, as some models may fit into hybrid categories. In addition, the classes are not fully exhaustive, as there may be models that do not align with any of the existing categories. However, for the purpose of this review, the taxonomy discussed above is sufficient.
Recurrent GNNs. Recurrent graph neural network modeling mainly includes the results of early studies on GNNs, where the node embeddings are generated by iteratively propagating neighbor information until a stable fixed point is reached. The iterative equation for the calculation of the embedding (or hidden state) of node $u$ has the form
$h_{u}^{(k)}=f\left(x_{u},\{x_{v},r_{(u,v)},h_{v}^{(k-1)}\mid v\in\mathcal{N}(u)\}\right)\>,$ (2)
where $k$ is the index related to iteration, $f$ is a parametric trainable function, $x_{u}$ is the attribute associated to node $u$, $r_{(u,v)}$ is the weight of the edge $(u,v)$ between the nodes $u$ and $v$, $\mathcal{N}(v)$ denotes the neighborhood of node $v$, and the initial hidden states $h^{(0)}$ are usually randomly initialized. As detailed in Ref. [Scarselli2009], which introduced recurrent GNNs, to ensure the convergence of Eq. (2), it is necessary that the iterative function $f$ is a contraction mapping.
Convolutional GNNs. Convolutional graph neural networks are closely related to recurrent graph neural networks. They were at first introduced by Kipf and Welling in 2017 in Ref. [graph_convolution_kipf]. Unlike recurrent GNNs, which apply the same iterative contraction mapping $f$ until an equilibrium is reached, convolutional GNNs use different parameters at each updating step, by stacking multiple graph convolutional layers to extract the node embeddings [cit_rev3].
A basic form for the equation for the $k^{\text{th}}$ layer of a convolutional graph neural network with $n$ nodes is
$H^{(k)}=f_{k}\left(\tilde{\mathcal{A}}H^{(k-1)}\Theta_{k-1}\right)\>,\quad\tilde{\mathcal{A}}=\hat{\mathcal{D}}^{-\frac{1}{2}}\hat{\mathcal{A}}\hat{\mathcal{D}}^{-\frac{1}{2}}\>,\quad\hat{\mathcal{A}}=\mathcal{A}+I\>,\quad\hat{\mathcal{D}}_{ii}=\sum_{j}\hat{\mathcal{A}}_{ij}\>,\quad\quad H^{(0)}=X\>,$ (3)
where $f_{k}$ represents the convolutional operation performed at layer $k\in\{1,\dots,K\}$, $\mathcal{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix, $I\in\mathbb{R}^{n\times n}$ is the identity matrix (used to add the self-loops in the adjacency matrix $\mathcal{A}$), $\hat{\mathcal{D}}\in\mathbb{R}^{n\times n}$ is the diagonal degree matrix associated to $\hat{\mathcal{A}}$, $H^{(k)}\in\mathbb{R}^{n\times d_{k}}$ are the $d_{k}$-dimensional node embeddings $h^{(k)}_{v}$ of the entire graph produced by the $k^{\text{th}}$ layer, $X\in\mathbb{R}^{n\times d_{0}}$ is the collection of all node attributes $x_{v}$ of the entire graph, and $\Theta_{k}\in\mathbb{R}^{d_{k-1}\times d_{k}}$ is a trainable weight matrix for the $k^{\text{th}}$ layer [survey_Lucheroni]. The mechanism of node embedding generation from Eq. (3) is illustrated in Fig. 6(b), where it is put in contrast with the mechanism of a recurrent GNNs, shown in Fig. 6(a).
(a) Recurrent GNNs use the same graph recurrent layer in the iterative generation of the node embeddings. Here $\Theta$ represents the weights of the parametric trainable function $f$.
(b) Convolutional GNNs use different graph convolutional layers in the generation of the node embeddings.
Figure 6: Comparison of the node embedding generation mechanisms in a recurrent GNN (left panel) and a convolutional GNN (right panel).
Attentional GNNs. In attentional GNNs, the aggregation process uses the attention mechanism [attention] to combine node features. The equations to compute the node embedding $h_{u}$ of layer $k+1$ from the embeddings of layer $k$ for each node $u$ are
$z_{u}^{(k)}$ $=W_{k}h_{u}^{(k)}\>,$ (4a)
$\epsilon_{uv}^{(k)}$ $=\text{LeakyReLU}\left(\vec{a}^{\,(k)^{\top}}\left(z_{u}^{(k)}\,||\,z_{v}^{(k)}\right)\right)\>,$ (4b)
$\alpha_{uv}^{(k)}$ $=\frac{\exp\left(\epsilon_{uv}^{(k)}\right)}{\displaystyle\sum\limits_{w\in\mathcal{N}(u)}\exp\left(\epsilon_{uw}^{(k)}\right)}\>,$ (4c)
$h_{u}^{(k+1)}$ $=f\left(\sum_{v\in\mathcal{N}(u)}{\alpha^{(k)}_{uv}z^{(k)}_{v}}\right)\>.$ (4d)
Eq. (4a) represents a learnable linear transformation of the node embedding $h_{u}^{(k)}$ using the trainable weight matrix $W_{k}$ of layer $k$. Eq. (4b) computes a pairwise masked attention score between nodes $u$ and $v$ by concatenating (||||| |) the vectors $z_{u}^{(k)}$ and $z_{v}^{(k)}$, calculating a dot product with a learnable weight vector $\vec{a}^{\,(k)}$, and applying the LeakyReLU nonlinearity. The quantity $\epsilon_{uv}$ indicates the importance of the features of node $v$ to node $u$. The term ”masked” refers to the fact that $\epsilon_{uv}$ are only computed for neighboring nodes. Eq. (4c) normalizes the attention score through the softmax function in order to make the $\alpha_{uv}$ comparable across different nodes. Finally, Eq. (4d) computes the next layer embedding $h_{u}^{(k+1)}$ by aggregating the information in neighboring nodes, weighted by the attention scores [GAT]. Graph attention networks can also further use a multi-head attention mechanism to stabilize the learning process of self-attention. In practice, the node updating operation in graph attention networks is a generalization of the traditional averaging or max-pooling of neighboring nodes, which allows each node to compute a weighted average of its neighbors and identify the most relevant ones [Brody2021HowAA].
As anticipated, the taxonomy discussed so far focuses only on the propagation mechanisms of the spatial component. However, spatio-temporal GNNs typically consist of a stack of spatial, temporal, or hybrid modules, each one with distinct roles.
The spatial module is responsible for propagating information between nodes, thereby enabling the analysis of cross-sectional inter-dependencies between different variables. As previously explained in the description of the mechanisms underlying the taxonomy, the spatial module aggregates information starting from immediate neighboring nodes and extends it outward, thereby capturing also the influence of distant nodes. Noticeably, this spatial operation does not account in principle for temporal information. In contrast, the temporal module, often based on architectures such as LSTM, GRU, or self-attention, focuses on the evolution of data over time, independently from cross-sectional node interactions.
When these different modules are stacked together, spatial and temporal information end up to be used simultaneously, although they are processed in separate modules. This occurs because the computation of each node state is influenced by both spatial and temporal information. This combined approach allows the spatio-temporal GNN model to effectively capture at the same time both spatial relationships and temporal dynamics.
4.4 Determination of the graph structure
A crucial issue when dealing with spatio-temporal GNNs for time series problems is the determination of the graph structure, i.e. the connectivity of the nodes. Namely, some time series datasets come by nature with a pre-defined graph structure (e.g., road network), while some others do not. When a natural pre-defined graph structure is available, it helps the model to better capture the underlying dynamics of the system. When it is not directly available, it must be in some way defined by the user (based on domain knowledge or some metrics), determined by some algorithm (such as visibility graphs [visibility_Gori], [visibility_graph], [art_14]), or learned by the model itself.
Once the graph structure has been determined, it is necessary to define the weighted graph adjacency matrix, which is a generalization of the graph adjacency matrix whose entries are given by the edge weights of the graph. The edge weights can be defined a priori by the user, or again, learned by the model itself based on a pre-defined architecture. In the first case, the edge weights are assigned based on some pre-defined metrics or criteria chosen by the user (e.g., spatial distance between the sites or similarity measures). In the latter case, the weights are continuously adjusted throughout training as the model learns from the data.
Each section of the following thematic analysis examines which approach is prevalent.
5 Thematic analysis
This section explores the diverse applications of GNN models in different groups. As said, the selected journal papers are divided in groups according to the application domain. The goal is to provide an overview of the various empirical studies, the most common approaches, and the results obtained for different applications.
In each subsection of the following thematic analysis, a list of the benchmark models that appear in the selected journal papers will be provided. They will be classified as: mathematical and statistical models coming from Econometrics, traditional machine learning models, and GNN-based models. As for mathematical and statistical models, they generate forecasts by analyzing a limited amount of data, and they usually struggle with cross-sectional large-scale and temporal long-term dependencies. Traditional machine learning models use more complex architectures, mainly neural networks, but they are not necessarily able to capture intricate spatio-temporal dependencies within the data. GNN-based models instead, attempt to represent these complex dependencies using various techniques, resulting in better, though not perfect, accuracy. In addition, information on the most common datasets will be given, along with the best results reported in the literature. All results presented in the tables of this SLR are taken directly from the tables of the selected papers.
The groups of papers are presented in alphabetical order, with the exception of the ”Generic” applications group (which cannot be directly attributed to a specific case study) and the ”Other topics” group (focusing on specific problems of other disciplines). In subsections with a large number of papers and high homogeneity of the datasets, more precise comparisons of the results and in-depth analyses will be conducted.
5.1 Energy
The first group of papers is ”Energy” (with 13 out of 156 papers), which includes all the studies related to energy systems. In particular, it includes research on electricity load, energy consumption, and power generation forecasting. This is somehow connected to the ”Environment” subsection (Subsec. 5.2). Indeed, green energy sources are associated with environmental phenomena such as wind speed and solar radiation inflow. However, this subsection focuses on energy applications, whereas Subsec. 5.2 refers to the study of environmental data that may or may not be related to energy applications. Wind speed forecasting, for instance, is connected to wind power generation due to the cubic relationship between wind speed and wind power [art_29]. However, wind speed forecasting may also be of interest for other reasons, and thus it is included in the more general Subsec. 5.2.
5.1.1 Overview
Energy forecasting is crucial for electricity power grid stability and operational efficiency. Accurate load and generation forecasts ensure that supply meets real-time demand, preventing outages and instability, and maintaining a reliable supply of power. In addition, effective forecasting allows for optimal scheduling of power plants and energy resources, reducing operational costs and improving overall system efficiency. The majority of econometric and machine learning models tend to focus on time series from a single site or consider multiple sites without explicitly capturing the relationships between them. GNN models may be an innovative and effective approach to addressing energy related problems.
An overview of the selected papers reveals a noticeable interest in the application of GNNs in the energy field since 2022. The most frequently studied fields include forecasting of wind ([art_20], [art_24], [art_27]) and solar power ([art_19], [art_21], [art_22], [art_26]), which are renewable energy sources characterized by their intermittency. The two most popular journals among the selected papers are Applied Energy by Elsevier, and Frontiers in Energy Research by Frontiers Media SA. The majority of the corresponding authors are affiliated with institutions in China and in the United States.
5.1.2 Datasets
The study of the selected papers in the ”Energy” group suggests that there are no particularly common or widely recognized benchmark datasets. Indeed, researchers often focus on specific datasets, typically collected within the university campus or in the country where they are located. For the same reason, the links to these datasets are not always available. In Tab. LABEL:t:link_dataset_energy are listed the links to the public datasets mentioned in the papers.
Table 3: List of public datasets in the ”Energy” group and their corresponding links.
Dataset Used by Link
Danish smart heat meters [art_15] https://doi.org/10.5281/zenodo.6563114
DKA Solar Centre [art_19] https://dkasolarcentre.com.au/
IEEE bus systems [art_23] https://doi.org/10.1109/UPEC.2015.7339813 (source paper)
JERICHO-E-usage [art_15] https://doi.org/10.6084/m9.figshare.c.5245457.v1
NREL [art_20], [art_22], [art_24] https://www.nrel.gov/
OPSD [art_16] https://open-power-system-data.org/
PV Switzerland [art_21], [art_22] https://doi.org/10.1109/IJCNN48605.2020.9207573 (source paper)
UMass Smart* Dataset [art_17] https://traces.cs.umass.edu/index.php/Smart/Smart
Three of the selected papers ([art_20], [art_22], [art_24]) use data from the National Renewable Energy Laboratory (NREL), a national laboratory of the U.S. Department of Energy (https://www.nrel.gov/). Additional datasets were used for exogenous weather variables and satellite imagery (such as Himawari-8, available at https://www.data.jma.go.jp/mscweb/en/himawari89/space_segment/spsg_sample.html in Ref. [art_19]), but they are of course highly dependent on the location of the study. The data granularity ranges from 1 minute to 1 hour, and the forecasting horizon from tens of minutes to days. They depend on the dataset, the domain of application, and the purpose of the forecast. Most datasets require pre-processing, including missing data interpolation, normalization, and satellite image processing.
5.1.3 Proposed models
As for the taxonomy of the proposed GNN models, 8 out of 13 papers use a convolutional GNN, and the other 5 use an attentional GNN. Regarding the graph structure, the papers typically provide only a brief description. For example, it is rarely highlighted whether the graph structure is static or dynamic, and the way in which it is constructed is often described summarily. In some papers, the graph structure reflects the geographic location of the elements (e.g., wind or solar power plants, as in Refs. [art_24] and [art_26]), whereas in others it is based on the correlation or similarity between time series (as in Ref. [art_19]). In some studies, a pre-defined structure is imposed and the weighted adjacency matrix is learned directly from the model, often using the attention mechanism as in Ref. [art_22].
The most common loss function used to train the models is the MSE. Many papers mention the programming languages Python and Matlab, and few of them mention the Python libraries PyTorch [PyTorch] and TensorFlow [2] for the implementation of the GNN model. In this group, only Ref. [art_17] provides the link to the source code of the proposed residential load forecasting with multiple correlation temporal graph neural networks (RLF-MGNN) model (available at https://codeocean.com/capsule/9294192/tree/v1).
5.1.4 Benchmark models
Tab. LABEL:t:benchmark_models_energy lists the benchmark models used in the ”Energy” group by at least two papers included in this review.
Table 4: List of benchmark models in the ”Energy” group divided per category.
Category Model Used by
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_15], [art_16], [art_22], [art_25]
Classical machine learning methods Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_15], [art_19]
Convolutional neural network (CNN) [art_15], [art_18], [art_21], [art_22]
Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [art_15], [art_17]
Encoder-decorer architectures (ED) [art_15], [art_19], [art_21], [art_22]
Feed-forward neural network (FNN) [art_15], [art_18], [art_19], [art_25], [art_27]
Gated recurrent unit (GRU) [art_15], [art_25]
Informer [Informer] [art_15], [art_16]
$k$-nearest neighbors (KNN) [art_18], [art_24]
Long-short term memory network (LSTM) [art_15], [art_16], [art_19], [art_20], [art_23], [art_24]
Recurrent neural network (RNN) [art_16], [art_17], [art_19], [art_20]
Support vector regression (SVR) [art_21], [art_22], [art_24], [art_27]
Transformer [attention] [art_16], [art_19]
GNN methods Spatio-temporal autoregressive model (ST-AR) [ST-AR] [art_21], [art_22]
The majority of the papers only use classical machine learning benchmarks. Some papers use simple mathematical and statistical models, such as linear regression (that we incorporated into ARIMA) or historical average. Only few papers use other GNN-based models as benchmarks, and there are no common benchmarks among them. This is due to the fact that this field has not been extensively explored in the past, and only a very limited number of GNN models have been specifically developed for energy forecasting.
5.1.5 Results
In all the papers, the authors claim that the proposed GNN model outperforms the benchmarks. The most common error metrics are the MAE, the RMSE, and the Normalized Root Mean Square Error (NRMSE).
It is not possible to compare the accuracy of the discussed models because there are no common benchmark datasets. The only two papers that study the same dataset are written by the same authors (Refs. [art_21] and [art_22]), and their most recent paper includes a comparison with the models proposed in the first one.
5.2 Environment
The group ”Environment” includes a significant number of the selected papers (27 out of 156), which suggests a considerable research interest on the application of GNNs in this field. However, this group is highly fragmented, as it includes a large number of subdomains and applications, all related to the study of environmental data.
5.2.1 Overview
The group of environmental studies is broad and includes several disciplines such as physics, climatology, oceanography, and atmospheric science. The breadth of the group is reflected in the number of case studies presented in the selected papers, which cover air quality prediction (in terms of PM2.5 concentration), sea temperature, wind speed, and other applications. Forecasting these quantities can be useful for several reasons. For instance, predicting wind speed or rainfall can assist in anticipating some extreme weather phenomena, and is also related to the prediction of green energy outputs. For example, in Refs. [art_29] and [art_34] the authors estimate the wind power forecasting error obtained by using wind speed forecasting models. As another example, predicting sea surface temperature can be useful for weather forecasting, fishing directions, and disaster warnings [art_28]. Moreover, monitoring and forecasting environmental data are crucial for evaluating the impact of human activities on the environment and tracking the progression of climate change.
The methods for forecasting environmental data are typically categorized as numerical models, statistical models, and machine learning methods. Numerical methods are based on atmospheric models, and are designed to quantify the interaction of different atmospheric variables. The accuracy of these models depends on the availability of data, and the mathematical simulations to get the forecasts may take days or even weeks, thus limiting the ability to make good short-term forecasts. Linear statistical methods leverage long-term dependencies in the recorded data by using regression-like models. However, they cannot capture non-linear relationships in the dataset, which limits their power. Machine learning models, on the other hand, are often a generalization of linear statistical methods that can capture more complex and non-linear relationships in the data.
An overview of the selected papers reveals that there has been a rapid increase in interest in the field among the GNN research community, and more than 20 of the selected papers were published in 2023. The most investigated subjects are sea temperature ([art_28], [art_33], [art_35], [art_36], [art_39], [art_41], [art_47]), wind speed ([art_29], [art_31], [art_34], [art_40], [art_44], [art_46]), and PM2.5 concentration forecasting ([art_2], [art_3], [art_4], [art_6], [art_7], [art_8]). All the selected papers are published in different journals, with the exception of the Engineering Applications of Artificial Intelligence by Elsevier, which has published two of the papers. The majority of the corresponding authors are affiliated with institutions in China.
5.2.2 Datasets
The analysis of the selected papers shows that in this group of papers there is no reference benchmark dataset. The publicly available datasets cited in the papers are listed in Tab. LABEL:t:link_dataset_environment.
Table 5: List of public datasets in the ”Environment” group and their corresponding links.
Dataset Used by Link
AgroData-10 [art_28] https://www.aoml.noaa.gov/phod/argo
CCMP wind data [art_31], [art_40], [art_44] http://data.remss.com
China Environmental Monitoring Station PM2.5 [art_2], [art_8] https://www.cnemc.cn/
China National Urban Air Quality [art_1], [art_5] https://github.com/Friger/GAGNN
Copernicus 3D thermohaline data [art_35] https://resources.marine.copernicus.eu/products
ECMWF East China Sea [art_45] https://github.com/Boltzxuann/GIPMN/tree/dataset
ECMWF South and East China Sea [art_32] http://apps.ecmwf.int/datasets/data/interim-full-daily
KnowAir [art_6] https://doi.org/10.1145/3397536.3422208 (source paper)
Korean Peninsula sea data [art_33] https://www.nifs.go.kr/risa/main.risa
NOAA sea datasets [art_28], [art_36], [art_39], [art_46], [art_47] https://www.ndbc.noaa.gov/
Norwegian offshore wind [art_29], [art_34] https://frost.met.no/index.html
Taiwan air quality dataset [art_3] https://data.gov.tw/en
UrbanAir [art_6] https://doi.org/10.1016/j.chemosphere.2018.12.128 (source paper)
US EPA PM2.5 [art_1], [art_4] https://aqs.epa.gov/aqsweb/airdata/download_files.html
Five of the selected papers ([art_28], [art_36], [art_39], [art_46], [art_47]) use National Oceanic and Atmospheric Administration (NOAA) sea datasets (https://www.noaa.gov/) from the U.S. Department of Commerce (US DOC). The US DOC website permits users to download daily, weekly and monthly mean optimum interpolation sea surface temperature (OISST) data from September 1981 for many geographical locations. In over half of the papers, exogenous variables are incorporated into the model. These include weather data (e.g., humidity, rainfall, pressure, temperature, wind) and, in the case of PM2.5 concentration forecasting, other pollutants (e.g., CO, NO2, O3, PM10, SO2).
The data granularity ranges from 1 minute to 1 month, depending on the case study, and the forecasting horizon from 10 minutes to 240 days. Most datasets require some pre-processing, including interpolation of missing data, removal of outliers and normalization.
5.2.3 Proposed models
Regarding the taxonomy of the proposed GNN models, the majority of papers present convolutional GNN (17 out of 27 papers), followed by 6 attentional GNNs and 3 hybrid architectures. The description in one paper (Ref. [art_5]), however, does not specify the classification of its model within the proposed taxonomy. As for the graph structure, not all of the papers describe it with sufficient precision, and some of them are vague about the graph construction and its static or dynamic nature. In almost half of the papers, the edge weights are learned directly from the model, often using the graph attention mechanism. When the graph structure is defined manually, the most popular criterion is the spatial distance between the sites ($d$), even in its exponential ($e^{-d}$) or inverse proportion ($1/d$) formulation. Finally, there are a few papers where more than one graph structure is fused together in order to capture more complex dynamics, as in Ref. [art_2].
Not all papers specify the loss function used to train the models. Among those that do, MAE and MSE are used in roughly equal proportions. Many papers mention the programming language Python and the library PyTorch for the implementation of the models. Few papers also provide a link to the source code for the proposed model, as shown in Tab. LABEL:t:link_codici_environment.
Table 6: List of source codes of the ”Environment” models in the review.
Model Link
Group-aware graph neural network (GAGNN) [art_1] https://github.com/Friger/GAGNN
Gridded information propagation and mixing network (GIPMN) [art_45] https://github.com/Boltzxuann/GIPMN/tree/dataset
Hierarchical graph recurrent network (HiGRN) [art_28] https://github.com/Neoyanghc/HiGRN
Spatio-temporal FFTransformer (ST-FFTransformer) [art_29] https://github.com/LarsBentsen/FFTransformer
5.2.4 Benchmark models
Tab. LABEL:t:benchmark_models_environment presents the benchmark models employed in the ”Environment” group by at least two articles examined in this review.
Table 7: List of benchmark models in the ”Environment” group divided per category.
Category Model Used by
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_28], [art_33], [art_34], [art_38]
Historical average (HA) [art_3], [art_6], [art_28], [art_38]
Naive [art_29], [art_34]
Vector autoregression (VAR) [art_28], [art_34], [art_38]
Classical machine learning methods Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_33], [art_46]
Combined FC-LSTM and convolution neural network (CFCC-LSTM) [CFCC-LSTM] [art_35], [art_39]
Hybrid convolutional neural network and gated recurrent unit (CGRU) [CGRU] [art_31], [art_40], [art_44]
Convolutional neural network (CNN) [art_30], [art_46]
Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [art_4], [art_41]
Convolutional long-short term memory network (ConvLSTM) [ConvLSTM] [art_28], [art_41]
Feed-forward neural network (FNN) [art_2], [art_7], [art_38]
Gated recurrent unit encoder–decoder (GED) [GED] [art_36], [art_39], [art_47]
Gated recurrent unit (GRU) [art_2], [art_6], [art_7], [art_8], [art_38], [art_43], [art_46]
Informer [Informer] [art_30], [art_47]
Long-short term memory network (LSTM) [art_1], [art_2], [art_3], [art_4], [art_5], [art_6], [art_7], [art_8], [art_28], [art_30], [art_32], [art_33], [art_35], [art_36], [art_38], [art_39], [art_43], [art_46], [art_47]
Predictive deep convolutional neural network (PDCNN) [PDCNN] [art_31], [art_40]
Support vector regression (SVR) [art_3], [art_6], [art_35], [art_36], [art_38]
Temporal convolutional network (TCN) [TCN] [art_6], [art_30]
Temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM] [art_6], [art_31], [art_32], [art_35], [art_40], [art_44]
Transformer [attention] [art_30], [art_33]
Hybrid wavelet random forest and deep belief network (Wavelet-DBN-RF) [art_31], [art_40], [art_44]
Extreme gradient boosting (XGBoost) [art_1], [art_2], [art_5], [art_46]
GNN methods Deeper graph convolutional network (DeeperGCN) [DeeperGCN] [art_1], [art_5]
Graph convolutional neural network and long-short term memory (GC-LSTM) [GC-LSTM] [art_1], [art_2], [art_5], [art_6], [art_7]
Graph convolutional network (GCN) [art_38], [art_41]
Graph WaveNet (GWN) [GWN] [art_1], [art_6], [art_44], [art_47]
Hierarchical graph convolution network (HGCN) [HGCN] [art_1], [art_28]
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_6], [art_32], [art_44]
Multi long-short term memory network (Multi-LSTM) [Multi-LSTM] [art_31], [art_40], [art_44]
Superposition graph neural network (SGNN) [art_24] [art_31], [art_40]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_3], [art_6], [art_28], [art_35], [art_47]
Spatio-temporal correlation graph neural network (STGN) [art_40] [art_31], [art_44]
Spatio-Temporal U-Network (ST-UNet) [ST-UNet] [art_1], [art_5]
Temporal graph convolutional network (T-GCN) [T-GCN] [art_6], [art_30], [art_35]
The majority of the benchmarks used fall within the category of classical machine learning, although there are several GNN benchmarks as well. The two most common benchmark models are LSTM and GRU. These models are widely used by environmental data forecasters due to their ability to capture temporal dependencies, and they are considered a reliable benchmark in the field. As for GNN models, the most common benckmarks are the graph convolutional neural network and long-short term memory (GC-LSTM) [GC-LSTM] (which appears only in quality air forecasting problems) and spatio-temporal graph convolutional network (STGCN) [STGCN]. The GC-LSTM is a model that integrates graph convolutional networks and long-short-term memory networks, proposed by Qi et al. in 2019 to forecast the spatio-temporal variation of PM2.5 concentration. The STGCN model instead was proposed in the traffic domain by Yu et al. in 2018. This model is composed of two spatio-temporal convolutional blocks and a fully-connected output layer. Each convolutional block contains two temporal gated convolution layers and one spatial graph convolution layer in the middle. The code for the STGCN model is available at https://github.com/VeritasYin/STGCN_IJCAI-18. Finally, it is important to point out that the multivariate time series forecasting with graph neural network (MTGNN) model proposed by Wu et al. in 2020 in Ref. [MTGNN] is often the second-best performing model in the papers using it, thus it deserves more attention and further investigation in future research.
5.2.5 Results
In each paper, the proposed model significantly improves on the benchmarks. The accuracy of the models is measured in terms of MAE, RMSE, and MAPE. Among the papers, there is a high diversity in the choice of the proposed datasets. However, there are two sets of papers using the same dataset, whose results can be compared.
The first dataset is the CCMP wind speed dataset (available at http://data.remss.com), which contains 3,942 geographical grid points from the east and southeast sea areas of China. The three papers using it ([art_31], [art_40], [art_44]) select 120 points from the area and use data records from January 2010 to April 2019 at 6 hours intervals, with a total of 13,624 samples at each node, for 10 different forecasting horizons. All papers present results for the complete set of 120 nodes and a subset of nodes. However, since their results are similar, only the average results for all the 120 nodes are reported in Tab. LABEL:t:risultati_ccmp for the sake of simplicity.
Table 8: Comparison of the average accuracy of the different models on the CCMP wind speed dataset for time horizons from 6 hours to 7 days ahead, expressed in terms of MAE and RMSE. The smallest errors are underlined. Numbers from the papers.
Model Metrics 6 h 12 h 18 h 1 d 2 d 3 d 4 d 5 d 6 d 7 d
CGRU MAE 1.658 1.795 1.937 2.038 2.290 2.401 2.478 2.525 2.556 2.594
RMSE 2.100 2.277 2.461 2.594 2.911 3.042 3.141 3.198 3.245 3.285
DAGLN [art_31] MAE 1.068 1.219 1.346 1.443 1.716 1.852 1.935 1.984 2.019 2.067
RMSE 1.423 1.625 1.793 1.917 2.255 2.428 2.512 2.570 2.607 2.656
DASTGN [art_44] MAE 1.039 1.184 1.297 1.387 1.662 1.806 1.900 1.945 1.995 2.028
RMSE 1.384 1.572 1.721 1.829 2.170 2.338 2.449 2.498 2.554 2.600
DLinear MAE 1.145 1.306 1.432 1.528 1.798 1.937 2.022 2.080 2.124 2.155
RMSE 1.528 1.745 1.913 2.037 2.366 2.512 2.604 2.666 2.713 2.748
EMD-SVRCKH MAE 1.111 1.311 1.432 1.526 1.788 1.919 1.998 2.051 2.090 2.119
RMSE 1.489 1.755 1.909 2.027 2.344 2.495 2.587 2.650 2.695 2.730
Graph WaveNet MAE 1.101 1.265 1.386 1.487 1.762 1.901 1.951 1.992 2.089 2.124
RMSE 1.467 1.685 1.852 1.982 2.304 2.466 2.554 2.582 2.684 2.717
Historical Inertia MAE 1.139 1.301 1.405 1.511 1.773 1.912 2.007 2.081 2.109 2.114
RMSE 1.494 1.704 1.843 1.977 2.304 2.467 2.571 2.637 2.678 2.687
MTGNN MAE 1.091 1.237 1.351 1.454 1.724 1.864 1.935 1.990 2.036 2.059
RMSE 1.454 1.650 1.801 1.936 2.267 2.436 2.515 2.579 2.632 2.657
Multi LSTMs MAE 1.312 1.645 1.914 2.129 2.610 2.748 2.793 2.818 2.842 2.867
RMSE 1.663 2.079 2.415 2.679 3.257 3.420 3.477 3.510 3.540 3.573
PDCNN MAE 2.025 2.026 2.385 2.551 2.959 3.011 3.012 2.971 2.918 2.877
RMSE 2.561 2.562 3.033 3.243 3.750 3.824 3.833 3.789 3.727 3.675
SGNN MAE 2.542 2.562 2.573 2.578 2.583 2.618 2.666 2.695 2.712 2.716
RMSE 3.222 3.246 3.264 3.270 3.275 3.328 3.392 3.431 3.453 3.458
STGN [art_40] MAE 1.084 1.281 1.407 1.523 1.809 1.942 2.046 2.073 2.137 2.189
RMSE 1.430 1.697 1.863 2.013 2.358 2.514 2.631 2.649 2.717 2.771
TPA-LSTM MAE 1.203 1.552 1.803 1.984 2.432 2.667 2.836 2.959 3.027 3.045
RMSE 1.562 1.997 2.319 2.544 3.082 3.383 3.598 3.745 3.825 3.842
Wavelet-DBN-RF MAE 1.451 1.597 1.749 1.836 2.087 2.162 2.200 2.217 2.232 2.243
RMSE 1.855 2.038 2.226 2.328 2.617 2.703 2.744 2.766 2.788 2.798
For all the time horizons, the most recent dynamic adaptive spatio-temporal graph neural network (DASTGN) model [art_44] has the highest accuracy. Other two well performing models are the data-based adaptive graph learning network (DAGLN) [art_31] and the multivariate time series forecasting with graph neural network (MTGNN) [MTGNN].
The second dataset shared by two papers ([art_1], [art_5]) is the one of China National Urban Air Quality (available at https://github.com/Friger/GAGNN). It contains air quality index data (AQI) of 209 cities, collected from January 2017 to April 2019 at 1 hour granularity, for a total of 20,370 samples. Both the papers also integrate exogenous weather variables (humidity, wind direction, wind speed, rainfall, air pressure, temperature, and visibility) collected from Envicloud (http://www.envicloud.cn/). The average accuracy of the models used in the two papers is compared in Tab. LABEL:t:risultati_chinese_aqi.
Table 9: Comparison of the average accuracy of the different models on the Chinese AQI dataset for time horizons from 1 to 6 hours ahead, expressed in terms of MAE and RMSE. The smallest errors are underlined. Numbers from the papers.
Model Metrics 1h 2h 3h 4h 5h 6h
AirFormer MAE 5.95 9.31 11.87 13.87 15.52 16.95
RMSE 11.49 17.23 21.32 24.31 26.72 28.71
DeeperGCN MAE 6.54 9.74 11.77 13.40 15.29 16.41
RMSE 13.67 18.93 21.14 23.83 26.25 28.02
FGA MAE 5.87 9.14 11.71 13.75 15.42 16.80
RMSE 11.36 17.01 21.05 24.09 26.55 28.52
GAGNN [art_1] MAE 5.56 8.59 10.80 12.52 13.91 15.10
RMSE 10.81 16.17 19.84 22.51 24.63 26.37
GC-LSTM MAE 5.95 9.16 11.58 13.46 15.00 16.31
RMSE 11.91 16.98 20.82 23.69 25.97 27.82
GWNet MAE 5.76 9.64 12.79 15.30 17.28 18.81
RMSE 11.27 17.57 22.31 25.75 28.52 30.48
HGCN MAE 5.70 9.09 11.73 13.84 15.55 16.95
RMSE 11.18 17.09 21.33 24.52 26.99 28.94
HighAir MAE 5.50 8.52 10.81 12.50 14.00 15.09
RMSE 10.80 16.10 19.85 22.70 24.91 26.40
INNGNN [art_5] MAE 5.48 8.49 10.67 12.34 13.72 14.91
RMSE 10.70 16.03 19.66 22.29 24.37 26.11
LSTM MAE 6.50 10.26 13.18 15.52 17.40 18.91
RMSE 13.85 19.26 23.52 26.83 29.46 31.55
MegaCRN MAE 5.38 8.76 10.80 12.73 14.46 16.03
RMSE 10.64 16.46 19.92 22.82 25.45 27.60
SHARE MAE 5.84 9.07 11.49 13.35 14.74 15.79
RMSE 11.27 16.84 20.77 23.60 25.80 27.38
ST-UNet MAE 5.95 9.30 11.58 13.38 14.82 16.02
RMSE 11.74 18.01 21.34 23.90 25.94 27.64
XGBoost MAE 6.85 10.89 13.99 16.27 18.14 19.56
RMSE 14.25 19.80 24.72 28.14 30.63 33.44
The hybrid interpretable neural network and graph neural network model (INNGNN) [art_5] is the most accurate model for all time horizons except 1 hour, where it is beaten by the meta-graph convolutional recurrent network model (MegaCRN) [MegaCRN]. The authors attribute this result to the fact that in the short term the AQI is strongly influenced by nearby cities, and the MegaCRN model is more focused on local features.
5.3 Finance
”Finance” is another emerging field of study, though not yet widely explored, with 9 out of 156 selected papers. The difficulty of predicting financial data is a well-established fact. This is due to the inherently complex nature of markets, the influence of geopolitical events on them, and the unpredictable behavior of humans which is often irrational and difficult to predict. This complexity makes financial time series highly volatile, and accurate forecasts and data modeling are essential to develop effective trading strategies. GNNs are taken into consideration to model the inter-dependencies between variables together with series dynamics.
5.3.1 Overview
The ”Finance” group is relatively narrow, in part due to the limited number of published papers. The majority of studies are concentrated on stock price prediction ([art_59], [art_60], [art_62], [art_63], [art_65]), with the aim of forecasting the future trend of the price of a company. Sometimes the stock prediction problem is seen as a classification task. Indeed, due to the stochastic nature and volatility of financial markets, predicting price movements (”up”, ”neutral”, ”down”) can be more feasible than accurately forecasting specific prices. For example, assuming a daily granularity, a rise (”up”) can be considered to occur when the next trading day’s closing price ($p_{t}$) is at least 0.55% higher than the previous day’s closing price ($p_{t-1}$), whereas if $p_{t}$ is more than 0.50% lower than $p_{t-1}$, the price movement can be classified as a ”down”. Otherwise, it is classified as ”neutral”. All the selected papers addressing stock prices approach the problem from a classification perspective. Another distinctive aspect in this group of papers is that researchers usually also evaluate how their predictions affect the stability of investment returns over a specified period. This is commonly done through trading simulations in the test set, evaluated with metrics such as the Sharpe ratio and overall returns. These simulations require modeling rather than forecasting.
An examination of the selected papers indicates an interest in the application of GNNs in Finance from 2022. Two of the papers were published in IEEE Access by IEEE, while the remaining ones appeared in various other journals. Most of the corresponding authors are affiliated with academic institutions in China.
5.3.2 Datasets
The number of datasets in this group is limited. The links to the public datasets referenced in the selected papers are listed in Tab. LABEL:t:link_dataset_finance.
Table 10: List of public datasets in the ”Finance” group and their corresponding links.
Dataset Used by Link
China A-Shares [art_63], [art_65], [art_66] https://finance.sina.com.cn/stock
CSI 100 index [art_62] https://tushare.pro/
CSI 300 index [art_59], [art_62] https://global.csmar.com/
S&P 100 index [art_62] https://finance.yahoo.com/
S&P 500 index [art_59] https://finance.yahoo.com/
Russell 1K index [art_62] https://finance.yahoo.com/
Russell 3K index [art_62] https://finance.yahoo.com/
All the listed datasets represent collections of stock indices. The most common datasets include the China Securities Index (CSI), the 100 and 500 Standard and Poor’s (S&P) indices, and the 100 and 300 China A-Shares indices. Many studies also include exogenous data, such as other market data, events, and news. As for the news, they can be processed by some language model like BERT (as in Refs. [art_59] and [art_66]). Most of the data have a granularity of 1 day (which is typical of financial data), and the forecasting horizon is also generally 1 day. Only one of the selected papers (Ref. [art_62]) analyses also longer time horizons for stock forecasting. As for pre-processing, half of the selected papers use normalization techniques, including min-max normalization, Z-score normalization, and logarithmic normalization.
5.3.3 Proposed models
In terms of the taxonomy of the proposed GNN models, 5 out of 9 papers utilize attentional GNNs, while the other 4 papers employ convolutional GNNs. As for the graph structure, some papers use models which learn the graph structure themselves ([art_62], [art_64]), while some others defines it a priori based on the relationships between companies and stocks ([art_63], [art_67], [art_68]).
Since the majority of papers deal with a stock movement classification problem, the most common loss function is the cross entropy. The majority of the papers explicitly indicate that Python is the programming language utilized, with nearly half employing PyTorch and a few utilizing TensorFlow. Only Ref. [art_63] provides a link to the source code of the proposed knowledge graph and graph convolution neural network (KG-GCN) model, available at https://github.com/Gjl12321/KG_GCN_Stock_Price_Trend_Prediction_System.
5.3.4 Benchmark models
Tab. LABEL:t:benchmark_models_finance displays the benchmark models utilized in the ”Finance” group by a minimum of two of the selected papers.
Table 11: List of benchmark models in the ”Finance” group divided per category.
Category Model Used by
Mathematical and statistical methods Mean reversion (MR) [art_63], [art_65]
Classical machine learning methods Convolutional neural network (CNN) [art_59], [art_64]
Dual-stage attention-based recurrent neural network (DARNN) [art_63], [art_65]
Feed-forward neural network (FNN) [art_59], [art_68]
Gated recurrent unit (GRU) [art_60], [art_68]
Long-short term memory network (LSTM) [art_59], [art_62], [art_63], [art_64], [art_65], [art_67], [art_68]
Support vector machine (SVM) [art_64], [art_68]
GNN methods Graph convolutional network (GCN) [art_59], [art_60], [art_62], [art_64], [art_68]
Financial Graph Attention Network (FinGAT) [FinGAT] [art_60], [art_62]
Hierarchical graph attention network (HATS) [HATS] [art_62], [art_63], [art_65]
Spatiotemporal hypergraph convolution network (STHGCN) [STHGCN] [art_63], [art_65]
It is evident that there is no consensus regarding the most appropriate benchmark models, especially in the context of GNNs. The most common benchmark is the LSTM, which is considered an effective approach for processing long term information. The most commonly used GNN model is a vanilla graph convolutional network (GCN), followed by a more complex hierarchical graph attention network (HATS) proposed by Raehyun et al. in 2019 [HATS]. The HATS model has been proposed for predicting stock movements using relational data. It aggregates information from different types of relationships among data and incorporates this information into each stock representation. The source code for the HATS model is available at https://github.com/dmis-lab/hats.
5.3.5 Results
The accuracy of the proposed classification models is primarily evaluated in terms of accuracy, precision and F1 score. The only two papers with comparable results are Refs. [art_63] and [art_65], which study 758 frequently traded stocks collected from the A-share market in China from January 2013 to December 2019 (available at https://finance.sina.com.cn/stock). These two papers compare the results obtained by using 5, 10 and 20 trading days as records. For sake of simplicity, a comparison of the accuracy of the models in the two papers with 20 days only is provided in Tab. LABEL:t:risultati_a-share.
Table 12: Comparison of the average accuracy of the different models on the A-share market in China with 20 training days as records, expressed in terms of accuracy, precision, recall, and F1 score. The best results are underlined.
Model Accuracy Precision Recall F1 score
DARNN 38.41% 37.99% 39.24% 38.60%
GCN+LSTM 37.30% 39.28% 34.16% 36.54%
HATS 38.85% 38.70% 35.06% 36.78%
HGTAN 40.02% 41.77% 39.03% 40.32%
KG-GCN [art_63] 51.38% 65.72% 31.45% 39.27%
LSTM 35.03% 36.43% 34.23% 35.20%
MOM 35.73% 35.19% 32.82% 33.96%
MONEY [art_65] 39.90% 43.92% 40.61% 42.20%
MR 35.32% 38.03% 33.60% 35.68%
SFM 34.54% 26.93% 33.32% 29.49%
STHGCN 38.45% 37.22% 32.82% 34.87%
TGC 37.81% 36.96% 34.49% 35.67%
The two best models are the knowledge graph and graph convolution neural network (KG-GCN) [art_63] and the stock price movement prediction via convolutional network with adversarial hypergraph model (MONEY) [art_65], depending on the metrics considered.
5.4 Health
The fourth group of papers is ”Health”, which includes 8 out of 156 publications. These papers focus on critical aspects of health monitoring, disease modeling, and diagnostic tools, with particular emphasis on the spread and diagnosis of disease. All the papers working with functional magnetic resonance imaging (fMRI) data were excluded from the analysis, as the dataset was generated as image sequences rather than time series. This key difference in data structure makes them less relevant to the spatio-temporal focus of this review.
5.4.1 Overview
Applications of GNNs in the ”Health” group have been studied starting from 2021. The reason of this spread may be attributed to two key factors. The first reason is the growing presence of sensors in hospitals, which enable the collection of multivariate time series signals which can be used for the timely diagnosis of chronic diseases. The second reason is the spread of the COVID-19 pandemic, which has motivated researchers to investigate and forecast the spread of the virus. Epidemic forecasting is the most common topic among the selected papers. The prediction of epidemics presents two significant challenges. First, their seasonality can vary greatly over time, and they can manifest with different intensities and durations. Second, occasional pandemics can disrupt established patterns even for years. These complexities make the forecasting task more difficult, and machine learning and GNN models have been studied to address this issue. A common reference model for epidemic prediction is the SIR model [modello_SIR], an acronym referring to individuals which stands for susceptible (S), infectious (I), and recovered (R). The SIR model divides the population into individuals who are susceptible to the disease, those who are infected and can spread it, and those who have recovered and are now immune, and uses differential equations to describe how the disease spreads. In general, some papers use it as benchmark (e.g., Ref. [art_49]).
All the selected papers are published in different journals, but primarily in the field of computer science. The majority of the corresponding authors are affiliated with academic institutions based in China.
5.4.2 Datasets
Tab. LABEL:t:link_dataset_health provides a list of the public datasets utilized in the selected papers, along with links for accessing them.
Table 13: List of public datasets in the ”Health” group and their corresponding links.
Dataset Used by Link
Bonn dataset [art_14] http://dx.doi.org/10.1103/PhysRevE.64.061907 (source paper)
Brazilian COVID-19 data [art_52] https://github.com/lcaldeira/GrafoBrasilCovid
COVID-19 data [art_49] https://www.2019ncov.chinacdc.cn/
COVID-19 vaccinations [art_50] https://ourworldindata.org/us-states-vaccinations
eICU [art_189] https://doi.org/10.1038/sdata.2018.178 (source paper)
MIMIC-IV [art_189] https://physionet.org/content/mimiciv/2.0/
Nytimes COVID-19 data [art_50] https://github.com/nytimes/covid-19-data
Parkinson’s Progression Markers Initiative (PPMI) [art_92] http://www.ppmi-info.org/
Parkinson Speech data set (PS) [art_92] https://doi.org/10.1109/jbhi.2013.2245674 (source paper)
Spanish COVID-19 data [art_51] https://cnecovid.isciii.es/
Spikes and slow waves (SSW) dataset [art_14] http://dx.doi.org/10.1109/ICDM50108.2020.00067 (source paper)
US influenza [art_48] https://github.com/aI-area/DVGSN/tree/main/dataset
None of the papers in this section use the same dataset, so it is not possible to compare the results of the proposed models. Most of the datasets are related to the COVID-19 effects and have a daily granularity. In addition, some papers ([art_49], [art_52], [art_92], [art_189]) include exogenous variables that can be either context-related (e.g., geography, population, health conditions, vaccinations, economic factors), or specific to a single patient (including personal information such as age, gender, and medical conditions).
5.4.3 Proposed models
As for the taxonomy of the proposed models, 5 out of 8 papers use a convolutional GNN architecture. Two papers use a pure attentional model, while another paper presents a hybrid convolutional-attentional architecture. Different approaches are used to define the graph structure. Some papers base their graphs on the geographical distribution of the data (as in Refs. [art_49], [art_52] and [art_92]), and others use models that learn the graph structure directly, as in Refs. [art_48] and [art_50]. There is also a paper that uses a variant of the visibility algorithm [art_14]. In almost all the cases, the graph is static.
In terms of programming languages, this community uses only Python with PyTorch. As for loss functions, MSE is typically used for prediction tasks, and cross entropy is used for classification tasks. Tab. LABEL:t:link_codici_health shows the links to the related source codes.
Table 14: List of source codes of the ”Health” models in the review.
Model Link
Dynamic virtual graph significance networks (DVGSN) [art_48] https://github.com/aI-area/DVGSN
Deep learning of contagion dynamics on complex networks [art_51] https://doi.org/10.5281/zenodo.4974521
Sparse spectra graph convolutional network (SSGCNet) [art_14] https://github.com/anonymous2020-source-code/WNFG-SSGCNet-ADMM
Spatial-Temporal Graph Convolutional Network (STGCN) [art_52] https://github.com/lcaldeira/CovidPandemicForecasting
Time-aware Context-Gated Graph Attention Network (T-ContextGGAN) [art_189] https://github.com/OwlCitizen/TContext-GGAN
5.4.4 Benchmark models
Tab. LABEL:t:benchmark_models_health lists the benchmark models in the ”Health” group that are discussed in at least two papers.
Table 15: List of benchmark models in the ”Health” group divided per category.
Category Model Used by
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_48], [art_49], [art_52]
Classical machine learning methods Feed-forward neural network (FNN) [art_14], [art_48]
Long-short term memory network (LSTM) [art_49], [art_50]
There are no particularly relevant benchmark models, and none fall within the category of GNNs. This is because GNNs have not been widely studied in the context of health.
5.4.5 Results
In each paper, the authors claim that the proposed GNN-based model outperforms the benchmarks. The most common error metrics are the MAE and the MSE for forecasting problems and accuracy, sensitivity and specificity for classification problems. A comparison of the accuracy of the different models is not possible due to the lack of common benchmark datasets.
5.5 Mobility
The topic that includes the largest number of papers identified by the query (39 out of 156) is ”Mobility”, which regards the movement of people. It includes urban traffic, air travel, and bicycle demand, among other applications. Given the large number of articles in this group, a more detailed comparison of benchmarks, models, and results can be provided.
5.5.1 Overview
Early traffic prediction is essential for improving the efficiency of transportation systems, helping drivers plan their trips more effectively, and preventing urban congestion. The advent of smart cities infrastructures and transportation systems has contributed to the collection of rich data from road sensors that can be used for this purpose. However, making accurate traffic forecasts is challenging due to constantly changing traffic patterns over time and space, as well as the influence of external factors such as weather conditions and special events. Unlike classical statistical and machine learning models, GNNs can be particularly effective for traffic forecasting due to their ability to model complex relationships in spatial and temporal domain. The increasing number of papers published each year in this group reflects both the growing research interest in this field and the growing confidence in GNN models. The most commonly investigated areas regard urban traffic ([art_93], [art_95], [art_97], [art_98], [art_99], [art_100], [art_103], [art_105], [art_106], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_122], [art_123], [art_124], [art_125], [art_126], [3], [art_128], [art_130], [art_131], [art_132], [art_135]), especially urban traffic flow and urban traffic speed. Among the journals where the selected papers were published, Applied Intelligence by Springer is the most common, with 5 out of 39 papers. This is followed by ACM Transactions on Intelligent Systems and Technology and ACM Transactions on Knowledge Discovery from Data, both published by the Association for Computing Machinery, each with 4 out of 39 papers. Most of the corresponding authors are affiliated with Chinese institutions.
5.5.2 Datasets
Tab. LABEL:t:link_dataset_traffico provides a list of the public datasets mentioned in the selected papers, along with the links to access them.
Table 16: List of public datasets in the ”Mobility” group and their corresponding links.
Dataset Used by Link
CAAC flights [art_129] http://www.caac.gov.cn/index.html
CD-HW [art_126] https://doi.org/10.1109/ACCESS.2020.3027375 (source paper)
City Bike NYC [art_94], [art_114] https://citibikenyc.com/system-data
DiDi Beijing [art_102] https://github.com/didi/TrafficIndex
DiDi Chengdu [art_93], [art_98] https://gaia.didichuxing.com/
Huai’an Bus Transaction [art_101] https://www.pandabus.cn/panda/panda-minibus.html
Los-loop [art_103] https://github.com/lehaifeng/T-GCN/tree/master/data
METR-LA [art_93], [art_99], [art_100], [art_109], [art_110], [art_113], [art_116], [art_120], [art_122], [art_126], [art_128] https://paperswithcode.com/dataset/metr-la
Milan network [art_123] https://github.com/arunasubbiah/milan-telecom-data-modeling
NE-BJ [art_100] https://github.com/tsinghua-fib-lab/Traffic-Benchmark
NYC-Taxi [art_96] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
PeMS03 [art_108], [art_113], [art_115], [art_117], [art_125], [art_130] https://www.kaggle.com/datasets/elmahy/pems-dataset
PeMS04 (PeMSD4) [art_95], [art_97], [art_99], [art_104], [art_108], [art_113], [art_115], [art_117], [art_122], [art_125], [3], [art_130], [art_131], [art_132] https://paperswithcode.com/dataset/pems04
PeMS07 (PeMSD7) [art_108], [art_110], [art_113], [art_117], [art_125], [3], [art_130] https://paperswithcode.com/dataset/pems07
PeMS08 (PeMSD8) [art_95], [art_97], [art_99], [art_108], [art_115], [art_117], [art_122], [art_125], [art_130], [art_131], [art_132] https://paperswithcode.com/dataset/pems08
PEMS-BAY [art_100], [art_109], [art_110], [art_116], [art_120], [art_122] https://paperswithcode.com/dataset/pems-bay
Q-Traffic [art_124] https://github.com/JingqingZ/BaiduTraffic#Dataset
RCOTP [art_95] https://developer.ibm.com/exchanges/data/all/airline/
UVDS [art_128] https://doi.org/10.1007/978-3-030-73280-6_6 (source paper)
TaxiBJ [art_126] https://www.microsoft.com/en-us/research/publication/forecasting-citywide-crowd-flows-based-big-data/
Taxi Shenzhen [art_103], [art_135] https://github.com/lehaifeng/T-GCN/tree/master/data
TLC Taxi [art_119] https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page
The most popular datasets are PeMS ones (http://pems.dot.ca.gov/), provided by the California Transportation Agency Performance Measurement System, which include various versions differing in time range, spatial coverage, and the number of sensors, and METR-LA, collected from loop detectors in Los Angeles County highway. The number of sensors is in the hundreds, and so is the number of nodes in the graph. For a more detailed overview of the benchmark datasets in the traffic domain, see Ref. [review_traffico_con_dataset]. Other common public traffic datasets can be found at https://paperswithcode.com/task/traffic-prediction. In terms of external data, the most commonly used are weather conditions, time information (time of day and day of week) and events (e.g., holidays), as well as information about the road network (e.g., points of interest).
The majority of the papers utilize data with a 5-minute granularity, and provide forecasts for horizons ranging from 5 minutes to 1 hour ahead. In addition to imputation of missing values, in almost half of the cases the pre-processing of the data includes data normalization (min-max or Z-score normalization) and the selection of time windows regarding recent, daily and weekly data (as in 8 out of 39 papers).
5.5.3 Proposed models
As for the proposed models, 18 out of 39 belong to the pure convolutional GNN paradigm, followed by 11 attentional GNNs, 9 hybrid convolutional-attention architectures, and 1 recurrent GNN. Regarding the definition of the graph structure, most of the papers discuss it in detail. Half of the papers are based on a pre-defined adjacency matrix, and among the remaining papers that propose models that learn the adjacency matrix on their own, a further half is based on a pre-initialization of the matrix, as for example in Ref. [art_108]. Here, more than in any other group, there are multiple ways to define the adjacency matrix, taking into account either the structure of the road network, the actual flow of people, or a combination of different information. More in detail, some models use the road connectivity ([art_104], [art_120], [art_123], [art_135]), others their spatial distance ([art_93], [art_96], [art_99], [art_110], [art_128], [art_132]), and some others incorporate extra information about the so-called points of interest (such as Refs. [art_94], [art_96]). Other models instead consider some similarity measures between the time series ([art_97], [art_117], [art_122], [art_123]), such as the correlation coefficient or cosine similarity. Still others directly use the number of people traveling on the road at a given time step ([art_94], [art_102], [art_129], [art_133]).
The two most common loss functions used to train the models are the MSE and the MAE. Few papers consider a combination of loss functions, especially when dealing with a model that has to learn the graph structure by itself (Refs. [art_95], [art_126], [art_130]). Not all papers specify the language or library used for the code. Among those that do, Python is the most common language, with the majority using PyTorch, while only a few researchers use TensorFlow. In addition, a few papers provide a link to the source code for the proposed model. The papers that provide such links are listed in Tab. LABEL:t:link_codici_traffico.
Table 17: List of source codes of the ”Mobility” models in the review.
Model Link
3-Dimensional Graph Convolution Network (3DGCN) [art_94] https://github.com/FIBLAB/3D-DGCN
Adaptive generalized PageRank graph neural network (AGP-GNN) [art_116] https://github.com/guoxiaoyuatbjtu/agp-gnn
Automated dilated spatio-temporal synchronous graph network (Auto-DSTSGN) [art_108] https://github.com/jinguangyi-uto-DSTSGN
Dynamic multi-view graph neural network for citywide traffic inference (CTVI+) [art_97] https://github.com/dsj96/TKDD
Dynamic Graph Convolutional Recurrent Network (DGCRN) [art_100] https://github.com/tsinghua-fib-lab/Traffic-Benchmark
Spatio-temporal Causal Graph Attention Network (STCGAT) [art_125] https://github.com/zsqZZU/STCGAT/tree/v1.0.0
Spatial-Temporal Graph Convolutional Neural Network with LSTM layers (STGCN-LSTM) [3] https://github.com/ant-agafonov/stgcn-lstm
Spatio-Temporal Graph Mixformer (STGM) [art_110] https://github.com/mouradost/stgm
5.5.4 Benchmark models
Tab. LABEL:t:benchmark_models_traffico lists the benchmark models used in the papers in the ”Mobility” group. The benchmarks are categorized into mathematical and statistical methods, classical machine learning models, and GNN models, and only those appearing in more than two papers are included.
Table 18: List of benchmark models in the ”Mobility” group divided per category.
Category Model Used by
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_93], [art_94], [art_95], [art_96], [art_98], [art_100], [art_101], [art_103], [art_106], [art_107], [art_109], [art_116], [art_123], [art_126], [art_129], [art_132], [art_135]
Historical average (HA) [art_93], [art_94], [art_96], [art_98], [art_100], [art_101], [art_102], [art_103], [art_109], [art_110], [art_120], [art_123], [art_125], [art_131], [art_132], [art_135]
Vector autoregression (VAR) [art_99], [art_100], [art_126], [art_131], [art_132]
Classical machine learning methods Feed-forward neural network (FNN) [art_93], [art_97], [art_100], [art_105], [art_114], [art_120], [art_121], [art_124], [art_126]
Gated recurrent unit (GRU) [art_98], [art_99], [art_101], [art_106], [art_121], [art_123], [art_124], [3], [art_131], [art_132], [art_135]
Long-short term memory network (LSTM) [art_93], [art_96], [art_98], [art_99], [art_100], [art_106], [art_107], [art_108], [art_109], [art_113], [art_115], [art_116], [art_120], [art_121], [art_122], [art_125], [art_126], [3], [art_130], [art_131], [art_132]
Long-short term time series network (LSTNet) [LSTNet] [art_102], [art_110]
Seq2Seq [Seq2seq] [art_105], [art_124]
Spatial and temporal normalization for multivariate time series forecasting (ST-Norm) [ST-Norm] [art_116], [art_117]
Support vector regression (SVR) [art_99], [art_100], [art_103], [art_107], [art_109], [art_119], [art_123], [art_129], [art_130], [art_135]
Extreme gradient boosting (XGBoost) [XGBoost] [art_97], [art_114]
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_95], [art_98], [art_100], [art_101], [art_110], [art_117], [art_124], [art_125], [art_130], [art_131]
Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_93], [art_95], [art_98], [art_99], [art_100], [art_107], [art_108], [art_110], [art_113], [art_115], [art_117], [art_121], [art_125], [art_126], [3], [art_130], [art_131], [art_132]
Dual-stage attention-based recurrent neural network (DA-RNN) [DA-RNN] [art_96], [art_105]
Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_93], [art_94], [art_95], [art_96], [art_100], [art_101], [art_103], [art_107], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_122], [art_124], [art_125], [art_126], [art_128], [art_129], [art_130], [art_131], [art_132]
Dynamic spatial-temporal aware graph neural network (DSTAGNN) [DSTAGNN] [art_95], [art_125]
GeoMAN [GeoMAN] [art_96], [art_132]
Graph convolutional network (GCN) [art_104], [art_135]
Graph multi-attention network (GMAN) [GMAN] [art_100], [art_109], [art_110], [art_116], [art_122], [art_128]
Graph WaveNet (GWN) [GWN] [art_93], [art_95], [art_100], [art_101], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_122], [art_124], [art_128]
Multi-range attentive bicomponent graph convolutional network (MRA-BGCN) [MRA-BGCN] [art_109], [art_116]
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_100], [art_109], [art_110], [art_113], [art_116], [art_117], [art_122], [art_124], [art_128]
Multi-view graph convolutional network (MVGCN) [MVGCN] [art_94], [art_96]
Spatial-temporal fusion graph neural network (STFGNN) [STFGNN] [art_98], [art_101], [art_108], [art_115], [art_117], [art_125], [art_130]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_93], [art_94], [art_95], [art_96], [art_99], [art_100], [art_101], [art_106], [art_107], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_121], [art_122], [art_124], [art_126], [3], [art_128], [art_129], [art_130], [art_131], [art_132]
Spatial temporal graph neural network (STGNN) [STGNN] [art_94], [art_103]
Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_99], [art_108], [art_117], [art_125]
Deep-meta-learning based model (ST-MetaNet) [ST-MetaNet] [art_96], [art_100], [art_109], [art_116]
Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_98], [art_99], [art_100], [art_108], [art_115], [art_117], [art_125], [art_130], [art_131]
Temporal graph convolutional network (T-GCN) [T-GCN] [art_103], [art_107], [art_113], [art_121], [art_123]
As for the basic mathematical and statistical methods, two widely used benchmarks are the ARIMA model and historical average model (HA), which uses the average of past observations as the forecast for future values. Despite their apparent simplicity, these models are sometimes shown to be very powerful benchmarks, with an accuracy comparable to that of more complex models. As for the traditional machine learning models, LSTMs are widely recognized as a type of recurrent neural network capable of handling and long-term dependencies in traffic data. As for GNN-based models instead, the three most used models are, in order, the spatio-temporal graph convolutional network (STGCN) [STGCN], the diffusion convolutional recurrent neural network (DCRNN) [DCRNN], and the attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN], all developed specifically for the traffic forecasting problem.
The STGCN model was proposed by Yu et al. in 2018, and it integrates graph convolution and gated temporal convolution through spatio-temporal convolutional blocks. The source code is available at https://github.com/VeritasYin/STGCN_IJCAI-18. The DCRNN model was proposed by Li et al. in 2017, and it aims to capture the spatial dependency using bidirectional random walks on the graph, and the temporal dependency using the encoder-decoder architecture with scheduled sampling. The code can be accessed at https://github.com/liyaguang/DCRNN. Finally, the ASTGCN model was proposed by Guo et al. in 2019, and it models in parallel recent, daily-periodic and weekly-periodic dependencies as three independent components with a spatio-temporal attention mechanism to capture the dynamic spatio-temporal correlations, and a spatio-temporal convolution which simultaneously employs graph convolutions to capture the spatial patterns. The code is available at https://github.com/Davidham3/ASTGCN-2019-mxnet. For a more comprehensive list of source codes for the most dated benchmark GNN models in the traffic domain see Ref. [review_traffico_jiang].
5.5.5 Results
In all papers, the authors claim that the proposed GNN-based models are able to outperform the other benchmarks. However, there are a limited number of situations in which some simple statistical and mathematical models demonstrate comparable accuracy, as it can be observed in Tabs. LABEL:t:risultati_metr-la and LABEL:t:risultati_pems-bay.
For the and comparison and evaluation of the models, the most commonly used error metrics are MAE, RMSE, and MAPE. Few papers also use different metrics, such as the coefficient of determination $R^{2}$. However, despite the use of the same metrics and the availability of widely used datasets, comparing models is not trivial because different papers often focus on different time windows of a dataset.
[…]
Table 19: Comparison of the average accuracy of the different models on the METR-LA dataset for 15, 30 and 60 minute time horizons, expressed in terms of MAE, MAPE and RMSE. The results of the models marked with are taken from papers in the ”Generic” group. The smallest errors are underlined. Numbers from the papers.
Time horizon 15 min 30 min 60 min
Metrics MAE MAPE RMSE MAE MAPE RMSE MAE MAPE RMSE
AGCRN 2.87 7.70 5.58 3.23 9.00 6.58 3.62 10.38 7.51
AGP-GNN [art_116] 2.71 7.21 5.25 3.04 8.15 6.20 3.41 9.70 7.20
ARIMA 3.99 9.60 8.12 5.15 12.70 10.45 6.90 17.40 13.23
ASTGCN 4.86 9.21 9.27 5.43 10.13 10.61 6.51 11.64 12.52
DCRNN 2.77 7.30 5.38 3.15 8.80 6.45 3.60 10.50 7.60
DGCRN [art_100] 2.62 6.63 5.01 2.99 8.02 6.05 3.44 9.73 7.19
DST-GCNN [art_126] 2.68 7.20 5.35 3.01 8.50 6.23 3.41 10.30 7.47
DyGCN-LSTM [art_113] 2.64 6.35 4.64 3.24 8.04 5.24 3.53 9.30 6.53
FC-GAGA 2.75 7.25 5.34 3.10 8.57 6.30 3.51 10.14 7.31
FNN 3.99 9.90 7.94 4.23 12.90 8.17 4.49 14.00 8.69
GA-LSTM [art_120] 2.30 7.01 4.97 2.98 8.50 6.00 3.55 9.97 7.23
GGRU 2.71 6.99 5.24 3.12 8.56 6.36 3.64 10.62 7.65
GMAN 2.77 7.25 5.48 3.07 8.35 6.34 3.40 9.72 7.22
GWN 2.69 6.90 5.15 3.07 8.37 6.22 3.53 10.01 7.37
HA 4.16 13.00 7.80 4.16 13.00 7.80 4.16 13.00 7.80
LSTM 3.44 9.60 6.30 3.77 10.90 7.23 4.37 13.20 8.69
MCFGNN [art_86] 2.31 5.86 5.29 2.66 6.89 6.27 3.05 7.65 7.20
MRA-BGCN 2.67 6.80 5.12 3.06 8.30 6.17 3.49 10.00 7.30
MS-GAT 2.65 6.43 2.99 4.83 8.31 5.91 6.48 9.48 8.11
MTGNN 2.69 6.86 5.18 3.05 8.19 6.17 3.49 9.87 7.23
MTGODE [art_75] 2.66 6.87 5.10 3.00 8.19 6.05 3.39 9.80 7.05
MVST-GNN [art_109] 2.70 5.19 6.97 3.06 8.12 6.08 3.40 9.60 6.98
SDGCN 2.76 7.16 5.38 3.15 8.53 6.46 3.61 9.87 7.43
STAG-GCN [art_93] 2.67 7.00 5.23 3.07 8.26 6.15 3.50 9.93 7.24
STFGNN 2.57 6.51 4.73 2.83 7.46 5.46 3.18 8.81 6.40
STGCN 2.88 7.62 5.74 3.47 9.57 7.24 4.59 12.70 9.40
STGODE 3.47 8.76 6.76 4.36 11.14 8.47 5.50 14.32 10.33
STG-NCDE 3.77 8.54 9.47 4.84 10.63 12.04 6.35 13.49 14.94
STGRAT 2.60 6.61 5.07 3.01 8.15 6.21 3.49 10.01 7.42
ST-MetaNet 2.69 6.91 5.17 3.10 8.57 6.28 3.59 10.63 7.52
STSGCN 3.31 8.06 7.62 4.13 10.29 9.77 5.06 12.91 11.66
SVR 3.99 9.30 8.45 5.05 12.10 10.87 6.72 16.70 13.76
TF-GAN [art_122] 2.63 6.55 4.94 3.06 8.36 6.20 3.32 9.48 7.11
T-GCN 3.03 7.81 5.26 3.52 9.45 6.12 4.30 11.80 7.31
VAR 4.42 10.20 7.89 5.41 12.70 9.13 6.52 15.80 10.11
WaveNet 2.99 8.04 5.89 3.59 10.25 7.28 4.45 13.62 8.93
The results on the METR-LA dataset are not uniform, and the best performing models vary depending on the metric and the time horizon. However, especially for longer time horizons, the Dynamic Graph Convolution LSTM Network (DyGCN-LSTM) [art_113] appears to be the most accurate one (excluding the ”Generic” models, among which the best model is the multichannel fusion graph neural network (MCFGNN) [art_86]).
[…]
Table 20: Comparison of the average accuracy of the different models on the PEMS-BAY dataset for 15, 30 and 60 minute time horizons, expressed in terms of MAE, MAPE and RMSE. The smallest errors are underlined. Numbers from the papers.
Time horizon 15 min 30 min 60 min
Metrics MAE MAPE RMSE MAE MAPE RMSE MAE MAPE RMSE
AGCRN 1.37 2.94 2.87 1.69 3.87 3.85 1.96 4.64 4.54
AGP-GNN [art_116] 1.33 2.79 2.86 1.61 3.66 3.67 1.86 4.30 4.34
ARIMA 1.62 3.50 3.30 2.33 5.40 4.76 3.38 8.30 6.50
ASTGCN 1.52 3.22 3.13 2.01 4.48 4.27 2.61 6.00 5.42
DCRNN 1.38 2.90 2.95 1.74 3.9 3.97 2.07 4.90 4.74
DGCRN [art_100] 1.28 2.66 2.69 1.59 3.55 3.63 1.89 4.43 4.42
FC-GAGA 1.36 2.87 2.86 1.68 3.80 3.80 1.97 4.67 4.52
FNN 2.20 5.19 4.42 2.30 5.43 4.63 2.46 5.89 4.98
GMAN 1.34 2.81 2.82 1.62 3.62 3.72 1.86 4.31 4.32
GWN 1.30 2.73 2.74 1.63 3.67 3.70 1.95 4.63 4.52
HA 2.88 6.80 5.59 2.88 6.80 5.59 2.88 6.80 5.59
LSTM 2.05 4.80 4.19 2.20 5.20 4.55 2.37 5.70 4.96
MRA-BGCN 1.29 2.90 2.72 1.61 3.80 3.67 1.91 4.60 4.46
MTGNN 1.32 2.77 2.79 1.65 3.69 3.74 1.94 4.53 4.49
MTGODE [art_75] 1.29 2.72 2.73 1.61 3.61 3.66 1.88 4.39 4.31
MVST-GNN [art_109] 1.30 2.73 2.74 1.57 3.56 3.54 1.80 4.19 4.26
SDGCN 1.35 2.89 2.88 1.69 3.87 3.91 1.99 4.69 4.59
STGCN 1.36 2.90 2.96 1.81 4.17 4.27 2.49 5.79 5.69
STGODE 1.43 2.99 2.88 1.84 3.84 3.90 2.30 4.61 4.89
STG-NCDE 1.38 2.91 2.93 1.71 3.91 3.84 2.03 4.82 4.58
STGRAT 1.29 2.67 2.71 1.61 3.63 3.69 1.95 4.64 4.54
STID 1.30 2.73 2.81 1.62 3.68 3.72 1.89 4.47 4.4
ST-MetaNet 1.36 2.82 2.90 1.76 4.00 4.02 2.20 5.45 5.06
ST-Norm 1.34 2.82 2.88 1.67 3.75 3.83 1.96 4.62 4.52
STSGCN 1.44 3.04 3.01 1.83 4.17 4.18 2.26 5.40 5.21
SVR 1.85 3.80 3.59 2.48 5.50 5.18 3.28 8.00 7.08
VAR 1.74 3.60 3.16 2.32 5.00 4.25 2.93 6.50 5.44
WaveNet 1.39 2.91 3.01 1.83 4.16 4.21 2.35 5.87 5.43
The results on the the PEMS-BAY dataset indicate that the dynamic graph convolutional recurrent network (DGCRN) [art_100] is the most accurate model for the 15-minute forecasting horizon, while the multi-view spatial-temporal graph neural network (MVST-GNN) [art_109] is the best model for the longer horizons.
[…]
Table 21: Comparison of the average accuracy of the different models on the PeMS03 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE. The results of the models marked with are taken from papers in the ”Generic” group. The smallest errors are underlined. Numbers from the papers.
Time horizon 60 min
Metrics MAE MAPE RMSE
AGCRN 17.69 19.40 29.66
ASTGCN 17.69 19.40 29.66
Auto-DSTSGN [art_108] 14.59 14.22 25.17
AutoSTG 16.27 16.10 27.63
DCGCN [art_117] 15.29 - 25.98
DCRNN 18.18 18.91 30.31
Graph WaveNet 19.85 19.31 32.94
LSGCN 17.94 - 29.85
LSTM 21.33 23.33 35.11
STAGCN [art_130] 15.40 14.48 26.23
STCGNN [art_115] 15.81 14.59 27.23
STFGCN 16.77 16.30 28.34
STGCN 17.49 17.15 30.12
STGODE 16.53 16.68 27.79
STSGCN 17.48 16.78 29.21
SVR 21.97 21.51 35.29
[…]
Table 22: Comparison of the average accuracy of the different models on the PeMS04 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE. The results of the models marked with are taken from papers in the ”Generic” group. The smallest errors are underlined. Numbers from the papers.
Time horizon 60 min
Metrics MAE MAPE RMSE
AGCRN 19.83 12.97 32.26
ARIMA 33.73 24.18 48.80
ASTGCN 22.93 16.56 35.22
Auto-DSTSGN [art_108] 18.85 13.21 30.48
Autoformer 23.76 18.01 36.59
AutoSTG 20.38 14.12 32.51
Crossformer 20.40 14.62 32.79
DCGCN [art_117] 20.28 - 31.65
DCRNN 19.71 13.54 31.43
DGCRN 19.04 12.80 30.82
DMGF-Net [art_99] 20.59 13.63 32.43
DSA-NET 22.79 16.03 35.77
DSTGN [art_73] 18.61 12.31 30.79
DSTIGNN [art_74] 18.41 12.45 29.97
FEDFormer 22.86 16.04 35.07
GMAN 20.93 14.06 33.34
GODERN-FS [art_88] 19.17 12.74 30.96
GRU 23.68 16.44 39.27
GWN 19.36 13.31 31.72
HA 38.03 27.88 59.24
LSGCN 21.53 13.18 33.86
LSTM 23.81 18.12 36.62
MTGNN 24.89 17.29 39.66
MTGODE 19.55 13.08 32.99
SCINet 19.29 11.89 31.28
SDGL [art_78] 18.65 12.38 31.30
STAGCN [art_130] 19.02 12.46 30.75
STCGNN [art_115] 19.39 12.71 31.17
STFGCN 19.83 13.02 31.88
STG2SEQ 25.20 18.77 38.48
STGODE 20.84 13.77 32.82
STIDGCN 18.47 12.42 29.90
STGCN 21.16 13.83 34.89
STSGCN 21.08 13.88 33.83
STGODE 20.84 13.76 32.84
STSGCN 21.19 13.90 33.65
SVR 28.70 19.20 44.56
TCN 23.22 15.59 37.26
VAR 23.51 17.85 36.39
[…]
Table 23: Comparison of the average accuracy of the different models on the PeMS07 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE. The smallest errors are underlined. Numbers from the papers.
Time horizon 60 min
Metrics MAE MAPE RMSE
AGCRN 22.37 9.12 36.55
ASTGCN 28.05 13.92 42.57
Auto-DSTSGN [art_108] 20.08 8.57 33.02
AutoSTG 23.22 9.95 36.47
DCGCN [art_117] 22.06 - 34.66
DCRNN 25.30 11.66 38.58
GWN 26.85 12.12 42.78
LSGCN 27.31 - 41.46
LSTM 29.98 13.20 45.94
STAGCN [art_130] 21.10 8.92 34.10
STFGCN 22.07 9.21 35.80
STGCN 25.38 11.08 38.78
STGODE 22.59 - 37.54
STSGCN 24.26 10.21 39.03
SVR 32.49 14.26 50.22
[…]
Table 24: Comparison of the average accuracy of the different models on the PeMS08 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE. The results of the models marked with are taken from papers in the ”Generic” group. The smallest errors are underlined. Numbers from the papers.
Time horizon 60 min
Metrics MAE MAPE RMSE
AGCRN 15.95 10.09 25.22
ARIMA 31.09 22.73 44.32
ASTGCN 18.61 13.08 28.16
Auto-DSTSGN [art_108] 14.74 9.45 23.76
Autoformer 21.04 15.98 31.33
AutoSTG 16.37 10.36 25.46
Crossformer 16.25 11.14 26.14
DCGCN [art_117] 15.68 - 24.39
DCRNN 15.26 9.96 24.28
DMGF-Net [art_99] 16.48 10.56 25.73
DSA-NET 17.14 11.32 26.96
DSTIGNN [art_74] 13.88 9.04 23.64
FEDformer 19.55 13.58 29.30
GMAN 16.97 11.32 26.70
GODERN-FS [art_88] 15.59 9.95 24.79
GRU 22.00 13.33 36.24
GWN 15.07 9.51 23.85
HA 34.86 24.07 52.04
LSGCN 17.73 11.20 26.76
LSTM 22.20 14.20 34.06
MTGNN 18.28 12.15 30.05
MTGODE 15.61 10.15 25.96
SCINet 15.78 9.97 24.60
SDGL [art_78] 14.93 9.61 24.13
STAGCN [art_130] 15.36 9.80 24.32
STCGNN [art_115] 15.55 10.02 24.61
STFGCN 16.64 10.60 26.22
STIDGCN 14.10 9.15 23.72
STG2SEQ 20.17 17.32 30.71
STGCN 17.50 11.29 27.09
STGODE 16.79 10.58 26.01
STSGCN 17.13 10.96 26.80
SVR 23.25 14.64 36.16
TCN 22.72 14.03 35.79
VAR 22.07 14.04 31.02
5.6 Predictive monitoring
In modern industry, sensor-based monitoring of processes has become essential. This pairs with the rapid growth of the Internet of Things (IoT), which aims at a smarter management across various applications [art_9]. In this framework, multi-sensor systems can be used to detect anomalies in complex scenarios and to monitor the overall status of a system. Predictive monitoring is a task that involves the continuous observation and analysis of a system, in order to forecast its future states and verify whether the predicted outcomes meet certain standards. This predictive monitoring task is important because it significantly improves efficiency, reliability, and cost of industrial operations. Related approaches include several techniques, such as anomaly detection, fault diagnosis and remaining useful life estimation.
Despite the specificity of the field, 16 out of 156 papers address the topic of predictive monitoring. This includes anomaly detection ([art_9], [art_10]), fault diagnosis ([art_53], [art_54], [art_55], [art_56], [art_57], [art_58], [art_192]) and RUL estimation ([art_152], [art_153], [art_154], [art_155], [art_156], [art_157], [art_158]). The use of GNNs in these analyses is justified by the necessity of using a model capable of capturing the spatio-temporal correlations among the many sensors present in the system.
5.6.1 Overview
Predictive monitoring includes several techniques, such as anomaly detection, fault diagnosis and remaining useful life estimation. Anomaly detection is a technique that aims to identify abnormal data that are not due to random deviations, and are rather generated by a different underlying mechanism [ref_anomaly_detection_1]. It is a prerequisite part of fault diagnosis [ref_anomaly_detection_2]. Fault diagnosis is the process of determining whether a fault has occurred in a system, including the identification of the time, location, type, and severity of the fault. It is usually considered to be a classification problem. As for remaining useful life (RUL) estimation, it aims to determine how long a machine will operate before it needs to be repaired or replaced, and it is useful for scheduling maintenance interventions.
The selected papers implement these three techniques, with anomaly detection being less prevalent than the other two. The main applications studied are industrial (e.g., bearings fault diagnosis [art_54], [art_55], [art_57], [art_155], [art_156]) and mechanical (e.g, engines RUL estimation [art_152], [art_153], [art_154], [art_157]). There are also two papers dealing with energy related tasks, specifically fault diagnosis of photovoltaic systems [art_53] and energy networks [art_56]. They are included in this section due to their specific focus on fault classification rather than on energy-related time series forecasting. It shoud also be underlined that the majority of the papers aim at solving a classification task, which is typical of anomaly detection and fault diagnosis. As for RUL estimation, the quantity to be predicted is the RUL itself. For this reason, we do not refer to forecasting time horizons as in the other sections, since the goal is to estimate the RUL rather than to forecast data over future time intervals.
The majority of the collected papers were published in 2023 by authors affiliated with Chinese universities. The most popular journal for these topics is IEEE Transactions on Instrumentation and Measurement by the Institute of Electrical and Electronics Engineers by IEEE, with 5 out of 15 papers. All other journals contain only one of the selected articles. It can be observed that the majority of the articles are published in IEEE journals.
5.6.2 Datasets
Tab. LABEL:t:link_dataset_predictive_monitoring contains a list of the public datasets used in the selected papers, with the links to access them.
Table 25: List of public datasets in the ”Predictive monitoring” group and their corresponding links.
Dataset Used by Link
Air Quality China [art_9] https://www.aqistudy.cn/historydata/
Air Quality - TPS [art_9] https://www.kaggle.com/amritpal333/tps-july-2021-original-dataset-clean
C-MAPSS [art_152], [art_153], [art_154], [art_157] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
CWRU bearing dataset [art_55], [art_192] http://csegroups.case.edu/bearingdatacenter/home
JNU bearing dataset [art_55] http://mad-net.org:8765/explore.html?id=9&t=0.9355271549540183
NASA battery dataset [art_158] http://ti.arc.nasa.gov/project/prognostic-data-repository
NASA prognostics data [art_58] https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/#bearing
N-CMAPSS [art_152], [art_154] https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/
MFPT [art_192] https://api.semanticscholar.org/CorpusID:26870683 (source paper)
NREL PV [art_53] https://www.nrel.gov/
PRONOSTIA [art_155], [art_156] https://hal.science/hal-00719503/document (source paper)
SEU bearing dataset [art_55] https://github.com/cathysiyu/Mechanical-datasets
XJTUGearbox and XJTUSpurgear [art_192] http://dx.doi.org/10.1016/j.ymssp.2021.108653 (source paper)
The two most popular datasets are C-MAPSS and N-CMAPSS, two well-known datasets for RUL prediction provided by NASA. C-MAPSS describes the deterioration of aircraft engines by recording temperature, pressure, and fan speed during the cruise phase (https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf). N-CMAPSS is an extended version of C-MAPSS, whose records cover climbing, cruising, and descending flight conditions (https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/). Practically no papers use exogenous variables, with the exception of Ref. [art_53], which evaluates the impact of meteorological variables on fault diagnosis of photovoltaic systems.
The granularity of the studied time series is highly variable, ranging from 1 hour to fractions of seconds. Many papers report the frequency of the sensors, with values ranging from 1 Hz to 50 kHz. Most datasets are normalized, by either using min-max normalization or Z-score normalization. In some papers, a data augmentation procedure is employed to mitigate the natural imbalance between the number of normal and anomalous conditions.
5.6.3 Proposed models
Half of the proposed GNN models (9 out of 16 papers) fall under the category of pure convolutional models, followed by 4 attentional GNNs, 2 hybrid models, and one whose description does not clarify which class of the taxonomy it belongs to (Ref. [art_155]). Almost all the models have a graph learning module, so that the graph structure is rarely defined a priori by the researchers. Three papers ([art_55], [art_57], [art_154]) use the visibility graph algorithm proposed by Lacasa et al. in Ref. [visibility_graph] for converting the time series into a graph.
The most common loss functions in this group are the cross entropy function for classification problems, and MSE or RMSE for RUL forecasting problems. Most of the papers mention the Python language and the PyTorch library, but only two papers provide a link to the source code of their model, and in both cases, it is for RUL prediction. The source code for the convolution-graph attention network (ConvGAT) [art_153] is available at https://github.com/CUG-FDGroup, while the one for the graph network with causal connectivity and temporal convolutional neural network feature extractor proposed in Ref. [art_155] is available at https://github.com/mylonasc/gnn-tcnn.
5.6.4 Benchmark models
Tab. LABEL:t:benchmark_models_predictive_monitoring presents the benchmark models employed in the ”Predictive monitoring” group by at least two articles examined in this review.
Table 26: List of benchmark models in the ”Predictive monitoring” group divided per category.
Category Model Used by
Classical machine learning methods Attention-based long-short term memory network [art_153], [art_157]
Feature-attention based bidirectional GRU and CNN model (AGCNN) [AGCNN] [art_153], [art_154], [art_157]
Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_152], [art_153], [art_154], [art_157]
Convolutional neural network (CNN) [art_9], [art_56], [art_152], [art_153], [art_154], [art_156], [art_157]
Decision tree (DT) [art_9], [art_56]
Feed-forward neural network (FNN) [art_55], [art_155]
Gradient boosting (GB) [art_56], [art_154]
Gated recurrent unit (GRU) [art_53], [art_56], [art_156]
Logistic regression (LR) [art_9], [art_56]
Long-short term memory network (LSTM) [art_55], [art_56], [art_154], [art_157], [art_158]
Random forest (RF) [art_56], [art_154]
Support vector machine (SVM) [art_9], [art_56]
Transformer [attention] [art_152], [art_154]
GNN methods Graph attention network (GAT) [GAT] [art_54], [art_56], [art_192]
Graph convolutional network (GCN) [art_54], [art_55], [art_156], [art_192]
Hierarchical attention graph convolutional network (HAGCN) [HAGCN] [art_152], [art_153]
Spatio-temporal fusion attention (STFA) [art_157] [art_153], [art_154]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_58], [art_152]
In this group of papers there is a high variety of benchmarks used, and only few papers share a benchmark model, especially in the GNN category. The most popular benchmarks are two classic well-known machine learning models, the CNN and the LSTM network. Not surprisingly, there are no classical statistical or mathematical benchmark models, since they are more frequently used in forecasting problems compared to classification problems.
Among the GNN benchmarks, there are two models that have been specifically developed for remaining useful life prediction, namely the hierarchical attention graph convolutional network (HAGCN) [HAGCN] and the spatio-temporal fusion attention (STFA) [art_157]. In the HAGCN model, proposed by Li et al. in 2021, the spatial dependencies are modeled by a hierarchical graph representation layer, and a bidirectional long-short term memory network is used for modeling temporal dependencies of sensor measurements. As for the STFA model, proposed by Kong et al. in 2022, it integrates a priori knowledge about the equipment’s structure with a spatio-temporal deep learning architecture which employs LSTM cells and an attention mechanism. Neither of the two original papers includes a link to the source code.
5.6.5 Results
The results demonstrate that the proposed GNN models outperform the benchmarks. The accuracy of the models is evaluated using different metrics, depending on whether the context is classification or forecasting. For classification tasks, several studies present confusion matrices and calculate accuracy percentages. For RUL forecasting problems instead, the most commonly used metrics are the RMSE and the score function.
The only results that can be compared across papers are those for the C-MAPSS dataset. Tab. LABEL:t:risultati_c-mapss shows the accuracy of the different models, expressed in terms of RMSE and score function.
Table 27: Comparison of the accuracy of the different models on the C-MAPSS dataset on sub-datasets FD001–FD004, expressed in terms of RMSE and score function. The smallest errors are underlined. Numbers from the papers.
Sub-dataset FD001 FD002 FD003 FD004
Metrics RMSE Score RMSE Score RMSE Score RMSE Score
1D CNN 12.61 273.70 22.36 10412.00 12.64 284.10 23.31 12466.00
2D CNN 18.45 1286.70 30.29 13570.00 19.82 1596.20 29.16 7886.40
AGCNN 12.42 225.51 19.43 1492.76 13.39 227.09 21.50 3392.60
Attention-based LSTM 14.53 322.44 - - - - 27.08 5649.14
ATT-LSTM 13.95 320.00 17.65 2102.00 12.72 223.00 20.21 3100.00
BiGRU-AS 13.68 284.00 20.81 2454.00 15.53 428.00 27.31 4708.00
Bi-LSTM 13.65 295.00 23.18 4130.00 13.74 317.00 24.86 5430.00
BiLSTM-ED 14.74 273.00 22.07 3099.00 17.48 574.00 23.49 3202.00
BLCNN 13.18 302.28 19.09 1557.56 13.75 381.37 20.97 3858.78
CDSG [art_154] 11.26 188.00 18.13 1740.00 12.03 218.00 19.73 2332.00
ChebyNet-EdgePool 15.21 450.22 16.28 1229.22 14.67 421.02 16.26 1159.74
CNN-LSTM 14.40 290.00 27.23 9869.00 14.32 316.00 26.69 6594.00
ConvGAT [art_153] 11.34 197.43 14.12 771.61 10.97 235.26 15.51 1231.17
C-Transformer 13.79 475.46 16.11 2214.59 17.10 939.10 19.77 3237.37
DBN 15.21 417.59 27.12 9031.64 14.71 442.43 29.88 7954.51
DCNN 12.61 273.70 22.36 10412.00 12.64 284.1 23.31 12466.00
Earlier CNN 18.45 1287.00 30.29 13570.00 19.82 1596.00 29.16 7886.00
ELM 17.27 523.00 37.28 498149.97 18.90 573.78 38.43 121414.00
GA+RBM+LSTM 12.56 231.00 22.73 3366.00 12.10 251.00 22.66 2840.00
GAT-EdgePool 13.53 309.59 15.07 1380.41 14.66 470.74 17.60 1726.53
GAT-TopkPool 13.21 303.18 17.25 5338.80 15.36 507.52 21.44 2971.93
GB 15.67 474.01 29.09 87280.06 16.84 576.72 29.01 17817.92
GGCN 11.82 186.70 17.24 1493.70 12.21 245.19 17.36 1371.50
HAGCN 11.93 222.30 15.05 1144.10 11.53 240.30 15.74 1218.60
HDNN 13.02 245.00 15.24 1282.42 12.22 287.72 18.16 1527.42
IMDSSN 12.14 206.11 17.40 1775.15 12.35 229.54 19.78 2852.81
LSTM 16.14 338.00 24.49 4450.00 16.18 852.00 28.17 5550.00
LSTMBS 14.89 481.00 26.86 7982.00 15.11 493.00 27.11 5200.00
MODBN 15.04 334.23 25.05 5585.00 12.51 421.91 28.66 6557.00
MS-CNN 11.44 196.22 19.35 3747.00 11.67 241.89 22.22 4844.00
RF 17.91 479.75 29.59 70456.86 20.27 711.13 31.12 46567.63
SBI 13.58 228.00 19.59 2650.00 19.16 1727.00 22.15 2901.00
SMDN 13.40 272.00 - - - - 23.40 4302.00
STFA [art_157] 11.35 194.44 19.17 2493.09 11.64 224.53 21.41 2760.13
Transformer 13.52 287.07 19.32 1436.74 13.44 263.64 20.38 2784.62
In general, the best results on the different sub-datasets are achieved by the GNN-based comprehensive dynamic structure GNN (CDSG) [art_154] and convolution-graph attention network (ConvGAT) [art_153] models.
5.7 Generic
This subsection includes all the papers that do not explicitly address a specific problem. They still present the results on some benchmark case studies, but the related datasets may cover different groups. In particular, the focus is on papers that do not discuss or mention any specific case study prior to the experimental section. The reason for this choice is that they are considered to explore broader methodologies or frameworks that are applied to diverse datasets without a declared focus on a particular case.
5.7.1 Overview
A significant number of papers (24 out of 156) have been categorized under the ”Generic” group. The majority of these works commit to address a multivariate time series forecasting problem which is not necessarily limited to a specific field. This ”Generic” subset includes one paper from 2021, 4 from 2022, and 19 from 2023, highlighting the recent interest in applying GNNs in more general contexts. These papers are distributed across various journals, with the exception of Expert Systems with Applications by Elsevier, IEEE Transactions on Knowledge and Data Engineering by IEEE, and Information Sciences by Elsevier, each publishing 2 of the selected papers. The majority of authors are affiliated with institutions in China.
5.7.2 Datasets
The selected papers make use of the datasets listed in Tab. LABEL:t:link_dataset_generic in their experiments.
Table 28: List of public datasets in the ”Generic” papers and their corresponding links.
Dataset Used by Link
Beijing Traffic [art_84] https://github.com/BuaaPercy/Traffic-DataSet-of-Beijing-Road
CCMP wind data [art_73] http://data.remss.com
CI earthquakes [art_160] https://doi.org/10.5194/adgeo-43-31-2016 (source paper)
Chickenpox cases [art_160] https://doi.org/10.48550/arXiv.2102.08100 (source paper)
CW earthquakes [art_89] https://doi.org/10.5194/adgeo-43-31-2016 (source paper)
Electricity consumption [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_88], [art_90] https://github.com/laiguokun/multivariate-time-series-data
Electricity Consuming Load (ECL) [art_82], [art_85] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014
Electricity transformer Temperature (ETT) [art_76], [art_82], [art_85] https://github.com/zhouhaoyi/ETDataset
Energy [art_77] http://dx.doi.org/10.1016/j.enbuild.2017.01.083 (source paper)
Enron [art_202] https://doi.org/10.1073/pnas.1800683115 (source paper)
Eu-Core [art_202] https://doi.org/10.1145/ (source paper)
Exchange-rate [art_73], [art_74], [art_76], [art_77], [art_78], [art_79], [art_80], [art_81], [art_85], [art_87], [art_89], [art_90] https://github.com/laiguokun/multivariate-time-series-data
Facebook [art_202] http://networkrepository.com/socfb
GEFCOM 2012 [art_89] https://doi.org/10.1016/j.ijforecast.2013.07.001 (source paper)
Hypertext [art_202] http://networkrepository.com/ia-infect-hyper (source paper)
ILI [art_76], [art_85], [art_89] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
Jena Weather [art_76] https://www.bgc-jena.mpg.de/wetter/
Loan [art_202] https://ai.ppdai.com/mirror/showCompetitionRisk
METR-LA [art_72], [art_75], [art_79], [art_80], [art_86], [art_160] https://paperswithcode.com/dataset/metr-la
MIMIC-III Heart Failure [art_71] https://doi.org/10.1038/sdata.2016.35 (source paper)
MIMIC-III Infection [art_71] https://doi.org/10.1038/sdata.2016.35 (source paper)
MOOC network [art_70] http://snap.stanford.edu/jodie/mooc.csv
Nasdaq [art_77], [art_80] https://doi.org/10.48550/arXiv.1704.02971 (source paper)
NOAA Weather [art_82] https://www.ncei.noaa.gov/data/local-climatological-data/
PeMS03 [art_74] https://www.kaggle.com/datasets/elmahy/pems-dataset
PeMS04 (PeMSD4) [art_73], [art_74], [art_78], [art_79], [art_88] https://paperswithcode.com/dataset/pems04
PeMS08 (PeMSD8) [art_73], [art_74], [art_78], [art_79], [art_88] https://paperswithcode.com/dataset/pems08
PEMS-BAY [art_72], [art_75], [art_160] https://paperswithcode.com/dataset/pems-bay
PhysioNet [art_71] https://archive.physionet.org/challenge/2012/papers/
Reddit network [art_70] http://snap.stanford.edu/jodie/reddit.csv
Solar-Energy [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_88], [art_90] https://github.com/laiguokun/multivariate-time-series-data
TAIEX [art_84] https://finance.yahoo.com/
TaoBao [art_202] https://tianchi.aliyun.com/competition/entrance/
Traffic [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_85], [art_87], [art_90] https://github.com/laiguokun/multivariate-time-series-data
US stock market price [art_89] https://doi.org/10.48550/arXiv.1810.09936 (source paper)
Wikipedia network [art_70] http://snap.stanford.edu/jodie/wikipedia.csv
Wind-Speed [art_76] https://www.kaggle.com/datasets/fedesoriano/wind-speed-prediction-dataset
It should be noted that there are multiple recurring datasets that can be considered as benchmarks for the community. Some of these have already been discussed in the previous sections. If the results for these datasets have already been shown in previous sections, they will not be repeated again here. Instead, they will be just marked with a star together with the benchmark models in the tables of the previous sections.
The four most common datasets are Electricity consumption, Exchange-rate, Solar-Energy and Traffic, all available at https://github.com/laiguokun/multivariate-time-series-data. The Electricity consumption dataset contains the hourly electricity consumption in kWh for 321 customers recorded from 2012 to 2014. The Traffic dataset describes the road occupancy rates measured by different sensors on the San Francisco Bay area road network from 2015 to 2016, with an hourly granularity. In Solar energy it is recorded the solar power production from 137 photovoltaic plants in Alabama State in 2006, with a 10 minutes granularity. Finally, Exchange rate contains daily exchange rates of 8 countries including Australia, British, Canada, Switzerland, China, Japan, New Zealand and Singapore, from 1990 to 2016.
In general, the data granularity varies significantly, ranging from fractions of seconds to one week, indicating that there is no particular focus on a specific granularity. Individual papers often cover multiple granularities, and many papers take into account datasets with granularity from 5 minutes to 1 day. As for the forecasting horizon, almost all papers focus on multi-step forecasting, usually considering 12 different forecasting horizons at the same time. Only a few papers mention data pre-processing, and usually utilize min-max or Z-score normalization.
5.7.3 Proposed models
Among the 24 ”Generic” papers, 19 use a purely convolutional GNN approach, and two employ a purely attentional model. One paper lacks sufficient detail to determine its classification within the taxonomy [art_85], and another declares it is using an aggregating approach that is simpler than either convolutional or attentional approaches [art_81]. One more paper proposes a multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) model, where the continuous dynamics of simplified graph propagation is described by an ordinary differential equation (ODE) [art_75].
As for the definition of the graph structure, the majority of these models use self-learning techniques. This is a straightforward choice, since all the papers in this section want to ensure the applicability of the model in different contexts, and the use of pre-defined rules for the definition of the graph limits the model’s universal applicability. Half of the papers employ a dynamic graph structure, which adapts to the varying temporal patterns observed in the data.
The two most commonly utilized loss functions are the MSE and the MAE. Since the graph structure is typically learned by the models themselves, many papers include a regularization term in the loss function for the optimization of the graph structure (e.g., Refs. [art_73], [art_78], [art_79], [art_86]).
The majority of the papers specify the language and libraries used for the code, namely Python with PyTorch, PyTorch Geometric [PyTorch_Geometric], and Torch Spatiotemporal [Cini_Torch_Spatiotemporal_2022]. However, only slightly less than half of the papers provide a link to the source code. The links to the code repositories are listed in Tab. LABEL:t:link_codici_generic.
Table 29: List of source codes of the ”Generic” models in the review.
Model Link
Adaptive dependency learning neural network (ADLNN) [art_83] https://github.com/AbishekSriramulu/ADLGNN.git
Graph neural network with neural Granger causality (CauGNN) [art_77] https://github.com/RRRussell/CauGNN
Gformer [art_82] https://github.com/wxh453751461/Gformer
Multi-scale adaptive graph neural network (MAGNN) [art_80] https://github.com/shangzongjiang/MAGNN
Multivariate time series deep spatiotemporal forecasting model with a graph neural network (MDST-GNN) [art_90] https://github.com/yiminghzc/MDST-GNN
Multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) [art_75] https://github.com/GRAND-Lab/MTGODE
Static and dynamic graph learning network (SDGL) [art_78] https://github.com/ZhuoLinLi-shu/SDGL
Sparse graph learning from spatiotemporal time series [art_72] https://github.com/andreacini/sparse-graph-learning
Temporal decomposition enhanced graph neural network for multivariate time series forecasting (TDG4MSF) [art_81] https://github.com/TYUT-Theta/MHZN.git
Temporal graph convolution and attention (T-GAN) [art_202] https://github.com/malei666666/T_GAN
Two GNN models with different graph structures [art_160] https://github.com/StefanBloemheuvel/graph_comparison
5.7.4 Benchmark models
Tab. LABEL:t:benchmark_models_generic displays the benchmark models utilized by at least two of the selected ”Generic” papers.
Table 30: List of benchmark models in the ”Generic” group divided per category.
Category Model Used by
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_73], [art_74], [art_78], [art_79], [art_80], [art_81], [art_83], [art_84], [art_88], [art_89]
Gaussian process (GP) [art_73], [art_74], [art_78], [art_80], [art_81], [art_83]
Historical average (HA) [art_78], [art_88]
Vector autoregression (VAR) [art_73], [art_74], [art_77], [art_78], [art_79], [art_81], [art_84], [art_88]
Classical machine learning methods Autoformer [Autoformer] [art_74], [art_76], [art_85]
Dual self-attention network for multivariate time series forecasting (DSANet) [DSANet] [art_73], [art_78]
Feed-forward neural network (FNN) [art_84], [art_89]
Gated recurrent unit (GRU) [art_73], [art_78], [art_82], [art_88], [art_90]
Hybrid framework based on fully Dilated CNN (HyDCNN) [HyDCNN] [art_75], [art_83]
Informer [Informer] [art_74], [art_76], [art_82], [art_85]
LogTrans [LogTrans] [art_82], [art_85]
Long-short term memory network (LSTM) [art_71], [art_74], [art_79], [art_82], [art_84], [art_85], [art_86], [art_88], [art_89]
Long-short term time series network (LSTNet) [LSTNet] [art_73], [art_74], [art_75], [art_77], [art_78], [art_80], [art_81], [art_82], [art_83], [art_85], [art_87], [art_89], [art_90]
Multi-Level Construal Neural Network (MLCNN) [MLCNN] [art_77], [art_87]
Neural basis expansion analysis for interpretable time series forecasting (N-BEATS) [N-BEATS] [art_69], [art_85]
Recurrent neural network with fully connected gated recurrent units (RNN-GRU) [art_73], [art_77], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87]
Reformer [Reformer] [art_76], [art_82], [art_85]
SCINet [SCINet] [art_74], [art_90]
Temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM] [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_90]
Transformer [attention] [art_71], [art_77]
Hybrid of the multilayer perception and autoregressive model (VAR-MLP) [VAR-MLP] [art_73], [art_74], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_90]
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_73], [art_74], [art_78], [art_79], [art_80], [art_88]
Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_78], [art_79]
Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_73], [art_74], [art_75], [art_78], [art_79], [art_84], [art_86], [art_87], [art_88]
Graph auto encoder (GAE) [GAE] [art_70], [art_202]
Graph multi-attention network (GMAN) [GMAN] [art_73], [art_75], [art_86], [art_88]
Graph for time series (GTS) [GTS] [art_72], [art_78]
Graph WaveNet (GWN) [GWN] [art_74], [art_75], [art_79], [art_80], [art_86], [art_87]
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_72], [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_86], [art_87], [art_88], [art_90]
Node2vec [node2vec] [art_70], [art_202]
Spatial-temporal fusion graph neural network (STFGNN) [STFGNN] [art_73], [art_86]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_73], [art_75], [art_78], [art_79], [art_84], [art_86], [art_88]
Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_75], [art_88]
Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_73], [art_74], [art_78]
There is a wide variety of benchmark models in this ”Generic” group. As for mathematical and statistical models, autoregressions and ARIMA models are very popular for their simplicity and interpretability. Among classical machine learning models, the most widely used models are LSTMs, temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM], and long-short term time series network (LSTNet) [LSTNet]. The TPA-LSTM model, introduced by Shih et al. in 2019, employs a set of filters to capture time-invariant temporal patterns, and an attention mechanism to identify relevant time series for multivariate forecasting. The source code for this benchmark model is available at https://github.com/gantheory/TPA-LSTM. The LSTNet model, proposed by Lai et al. in 2018, combines a convolution neural network and a recurrent neural network to capture short-term local dependencies among variables and identify long-term patterns. Additionally, it incorporates an autoregressive model that enhances the robustness of the deep learning approach to time series with significant scale fluctuations. The source code for the model is accessible at https://github.com/laiguokun/LSTNet. Among GNN models, the most common ones are DCRNN [DCRNN] and MTGNN [MTGNN], already discussed in the previous sections.
5.7.5 Results
The most common error metrics used in the evaluation of the models in the ”Generic” group are, in order, the mean absolute error (MAE), correlation coefficient (CORR), root mean square error (RMSE), root relative squared error (RRSE), mean absolute percentage error (MAPE), and mean squared error (MSE). Differently from other metrics, where lower values indicate higher accuracy, the correlation coefficient (CORR) quantifies the strength of the the linear relationship between predicted and actual values, with higher values indicating a better performance. This Results subsection compares the results obtained for the Electricity consumption, Exchange rate, Solar energy and Traffic datasets. The results of the proposed models and benchmarks on METR-LA and PEMS datasets were already reported with a in Subsec. 5.5.5.
[…]
Table 31: Comparison of the average accuracy of the different models on the Electricity consumption dataset for 3, 6, 12, and 24 steps time horizons, expressed in terms of RRSE and CORR. The numbers corresponding to the highest accuracy are underlined.
Time horizon 3 steps 6 steps 12 steps 24 steps
Metrics RRSE CORR RRSE CORR RRSE CORR RRSE CORR
ADLGNN [art_83] 0.0719 0.9506 0.0809 0.9386 0.0887 0.9312 0.0930 0.9294
AGCRN 0.0766 0.9408 0.0894 0.9309 0.0921 0.9222 0.0967 0.9183
AGLG-GRU [art_79] 0.0738 0.9434 0.0864 0.9302 0.0912 0.9283 0.0947 0.9274
AR 0.0995 0.8845 0.1035 0.8632 0.105 0.8591 0.1054 0.8595
ARIMA 0.0917 0.8902 0.1002 0.8834 0.1028 0.8604 0.1042 0.8487
Autoformer 0.1258 0.9147 0.1344 0.9001 0.1357 0.8921 0.1554 0.8704
Crossformer 0.0742 0.9452 0.0855 0.9351 0.0901 0.9280 0.0967 0.9205
DSTGN [art_73] 0.0713 0.9518 0.0821 0.9424 0.0887 0.9357 - -
DSTIGNN [art_74] 0.0733 0.9515 0.082 0.9412 0.0899 0.9338 0.0954 0.9289
ESG 0.0718 0.9494 0.0844 0.9372 0.0898 0.9321 0.0962 0.9279
FEDformer 0.0889 0.9321 0.1006 0.9191 0.1154 0.908 0.1202 0.9012
GP 0.1500 0.8670 0.1907 0.8334 0.1621 0.8394 0.1273 0.8818
GTS 0.0790 0.9291 0.0884 0.9187 0.0957 0.9135 0.0951 0.9098
GWN 0.0746 0.9459 0.0922 0.9310 0.0909 0.9267 0.0962 0.9226
HyDCNN 0.0832 0.9354 0.0898 0.9329 0.0921 0.9285 - -
Informer 0.1337 0.8903 0.1532 0.8705 0.1635 0.8527 0.1834 0.8399
LSTNet 0.0864 0.9283 0.0931 0.9135 0.1007 0.9077 0.1007 0.9119
MAGNN [art_80] 0.0745 0.9476 0.0876 0.9323 0.0908 0.9282 0.0963 0.9217
MDST-GNN [art_90] 0.0738 0.9454 0.0833 0.9346 0.0884 0.9264 0.0922 0.9222
MTGNN 0.0745 0.9474 0.0878 0.9316 0.0916 0.9278 0.0953 0.9234
MTGODE [art_75] 0.0736 0.9430 0.0809 0.9340 0.0891 0.9279 - -
MTHetGNN 0.0749 0.9456 0.0892 0.9307 0.0959 0.8783 0.0969 0.8782
MTNet 0.0840 0.9319 0.0901 0.9226 0.0934 0.9165 0.0969 0.9147
RNN-GRU 0.1102 0.8597 0.1144 0.8623 0.1183 0.8472 0.1295 0.8651
SARIMA 0.0906 0.9055 0.0999 0.8829 0.1026 0.8674 0.1036 0.8621
SCINet 0.0740 0.9494 0.0845 0.9387 0.0929 0.9305 0.0967 0.927
SDGL [art_78] 0.0698 0.9534 0.0805 0.9445 0.0889 0.9351 0.0935 0.9301
SDLGNN 0.0726 0.9502 0.082 0.9384 0.0896 0.9304 0.0947 0.9257
SDLGNN-Corr 0.0737 0.9475 0.0841 0.9346 0.0923 0.9263 0.0971 0.9227
STG-NCDE 0.6152 0.8739 0.6584 0.8663 0.7302 0.8728 - -
TDG4-MSF [art_81] 0.0731 0.9499 0.0828 0.9371 0.0894 0.9306 0.0969 0.9246
Theta 0.0975 0.8906 0.1029 0.8723 0.1040 0.8576 0.1041 0.8595
TPA-LSTM 0.0823 0.9439 0.0916 0.9337 0.0964 0.9250 0.1006 0.9133
TRMF 0.1802 0.8538 0.2039 0.8424 0.2186 0.8304 0.3656 0.7471
VAR-MLP 0.1393 0.8708 0.1620 0.8389 0.1557 0.8192 0.1274 0.8679
The accuracy of the various models varies depending on the studied forecasting horizon but, overall, one can notice that the GNN models perform better than the others, with the static and dynamic graph learning network (SDGL) [art_78] being the most accurate.
Tab. LABEL:t:risultati_exchange_rate presents the accuracy of different models applied to the Exchange rate dataset, with daily data from 1990 to 2016. The evaluation covers horizons of 3, 6, 12, and 24 steps ahead, with accuracy measured by RRSE and CORR.
Table 32: Comparison of the average accuracy of the different models on the Exchange rate dataset for 3, 6, 12, and 24 steps time horizons, expressed in terms of RRSE and CORR. The numbers corresponding to the highest accuracy are underlined.
Time horizon 3 steps 6 steps 12 steps 24 steps
Metrics RRSE CORR RRSE CORR RRSE CORR RRSE CORR
AGCRN 0.0269 0.9717 0.0331 0.9615 0.0374 0.9531 0.0476 0.9334
AGLG-GRU [art_79] 0.0191 0.9792 0.0233 0.9701 0.0328 0.9548 0.0449 0.9372
AR 0.0228 0.9734 0.0279 0.9656 0.0353 0.9526 0.0445 0.9357
ARIMA 0.0198 0.9754 0.0261 0.9721 0.0344 0.9548 0.0445 0.9301
Autoformer 0.0291 0.9631 0.0379 0.9397 0.0441 0.9201 0.0501 0.9012
Crossformer 0.0226 0.9759 0.0269 0.9656 0.0358 0.9436 0.0489 0.9291
DSTGN [art_73] 0.0179 0.9782 0.0250 0.9715 0.0346 0.9562 - -
DSTIGNN [art_74] 0.0173 0.9818 0.0244 0.9723 0.0333 0.9571 0.0430 0.9407
ESG 0.0181 0.9792 0.0246 0.9717 0.0345 0.9564 0.0468 0.9392
FEDformer 0.0256 0.9701 0.0287 0.9555 0.0379 0.9385 0.0487 0.9190
GP 0.0239 0.8713 0.0272 0.8193 0.0394 0.8484 0.0580 0.8278
GTS 0.0180 0.9898 0.0260 0.9824 0.0333 0.9701 0.0442 0.9518
GWN 0.0251 0.9740 0.0300 0.9640 0.0381 0.951 0.0486 0.9294
Informer 0.0882 0.9563 0.1081 0.9321 0.1301 0.8911 0.1521 0.8021
LSTNet 0.0226 0.9735 0.0280 0.9658 0.0356 0.9511 0.0449 0.9354
MAGNN [art_80] 0.0183 0.9778 0.0246 0.9712 0.0343 0.9557 0.0474 0.9339
MDST-GNN [art_90] 0.0172 0.9811 0.0245 0.9727 0.0337 0.9578 0.0431 0.9392
MTGNN 0.0194 0.9786 0.0259 0.9708 0.0349 0.9551 0.0456 0.9372
MTHetGNN 0.0198 0.9769 0.0259 0.9701 0.0345 0.9539 0.0451 0.9360
MTNet 0.0212 0.9767 0.0258 0.9703 0.0347 0.9561 0.0442 0.9388
RNN-GRU 0.0192 0.9786 0.0264 0.9712 0.0408 0.9531 0.0626 0.9223
SARIMA 0.0197 0.9748 0.0253 0.9643 0.0338 0.9495 0.0444 0.9370
SCINet 0.0171 0.9787 0.0240 0.9704 0.0331 0.9553 0.0436 0.9396
SDGL [art_78] 0.0180 0.9808 0.0249 0.9730 0.0342 0.9583 0.0455 0.9402
TDG4-MSF [art_81] 0.0172 0.9825 0.0244 0.9718 0.0330 0.9569 0.0433 0.9386
Theta 0.0497 0.9738 0.0257 0.9656 0.0342 0.9510 0.0441 0.9323
TPA-LSTM 0.0174 0.9790 0.0241 0.9709 0.0341 0.9564 0.0444 0.9381
TRMF 0.0351 0.9142 0.0875 0.8123 0.0494 0.8993 0.0563 0.8678
VAR-MLP 0.0265 0.8609 0.0394 0.8725 0.0407 0.8280 0.0578 0.7675
Also in this case, GNN models are the best overall. The discrete graph structure learning for time series model (GTS), proposed in 2021 by Shang et al. in Ref. [GTS] and used as a benchmark in Ref. [art_78], demonstrates itself to have a superior performance. It is therefore a suitable benchmark for comparison.
Tab. LABEL:t:risultati_solar_energy summarizes the accuracy of the various models on the Solar energy dataset, covering the period from January 1, 2016, to December 31, 2016, at 10-minutes intervals. The analysis considers forecasting horizons of 3, 6, 12, and 24 steps ahead, and the results are expressed in terms of RRSE and CORR.
Table 33: Comparison of the average accuracy of the different models on the Solar energy dataset for 3, 6, 12, and 24 steps time horizons, expressed in terms of RRSE and CORR. The numbers corresponding to the highest accuracy are underlined.
Time horizon 3 steps 6 steps 12 steps 24 steps
Metrics RRSE CORR RRSE CORR RRSE CORR RRSE CORR
ADLGNN [art_83] 0.1708 0.9866 0.2188 0.9768 0.2897 0.9551 0.4128 0.906
AGCRN 0.1840 0.9841 0.2432 0.9708 0.3185 0.9487 0.4141 0.9087
AGLG-GRU [art_79] 0.1762 0.9842 0.2302 0.9682 0.3021 0.9532 0.4130 0.9084
AR 0.2435 0.9710 0.3790 0.9263 0.5911 0.8107 0.8699 0.5314
ARIMA 0.2328 0.9739 0.3413 0.9402 0.4531 0.8886 0.5810 0.8133
Autoformer 0.1935 0.9784 0.2604 0.9651 0.3959 0.9111 0.6064 0.818
Crossformer 0.1758 0.9801 0.2304 0.9679 0.3101 0.9437 0.4115 0.9001
DSTGN [art_73] 0.1787 0.9867 0.2358 0.9721 0.3079 0.9513 - -
DSTIGNN [art_74] 0.1684 0.9869 0.2165 0.9773 0.2863 0.9584 0.4064 0.9118
ESG 0.1708 0.9865 0.2278 0.9743 0.3073 0.9519 0.4101 0.9100
FEDformer 0.1901 0.9827 0.2444 0.9606 0.3502 0.9275 0.4851 0.8876
GP 0.2259 0.9751 0.3286 0.9448 0.5200 0.8518 0.7973 0.5971
GTS 0.1842 0.9842 0.2691 0.9645 0.3259 0.9481 0.4796 0.8678
GWN 0.1773 0.9846 0.2279 0.9743 0.3068 0.9527 0.4206 0.9055
HyDCNN 0.1806 0.9865 0.2335 0.9747 0.3094 0.9515 - -
Informer 0.2134 0.9715 0.2701 0.9549 0.4331 0.8985 0.7017 0.7921
LSTNet 0.1843 0.9843 0.2559 0.969 0.3254 0.9467 0.4643 0.887
MAGNN [art_80] 0.1771 0.9853 0.2361 0.9724 0.3015 0.9539 0.4108 0.9097
MDST-GNN [art_90] 0.1764 0.9855 0.2321 0.9735 0.3082 0.9519 0.4119 0.9103
MTGNN 0.1778 0.9852 0.2348 0.9726 0.3109 0.9509 0.4270 0.9031
MTGODE [art_75] 0.1693 0.9868 0.2171 0.9771 0.2901 0.9577 - -
MTHetGNN 0.1838 0.9845 0.2600 0.9681 0.3169 0.9486 0.4231 0.9031
MTNet 0.1847 0.9840 0.2398 0.9723 0.3251 0.9462 0.4285 0.9013
RNN-GRU 0.1932 0.9823 0.2628 0.9675 0.4163 0.9150 0.4852 0.8823
SARIMA 0.2227 0.9760 0.3228 0.9461 0.4512 0.8900 0.5714 0.8184
SCINet 0.1775 0.9853 0.2301 0.9739 0.2997 0.9550 0.4081 0.9112
SDGL [art_78] 0.1699 0.9866 0.2222 0.9762 0.2924 0.9565 0.4047 0.9119
SDLGNN 0.1720 0.9864 0.2249 0.9757 0.3024 0.9547 0.4184 0.9051
SDLGNN-Corr 0.1806 0.9848 0.2378 0.9722 0.3042 0.9534 0.4173 0.9067
STG-NCDE 0.2346 0.9748 0.2908 0.9605 0.5149 0.8639 - -
TDG4-MSF [art_81] 0.1746 0.9858 0.2348 0.9727 0.3082 0.9527 0.4031 0.9143
Theta 0.2442 0.9685 0.3327 0.9445 0.4488 0.8979 0.6092 0.8139
TPA-LSTM 0.1803 0.9850 0.2347 0.9742 0.3234 0.9487 0.4389 0.9081
TRMF 0.2473 0.9703 0.3470 0.9418 0.5597 0.8475 0.9005 0.5598
VAR-MLP 0.1922 0.9829 0.2679 0.9655 0.4244 0.9058 0.6841 0.7149
The results indicate that the dynamic spatiotemporal interactive graph neural network (DSTIGNN) model [art_74] and the static and dynamic graph learning network (SDGL) model [art_78] are the most accurate on the solar energy dataset, for shorter and longer time horizons, respectively.
Finally, Tab. LABEL:t:risultati_traffic illustrates the performance of different models on the Traffic dataset, with hourly data between 2015 and 2016. The accuracy of each model is assessed across 3, 6, 12, and 24 steps ahead horizons, with RRSE and CORR as the evaluation metrics.
Table 34: Comparison of the average accuracy of the different models on the Traffic dataset for 3, 6, 12, and 24 steps time horizons, expressed in terms of RRSE and CORR. The numbers corresponding to the highest accuracy are underlined.
Time horizon 3 steps 6 steps 12 steps 24 steps
Metrics RRSE CORR RRSE CORR RRSE CORR RRSE CORR
ADLGNN [art_83] 0.4047 0.9028 0.4201 0.8928 0.4299 0.8876 0.4416 0.8818
AGCRN 0.4379 0.8850 0.4635 0.8670 0.4694 0.8679 0.4707 0.8664
AGLG-GRU [art_79] 0.4173 0.8958 0.4722 0.8541 0.4427 0.8755 0.4526 0.8842
AR 0.5991 0.7752 0.6218 0.7568 0.6252 0.7544 0.6300 0.7519
ARIMA 0.5841 0.7959 0.6194 0.7655 0.6197 0.7652 0.6248 0.7610
GP 0.6082 0.7831 0.6772 0.7406 0.6406 0.7671 0.5995 0.7909
GTS 0.4665 0.8695 0.4779 0.8582 0.4792 0.8589 0.4766 0.8573
GWN 0.4484 0.8801 0.4689 0.8674 0.4725 0.8646 0.4741 0.8646
HyDCNN 0.4198 0.8915 0.4290 0.8855 0.4352 0.8858 0.4423 0.8819
LSNet 0.4777 0.8721 0.4893 0.8690 0.4950 0.8614 0.4973 0.8588
MAGNN [art_80] 0.4097 0.8992 0.4555 0.8753 0.4423 0.8815 0.4434 0.8813
MDST-GNN [art_90] 0.4162 0.8958 0.4461 0.8803 0.4377 0.8841 0.4452 0.8792
MTGNN 0.4162 0.8963 0.4754 0.8667 0.4461 0.8794 0.4535 0.8810
MTGODE [art_75] 0.4127 0.900 0.4259 0.8945 0.4329 0.8899 - -
MTHetGNN 0.4826 0.8643 0.5198 0.8452 0.5147 0.8744 0.5250 0.8418
MTNet 0.4764 0.8728 0.4855 0.8681 0.4877 0.8644 0.5023 0.8570
RNN-GRU 0.5358 0.8511 0.5522 0.8405 0.5562 0.8345 0.5633 0.8300
SARIMA 0.5823 0.7967 0.5974 0.7837 0.6002 0.7811 0.6151 0.7697
SCINet 0.4216 0.8920 0.4414 0.8809 0.4495 0.8772 0.4453 0.8825
SDGL [art_78] 0.4142 0.9010 0.4475 0.8825 0.4584 0.8760 0.4571 0.8766
SDLGNN 0.4053 0.9017 0.4209 0.8925 0.4313 0.8868 0.4444 0.8801
SDLGNN-Corr 0.4227 0.8937 0.4378 0.8846 0.4576 0.8746 0.4579 0.8784
TDG4-MSF [art_81] 0.4029 0.9014 0.4196 0.8925 0.4294 0.8864 0.4366 0.8834
Theta 0.6071 0.7768 0.6241 0.7606 0.6271 0.7591 0.6012 0.783
TPA-LSTM 0.4487 0.8812 0.4658 0.8717 0.4641 0.8717 0.4765 0.8629
TRMF 0.6708 0.6964 0.6261 0.7430 0.5956 0.7748 0.6442 0.7278
VAR-MLP 0.5582 0.8245 0.6570 0.7695 0.6023 0.7929 0.6146 0.7891
The accuracy of the models varies by horizon and metric. However, the best one in terms of RRSE over all horizons is the temporal decomposition enhanced graph neural network for multivariate time series forecasting (TDG4MSF) [art_81].
5.8 Other topics
This final subsection includes 20 papers that fall outside the previously discussed groups. The case studies include cellular traffic prediction, recommendation systems for user preferences, and other forecasting or classification problems. Given the heterogeneity of the papers, only a general overview is provided here, and no comparisons of the approaches or discussions of benchmarks are presented.
5.8.1 Overview
Since GNNs have recently gained considerable popularity, many recent papers focus on applications that do not relate to the categories discussed so far. A frequently discussed topic in this ”Other topics” category is cellular traffic prediction ([art_138], [art_144], [art_148]), in terms of SMS messages, calls, internet connections, phone’s traffic statistics or locations, which are essential for the cellular network resource management system. Another topic is modulation classification ([art_12], [art_13], [4]), which is the process of determining the modulation used at the transmitter based on observations of the received signal [modulation_classification_definition]. A further area of investigation is the study of user preferences ([art_147], [art_151]) to enable personalized experiences and recommendations.
Other applications discussed in the papers of this section are encrypted traffic classification [art_118], urban spatio-temporal event prediction [art_136], prediction of the quality of operational processes [art_137], seismic intensity forecasting [art_139] (with source code available at https://github.com/StefanBloemheuvel/GCNTimeseriesRegression), prediction of spatio-temporal dynamics with complex structures [art_140], tailing dam monitoring [art_141], forecasting of educational video engagement [art_142], subsurface production forecasting [art_143], time series prediction for data centers maintenance [art_145], hydraulic runoff [art_146], water demand forecasting [art_149], and human activity recognition [art_159] (with source code available at https://github.com/riktimmondal/HAR-Sensor).
5.8.2 Datasets
Among the papers selected here, there are only three shared datasets, two for modulation classification and one for cellular traffic prediction. However, the datasets are handled in different ways, either in terms of pre-processing techniques, selected time windows, or train/test split percentages. For this reason, the results of these papers are not directly comparable among each other.
The two datasets for modulation classification used by Refs. [art_12], [art_13] and [4] are RML2016.10a (https://www.kaggle.com/datasets/raindrops12/rml201610a, source paper: http://dx.doi.org/10.1007/978-3-319-44188-7_16) and RML2016.10b (https://www.kaggle.com/datasets/marwanabudeeb/rml201610b, source paper: https://pubs.gnuradio.org/index.php/grcon/article/view/11). They include 220,000 and 1.2 million samples respectively, in many modulation types. Each modulation type has 20 levels of signal-to-noise ratios at 2 dB intervals from -20 dB to 18 dB.
As for the cellular traffic prediction dataset used in Refs. [art_138] and [art_144], it was proposed by Telecom Italia and MIT Media Lab during the ”Telecom Italia Big Data Challenge”, and it is available at https://dataverse.harvard.edu/dataverse/bigdatachallenge. The investigated geographical area is divided in a grid, and each cell records the number of SMS messages, calls, and wireless network traffic data in a 10 minutes interval.
6 Discussion
The aim of this SLR is to provide a comprehensive and detailed overview of the use of spatio-temporal GNN models for time series classification and forecasting in various fields, and to assess their performance in different application domains. Existing literature has been collected and synthesized in order to present an overview of datasets, models, and tables of results to support researchers in their future work. In the Introduction two sets of research questions were defined: a generic set and a specific set. Now, the following discussion will answer those research questions. Actually, most of the general research questions were already addressed in the overview presented in Sec. 3. As for the tools, the GNN community mostly uses Python and PyTorch, although some researchers use TensorFlow. Unfortunately, only few papers provide a link to the source code of the proposed model. Almost all papers are funded by public or private entities.
As for the specific questions, the answers are the following.
SQ1 (Applications). The three most investigated fields are, in order, ”Mobility”, ”Environment”, and ”Generic”. Regarding the first two groups, their datasets can be naturally translated into graphs, which explains why they have been studied more extensively over time. In contrast, the ”Generic” group has recently gained a significant interest for the applicability to broader contexts. As for the differences between the various fields, the main ones are related to the definition of the graph, which is not always explicit, and the benchmarks used, often related to the mindset of specific communities. It is not easy to compare results across different applications and to determine the most promising fields. The studies in the ”Generic” group appear promising across a wide range of domains.
SQ2 (Graph construction). Most of the selected papers focus on pre-defined graph structures (when available), with the objective of extracting the maximum amount of information and enhancing the interpretability of the model. However, there is a recent growing interest in models that learn the graph structure and the edge weights themselves. This trend is expected to become more prevalent in the future.
SQ3 (Taxonomy). As observed in Ref. [survey_spatiotemporal], there are two main approaches to the design of spatio-temporal GNNs: one that treats the spatial and temporal components in separate modules, and one that integrates and processes them together. The analysis of the collected papers reveals that the most common approach in literature is that of separate modules. Specifically, researchers often address the spatial and temporal aspects of the problem independently from each other, and focus on each module separately. The proposed taxonomy of GNN models refers only to the spatial component, and the review indicates that the convolutional and the attentional approaches are the prevalent ones. Approximately 62% of the models are purely convolutional, 25% are purely attentional, and 8% are hybrid convolutional-attentional. With regard to the temporal component, the recurrent structures of GRUs and the attention mechanism are widely used.
SQ4 (Benchmark models). As for the benchmark models, there are many options available, and their choice depends on the specific application. In fields such as energy, Finance, health, and predictive monitoring, the focus tends to be on simpler statistical and classical machine learning benchmarks, with only a limited use of GNN benchmarks. In the ”Generic” group, many recent machine learning benchmark models based on the Transformer architecture have emerged. Notably, in ”Mobility”, ”Environment”, and ”Generic” groups (which are also the most investigated fields), there are many reference GNN benchmark models. The most prevalent benchmarks are the graph convolutional neural network with long-short term memory (GC-LSTM) [GC-LSTM], spatio-temporal graph convolutional network (STGCN) [STGCN], attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN], diffusion convolutional recurrent neural network (DCRNN) [DCRNN], spatio-temporal graph convolutional network (STGCN) [STGCN], and multivariate time series forecasting with graph neural network (MTGNN) [MTGNN].
SQ5 (Benchmark datasets). The datasets mentioned in the selected papers are closely related to the specific case of study. Even though some benchmark datasets are listed in the ”Generic” group, there is no common standard dataset for the entire GNN research community. It would be beneficial to agreeing on, and start adopting some of the most commonly used datasets such as the traffic ones, with the goal of developing a shared benchmark dataset for the whole research community. Hopefully, this review could serve this purpose by providing comprehensive tables of the results of the models on different datasets to facilitate comparison between spatio-temporal GNN models.
SQ6 (Modeling paradigms). As for the modeling paradigm, most of the selected papers work with a homogeneous graph, which models the relationships among multiple entities of the same nature. This is due to the fact that many basic GNN algorithms were originally developed for homogeneous graphs, and because it is easier to identify relationships between quantities of the same type. Also, in many cases, the focus is on multivariate series where the variables of interest and the target quantities are inherently of the same nature.
SQ7 (Metrics). Regarding the error metrics, their choice is highly dependent on the specific case study. However, the most common ones are mean absolute error (MAE), mean squared error (MSE), and mean absolute percentage error (MAPE) for forecasting problems, and accuracy for classification problems.
7 Limits, challenges, and future research directions
In this section the limitations and challenges of spatio-temporal GNN modeling will be discussed, together with directions for further research.
Comparability. As highlighted in Ref. [cit_rev2], the evaluation of GNN models has improved thanks to the introduction of the Open Graph Benchmark (OGB) [NEURIPS2020_fb60d411], which provides a standardized evaluation framework and a variety of benchmark graph datasets. However, there is currently no standardized benchmark for spatio-temporal GNNs. As a result, each model is evaluated on its own selection of datasets, and this fragmentation makes it difficult to compare the results of different studies. To alleviate this problem, this review presented all the information gathered from the selected papers, including datasets, benchmarks, codes, and tables of results. It is intended to serve as a detailed overview and a foundation for further exploration, in the hope that researchers will begin to examine the data collected in the presented tables and, over time, identify relevant datasets and benchmark models.
Reproducibility. The limited availability of links to repositories, source code, and datasets makes it difficult to evaluate the progress of knowledge in spatio-temporal GNNs. Moreover, in many papers the authors do not give detailed information about the model and the graph structure, like definition of the graph, number of nodes, calculation of edge weights. This issues complicate the verification of results and the assessment of the reproducibility of experiments, which are essential to drive further research in this field. Without access to these fundamental resources, it is extremely difficult to assess the accuracy and significance of the proposed models.
Explainability. A crucial aspect of GNN models is explainability, intended as the ability to interpret and understand their decision-making processes. Despite their attempt to explicitly model spatial relationships between series, GNNs are often considered as ”black box” models. None of the selected papers adequately tackle the concept of explainability. However, recent literature is beginning to investigate this important topic. For instance, few papers studied perturbation-based explanation methods, which however have proven to be somewhat ineffective as reported in Ref. [interpret_3]. Other approaches, such as GNNExplainer [interpret_2], aim to identify a small subset of node features that have a crucial role in GNN’s predictions. Recently, some researchers have started to explore graph counterfactuals as a means of generating explanations, as reviewed in Ref. [interpret_1].
Poor information capacity. Another limitation of GNNs lies in the poor information behind the design of the graph structure and the model definition. A poorly constructed graph with unrepresentative or overly sparse connections, as well as a lack of physical constraints in the equations of the model, can significantly affect the performance. Although efforts have been made to address these issues, such as integrating physical constraints through differential equations, these approaches are not always effective. Therefore, it is essential to develop new techniques to overcome these limitations.
Heterogeneity. The majority of current GNN models are designed to deal with homogeneous graphs, where nodes and edges are all of the same type. As a result, it is challenging to use these GNNs with heterogeneous graphs, which have different types of nodes and edges or different inputs. Therefore, further research is necessary to develop GNN models that can effectively capture the interactions between various types of nodes and edges.
Scalability. Spatio-temporal GNNs are widely used for modeling and analyzing large and complex time series-based network structures. However, GNN models often require a significant amount of memory to compute the adjacency matrix and node embeddings, especially in the case of dynamic graph structures. This scalability challenge results in high computational costs, and in many cases the necessity to use significant GPU resources. Therefore, the development of more scalable GNN models is crucial to facilitate time series analysis even in environments with limited computing power.
8 Conclusions
This paper presented the results of a SLR on the application of spatio-temporal GNN models to time series classification and forecasting problems in different fields. Lately, GNNs have gained significant popularity due to their ability to process graph-structured data. This has led in more recent years to the development of spatio-temporal GNNs in the field of time series analysis, due to their ability to model dependencies between variables and across time points.
This SLR has brought forward on two set of questions: generic questions which can be answered at the level of bibliographic overview, and more specific questions regarding particular aspects of the proposed spatio-temporal GNN models. By answering these questions, it has emerged that the majority of the models in the selected papers is characterized by a convolutional spatial aggregation. However, several graph attentional models are also emerging. In addition, while the majority of the selected papers focus on models with a pre-defined graph structure, an increasing number of studies, particularly in the ”Generic” group, are beginning to develop models that learn the graph structure autonomously.
A first key point which results from the presented overview is that the current literature on spatio-temporal GNNs is very fragmented, as it was pictorially reported in Fig. 4. This can be attributed to the fact that the involved researchers come from different communities with focus on specific application domains. As a result, there seems to be a lack of standardised datasets or benchmarks. The objective of this review was also to collect information on datasets, proposed models, links to source codes, benchmarks, and results, in order to provide a foundation for future studies. A second major key point resulting from the proposed SLR seems hence to be that there is a need for the GNN research community to work on common datasets and develop comparable and reproducible models. This would enhance transparency and make it easier to evaluate advancements in the field and to share them more easily.
Appendix A List of selected journal papers
Tab. LABEL:t:selected_papers lists all the journal papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (e.g., classification or forecasting).
Table 35: List of selected journal papers, with year, group they belong to, case study, and nature of the task.
Ref. Year Group Case study Task
[art_15] 2023 Energy Heat load Forecasting
[art_16] 2023 Energy Electricity net load Forecasting
[art_17] 2023 Energy Residential load Forecasting
[art_18] 2023 Energy Electricity load Forecasting
[art_19] 2023 Energy Photovoltaic power Forecasting
[art_20] 2022 Energy Wind power Forecasting
[art_21] 2022 Energy Photovoltaic power Forecasting
[art_22] 2022 Energy Photovoltaic power Forecasting
[art_23] 2022 Energy Power system transient dynamics Forecasting
[art_24] 2020 Energy Wind power Forecasting
[art_25] 2022 Energy Urban buildings energy consumption Forecasting
[art_26] 2023 Energy Solar power Forecasting
[art_27] 2022 Energy Wind power Forecasting
[art_1] 2023 Environment Air quality index Forecasting
[art_2] 2023 Environment PM 2.5 concentration Forecasting
[art_3] 2023 Environment PM 2.5 concentration Forecasting
[art_4] 2023 Environment PM 2.5 concentration Forecasting
[art_5] 2023 Environment Air quality index Forecasting
[art_6] 2023 Environment PM 2.5 concentration Forecasting
[art_7] 2023 Environment PM 2.5 concentration Forecasting
[art_8] 2022 Environment PM 2.5 concentration Forecasting
[art_28] 2023 Environment Sea temperature Forecasting
[art_29] 2023 Environment Wind speed Forecasting
[art_30] 2023 Environment Water quality Forecasting
[art_31] 2023 Environment Wind speed Forecasting
[art_32] 2023 Environment Significant wave height Forecasting
[art_33] 2023 Environment Sea temperature Forecasting
[art_34] 2023 Environment Wind speed Forecasting
[art_35] 2023 Environment Sea temperature Forecasting
[art_36] 2023 Environment Sea temperature Forecasting
[art_38] 2022 Environment Frost Classification/forecasting
[art_39] 2022 Environment Sea temperature Forecasting
[art_40] 2021 Environment Wind speed Forecasting
[art_41] 2021 Environment Sea temperature Forecasting
[art_42] 2022 Environment Water quality Forecasting
[art_43] 2023 Environment Groundwater level Forecasting
[art_44] 2023 Environment Wind speed Forecasting
[art_45] 2023 Environment Rainfall Forecasting
[art_46] 2023 Environment Wind speed Forecasting
[art_47] 2023 Environment Sea temperature Forecasting
[art_59] 2022 Finance Stock prediction Classification
[art_60] 2023 Finance Stock prediction Classification
[art_62] 2023 Finance Stock prediction Classification
[art_63] 2023 Finance Stock prediction Classification
[art_64] 2022 Finance Investment prediction Forecasting
[art_65] 2023 Finance Stock prediction Classification
[art_66] 2022 Finance Financial prediction Classification
[art_67] 2023 Finance Trading Classification
[art_68] 2023 Finance Sale prediction Forecasting
[art_69] 2023 Generic Batch workloads Forecasting
[art_70] 2023 Generic Node classification for websites Classification
[art_71] 2023 Generic Clinical risk Classification
[art_72] 2023 Generic Sensor network Forecasting
[art_73] 2023 Generic Multivariate time series Forecasting
[art_74] 2023 Generic Multivariate time series Forecasting
[art_75] 2023 Generic Multivariate time series Forecasting
[art_76] 2023 Generic Long-term series forecasting Forecasting
[art_77] 2023 Generic Multivariate time series Forecasting
[art_78] 2023 Generic Multivariate time series Forecasting
[art_79] 2023 Generic Multivariate time series Forecasting
[art_80] 2023 Generic Multivariate time series Forecasting
[art_81] 2023 Generic Multivariate time series Forecasting
[art_82] 2023 Generic Multivariate long sequence time series Forecasting
[art_83] 2023 Generic Multivariate time series Forecasting
[art_84] 2022 Generic Multivariate time series Classification/forecasting
[art_85] 2023 Generic Multivariate time series Forecasting
[art_86] 2022 Generic Multivariate time series Forecasting
[art_87] 2022 Generic Multivariate time series Forecasting
[art_88] 2023 Generic Multi-channel time series Forecasting
[art_89] 2023 Generic Multivariate time series Forecasting
[art_90] 2022 Generic Multivariate time series Forecasting
[art_160] 2023 Generic Impact of graph construction Forecasting
[art_202] 2021 Generic Generic Classification/forecasting
[art_14] 2023 Health Epilepsy diagnosis Classification
[art_48] 2023 Health Epidemic prediction Forecasting
[art_49] 2023 Health Epidemic prediction Forecasting
[art_50] 2022 Health Epidemic prediction Forecasting
[art_51] 2021 Health Epidemic prediction Forecasting
[art_52] 2022 Health Epidemic prediction Forecasting
[art_92] 2023 Health Disease diagnosis Classification
[art_189] 2023 Health Clinical risk classification Classification
[art_93] 2022 Mobility Urban traffic Forecasting
[art_94] 2021 Mobility Crowd flow Forecasting
[art_95] 2023 Mobility Flight delays, urban traffic Forecasting
[art_96] 2022 Mobility Crowd flow Forecasting
[art_97] 2023 Mobility Urban traffic Forecasting
[art_98] 2022 Mobility Urban traffic Forecasting
[art_99] 2023 Mobility Urban traffic Forecasting
[art_100] 2023 Mobility Urban traffic Forecasting
[art_101] 2023 Mobility Bus station profiling Profiling
[art_102] 2021 Mobility Passenger demand Forecasting
[art_103] 2023 Mobility Urban traffic Forecasting
[art_104] 2022 Mobility Internet of vehicles Forecasting
[art_105] 2020 Mobility Urban traffic Forecasting
[art_106] 2023 Mobility Urban traffic Forecasting
[art_107] 2023 Mobility Metro flow Forecasting
[art_108] 2023 Mobility Urban traffic Forecasting
[art_109] 2023 Mobility Urban traffic Forecasting
[art_110] 2023 Mobility Urban traffic Forecasting
[art_113] 2023 Mobility Urban traffic Forecasting
[art_114] 2023 Mobility Bike station demand Forecasting
[art_115] 2023 Mobility Urban traffic Forecasting
[art_116] 2023 Mobility Urban traffic Forecasting
[art_117] 2023 Mobility Urban traffic Forecasting
[art_119] 2022 Mobility Taxi demand Forecasting
[art_120] 2022 Mobility Urban traffic Forecasting
[art_121] 2022 Mobility Metro passengers Forecasting
[art_122] 2022 Mobility Urban traffic Forecasting
[art_123] 2021 Mobility Urban traffic Forecasting
[art_124] 2022 Mobility Urban traffic Forecasting
[art_125] 2023 Mobility Urban traffic Forecasting
[art_126] 2020 Mobility Urban traffic Forecasting
[3] 2021 Mobility Urban traffic Forecasting
[art_128] 2022 Mobility Urban traffic Forecasting
[art_129] 2022 Mobility Flight delay Forecasting
[art_130] 2022 Mobility Urban traffic Forecasting
[art_131] 2023 Mobility Urban traffic Forecasting
[art_132] 2022 Mobility Urban traffic Forecasting
[art_133] 2022 Mobility Bicycle demand Forecasting
[art_135] 2022 Mobility Urban traffic Forecasting
[art_12] 2023 Other Modulation classification Classification
[art_13] 2022 Other Modulation classification Classification
[art_118] 2023 Other Encrypted traffic classification Classification
[art_136] 2022 Other Spatial temporal events Forecasting
[art_137] 2023 Other Quality indicators Forecasting
[art_138] 2023 Other Cellular traffic Forecasting
[art_139] 2023 Other Seismic data Forecasting
[art_140] 2023 Other Structural dynamics on irregular domains Forecasting
[art_141] 2023 Other Tailing dam monitoring Forecasting
[art_142] 2023 Other Video engagement Forecasting
[art_143] 2023 Other Subsurface production Forecasting
[art_144] 2023 Other Cellular traffic Forecasting
[art_145] 2023 Other Data centers maintenance Forecasting
[art_146] 2023 Other Hydraulic runoff Forecasting
[art_147] 2022 Other App popularity Forecasting
[art_148] 2021 Other Cellular traffic Forecasting
[art_149] 2022 Other Water demand Forecasting
[art_151] 2023 Other User preference Edge prediction
[art_159] 2021 Other Human activity recognition Classification
[4] 2023 Other Modulation classification Classification
[art_9] 2022 Predictive monitoring Anomaly detection Classification
[art_10] 2023 Predictive monitoring Anomaly detection Forecasting
[art_53] 2023 Predictive monitoring Fault diagnosis Classification
[art_54] 2023 Predictive monitoring Fault diagnosis Classification
[art_55] 2021 Predictive monitoring Fault diagnosis Classification
[art_56] 2023 Predictive monitoring Fault diagnosis Classification
[art_57] 2023 Predictive monitoring Fault diagnosis Classification
[art_58] 2023 Predictive monitoring Fault prediction Forecasting
[art_152] 2023 Predictive monitoring Remaining useful life Forecasting
[art_153] 2023 Predictive monitoring Remaining useful life Forecasting
[art_154] 2023 Predictive monitoring Remaining useful life Forecasting
[art_155] 2021 Predictive monitoring Remaining useful life Forecasting
[art_156] 2022 Predictive monitoring Remaining useful life Forecasting
[art_157] 2022 Predictive monitoring Remaining useful life Forecasting
[art_158] 2023 Predictive monitoring Remaining useful life Forecasting
[art_192] 2023 Predictive monitoring Fault diagnosis Classification

### v2

@@
 Abstract
-In recent years, spatio-temporal graph neural networks (GNNs) have attracted considerable interest in the field of time series analysis, due to their ability to capture dependencies among variables and across time points.
-The objective of the presented systematic literature review is hence to provide a comprehensive overview of the various modeling approaches and application domains of GNNs for time series classification and forecasting.
-A database search was conducted, and over 150 journal papers were selected for a detailed examination of the current state-of-the-art in the field.
+In recent years, spatio-temporal graph neural networks (GNNs) have attracted considerable interest in the field of time series analysis, due to their ability to capture, at once, dependencies among variables and across time points.
+The objective of this systematic literature review is hence to provide a comprehensive overview of the various modeling approaches and application domains of GNNs for time series classification and forecasting.
+A database search was conducted, and 281 journal papers were selected for a detailed examination of the current state-of-the-art in the field.
 This examination is intended to offer to the reader a comprehensive collection of proposed models, links to related source code, available datasets, benchmark models, and fitting results.
 All this information is hoped to assist researchers in future studies.
-To the best of our knowledge, this is the first systematic literature review presenting a detailed comparison of the results of current spatio-temporal GNN models in different domains.
-In addition, in its final part this review discusses current limitations and challenges in the application of spatio-temporal GNNs, such as comparability, reproducibility, explainability, poor information capacity, and scalability.
+To the best of our knowledge, this is the first systematic literature review presenting a detailed comparison of results of current spatio-temporal GNN models applied to different domains.
+In its final part, this review discusses current limitations and challenges in the application of spatio-temporal GNNs, such as comparability, reproducibility, explainability, poor information capacity, and scalability.
 1 Introduction
-In recent years, graph neural networks (GNNs) have emerged as a powerful class of artificial neural network models for processing data that can be represented as graphs.
-GNNs are in fact particularly well-suited for a wide range of practical and engineering applications where data naturally lend themselves to be represented as graph structures, such as applications to transportation networks, image analysis, and natural language processing.
-The intuitive concept behind this approach is that nodes in a graph can be put in correspondence to objects or concepts, whereas edges can be assumed to represent their mutual relationships.
+In recent years, graph neural networks (GNNs) have emerged as a powerful class of artificial neural network models aimed at processing data that can be represented as graphs [Zheng2024].
+GNNs are in fact particularly well-suited for a wide range of practical and engineering applications where data naturally lend themselves to be represented as graph structures, such as applications to transportation networks [review_traffico_jiang], image analysis [GNN_image], classification of molecular structure [GNN_molecule], and natural language processing [GNN_NLP].
+The intuitive concept behind the GNN approach is that the nodes of a graph can be put in correspondence to objects or concepts, whereas the edges can be assumed to represent their mutual relationships [Scarselli2009].
 Data for these quantities can then be processed at once by a GNN model.
-GNNs can be used for three different classes of problems: at graph level, at edge level, and at node level.
-In a graph-level problem, the goal is to predict a property of a graph based on its entire structure, rather than on individual nodes or edges.
+GNNs can be used at three different graph levels, related to three distinct classes of problems: at graph level, at edge level, and at node level [review_traffico_con_codici].
+In graph-level problems, the goal is to predict a property of a graph based on its entire structure, rather than on individual nodes or edges.
 For example, a molecule from a sample can be represented as a graph, and global properties of this molecule can be inferred from data coming from its entire structure in relation to the sample.
-In an edge-level problem, the goal is to predict the presence or absence of edges between pairs of nodes.
+In edge-level problems, the goal is to predict the presence or absence of edges between pairs of nodes.
 As an example, this task can be used in recommendation systems to predict potential connections between users and items on the basis of past interactions.
-In node-level problems, the goal is to predict the identity or role of each node within a graph, in order to solve either a classification or a regression task.
-Noticeably, regression tasks can be ”autoregressions”, that is, they can include time.
-This review examines then the specific use of GNNs on time series related tasks, with a particular focus on time series classification and forecasting.
-The idea behind the use of GNNs for time series problems is that GNNs can be made capable of capturing complex relationships, both inter-variable (connections between different variables within a multivariate series) but also inter-temporal (dependencies between different points in time) at once.
-This results in a spatio-temporal GNN approach, where the spatial dimension is related to the multivariate framework, and the temporal dimension is related to the temporal nature of the data.
+In node-level problems, the goal is to predict the role or value of each node within a graph, in order to solve either a classification or a regression task.
+Noticeably, regression tasks can sometimes be put in the form of temporal “autoregressions”, that is, they can include time.
+The inclusion of time in GNN modeling has proven to have far-reaching consequences.
+For this reason, this review examines the available journals’ literature on the specific use of GNNs on time series related tasks, with a particular focus on time series classification and forecasting.
+The idea behind the use of GNNs for time series problems is that GNNs can be made capable of capturing complex relationships, both inter-variable (connections between different variables within a multivariate series) but also inter-temporal (dependencies between different points in time) at once [review_Alippi].
+The inclusion of this capability in GNN models results in a spatio-temporal GNN approach.
+Within this approach, the spatial dimension is related to the multivariate framework, where nodes typically correspond to different features of a multivariate time series at a given time step $t$, and edges to their relationships, and the temporal dimension is related to the temporal nature of the data.
 Namely, spatio-temporal GNNs are a class of GNN models designed to handle data in both spatial and temporal dimensions at once.
-In this case, in a spatio-temporal graph, data in each node or edge can evolve over time, and the challenge is here to model the related dynamic interactions over time.
-Notwithstanding all this potential interest, to the best of our knowledge there is a lack of systematic literature reviews (SLRs) on spatio-temporal GNN models for time series applications.
-Most of the existing surveys are not SLRs, they are not recent, or only focus on a limited number of application domains.
-Individually, recent reviews focus for example on algorithmic features [cit_rev1], [cit_rev2], [cit_rev3], [survey_Lucheroni], GNN characterization [cit_rev8], specific fields [cit_rev4], and distributed training [cit_rev5], [cit_rev6].
-Only a SLR on spatio-temporal GNNs was indeed recently published by Zeghina et al.
-[survey_spatiotemporal], but due to the restrictive search query, it covers only 52 publications.
-The intention of the SLR presented here is to provide a broad and comprehensive overview of the applications of spatio-temporal GNNs for time series classification and forecasting in different areas.
-This SLR also aims to assess if the current popularity of spatio-temporal GNN models is indeed due to their effectiveness, and to evaluate their efficacy and accuracy in different fields.
-Additionally, it collects results and benchmarks to assist researchers in their work by synthesizing the available literature on GNN applications.
-To the best of our knowledge, this is the first review to present comprehensive tables with all the results from the various models and benchmarks proposed in a highly fragmented literature.
-It is thus hoped that this collected and distilled knowledge can become a valuable resource and reference for researchers.
-In this review two sets of questions will be posed.
-The answers to the first set want to provide a general overview of the publications in the field, while the answers to the second set focus on specific aspects of the proposed GNN models.
+In a spatio-temporal graph, data in each node or edge can evolve over time, and the challenge is then to model their related dynamic interactions over time.
+Notwithstanding all the potential interest for such important features, to the best of our knowledge there is a lack of systematic literature reviews (SLRs) on spatio-temporal GNN models for time series applications.
+Most of the existing surveys are not SLRs, or they are not recent, or only focus on a limited number of application domains and specific issues not necessarily related to time series.
+For example, recent reviews focus on algorithmic features [cit_rev2, cit_rev3, survey_Lucheroni, survey_Chikwendu], GNN characterization [cit_rev8, review_Alippi], specific fields [cit_rev4], and distributed training [cit_rev5].
+Only one SLR on spatio-temporal GNNs was actually recently published, by [survey_spatiotemporal].
+However, due to the restrictive search query used in that paper, that SLR covers 52 publications only.
+In addition, [survey_kumar_spatiotemporal_data] proposed a recent survey on spatio-temporal predictive modeling techniques for different domains.
+However, the authors approach the topic from a more general perspective, analyzing different machine learning models without placing too much emphasis on spatio-temporal GNNs.
+As an organized summary of these contributions, a list with a short description of the main reviews on spatio-temporal GNN models is provided in Tab.
+LABEL:t:tabella_altre_review.
+Table 1: Short description of main reviews on spatio-temporal GNN models.
+Review Description
+Wu et al.
+(2020) [cit_rev3] One of the first reviews on the use of GNNs in machine learning and data mining.
+It introduces a taxonomy that classifies GNN architectures into four categories: recurrent, convolutional, graph autoencoders, and spatial-temporal GNNs.
+The focus of this review is primarily on early GNN models, rather than applications on time series data.
+Zhou et al.
+(2020) [review_traffico_con_codici] The authors outline a general design pipeline for GNN models, covering key aspects such as the graph structure, pooling techniques, loss function design, and components of the propagation module (including recurrent and convolutional operators, and skip connections).
+They present some applications of GNNs, including physical systems modeling, chemical analysis, text processing, image classification, recommendation systems, and analysis of other types of data.
+Jin et al.
+(2023) [cit_rev4] Review on the applications of spatio-temporal GNNs to predictive learning in urban computing.
+In the specific, the authors describe four methods for the construction of pre-defined graphs: topology-based, distance-based, similarity-based and interaction-based approaches.
+Then, they discuss various design strategies for spatio-temporal GNNs and their combination with some advanced learning frameworks, such as adversarial learning, continuous learning, and physics informed learning.
+Li et al.
+(2023) [li2023graphneuralnetworkspatiotemporal] The authors aim to provide a systematic review of spatio-temporal mining using GNNs.
+After providing an overview of the graph structure and the spatial and temporal components of GNNs, the authors describe some applications in fields such as transportation, disaster management, environment, power systems, and neuroscience.
+However, unlike our review, they do not outline the selection process of the papers or include detailed and comprehensive tables that cover datasets, benchmarks and results.
+Lin et al.
+(2023) [cit_rev5] This review is specialized on distributed training of GNNs.
+It investigates optimization techniques used in this context, including parallel mini-batch generation, dynamic mini-batch allocation, mini-batch transmission pipelining, and parallel aggregation with edge partitioning.
+Longa et al.
+(2023) [cit_rev2] The authors offer an overview of GNNs for temporal graphs.
+They describe supervised and unsupervised learning tasks on temporal graphs, and present a taxonomy of temporal GNNs that groups the model in snapshot-based and event-based approaches.
+Chen et al.
+(2024) [survey_Lucheroni] One of the first reviews on the use of GNNs for time series modeling.
+The authors thoroughly describe a wide range of models for time series classification, forecasting, and anomaly detection, based either on recurrent/convolutional GNNs and attentional GNNs.
+Moreover, they provide a brief comparison of the models’ performances in traffic forecasting and anomaly detection in water treatment.
+Chikwendu et al.
+(2024) [survey_Chikwendu] This survey provides an overview of graph neural architecture search, and explores methods for representing nodes, edges, and sub-graphs, which add context and semantics to graphs.
+The authors also describe some applications, including computer vision, brain networks, recommendation system, biomedical application, natural language processing, and traffic.
+Feng et al.
+(2024) [feng2024comprehensivesurveydynamicgraph] This survey examines 81 dynamic GNN models.
+It introduces a novel taxonomy that categorizes GNNs into discrete-time dynamic graphs and continuous-time dynamic graphs based on their structures, features, and dynamic modeling methods.
+In addition, the authors perform numerical experiments which evaluate 12 significant models on 6 graph datasets, which are not inherently related to time series.
+Jin et al.
+(2024) [review_Alippi] The authors present a comprehensive review of GNNs for time series analysis, covering forecasting, classification, anomaly detection, and imputation.
+They categorize spatio-temporal GNNs according to their three main components: the spatial module, the temporal module, and the overall model architecture.
+Although they focus on the same topic as our SLR, they do not provide a detailed comparison of the results of existing models in the literature.
+Khemani et al.
+(2024) [Khemani2024] The review discusses general concepts of GNNs, including architectures, techniques, datasets, applications, and challenges, offering a broad analysis that emphasizes these aspects without focusing specifically on time series analysis.
+It provides a brief description of papers in the literature, detailing application areas, datasets used, models applied, and performance evaluation.
+Kumar et al.
+(2024) [survey_kumar_spatiotemporal_data] This SLR includes 200 papers from high-impact journals and prominent conferences on spatio-temporal predictive modeling techniques across different domains.
+Although its structure is similar to that of our review, its focus is broader and not specifically on spatio-temporal GNNs, as it also covers and discusses statistical and non-GNN machine learning models.
+Wu et al.
+(2024) [cit_rev8] This review offers a comprehensive survey of GNN models from the perspective of computation theory.
+In particular, the authors compare existing GNN models through various perspectives: time, memory, communication, parallelism, and execution.
+Zeghina et al.
+(2024) [survey_spatiotemporal] This SLR includes 52 papers on spatio-temporal GNNs across various application domains.
+Compared to our SLR, it covers a limited set of papers due to a restricted search query and it does not provide detailed and comprehensive tables of datasets, benchmarks, and model results.
+The intention of the SLR presented here is thus to provide a broad and comprehensive overview of the applications of spatio-temporal GNNs for time series classification and forecasting in different areas.
+Necessarily, this SLR also aims to assess if the current popularity of spatio-temporal GNN models is indeed due to their effectiveness, and to evaluate their efficacy and accuracy in different fields.
+Additionally, it collects results and benchmarks used in the discussed papers in order to assist researchers in their work by synthesizing the available literature on GNN applications.
+To the best of our knowledge, this is the first review presenting comprehensive tables with all the results from the various models and benchmarks proposed in a highly fragmented literature.
+This synthesis and comparison of the results is performed by gathering data from the existing literature, without conducting any numerical experiment on data.
+Moreover, given the large number of selected papers, it was not feasible to include a description of all the proposed models.
+To compensate this limitation, we provide a description of the most commonly used benchmarks and the most accurate models within each group.
+In case, for more detailed descriptions of experimental setups and data pre-processing techniques, the reader is referred to the original papers.
+It is hoped that this collected and distilled knowledge can anyway become a valuable resource and reference for researchers.
+During the development of this review, we formulated two sets of questions, one general and one specific, which are outlined below.
+The answers to the first set want to provide a general overview of the publications in the field, whereas the answers to the second set focus on specific aspects of the proposed GNN models.
 The general overview questions (GQs) are:
   1. GQ1)
-    Trend: What is the temporal trend of publications on GNN models? Is there a growing interest in GNNs?
+    Trend: What is the temporal trend of the number of publications on GNN models? Is there a growing interest in GNNs?
   2. GQ2)
@@
   3. GQ3)
-    Journals: In which journals are the papers on GNNs published? Are these journals chosen because of the specific domain of application or not?
+    Journals: In which journals are the papers on GNNs published? Are these journals chosen because of the specific domain of application, or not?
   4. GQ4)
@@
   3. SQ3)
-    Taxonomy: Among the various taxonomy classes, which are the most common? Are there some recurring mechanisms?
+    Taxonomy: Among the various taxonomy classes for GNNs, which are the most common? Are there some recurring mechanisms?
   4. SQ4)
-    Benchmark models: Which are the most common benchmark models? Are they classical ML models or also GNN models?
+    Benchmark models: Which are the most commonly used benchmark models? Are they non-GNN machine learning models or also GNN models?
   5. SQ5)
-    Benchmark datasets: Which are the most common benchmark datasets?
+    Benchmark datasets: Which are the most commonly used benchmark datasets?
   6. SQ6)
-    Modeling paradigms: What is the most common paradigm?
+    Modeling paradigms: Which is the most common modeling paradigm?
 Is it modeling complex interacting systems, or modeling a system with multiple interacting quantities?
-Specifically, do graphs typically condense the relationships among multiple entities, or do they describe different aspects of the same entity?
-In other words, is the graph structure homogeneous or heterogeneous?
+Specifically, do graphs typically condense the relationships among multiple entities, or they describe different aspects of the same entity?
+In other words, is the used graph structure homogeneous or heterogeneous?
   7. SQ7)
     Metrics: Which are the most common metrics used to assess the accuracy of a given model?
-The answers to both the general and the specific questions will be integrated into the following discussion in order to provide a comprehensive overview of the present state-of-the-art of spatio-temporal GNN models.
-A general overview of such GNN models will be assembled by highlighting similarities and differences between the approaches across different fields.
-However, in case, for a more detailed analysis of a specific discipline or other aspects of GNN models, the reader is invited to consult the specialized reviews mentioned before ([cit_rev1], [cit_rev2], [cit_rev3], [survey_Lucheroni], [cit_rev8], [cit_rev4], [cit_rev5], [cit_rev6]).
+As the discussion will progress, the answers to both the general and the specific questions will be integrated in it.
+This will hopefully provide a comprehensive overview of the state-of-the-art of spatio-temporal GNN models.
+At the same time, a general overview of such GNN models will be assembled by highlighting similarities and differences between the approaches across different fields.
+In case of need of an even more detailed analysis of a specific discipline, or other aspects of GNN models, the reader is invited to consult the specialized reviews mentioned before [cit_rev1, cit_rev2, cit_rev3, survey_Lucheroni, cit_rev8, cit_rev4, cit_rev5, cit_rev6].
 The paper is organized as follows.
@@
 8 concludes the review.
-Appendix  A provides a list of all journal papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (e.g., classification or forecasting).
+A provides a list of all journal papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (e.g., classification or forecasting).
+All links mentioned in the papers are included in this review.
+For the reader’s convenience, links that no longer work at the time of submission are marked with a .
 2 Methodology for the systematic literature review
-To conduct this SLR, four primary databases were consulted: Scopus, IEEE Xplore, Web of Science, and ACM.
+In order to conduct this SLR, four databases of primary importance were consulted: Scopus, IEEE Xplore, Web of Science, and ACM.
 These databases were chosen for their comprehensive coverage of academic and scientific literature, ensuring a broad and inclusive search across multiple disciplines.
 Prior to commencing the literature search, some exclusion criteria were defined.
-First, papers from the current year 2024 were excluded because it has not yet ended, hence the review only covers up to year 2023.
-Second, only journal articles were considered, in order to guarantee the inclusion of ”established” sources, whereas both conference papers and book chapters were excluded.
-The exclusion of conference papers is also made in order to keep the number of papers manageable.
-However, it is important to notice that relevant conference papers which provide useful benchmarks are anyway referenced throughout the text.
-Third, only articles published in English were included because of facility of fruition and comparison.
-Review articles are also not included, although they may be mentioned in the test.
-An advanced search query submitted to the databases was designed to capture the breadth of the topic while maintaining specificity.
-The query used is (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting).
-The first two elements of query aim to restrict the pool of GNN models to the time series domain, and the second part focuses the search on classification and forecasting applications.
-The selection process of the papers followed the multi-stage approach described below:
+First, papers from the current year 2025 were excluded because it has not yet ended, hence the review will only cover the literature up to the end of year 2024.
+Second, only journal articles were considered, whereas both book chapters and conference papers were excluded.
+Such a selection is made for two reasons.
+Whereas it is true that new research directions often emerge from conference papers, which can contain outstanding results, most often the results presented in this form are considered preliminary even by the authors themselves, frequently because felt as lacking the stability required for standard journal publication.
+In addition, the exclusion of conference papers is made for readability reasons, specifically to ensure that a manageable number of articles can be discussed in full within a limited number of pages.
+Nevertheless, conference papers that provide useful benchmarks and which are frequently referenced as accurate benchmark models in other journal articles, are anyway referenced throughout the text.
+An advanced search query to be submitted to the databases was designed to capture the breadth of the topic while maintaining specificity.
+Its general form is (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting)  [^1] Specific search queries for the different databases: Scopus: TITLE-ABS-KEY((("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting)) AND PUBYEAR < 2025 AND (EXCLUDE(DOCTYPE,"cr") OR EXCLUDE (DOCTYPE,"cp")).
+IEEE Xplore: (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting).
+Filters applied: Journals, 2000 - 2024.
+Web of Science: (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting).
+Filters applied: Article, 2000 - 2024.
+ACM: [[All: "graph neural network"] OR [All: gnn]] AND [All: "time series"] AND [[All: classification] OR [All: forecasting]] AND [E-Publication Date: (01/01/2000 TO 12/31/2024)].
+.
+The first two elements of the query aim to restrict the pool of GNN models to the time series domain, and the second part focuses the search on classification and forecasting applications.
+Overall, the selection process of the papers followed the multi-stage approach described below:
   1. 1.
@@
   5. 5.
-    Data extraction: relevant data were extracted from the included studies, and analyzed in order to better discuss the approach and present the most common datasets and models for each identified field.
-A total of 473 records were identified, 188 from Scopus, 31 from IEEE, 104 from Web of Science, and 150 from ACM.
-A scheme of this selection process is sketched in Fig.
+    Data extraction: relevant data and features were extracted from the included studies, and analyzed in order to better discuss the approach and present the most common datasets and models for each identified field.
+A graphical representation of this selection process is provided in Fig.
 1.
-First, 142 duplicates were removed, then the titles and abstracts of the remaining 331 records were manually evaluated to select the pertinent papers.
-Among these, 170 were discarded because they were considered outside the scope of this review, and 5 were discarded because they were not in English.
-Hence, 156 records were included in the review.
-Figure 1: PRISMA flowchart to summarize the identification of the studies and the selection process, adapted from Ref. [PRISMA_statement].
-Next, Sec. 3 provides an overview of the selected publications.
+A total of 879 records were identified, 355 from Scopus, 67 from IEEE Xplore, 194 from Web of Science, and 263 from ACM.
+After that, 253 duplicates were removed.
+Then, the titles and abstracts of the remaining 626 records were manually evaluated to select the pertinent papers.
+Among these, 333 were discarded because they were considered outside the scope of this review, and 12 were discarded because they were not in English.
+Hence, 281 records were included in the review.
+Figure 1: PRISMA flowchart summarizing the identification and selection process, adapted from [PRISMA_statement].
+The flowchart is structured vertically into three key phases: identification (collection of records from databases and removal of duplicates), screening (assessment for eligibility and exclusion of out-of-scope studies), and inclusion (inclusion of papers in the review).
+The boxes on the right indicate the number of studies excluded at each stage.
+As for conference papers, a total of 1,765 records were identified across the four included databases.
+Assuming the same proportion of duplicates and articles to be excluded as those found for journal papers, about 600 conference papers would remain, making the total number of papers excessive.
+As previously stated, these conference papers have been excluded from this review.
+Next, Sec.
+3 provides an overview of the selected journal publications.
 3 Overview of the publications
-This section provides a comprehensive overview of the 156 selected journal articles.
+This section provides a comprehensive overview of the 281 selected journal papers.
 After their selection, the papers were grouped according to their domain of application, as better explained in the thematic analysis of Sec.
 5.
-In the following, some bibliographic information about the reviewed articles is presented.
+In the following, some overall bibliographic information about the reviewed articles is presented.
 Fig.
-2 illustrates the number of publications over time across different groups.
+2 displays the number of publications over time across different groups.
 A cumulative positive trend can be observed over time, reflecting an increasing interest in spatio-temporal GNN models across various fields, since 2020.
-The three most common groups that have seen the greatest increase in publications are ”Environment”, ”Generic”, and ”Mobility”.
+The four most common groups that have seen the greatest increase in publications are “Environment”, “Generic”, “Mobility”, and “Predictive monitoring”.
 For more details, see also the group numerosity reported in Tab.
-1.
+2.
 Figure 2: Number of journal publications over time across the different groups.
 Table 2: Pivot table illustrating the number of publications by year and group.
-Field / Year 2020 2021 2022 2023 Total
-Energy 1 6 6 13
-Environment 2 4 21 27
-Finance 3 6 9
-Generic 1 4 19 24
-Health 1 2 5 8
-Mobility 2 4 15 18 39
-Other 2 4 14 20
-Predictive monitoring 2 3 11 16
-Total 3 12 41 100 156
-An inspection of the publication sources indicates that 95 different journals have published articles on GNN models for time series forecasting or classification.
+Field / Year 2020 2021 2022 2023 2024 Total
+Energy 1 - 6 6 20 33
+Environment - 2 4 21 16 43
+Finance - - 3 6 7 16
+Generic - 1 4 19 20 44
+Health - 1 2 5 7 15
+Mobility 2 4 15 15 30 66
+Other - 2 4 14 9 29
+Predictive monitoring - 2 3 11 19 35
+Total 3 12 41 97 128 281
+The catalog of the publication sources indicates that, in the studies period, 145 different journals have published articles on GNN models for time series forecasting or classification.
 Tab.
-2 lists the 15 journals which contain the highest number of published papers.
-Table 2: Top 15 journals with the highest number of published paper.
-Journal Number of papers
-IEEE Access 7
-ACM Transactions on Knowledge Discovery from Data 6
-Applied Intelligence 6
-IEEE Transactions on Instrumentation and Measurement 6
-ACM Transactions on Intelligent Systems and Technology 5
-Expert Systems with Applications 5
-Sensors 4
-Applied Energy 3
-Applied Soft Computing 3
-Energy 3
-Engineering Applications of Artificial Intelligence 3
-IEEE Transactions on Intelligent Transportation Systems 3
-Information Sciences 3
-International Journal of Data Science and Analytics 3
-Knowledge-Based Systems 3
+3 lists the journals with more than 5 published papers.
+Table 3: Journals with more than 5 published paper.
+Journal Papers
+IEEE Access 14
+ACM Trans. Knowl. Discov. Data 10
+Knowledge-Based Systems 9
+IEEE Transactions on Instrumentation and Measurement 8
+Applied Intelligence 8
+IEEE Transactions on Knowledge and Data Engineering 7
+Expert Systems with Applications 7
+ACM Trans. Intell. Syst. Technol. 7
+Sensors 6
+IEEE Transactions on Intelligent Transportation Systems 6
+IEEE Transactions on Industrial Informatics 6
+Applied Energy 6
 The diversity of the journals indicates a vast range of potential applications for spatio-temporal GNN models in different fields.
-This suggests that researchers often select journals based on the specific domain of application, allowing them to reach targeted audiences and contribute within their respective fields.
-Such diversity also underlines the interdisciplinary nature of GNNs research, as it spans multiple domains.
-Figure 3: Pie chart of the distribution of corresponding authors in different countries.
+This suggests that researchers often select journals based on the specific domain of application.
+This allows them to reach targeted audiences, and contribute within their respective fields.
+Such diversity also underlines the interdisciplinary nature of GNNs research, clearly spanning multiple domains.
+As for the publishers, there are around 40 different publishers, and the ones with more than 5 papers are listed in Tab.
+4.
+Table 4: Publishers with more than 5 published paper.
+Journal Papers
+Elsevier 82
+Institute of Electrical and Electronics Engineers Inc. 42
+Association for Computing Machinery 27
+Multidisciplinary Digital Publishing Institute (MDPI) 27
+IEEE 18
+IEEE Computer Society 13
+Springer 10
+Springer Science and Business Media 10
+John Wiley and Sons Inc 5
 In order to examine the degree of collaboration among researchers, a co-authorship network was generated by using the VOSviewer software [VOSviewer].
 Fig.
-4 illustrates authors as nodes and collaborations as connections, revealing some research clusters.
-Two of these clusters are enlarged and shown in Fig.
-4.
-(a) Visualization of the complete collaboration network.
-(b) Zoom in on two different clusters.
+3 depicts authors as nodes of the network, and collaborations as connections, revealing some research clusters.
+Fig.
+3 shows two of these clusters enlarged.
+(a) Map of the complete collaboration network.
+(b) Enlargement on two different clusters.
 Figure 3: Network of collaboration between researchers, created using VOSviewer software.
-In the figure, the color of the nodes is related to the average publication year of an author, so that the evolution of research groups over time can be evaluated.
+In Fig.
+3, the color of the nodes is related to the average publication year of an author, so that the evolution of research groups over time can be evaluated.
 Two key observations can be made.
 First, the number of clusters suggests that there are many distinct research groups, indicating a high degree of fragmentation within the community.
-Second, the color of the nodes indicates that some groups were active only at the beginning and have since disappeared, while others have recently formed, as indicated by their yellow color.
-Some authors have various collaborations over time, as also shown in Fig.
-4.
+Second, the color of the nodes indicates that some groups were active only at the beginning and have subsequently disappeared, while others have recently emerged, as indicated by their yellow color.
+Some authors have had various collaborations over time, as also shown in Fig.
+3.
 Fig.
-3 shows a pie chart of the distribution of corresponding authors among different countries.
+4 shows a pie chart of how corresponding authors are distributed among different countries.
 This visualization provides an overview of the global distribution of research activities.
 More than the 70% of the authors are affiliated with institutions in China.
+Figure 4: Pie chart of the distribution of corresponding authors in different countries.
 4 Graph neural networks
@@
 Definition 1 (Time series).
-A time series is a sequence of data points indexed in time order (that here is assumed to be regularly sampled, i.e., equispaced).
+A time series is a sequence of data points indexed by timestamps, that here are assumed to be equispaced.
 An equispaced univariate time series of length $T$ is a sequence of scalar observations collected over time, denoted as $\vec{x}_{t}\in\mathbb{R}^{T}$.
@@
 Definition 2 (Graph).
-A graph $\mathcal{G}$ is a pair of finite sets $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ is a set of $n$ nodes (also called vertices) and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges.
-In an undirected graph each edge is an unordered pair of nodes $\{v_{i},v_{j}\}$, while in a directed graph the edges have an orientation and correspond to ordered pairs $(v_{i},v_{j})$.
+A graph $\mathcal{G}$ is a pair of finite sets $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ is a set of $n$ nodes (also called vertices) and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of connecting edges.
+In an undirected graph each edge is an unordered pair of nodes $\{v_{i},v_{j}\}$, while in a directed graph the edges have an orientation, and correspond to ordered pairs $(v_{i},v_{j})$.
 When an edge $(v_{i},v_{j})$ or $\{v_{i},v_{j}\}$ exists, the nodes $v_{i}$ and $v_{j}$ are called adjacent.
@@
 Notice that the sets of nodes and edges can change over time, as well as the set of attributes $\mathcal{X}_{t}$.
+Hereafter, the time index will be dropped to indicate graphs not necessarily changing in time, i.e., not necessarily spatio-temporal.
 Definition 4 (Adjacency matrix).
@@
 4.2 Overview of graph neural networks
-Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, a GNN model (not necessarily changing in time, i.e., spatio-temporal) aims to generate an embedding (i.e., a vector of real entries) for each node in the graph.
-An overall graph embedding consists of representing a graph into a possibly different-dimensional space while preserving its structural information.
+Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, a GNN model (not necessarily spatio-temporal) aims to generate an embedding (i.e., a vector of real entries) for each node in the graph.
+An overall graph embedding consists of representing a graph in a space of chosen dimension, while preserving its structural information.
 The notion of GNN was initially proposed by Gori et al.
@@
 in 2009 in Ref.
-[Scarselli2009] as an architecture to implement a function that maps a graph $\mathcal{G}$ with $n$ nodes into an $m$-dimensional Euclidean space.
+[Scarselli2009] as an architecture designed to implement a function that maps a graph $\mathcal{G}$ with $n$ nodes into an $m$-dimensional Euclidean space.
+Noticeably, this definition was intended in general for graph data, not specifically time-dependent data, which is the focus of this SLR.
 In Ref.
-[Scarselli2009] the target node’s embedding was learned by first propagating forward at fixed weights neighbor information within the graph in an iterative manner, until a stable fixed point is reached (whose existence is guaranteed under certain assumptions by Banach’s fixed point theorem), and then propagating errors on targets back across the graph.
-The calculation of node embeddings is in general quite complex, and over time many other methods have been developed.
-These methods are based on different architectures and sophisticated mechanisms.
+[Scarselli2009] the target node’s embedding was learned in two steps.
+First, neighbor information is iteratively propagated with fixed weights within the graph until a stable fixed point is reached (whose existence is guaranteed under certain assumptions by Banach’s fixed point theorem).
+Second, errors on targets are propagated back across the graph.
+Notice that in this setup, propagation is performed in “internal time”, not in “clock time”.
+More in general, the calculation of node embeddings is quite complex, and over time many other methods have been developed.
+These methods are based on different architectures and sophisticated aggregation mechanisms.
 Based on these aggregating mechanisms, a taxonomy of GNN models with different classes has been defined.
@@
 [survey_Lucheroni].
-More details will be given in Subsec.
+More details about that will be given in Subsec.
 4.3.
 4.3 Spatio-temporal GNNs and their taxonomy
-The majority of research on spatio-temporal GNNs is focused on multivariate time series, as they can be naturally abstracted into spatio-temporal graphs, as depicted in Fig.
+A large part of the research on spatio-temporal GNNs is focused on multivariate time series, as they can be naturally abstracted into spatio-temporal graphs, as depicted in Fig.
 5.
-In this context, nodes correspond to different variables at a given time step, and edges represent relationships between variables.
+In this context, nodes usually correspond to different variables at a given time step, and edges represent relationships between variables.
 Figure 5: Representation of a spatio-temporal graph. Here, each node has a dynamic feature, given by the points of the time series.
-This spatio-temporal modeling assumes that the feature information of a node depends on both its own historical values and on the historical data from its neighboring nodes [survey_spatiotemporal].
-Once the the graph structure is given, information is propagated within it until the output is obtained, as better explained in the following.
-When it comes to the development of a GNN model, there are two primary approaches: one that handles spatial and temporal substructures in separate blocks, and another that integrates and processes these two substructures together [survey_spatiotemporal].
-This is implemented by means of modules that are purely spatial, purely temporal, or hybrid combinations of the two.
-These modules are then organized into a series of blocks to create the final spatio-temporal GNN model.
-In the taxonomy of spatio-temporal GNNs used in this review, the focus is exclusively on the description of the spatial module that propagates the information within the nodes at fixed time $t$ (not explicitly included in the following equations for simplicity), which is in the reviewed literature a fundamental aspect that differentiates one GNN model from another.
-Based on the mechanism at the basis of the information propagation in the graph at a given time step, spatio-temporal GNNs can be further classified as recurrent GNNs, convolutional GNNs, and attentional GNNs, as explained in the following.
+This specific form of spatio-temporal modeling assumes that the feature information of a node depends on both its own historical values and on the historical data from its neighboring nodes [survey_spatiotemporal].
+Notice that once the the graph structure is given, information is propagated (in “internal time”) within it until the output is obtained, as better explained in the following.
+Our discussion will build up from this perspective.
+When it comes to the development of a spatio-temporal GNN model, there are two primary approaches: one that handles spatial and temporal substructures in separate blocks, and another that integrates and processes these two substructures at once [survey_spatiotemporal].
+These approaches are implemented by means of modules that are purely spatial, purely temporal, or hybrid combinations of the two.
+These modules can then be organized into a series of blocks, thus creating the sought spatio-temporal GNN model.
+In the taxonomy of spatio-temporal GNNs used in this review, the focus is exclusively on the description of the spatial module that propagates the information within the nodes at fixed time $t$ (not explicitly included in the following equations for simplicity), which, in the reviewed literature, is a fundamental aspect that differentiates one GNN model from another.
+Based on the mechanism at the basis of the information propagation in the graph at a given time step, spatio-temporal GNNs can be further classified as recurrent GNNs, convolutional GNNs, and attentional GNNs, as discussed in the following.
 These classes are not mutually exclusive, as some models may fit into hybrid categories.
-In addition, the classes are not fully exhaustive, as there may be models that do not align with any of the existing categories.
-However, for the purpose of this review, the taxonomy discussed above is sufficient.
+In addition, these classes are not fully exhaustive of the variety of available models, as there may be models that do not align with any of the existing categories.
+However, for the purpose of this review, the taxonomy discussed above will be considered sufficient.
 Recurrent GNNs.
-Recurrent graph neural network modeling mainly includes the results of early studies on GNNs, where the node embeddings are generated by iteratively propagating neighbor information until a stable fixed point is reached.
+Recurrent GNN modeling mainly proceeds from the results of early studies on GNNs, where the node embeddings are generated by iteratively propagating neighbor information until a stable fixed point is reached.
 The iterative equation for the calculation of the embedding (or hidden state) of node $u$ has the form
 $h_{u}^{(k)}=f\left(x_{u},\{x_{v},r_{(u,v)},h_{v}^{(k-1)}\mid v\in\mathcal{N}(u)\}\right)\>,$ (2)
-where $k$ is the index related to iteration, $f$ is a parametric trainable function, $x_{u}$ is the attribute associated to node $u$, $r_{(u,v)}$ is the weight of the edge $(u,v)$ between the nodes $u$ and $v$, $\mathcal{N}(v)$ denotes the neighborhood of node $v$, and the initial hidden states $h^{(0)}$ are usually randomly initialized.
-As detailed in Ref.
-[Scarselli2009], which introduced recurrent GNNs, to ensure the convergence of Eq.
-(2), it is necessary that the iterative function $f$ is a contraction mapping.
+where $k$ is the index related to iteration (hence running on “internal time”), $f$ is a parametric trainable function, $x_{u}$ is the attribute associated to node $u$, $r_{(u,v)}$ is the weight of the edge $(u,v)$ between the nodes $u$ and $v$, $\mathcal{N}(v)$ denotes the neighborhood of node $v$, and the initial hidden states $h^{(0)}$ are usually randomly initialized.
+As detailed in the seminal Ref.
+[Scarselli2009], which introduced recurrent GNNs, in order to ensure the convergence of Eq.
+(2) it is necessary that the iterative function $f$ is a contraction mapping.
 Convolutional GNNs.
-Convolutional graph neural networks are closely related to recurrent graph neural networks.
-They were at first introduced by Kipf and Welling in 2017 in Ref.
+Convolutional GNNs are closely related to recurrent GNNs.
+They were introduced by Kipf and Welling in 2017 in Ref.
 [graph_convolution_kipf].
-Unlike recurrent GNNs, which apply the same iterative contraction mapping $f$ until an equilibrium is reached, convolutional GNNs use different parameters at each updating step, by stacking multiple graph convolutional layers to extract the node embeddings [cit_rev3].
-A basic form for the equation for the $k^{\text{th}}$ layer of a convolutional graph neural network with $n$ nodes is
-$H^{(k)}=f_{k}\left(\tilde{\mathcal{A}}H^{(k-1)}\Theta_{k-1}\right)\>,\quad\tilde{\mathcal{A}}=\hat{\mathcal{D}}^{-\frac{1}{2}}\hat{\mathcal{A}}\hat{\mathcal{D}}^{-\frac{1}{2}}\>,\quad\hat{\mathcal{A}}=\mathcal{A}+I\>,\quad\hat{\mathcal{D}}_{ii}=\sum_{j}\hat{\mathcal{A}}_{ij}\>,\quad\quad H^{(0)}=X\>,$ (3)
-where $f_{k}$ represents the convolutional operation performed at layer $k\in\{1,\dots,K\}$, $\mathcal{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix, $I\in\mathbb{R}^{n\times n}$ is the identity matrix (used to add the self-loops in the adjacency matrix $\mathcal{A}$), $\hat{\mathcal{D}}\in\mathbb{R}^{n\times n}$ is the diagonal degree matrix associated to $\hat{\mathcal{A}}$, $H^{(k)}\in\mathbb{R}^{n\times d_{k}}$ are the $d_{k}$-dimensional node embeddings $h^{(k)}_{v}$ of the entire graph produced by the $k^{\text{th}}$ layer, $X\in\mathbb{R}^{n\times d_{0}}$ is the collection of all node attributes $x_{v}$ of the entire graph, and $\Theta_{k}\in\mathbb{R}^{d_{k-1}\times d_{k}}$ is a trainable weight matrix for the $k^{\text{th}}$ layer [survey_Lucheroni].
+Unlike recurrent GNNs, which apply the same iterative contraction mapping $f$ until an equilibrium is reached, convolutional GNNs can use different parameters at each updating step, by stacking multiple graph convolutional layers to extract the node embeddings [cit_rev3].
+Noticeably, as the number of convolutional layers increases, the convolutional perceptual field expands, and a more abstract representation of information is thus obtained [art_461].
+A basic form for the equation for the $k^{\text{th}}$ layer of a convolutional GNN with $n$ nodes is
+$\begin{gathered}H^{(k)}=f_{k}\left(\tilde{\mathcal{A}}H^{(k-1)}\Theta_{k-1}\right)\>,\quad\tilde{\mathcal{A}}=\hat{\mathcal{D}}^{-\frac{1}{2}}\hat{\mathcal{A}}\hat{\mathcal{D}}^{-\frac{1}{2}}\>,\quad\hat{\mathcal{A}}=\mathcal{A}+I\>,\quad\\ \hat{\mathcal{D}}_{ii}=\sum_{j}\hat{\mathcal{A}}_{ij}\>,\quad\quad H^{(0)}=X\>,\end{gathered}$ (3)
+where $f_{k}$ represents the convolutional operation performed at layer $k\in\{1,\dots,K\}$ (with $K$ typically kept small to prevent issues like over-smoothing [oversmoothing]), $\mathcal{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix, $I\in\mathbb{R}^{n\times n}$ is the identity matrix (used to add self-loops in the adjacency matrix $\mathcal{A}$), $\hat{\mathcal{D}}\in\mathbb{R}^{n\times n}$ is the diagonal degree matrix associated to $\hat{\mathcal{A}}$, $H^{(k)}\in\mathbb{R}^{n\times d_{k}}$ are the $d_{k}$-dimensional node embeddings $h^{(k)}_{v}$ of the entire graph produced by the $k^{\text{th}}$ layer, $X\in\mathbb{R}^{n\times d_{0}}$ is the collection of all node attributes $x_{v}$ of the entire graph, and $\Theta_{k}\in\mathbb{R}^{d_{k-1}\times d_{k}}$ is a trainable weight matrix for the $k^{\text{th}}$ layer [survey_Lucheroni].
 The mechanism of node embedding generation from Eq.
@@
 Eq.
-(4b) computes a pairwise masked attention score between nodes $u$ and $v$ by concatenating (||||| |) the vectors $z_{u}^{(k)}$ and $z_{v}^{(k)}$, calculating a dot product with a learnable weight vector $\vec{a}^{\,(k)}$, and applying the LeakyReLU nonlinearity.
+(4b) computes a pairwise masked attention score between nodes $u$ and $v$.
+This is done by applying the concatenation operation ||||| | to the vectors $z_{u}^{(k)}$ and $z_{v}^{(k)}$, then calculating a dot product of the concatenation with a learnable weight vector $\vec{a}^{\,(k)}$, and finally applying the LeakyReLU nonlinearity.
 The quantity $\epsilon_{uv}$ indicates the importance of the features of node $v$ to node $u$.
-The term ”masked” refers to the fact that $\epsilon_{uv}$ are only computed for neighboring nodes.
+The term “masked” refers to the fact that $\epsilon_{uv}$ are computed for neighboring nodes only.
 Eq.
@@
 Finally, Eq.
-(4d) computes the next layer embedding $h_{u}^{(k+1)}$ by aggregating the information in neighboring nodes, weighted by the attention scores [GAT].
-Graph attention networks can also further use a multi-head attention mechanism to stabilize the learning process of self-attention.
+(4d) computes the embedding $h_{u}^{(k+1)}$ of the next layer by aggregating the information in neighboring nodes, weighted by the attention scores [GAT].
+Graph attention networks can in addition use a multi-head attention mechanism to stabilize the learning process of self-attention.
 In practice, the node updating operation in graph attention networks is a generalization of the traditional averaging or max-pooling of neighboring nodes, which allows each node to compute a weighted average of its neighbors and identify the most relevant ones [Brody2021HowAA].
 As anticipated, the taxonomy discussed so far focuses only on the propagation mechanisms of the spatial component.
-However, spatio-temporal GNNs typically consist of a stack of spatial, temporal, or hybrid modules, each one with distinct roles.
+However, spatio-temporal GNNs typically consist of a stack of spatial, temporal, or also hybrid modules, each one with distinct roles.
+The function of these modules is detailed hereafter.
 The spatial module is responsible for propagating information between nodes, thereby enabling the analysis of cross-sectional inter-dependencies between different variables.
-As previously explained in the description of the mechanisms underlying the taxonomy, the spatial module aggregates information starting from immediate neighboring nodes and extends it outward, thereby capturing also the influence of distant nodes.
-Noticeably, this spatial operation does not account in principle for temporal information.
-In contrast, the temporal module, often based on architectures such as LSTM, GRU, or self-attention, focuses on the evolution of data over time, independently from cross-sectional node interactions.
+As previously discussed in the description of the mechanisms underlying the assumed taxonomy, the spatial module aggregates information starting from immediate neighboring nodes and extends it outward, thereby capturing also the influence of distant nodes.
+This spatial operation cannot directly account for temporal information.
+In contrast, the temporal module, often based on architectures such as LSTM, GRU, or self-attention, focuses on the evolution of data over time, independently from cross-sectional node interactions [art_384].
 When these different modules are stacked together, spatial and temporal information end up to be used simultaneously, although they are processed in separate modules.
 This occurs because the computation of each node state is influenced by both spatial and temporal information.
-This combined approach allows the spatio-temporal GNN model to effectively capture at the same time both spatial relationships and temporal dynamics.
+This combined approach allows the spatio-temporal GNN model to be a powerful architecture able to effectively capture at the same time both spatial relationships and temporal dynamics.
 4.4 Determination of the graph structure
-A crucial issue when dealing with spatio-temporal GNNs for time series problems is the determination of the graph structure, i.e.
-the connectivity of the nodes.
-Namely, some time series datasets come by nature with a pre-defined graph structure (e.g., road network), while some others do not.
+A crucial issue when dealing with spatio-temporal GNNs for time series problems is the determination of the graph structure, i.e., the connectivity of the nodes.
+Namely, some time series datasets come by their own nature with a pre-defined graph structure (e.g., a road network), whereas some others do not.
 When a natural pre-defined graph structure is available, it helps the model to better capture the underlying dynamics of the system.
-When it is not directly available, it must be in some way defined by the user (based on domain knowledge or some metrics), determined by some algorithm (such as visibility graphs [visibility_Gori], [visibility_graph], [art_14]), or learned by the model itself.
-Once the graph structure has been determined, it is necessary to define the weighted graph adjacency matrix, which is a generalization of the graph adjacency matrix whose entries are given by the edge weights of the graph.
+When it is not directly available, it must be defined by the user in some way (based on domain knowledge or some metrics), or determined by some algorithm (such as visibility graphs [visibility_Gori, visibility_graph, art_14]), or learned by the model itself.
+Once the graph structure is determined, it is necessary to define the weighted graph adjacency matrix, which is a generalization of the graph adjacency matrix $\mathcal{A}$ whose entries are given by the edge weights of the graph.
 The edge weights can be defined a priori by the user, or again, learned by the model itself based on a pre-defined architecture.
-In the first case, the edge weights are assigned based on some pre-defined metrics or criteria chosen by the user (e.g., spatial distance between the sites or similarity measures).
-In the latter case, the weights are continuously adjusted throughout training as the model learns from the data.
-Each section of the following thematic analysis examines which approach is prevalent.
+In the first case, the edge weights are assigned based on some pre-defined metrics or criteria chosen by the user, for example by the spatial distance between the sites, or by similarity measures.
+In the latter case, the weights are continuously adjusted throughout training as the model keeps learning from the data.
+In the following, we present a concise overview of the most common techniques for defining graph structures.
+Pre-defined graphs are typically generated based on prior knowledge and specific metrics.
+These methods for the construction of the graphs typically rely on distance or similarity information.
+As for location-based techniques, they utilize geographical coordinates (latitude and the longitude) of the nodes to calculate the distance between them (e.g., using haversine or geodesic distance).
+This distance information is then used to determine the edge weights with techniques such as the thresholded Gaussian kernel (which applies a negative exponential function to the distance), weighted $k$-NN, or reversed MinMax.
+As for similarity-based techniques instead, they rely on similarity (or dissimilarity) of signal data.
+These techniques include methods such as correlation, cosine similarity, dynamic time warping (DTW), maximum information coefficient (MIM), or transfer entropy.
+In general, no single approach stands out as a clear winner, and performance should be evaluated on each dataset.
+For references and a more detailed explanation of these and other techniques for constructing pre-defined structures, see Ref.
+[art_160].
+As for automatic graph structure learning techniques, a distinction can be made between models that rely on predefined graph structures (sometimes referred to as supplementary graph structure learning models), and those that learn the graph structure entirely from scratch.
+In the first category, there are models such as Graph WaveNet (GWN) [GWN], which constructs an adaptive matrix as a complement to a predefined graph structure [art_405].
+These techniques still rely on some form of prior assumptions.
+To address this limitation, models that learn the graph structure from scratch have been proposed.
+One of them is the MTGNN model [MTGNN], which extracts uni-directional relationships from node embeddings as it learns, and achieves state-of-the-art results across different datasets.
+It is one of the first models to use self-learned graph structures [art_417].
+When the graph structure is learned, some authors introduce a graph regularization term to the training loss function to optimize the learned graph adjacency matrix.
+This regularization term is typically designed to control the sparsity of the matrix, and to promote graph proximity, encouraging larger entries when the series in the corresponding nodes are “close enough” [art_448].
+Some researchers argue that models which generate graphs without prior knowledge may overlook local conditions when constructing adjacency matrices [art_397].
+Therefore, it is important to conduct further research in this area to identify the most effective techniques for generating informative graphs.
+For a more detailed explanation of graph learning models, see Refs.
+[art_396, art_405].
+Another key aspect in the determination of the graph structure is the choice between using a static or dynamic adjacency matrix, i.e.
+whether to use the same matrix for all graphs or not.
+When working with graphs defined in advance by specific metrics, it is common to use either a static matrix or a dynamic matrix generated through a sliding window approach.
+In contrast, when graphs are learned by the model, dynamic adjacency matrices are typically employed.
+The use of a dynamic adjacency matrix has the advantage of capturing the dynamic nature of the problem [art_447].
+However, the learned dynamic graphs often show significant differences between adjacent time steps, whereas real-world graph structures typically evolve smoothly over time.
+Therefore, this approach may not accurately capture the stable evolution observed in the real world.
+Consequently, some researchers are exploring an intermediate approach, assuming that adjacent graph structures follow a basic pattern.
+Instead of defining static or dynamic adjacency matrices, they fuse a static backbone graph with a dynamic adjacency matrix.
+The static network captures the common graph structure over time, while the dynamic network focuses on learning the changes over time.
+An example of this intermediate approach can be found in Ref.
+[art_415], which also provides a comparison between this intermediate approach and the classical ones.
+In general, there is no rule that can determine in advance the best approach to follow.
+In fact, it depends on both the application domain and the specific dataset.
+For this reason, we do not explore this aspect further.
+Each subsection of the thematic analysis in Sec.
+5 discusses the most prevalent approach for each group of papers.
+4.5 Model evaluation
+Evaluating the performance of a time series forecasting or classification model is essential to assess its reliability.
+Moreover, experimental results often show that even when datasets involve the same forecasting task, the transferability of model architectures is limited [art_408].
+Therefore, a proper evaluation of the model is a crucial aspect.
+A first key aspect to consider is its accuracy.
+The accuracy of a model can be measured by many different metrics, which vary between classification and forecasting tasks.
+Each metric provides its own insight into different aspects of prediction quality, so multiple metrics are typically used for a more comprehensive evaluation.
+In addition to presenting a table of the error metrics computed for the entire graph, some authors provide a spatial visualization of these errors across the various nodes, as for example in Refs.
+[art_28, art_449].
+As for forecasting problems, commonly used evaluation metrics include the mean absolute error (MAE), mean squared error (MSE), root mean squared error (RMSE), mean absolute percentage error (MAPE), root relative squared error (RRSE), the coefficient of determination (R2), and the correlation coefficient (CORR).
+Let $N$ denote the number of observations, $y_{i}$ the true value associated with observation $i$, $\hat{y}_{i}$ the corresponding output of the model, and the bar denote the average.
+The formulas for these metrics (for a single-step forecasting horizon) are provided in Tab.
+5.
+Table 5: Evaluation metrics for single-horizon forecasting tasks and their formulas.
+Metrics Formula
+Mean absolute error (MAE) $\frac{1}{N}\sum_{i=1}^{N}|y_{i}-\hat{y}_{i}|$
+Mean squared error (MSE) $\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}$
+Root mean squared error (RMSE) $\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}$
+Mean absolute percentage error (MAPE) $\frac{1}{N}\sum_{i=1}^{N}\left|\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right|\cdot 100$
+Root relative squared error (RRSE) $\sqrt{\frac{\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}}$
+Coefficient of determination (R2) $1-\frac{\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}$
+Correlation coefficient (CORR) $\frac{\sum_{i=1}^{N}(y_{i}-\bar{y})(\hat{y}_{i}-\bar{\hat{y}})}{\sqrt{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}\sum_{i=1}^{N}(\hat{y}_{i}-\bar{\hat{y}})^{2}}}$
+Similar formulas can be applied to multi-horizon forecasting problems.
+All the metrics listed, except for the last one, are considered better when their values are lower, as they indicate smaller discrepancies between the predicted and actual values.
+In contrast, CORR quantifies the linear correlation between actual values and forecasts, and a high correlation is preferable.
+As for classification problems, common evaluation metrics include accuracy, precision, recall, and F1 score.
+Given $TP$, $TN$, $FP$, and $FN$ the number of true positives, true negatives, false positives, and false negatives, respectively, the formulas for the binary classification problem are reported in Tab.
+6.
+Table 6: Evaluation metrics for binary classification tasks and their formulas.
+Metrics Formula
+Accuracy $\frac{TP+TN}{TP+TN+FP+FN}$
+Precision $\frac{TP}{TP+FP}$
+Recall $\frac{TP}{TP+FN}$
+F1 score $\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
+Similar formulas can be applied for the multi-class classification problem.
+Besides accuracy, another important aspect to consider is the complexity of the model.
+Finding an optimal balance between accuracy and complexity is crucial for developing models that are not only accurate but also efficient.
+For this reason, some authors include the training and inference computational times, as well as the number of parameters of their models, allowing for a comparison of these aspects across different models.
+However, not all authors do so, and even if they do it, the characteristics of the hardware and other settings used by each author must be taken into account when comparing computational times.
+Consequently, it was not possible to provide an accurate and detailed comparison of the relationship between complexity and performance across the different models in this review.
+Although creating a plot to compare accuracy and computational times of the different models would be very useful (as in Ref.
+[art_407]), it is not possible to provide such a comparison here.
+The authors of the various papers should be encouraged to incorporate such comparisons in their work.
 5 Thematic analysis
-This section explores the diverse applications of GNN models in different groups.
-As said, the selected journal papers are divided in groups according to the application domain.
-The goal is to provide an overview of the various empirical studies, the most common approaches, and the results obtained for different applications.
-In each subsection of the following thematic analysis, a list of the benchmark models that appear in the selected journal papers will be provided.
-They will be classified as: mathematical and statistical models coming from Econometrics, traditional machine learning models, and GNN-based models.
-As for mathematical and statistical models, they generate forecasts by analyzing a limited amount of data, and they usually struggle with cross-sectional large-scale and temporal long-term dependencies.
-Traditional machine learning models use more complex architectures, mainly neural networks, but they are not necessarily able to capture intricate spatio-temporal dependencies within the data.
-GNN-based models instead, attempt to represent these complex dependencies using various techniques, resulting in better, though not perfect, accuracy.
+This section explores the applications of the collected GNN models across different thematic groups.
+The goal is to provide an overview of available empirical studies, the most common approaches, and the results obtained concerning the discussed themes.
+In each following subsection, one per theme, a list of the benchmark models that appear in the selected journal papers will be provided.
+Within each theme, they will be classified as: mathematical and statistical models coming from Econometrics, non-GNN machine learning models, and GNN-based models.
+Mathematical and statistical models are characterized by the fact that they generate forecasts by analyzing a limited amount of data, and they usually struggle with cross-sectional large-scale and temporal long-term dependencies.
+Non-GNN machine learning models use more complex and highly nonlinear architectures, mainly neural networks, but they are not necessarily able to capture intricate spatio-temporal dependencies within the data.
+GNN-based models, conversely, attempt to represent these complex dependencies using various techniques, resulting in better, though not perfect, accuracy.
 In addition, information on the most common datasets will be given, along with the best results reported in the literature.
-All results presented in the tables of this SLR are taken directly from the tables of the selected papers.
-The groups of papers are presented in alphabetical order, with the exception of the ”Generic” applications group (which cannot be directly attributed to a specific case study) and the ”Other topics” group (focusing on specific problems of other disciplines).
+All results presented in the tables of this SLR are taken directly from the tables of the selected papers, as we did not conduct any further experiments to obtain additional results.
+Note that in the tables which list benchmark models we have included references to any GNN model found.
+In the results tables, instead, we have explicitly referenced only those models proposed in the selected papers, in order to highlight them.
+Further information about the models used as benchmarks can be found either in the benchmark tables or in the papers where they are used as benchmarks.
+The thematic groups of papers are presented in alphabetical order, with the exception of the “Generic” applications group (which cannot be directly attributed to a specific case study) and the “Other topics” group (focusing on specific problems of other disciplines).
 In subsections with a large number of papers and high homogeneity of the datasets, more precise comparisons of the results and in-depth analyses will be conducted.
 5.1 Energy
-The first group of papers is ”Energy” (with 13 out of 156 papers), which includes all the studies related to energy systems.
+The first thematic group is “Energy” (with 33 out of 281 papers), which includes all the studies related to energy systems.
 In particular, it includes research on electricity load, energy consumption, and power generation forecasting.
-This is somehow connected to the ”Environment” subsection (Subsec.
+These topics are somehow connected to those of the “Environment” subsection discussed later (Subsec.
 5.2).
-Indeed, green energy sources are associated with environmental phenomena such as wind speed and solar radiation inflow.
-However, this subsection focuses on energy applications, whereas Subsec.
-5.2 refers to the study of environmental data that may or may not be related to energy applications.
+For example, green energy sources are associated with environmental phenomena such as wind speed and solar radiation inflow.
+Our rationale for this selection is that here we focus on energy applications, whereas Subsec.
+5.2 will refer to the study of environmental data that may or may not be related to energy applications.
 Wind speed forecasting, for instance, is connected to wind power generation due to the cubic relationship between wind speed and wind power [art_29].
@@
 In addition, effective forecasting allows for optimal scheduling of power plants and energy resources, reducing operational costs and improving overall system efficiency.
-The majority of econometric and machine learning models tend to focus on time series from a single site or consider multiple sites without explicitly capturing the relationships between them.
+In the papers within this theme, the majority of econometric and non-GNN machine learning models tend to focus on time series from a single site or consider multiple sites without explicitly capturing the relationships between them.
 GNN models may be an innovative and effective approach to addressing energy related problems.
-An overview of the selected papers reveals a noticeable interest in the application of GNNs in the energy field since 2022.
-The most frequently studied fields include forecasting of wind ([art_20], [art_24], [art_27]) and solar power ([art_19], [art_21], [art_22], [art_26]), which are renewable energy sources characterized by their intermittency.
-The two most popular journals among the selected papers are Applied Energy by Elsevier, and Frontiers in Energy Research by Frontiers Media SA.
-The majority of the corresponding authors are affiliated with institutions in China and in the United States.
+An overview of the selected papers reveals a relevant interest in the application of GNNs in the energy field starting from 2022.
+The most frequently studied fields include forecasting of wind [art_20, art_24, art_27, art_332, art_333] and solar power [art_19, art_21, art_22, art_26, art_349], which are renewable energy sources characterized by their intermittency.
+The most popular journal among the selected papers is Applied Energy by Elsevier, with 5 published papers.
+The majority of the corresponding authors are affiliated with institutions in China.
 5.1.2 Datasets
-The study of the selected papers in the ”Energy” group suggests that there are no particularly common or widely recognized benchmark datasets.
+The study of the selected papers in the “Energy” group suggests that there are no particularly common or widely recognized benchmark datasets within this field.
 Indeed, researchers often focus on specific datasets, typically collected within the university campus or in the country where they are located.
@@
 LABEL:t:link_dataset_energy are listed the links to the public datasets mentioned in the papers.
-Table 3: List of public datasets in the ”Energy” group and their corresponding links.
+Table 7: List of public datasets in the “Energy” group and their corresponding links.
 Dataset Used by Link
+Commercial and Residential Load [art_348] https://www.osti.gov/biblio/1788456
 Danish smart heat meters [art_15] https://doi.org/10.5281/zenodo.6563114
@@
 IEEE bus systems [art_23] https://doi.org/10.1109/UPEC.2015.7339813 (source paper)
+Individual Household Electric Power Consumption [art_344] https://archive.ics.uci.edu/ml/datasets/individual+household+electric+power+consumption
 JERICHO-E-usage [art_15] https://doi.org/10.6084/m9.figshare.c.5245457.v1
@@
 OPSD [art_16] https://open-power-system-data.org/
+PJM Hourly Energy Consumption [art_390] https://www.pjm.com/markets-and-operations/data-dictionary
+PVGIS [art_349] http://re.jrc.ec.europa.eu/pvgis/apps4/pvest.php
 PV Switzerland [art_21], [art_22] https://doi.org/10.1109/IJCNN48605.2020.9207573 (source paper)
-UMass Smart* Dataset [art_17] https://traces.cs.umass.edu/index.php/Smart/Smart
-Three of the selected papers ([art_20], [art_22], [art_24]) use data from the National Renewable Energy Laboratory (NREL), a national laboratory of the U.S.
+Residential heat and electricity demand [art_341] https://ieee-dataport.org/open-access/8-years-hourly-heat-and-electricity-demand-residential-building
+SGCC dataset [art_342] http://www.sgcc.com.cn/
+Southern Power Grid [art_344] https://www.csg.cn/
+UMass Smart* Dataset [art_17] https://traces.cs.umass.edu/docs/traces/smartstar/
+Three of the selected papers [art_20, art_22, art_24] use data from the National Renewable Energy Laboratory (NREL), a national laboratory of the U.S.
 Department of Energy (https://www.nrel.gov/).
@@
 [art_19]), but they are of course highly dependent on the location of the study.
-The data granularity ranges from 1 minute to 1 hour, and the forecasting horizon from tens of minutes to days.
-They depend on the dataset, the domain of application, and the purpose of the forecast.
+The data time granularity ranges from 1 minute to 1 day, and the forecasting horizon from tens of minutes to days.
+These quantities depend on the dataset, the domain of application, and the purpose of the forecast.
 Most datasets require pre-processing, including missing data interpolation, normalization, and satellite image processing.
-5.1.3 Proposed models
-As for the taxonomy of the proposed GNN models, 8 out of 13 papers use a convolutional GNN, and the other 5 use an attentional GNN.
-Regarding the graph structure, the papers typically provide only a brief description.
-For example, it is rarely highlighted whether the graph structure is static or dynamic, and the way in which it is constructed is often described summarily.
+5.1.3 Types of models
+As for the taxonomy of the proposed GNN models, 17 out of 33 papers use a convolutional GNN, and the other 9 use an attentional GNN.
+Six papers from 2024 use both convolutional and attentional approaches, and one paper [art_338] uses the original recurrent approach on a graph defined via the visibility algorithm.
+As to the graph structure used, the papers typically provide only a brief description.
+For example, only few papers highlight whether the graph structure is static or dynamic, and the way in which it is constructed is often described summarily.
 In some papers, the graph structure reflects the geographic location of the elements (e.g., wind or solar power plants, as in Refs.
-[art_24] and [art_26]), whereas in others it is based on the correlation or similarity between time series (as in Ref.
+[art_24, art_26, art_349]), whereas in others it is based on the correlation or similarity between time series (as in Ref.
 [art_19]).
@@
 [art_22].
-The most common loss function used to train the models is the MSE.
-Many papers mention the programming languages Python and Matlab, and few of them mention the Python libraries PyTorch [PyTorch] and TensorFlow [2] for the implementation of the GNN model.
+The majority of papers in this group rely on a static graph.
+The most common loss function used to train these models is the MSE.
+Many papers mention the use of the programming languages Python and Matlab, and few of them mention the use of the Python libraries TensorFlow [2], PyTorch [PyTorch], and PyTorch Geometric [PyTorch_Geometric] for the implementation of the GNN model.
 In this group, only Ref.
@@
 5.1.4 Benchmark models
-Tab. LABEL:t:benchmark_models_energy lists the benchmark models used in the ”Energy” group by at least two papers included in this review.
-Table 4: List of benchmark models in the ”Energy” group divided per category.
+Tab. LABEL:t:benchmark_models_energy lists the benchmark models used in the “Energy” group by at least two papers included in this review.
+Table 8: List of benchmark models in the “Energy” group divided per category.
 Category Model Used by
-Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_15], [art_16], [art_22], [art_25]
-Classical machine learning methods Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_15], [art_19]
-Convolutional neural network (CNN) [art_15], [art_18], [art_21], [art_22]
-Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [art_15], [art_17]
+Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_15], [art_16], [art_22], [art_25], [art_333], [art_337], [art_338], [art_341], [art_343], [art_349]
+Historical average (HA) [art_350], [art_390]
+Vector ARIMA [art_332], [art_333], [art_343], [art_350], [art_390]
+Non-GNN machine learning methods Autoformer [Autoformer] [art_333], [art_341]
+Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_15], [art_19], [art_339]
+Convolutional neural network (CNN) [art_15], [art_18], [art_21], [art_22], [art_343], [art_347]
+Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [art_15], [art_17], [art_351]
 Encoder-decorer architectures (ED) [art_15], [art_19], [art_21], [art_22]
-Feed-forward neural network (FNN) [art_15], [art_18], [art_19], [art_25], [art_27]
-Gated recurrent unit (GRU) [art_15], [art_25]
+Feed-forward neural network (FNN) [art_15], [art_18], [art_19], [art_25], [art_27], [art_351]
+Gated recurrent unit (GRU) [art_15], [art_25], [art_334], [art_338], [art_343], [art_347]
 Informer [Informer] [art_15], [art_16]
-$k$-nearest neighbors (KNN) [art_18], [art_24]
-Long-short term memory network (LSTM) [art_15], [art_16], [art_19], [art_20], [art_23], [art_24]
+$k$-nearest neighbors (KNN) [art_18], [art_24], [art_342], [art_343]
+Long-short term memory network (LSTM) [art_15], [art_16], [art_19], [art_20], [art_23], [art_24], [art_333], [art_334], [art_335], [art_340], [art_341], [art_343], [art_344], [art_347], [art_348], [art_349], [art_353]
+Long-short term time series network (LSTNet) [LSTNet] [art_15], [art_332], [art_334]
 Recurrent neural network (RNN) [art_16], [art_17], [art_19], [art_20]
-Support vector regression (SVR) [art_21], [art_22], [art_24], [art_27]
-Transformer [attention] [art_16], [art_19]
-GNN methods Spatio-temporal autoregressive model (ST-AR) [ST-AR] [art_21], [art_22]
-The majority of the papers only use classical machine learning benchmarks.
-Some papers use simple mathematical and statistical models, such as linear regression (that we incorporated into ARIMA) or historical average.
-Only few papers use other GNN-based models as benchmarks, and there are no common benchmarks among them.
-This is due to the fact that this field has not been extensively explored in the past, and only a very limited number of GNN models have been specifically developed for energy forecasting.
+Seq2Seq [Seq2seq] [art_335], [art_343]
+Support vector regression (SVR) [art_21], [art_22], [art_24], [art_27], [art_338], [art_342], [art_343], [art_344], [art_348]
+Transformer [attention] [art_16], [art_19], [art_335], [art_343]
+Extreme gradient boosting (XGBoost) [art_25], [art_334], [art_335], [art_341]
+GNN methods Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_349], [art_350], [art_390]
+Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_332], [art_340], [art_341]
+Spatio-temporal autoregressive model (ST-AR) [ST-AR] [art_21], [art_22]
+Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_339], [art_341], [art_348], [art_351], [art_353]
+The majority of the papers only use non-GNN machine learning benchmarks.
+Some papers use simple mathematical and statistical models, such as the ARIMA model and the historical average.
+Only few papers use GNN-based models as benchmarks, and the most popular benchmark among them is the spatio-temporal graph convolutional network (STGCN).
+Before 2024, the only benchmark used by two papers was the spatio-temporal autoregressive model (ST-AR).
+This is due to the fact that this thematic field has not been extensively explored in the past, and only a very limited number of GNN models have been specifically developed for energy forecasting.
 5.1.5 Results
-In all the papers, the authors claim that the proposed GNN model outperforms the benchmarks. The most common error metrics are the MAE, the RMSE, and the Normalized Root Mean Square Error (NRMSE).
-It is not possible to compare the accuracy of the discussed models because there are no common benchmark datasets.
+In general, the authors of the papers in this group claim that their proposed GNN model outperforms the benchmarks.
+However, in some cases, it turns out that basic statistical methods may achieve good results as well [art_390].
+The most common error metrics used are the MAE, the RMSE, and the Normalized Root mean squared Error (NRMSE).
+Some papers calculate these error metrics across different seasons to evaluate the model’s performance under different conditions.
+An example is provided in Ref.
+[art_349], where the authors also use the GNNExplainer [interpret_2] to assess the influence of different variables.
+It is not possible to compare the accuracy of the models included in this group, because no common benchmark datasets were used.
 The only two papers that study the same dataset are written by the same authors (Refs.
-[art_21] and [art_22]), and their most recent paper includes a comparison with the models proposed in the first one.
+[art_21, art_22]), and their most recent paper includes a comparison with the models proposed in the first one.
 5.2 Environment
-The group ”Environment” includes a significant number of the selected papers (27 out of 156), which suggests a considerable research interest on the application of GNNs in this field.
+The thematic group “Environment” includes a relevant number of selected papers (43 out of 281).
+This suggests a considerable research interest on the application of GNNs in this field.
 However, this group is highly fragmented, as it includes a large number of subdomains and applications, all related to the study of environmental data.
 5.2.1 Overview
-The group of environmental studies is broad and includes several disciplines such as physics, climatology, oceanography, and atmospheric science.
-The breadth of the group is reflected in the number of case studies presented in the selected papers, which cover air quality prediction (in terms of PM2.5 concentration), sea temperature, wind speed, and other applications.
+The group of papers related to environmental topics is broad, and includes papers on disciplines such as physics, climatology, oceanography, and atmospheric science.
+The breadth of the group is reflected in the number of case studies, which cover air quality prediction (in terms of PM2.5 concentration), sea temperature, wind speed, and other applications.
 Forecasting these quantities can be useful for several reasons.
@@
 Moreover, monitoring and forecasting environmental data are crucial for evaluating the impact of human activities on the environment and tracking the progression of climate change.
-The methods for forecasting environmental data are typically categorized as numerical models, statistical models, and machine learning methods.
+The methods for forecasting environmental data are typically categorized as numerical methods, statistical methods, and machine learning methods.
 Numerical methods are based on atmospheric models, and are designed to quantify the interaction of different atmospheric variables.
-The accuracy of these models depends on the availability of data, and the mathematical simulations to get the forecasts may take days or even weeks, thus limiting the ability to make good short-term forecasts.
+The accuracy of these models depends on the availability of data, and the mathematical simulations needed to get the forecasts may take days or even weeks, thus limiting the ability to make good short-term forecasts.
 Linear statistical methods leverage long-term dependencies in the recorded data by using regression-like models.
 However, they cannot capture non-linear relationships in the dataset, which limits their power.
-Machine learning models, on the other hand, are often a generalization of linear statistical methods that can capture more complex and non-linear relationships in the data.
-An overview of the selected papers reveals that there has been a rapid increase in interest in the field among the GNN research community, and more than 20 of the selected papers were published in 2023.
-The most investigated subjects are sea temperature ([art_28], [art_33], [art_35], [art_36], [art_39], [art_41], [art_47]), wind speed ([art_29], [art_31], [art_34], [art_40], [art_44], [art_46]), and PM2.5 concentration forecasting ([art_2], [art_3], [art_4], [art_6], [art_7], [art_8]).
-All the selected papers are published in different journals, with the exception of the Engineering Applications of Artificial Intelligence by Elsevier, which has published two of the papers.
+Machine learning methods, on the other hand, are often advanced generalizations of linear statistical methods that can capture more complex and non-linear relationships in the data.
+An overview of the selected papers reveals that there has been a rapid increase in interest on this theme among the GNN research community too, and more than 35 of the selected papers were published from 2023.
+The most investigated data are related to sea temperature (Refs.
+[art_28, art_33, art_35, art_36, art_39, art_41, art_47]), wind speed (Refs.
+[art_29, art_31, art_34, art_40, art_44, art_46, art_366]), and PM2.5 concentration forecasting (Refs.
+[art_2, art_3, art_4, art_6, art_7, art_8, art_354, art_361, art_371]).
+Almost every selected paper is published in a different journal, with the exception of Energy, Engineering Applications of Artificial Intelligence, Journal of Cleaner Production, Knowledge-Based Systems, Science of the Total Environment by Elsevier, each with two published papers.
 The majority of the corresponding authors are affiliated with institutions in China.
 5.2.2 Datasets
-The analysis of the selected papers shows that in this group of papers there is no reference benchmark dataset.
+The analysis of the selected papers shows that in this thematic group there is no standardized benchmark dataset.
 The publicly available datasets cited in the papers are listed in Tab.
 LABEL:t:link_dataset_environment.
-Table 5: List of public datasets in the ”Environment” group and their corresponding links.
+Table 9: List of public datasets in the “Environment” group and their corresponding links.
 Dataset Used by Link
 AgroData-10 [art_28] https://www.aoml.noaa.gov/phod/argo
+Carbon Monitor [art_360] https://carbonmonitor.org/
 CCMP wind data [art_31], [art_40], [art_44] http://data.remss.com
-China Environmental Monitoring Station PM2.5 [art_2], [art_8] https://www.cnemc.cn/
+China Environmental Monitoring Station PM2.5 [art_2], [art_8], [art_371] https://www.cnemc.cn/
 China National Urban Air Quality [art_1], [art_5] https://github.com/Friger/GAGNN
@@
 ECMWF South and East China Sea [art_32] http://apps.ecmwf.int/datasets/data/interim-full-daily
-KnowAir [art_6] https://doi.org/10.1145/3397536.3422208 (source paper)
+HadCRUT5 [art_370] https://doi.org/10.1029/2019JD032361 (source paper)
+Jinan PM2.5 [art_354] https://github.com/liaohaibing/MGST_GNN/tree/master
+KnowAir [art_6] https://github.com/shuowang-ai/PM2.5-GNN
 Korean Peninsula sea data [art_33] https://www.nifs.go.kr/risa/main.risa
-NOAA sea datasets [art_28], [art_36], [art_39], [art_46], [art_47] https://www.ndbc.noaa.gov/
+NOAA sea datasets [art_28], [art_36], [art_39], [art_46], [art_47], [art_365] https://www.ndbc.noaa.gov/
 Norwegian offshore wind [art_29], [art_34] https://frost.met.no/index.html
+Overbetuwe groundwater data [art_359] https://github.com/mlttac/GroundwaterFlowGNN
 Taiwan air quality dataset [art_3] https://data.gov.tw/en
@@
 US EPA PM2.5 [art_1], [art_4] https://aqs.epa.gov/aqsweb/airdata/download_files.html
-Five of the selected papers ([art_28], [art_36], [art_39], [art_46], [art_47]) use National Oceanic and Atmospheric Administration (NOAA) sea datasets (https://www.noaa.gov/) from the U.S.
+US Geological Survey [art_355] https://earthquake.usgs.gov/earthquakes/search/
+Yangtze River PM2.5 [art_354] https://github.com/liaohaibing/MGST_GNN/tree/master
+Six of the selected papers [art_28, art_36, art_39, art_46, art_47, art_365] use National Oceanic and Atmospheric Administration (NOAA) sea datasets (https://www.noaa.gov/) from the U.S.
 Department of Commerce (US DOC).
@@
 In over half of the papers, exogenous variables are incorporated into the model.
-These include weather data (e.g., humidity, rainfall, pressure, temperature, wind) and, in the case of PM2.5 concentration forecasting, other pollutants (e.g., CO, NO2, O3, PM10, SO2).
-The data granularity ranges from 1 minute to 1 month, depending on the case study, and the forecasting horizon from 10 minutes to 240 days.
-Most datasets require some pre-processing, including interpolation of missing data, removal of outliers and normalization.
-5.2.3 Proposed models
-Regarding the taxonomy of the proposed GNN models, the majority of papers present convolutional GNN (17 out of 27 papers), followed by 6 attentional GNNs and 3 hybrid architectures.
-The description in one paper (Ref.
-[art_5]), however, does not specify the classification of its model within the proposed taxonomy.
+These include weather data (e.g., humidity, rainfall, pressure, temperature, wind) and, in the case of PM2.5 concentration forecasting, concentration of pollutants (like CO, NO2, O3, PM10 and SO2).
+Overall, many papers also consider exogenous variables, in particular weather data, as they have the greatest influence on environmental data.
+The data time granularity ranges from 1 minute to 1 month, depending on the case study, and the forecasting horizon ranges from 10 minutes to more than one year.
+Most datasets require some type of pre-processing, including interpolation of missing data, removal of outliers and normalization.
+5.2.3 Types of models
+Regarding the taxonomy of the proposed GNN models, the majority of papers present convolutional GNNs (28 out of 43 papers), followed by 9 attentional GNNs, and by 3 hybrid architectures.
+The description in 3 of the papers (Refs.
+[art_5, art_354, art_361]), however, does not allow us to specify the classification of the model within the proposed taxonomy.
 As for the graph structure, not all of the papers describe it with sufficient precision, and some of them are vague about the graph construction and its static or dynamic nature.
 In almost half of the papers, the edge weights are learned directly from the model, often using the graph attention mechanism.
-When the graph structure is defined manually, the most popular criterion is the spatial distance between the sites ($d$), even in its exponential ($e^{-d}$) or inverse proportion ($1/d$) formulation.
+When the graph structure is defined manually, the most popular criterion used is the spatial distance between the sites ($d$), even in its exponential ($e^{-d}$) or inverse proportion ($1/d$) formulation, or similarity measures, such as cosine similarity and correlation of the time series.
 Finally, there are a few papers where more than one graph structure is fused together in order to capture more complex dynamics, as in Ref.
 [art_2].
-Not all papers specify the loss function used to train the models.
-Among those that do, MAE and MSE are used in roughly equal proportions.
-Many papers mention the programming language Python and the library PyTorch for the implementation of the models.
+Almost all the papers in this thematic group deal with multivariate time series, with the exception of Ref.
+[art_370], which studies a univariate time series of global average temperature.
+In this work, the authors treat each time step $t$ as a node $v_{t}$ in the graph, and the adjacency matrix is constructed in such a way that $v_{t}$ is influenced by $v_{t-1}$ and $v_{t+1}$.
+Not all papers openly write down the loss function used to train the models.
+Among those that do, the most popular loss functions are MSE and MAE.
+Many papers mention the use of the programming language Python and of the library PyTorch for the implementation of the models.
 Few papers also provide a link to the source code for the proposed model, as shown in Tab.
 LABEL:t:link_codici_environment.
-Table 6: List of source codes of the ”Environment” models in the review.
+Table 10: List of source codes of the “Environment” models in the review.
 Model Link
+Dynamic Spatial-Temporal Graph Convolutional Recurrent Network (DSTGCRN) [art_360] https://github.com/PalaceTony/DSTGCRN
+Dynamic spatiotemporal graph neural network (DST-GNN) [art_354] https://github.com/liaohaibing/MGST_GNN/tree/master
 Group-aware graph neural network (GAGNN) [art_1] https://github.com/Friger/GAGNN
 Gridded information propagation and mixing network (GIPMN) [art_45] https://github.com/Boltzxuann/GIPMN/tree/dataset
+GNNCoder [art_356] https://doi.org/10.5281/zenodo.13358007
+GroundwaterFlowGNN [art_359] https://github.com/mlttac/GroundwaterFlowGNN
 Hierarchical graph recurrent network (HiGRN) [art_28] https://github.com/Neoyanghc/HiGRN
+Physics-aware graph networks (PaGN) [art_365] https://github.com/eunviho/novel_pagn
 Spatio-temporal FFTransformer (ST-FFTransformer) [art_29] https://github.com/LarsBentsen/FFTransformer
 5.2.4 Benchmark models
-Tab. LABEL:t:benchmark_models_environment presents the benchmark models employed in the ”Environment” group by at least two articles examined in this review.
-Table 7: List of benchmark models in the ”Environment” group divided per category.
+Tab. LABEL:t:benchmark_models_environment presents the benchmark models employed in the “Environment” group by at least two articles examined in this review.
+Table 11: List of benchmark models in the “Environment” group divided per category.
 Category Model Used by
-Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_28], [art_33], [art_34], [art_38]
+Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_28], [art_33], [art_34], [art_38], [art_363], [art_370], [art_371]
 Historical average (HA) [art_3], [art_6], [art_28], [art_38]
-Naive [art_29], [art_34]
-Vector autoregression (VAR) [art_28], [art_34], [art_38]
-Classical machine learning methods Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_33], [art_46]
-Combined FC-LSTM and convolution neural network (CFCC-LSTM) [CFCC-LSTM] [art_35], [art_39]
+Naive [art_29], [art_34], [art_366]
+Vector autoregression (VAR) [art_28], [art_34], [art_38], [art_360]
+Non-GNN machine learning methods Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_33], [art_46]
+Combined FC-LSTM and convolution neural network (CFCC-LSTM) [CFCC-LSTM] [art_35], [art_39], [art_370]
+DLinear [DLinear] [art_31], [art_364]
 Hybrid convolutional neural network and gated recurrent unit (CGRU) [CGRU] [art_31], [art_40], [art_44]
 Convolutional neural network (CNN) [art_30], [art_46]
-Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [art_4], [art_41]
-Convolutional long-short term memory network (ConvLSTM) [ConvLSTM] [art_28], [art_41]
-Feed-forward neural network (FNN) [art_2], [art_7], [art_38]
+Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [art_4], [art_41], [art_368], [art_371]
+Convolutional long-short term memory network (ConvLSTM) [ConvLSTM] [art_28], [art_41], [art_355]
+Feed-forward neural network (FNN) [art_2], [art_7], [art_38], [art_360], [art_365], [art_370]
 Gated recurrent unit encoder–decoder (GED) [GED] [art_36], [art_39], [art_47]
-Gated recurrent unit (GRU) [art_2], [art_6], [art_7], [art_8], [art_38], [art_43], [art_46]
-Informer [Informer] [art_30], [art_47]
-Long-short term memory network (LSTM) [art_1], [art_2], [art_3], [art_4], [art_5], [art_6], [art_7], [art_8], [art_28], [art_30], [art_32], [art_33], [art_35], [art_36], [art_38], [art_39], [art_43], [art_46], [art_47]
+Gated recurrent unit (GRU) [art_2], [art_6], [art_7], [art_8], [art_38], [art_43], [art_46], [art_361], [art_364], [art_365], [art_367], [art_370]
+Informer [Informer] [art_30], [art_47], [art_354], [art_364], [art_371]
+Long-short term time series network (LSTNet) [LSTNet] [art_355], [art_373]
 Predictive deep convolutional neural network (PDCNN) [PDCNN] [art_31], [art_40]
-Support vector regression (SVR) [art_3], [art_6], [art_35], [art_36], [art_38]
-Temporal convolutional network (TCN) [TCN] [art_6], [art_30]
+Random forest (RF) [art_6], [art_364], [art_370]
+Support vector regression (SVR) [art_3], [art_6], [art_35], [art_36], [art_38], [art_360], [art_370], [art_371]
+Recurrent neural network (RNN) [art_46], [art_365], [art_367]
+Temporal convolutional network (TCN) [TCN] [art_6], [art_30], [art_355], [art_356], [art_368], [art_371], [art_373]
 Temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM] [art_6], [art_31], [art_32], [art_35], [art_40], [art_44]
 Transformer [attention] [art_30], [art_33]
-Hybrid wavelet random forest and deep belief network (Wavelet-DBN-RF) [art_31], [art_40], [art_44]
-Extreme gradient boosting (XGBoost) [art_1], [art_2], [art_5], [art_46]
-GNN methods Deeper graph convolutional network (DeeperGCN) [DeeperGCN] [art_1], [art_5]
-Graph convolutional neural network and long-short term memory (GC-LSTM) [GC-LSTM] [art_1], [art_2], [art_5], [art_6], [art_7]
-Graph convolutional network (GCN) [art_38], [art_41]
-Graph WaveNet (GWN) [GWN] [art_1], [art_6], [art_44], [art_47]
+Hybrid wavelet random forest and deep belief network (Wavelet-DBN-RF) [Wavelet-DBN-RF] [art_31], [art_40], [art_44]
+Extreme gradient boosting (XGBoost) [art_1], [art_2], [art_5], [art_46], [art_364]
+GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_6], [art_360], [art_365], [art_368]
+Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_363], [art_368]
+Deeper graph convolutional network (DeeperGCN) [DeeperGCN] [art_1], [art_5]
+Graph convolutional neural network and long-short term memory (GC-LSTM) [GC-LSTM] [art_1], [art_2], [art_5], [art_6], [art_7], [art_354], [art_361]
+Graph convolutional network (GCN) [art_38], [art_41], [art_357], [art_367]
+Graph WaveNet (GWN) [GWN] [art_1], [art_6], [art_44], [art_47], [art_363], [art_368]
 Hierarchical graph convolution network (HGCN) [HGCN] [art_1], [art_28]
-Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_6], [art_32], [art_44]
+Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_6], [art_32], [art_44], [art_355], [art_360], [art_373]
 Multi long-short term memory network (Multi-LSTM) [Multi-LSTM] [art_31], [art_40], [art_44]
 Superposition graph neural network (SGNN) [art_24] [art_31], [art_40]
-Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_3], [art_6], [art_28], [art_35], [art_47]
+Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_3], [art_6], [art_28], [art_35], [art_47], [art_363], [art_368], [art_371]
 Spatio-temporal correlation graph neural network (STGN) [art_40] [art_31], [art_44]
 Spatio-Temporal U-Network (ST-UNet) [ST-UNet] [art_1], [art_5]
+Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_368], [art_371]
 Temporal graph convolutional network (T-GCN) [T-GCN] [art_6], [art_30], [art_35]
-The majority of the benchmarks used fall within the category of classical machine learning, although there are several GNN benchmarks as well.
-The two most common benchmark models are LSTM and GRU.
-These models are widely used by environmental data forecasters due to their ability to capture temporal dependencies, and they are considered a reliable benchmark in the field.
-As for GNN models, the most common benckmarks are the graph convolutional neural network and long-short term memory (GC-LSTM) [GC-LSTM] (which appears only in quality air forecasting problems) and spatio-temporal graph convolutional network (STGCN) [STGCN].
-The GC-LSTM is a model that integrates graph convolutional networks and long-short-term memory networks, proposed by Qi et al.
+The majority of the benchmarks used in this thematic group fall within the category of non-GNN machine learning methods, although there are several GNN benchmarks as well.
+The two most common benchmark models are LSTMs and GRUs.
+These models are widely used by environmental data forecasters due to their ability to capture temporal dependencies, and are considered a reliable benchmark in the field.
+As for GNN models, the most common benckmarks are the graph convolutional neural network and long-short term memory newtork (GC-LSTM) [GC-LSTM] (which appears only in quality air forecasting problems), and the spatio-temporal graph convolutional network (STGCN) [STGCN].
+The GC-LSTM network is a model that integrates graph convolutional networks and long-short-term memory networks.
+It was proposed by Qi et al.
 in 2019 to forecast the spatio-temporal variation of PM2.5 concentration.
-The STGCN model instead was proposed in the traffic domain by Yu et al.
-in 2018.
-This model is composed of two spatio-temporal convolutional blocks and a fully-connected output layer.
+In the original paper, the GC-LSTM model works with a pre-defined graph based on the geographical distance between air quality stations.
+The input layer of the GC-LSTM model receives historical air quality observations, meteorological variables, spatial predictors, and temporal predictors.
+This information is processed through a graph convolution operation to extract spatial features.
+After that, the temporal feature extraction is performed using a LSTM network, which takes as input the graph convolutional features concatenated with the original signals.
+The STGCN model, on its part, was proposed in the traffic domain by Yu et al.
+in 2018, and works with pre-defined adjacency matrices computed based on the distances among stations in the traffic network.
+The STGCN model is composed of two spatio-temporal convolutional blocks and a fully-connected output layer.
 Each convolutional block contains two temporal gated convolution layers and one spatial graph convolution layer in the middle.
 The code for the STGCN model is available at https://github.com/VeritasYin/STGCN_IJCAI-18.
-Finally, it is important to point out that the multivariate time series forecasting with graph neural network (MTGNN) model proposed by Wu et al.
+Finally, it is important to point out that the so-called multivariate time series forecasting with graph neural network (MTGNN) model proposed by Wu et al.
 in 2020 in Ref.
-[MTGNN] is often the second-best performing model in the papers using it, thus it deserves more attention and further investigation in future research.
+[MTGNN] is often reported in the papers using it as the second-best performing model, thus it probably deserves special attention and further investigation in future research.
 5.2.5 Results
-In each paper, the proposed model significantly improves on the benchmarks.
+In each paper of this group, it is stated that the proposed model is significantly better than the benchmarks.
 The accuracy of the models is measured in terms of MAE, RMSE, and MAPE.
 Among the papers, there is a high diversity in the choice of the proposed datasets.
-However, there are two sets of papers using the same dataset, whose results can be compared.
+However, there are two sets of papers using the same dataset, whose results are compared hereafter.
 The first dataset is the CCMP wind speed dataset (available at http://data.remss.com), which contains 3,942 geographical grid points from the east and southeast sea areas of China.
-The three papers using it ([art_31], [art_40], [art_44]) select 120 points from the area and use data records from January 2010 to April 2019 at 6 hours intervals, with a total of 13,624 samples at each node, for 10 different forecasting horizons.
+The three papers using it (Refs.
+[art_31, art_40, art_44]) select 120 points from the area and use data records from January 2010 to April 2019 at 6 hours intervals, with a total of 13,624 samples at each node, for 10 different forecasting horizons.
 All papers present results for the complete set of 120 nodes and a subset of nodes.
-However, since their results are similar, only the average results for all the 120 nodes are reported in Tab.
+Since these results are similar, only the average results for all the 120 nodes are reported in Tab.
 LABEL:t:risultati_ccmp for the sake of simplicity.
-Table 8: Comparison of the average accuracy of the different models on the CCMP wind speed dataset for time horizons from 6 hours to 7 days ahead, expressed in terms of MAE and RMSE.
+Table 12: Comparison of the average accuracy of the different models on the CCMP wind speed dataset for different time horizons, expressed in terms of MAE and RMSE.
 The smallest errors are underlined.
-Numbers from the papers.
+Numbers from the original papers.
 Model Metrics 6 h 12 h 18 h 1 d 2 d 3 d 4 d 5 d 6 d 7 d
@@
 RMSE 1.855 2.038 2.226 2.328 2.617 2.703 2.744 2.766 2.788 2.798
-For all the time horizons, the most recent dynamic adaptive spatio-temporal graph neural network (DASTGN) model [art_44] has the highest accuracy.
-Other two well performing models are the data-based adaptive graph learning network (DAGLN) [art_31] and the multivariate time series forecasting with graph neural network (MTGNN) [MTGNN].
-The second dataset shared by two papers ([art_1], [art_5]) is the one of China National Urban Air Quality (available at https://github.com/Friger/GAGNN).
+For all the time horizons considered in the papers, the dynamic adaptive spatio-temporal graph neural network (DASTGN) model [art_44] proposed by Gao et al.
+in 2023 has the highest accuracy.
+The DASTGN model is composed of a graph generation module and a spatio-temporal convolution module.
+The dynamic adaptive graph generation module generates a static graph to model long-term stable dependencies, and a dynamic graph to model the short-term dynamics.
+These graphs are then merged into an optimal dynamic graph.
+The spatio-temporal convolution module consists of an input module, a stack of spatio-temporal layers connected with residual and skip connections, and an output module that generates wind speed forecasts.
+Other two well performing models are the data-based adaptive graph learning network (DAGLN) [art_31], and the multivariate time series forecasting with graph neural network (MTGNN) [MTGNN].
+The DAGLN model was proposed by Ren at al.
+in 2023, and uses an encoder-decodes architecture.
+The encoder includes an adaptive graph learning module that learns the optimal adjacency matrix starting from a pre-defined one obtained by using the DTW algorithm, a graph convolution with node adaptive parameter learning algorithm for personalized learning of parameter space, and a GRU for temporal modeling.
+Then, this information is fed into the decoder that operates a convolutional and a linear operation to produce the wind speed forecast.
+The MTGNN model, proposed by Wu et al.
+in 2020, is a well-known state-of-the-art model for multivariate time series forecasting.
+It incorporates a graph learning layer that learns a graph adjacency matrix from scratch, followed by a stack of graph convolution and temporal convolution modules to capture spatial and temporal dependencies, and an output module.
+To avoid the problem of vanishing gradients, residual and skip connections are added between spatial and temporal convolutional modules.
+The second dataset shared by two papers (Refs.
+[art_1, art_5]) is called China National Urban Air Quality dataset (available at https://github.com/Friger/GAGNN).
 It contains air quality index data (AQI) of 209 cities, collected from January 2017 to April 2019 at 1 hour granularity, for a total of 20,370 samples.
-Both the papers also integrate exogenous weather variables (humidity, wind direction, wind speed, rainfall, air pressure, temperature, and visibility) collected from Envicloud (http://www.envicloud.cn/).
+Both these papers integrate exogenous weather variables (humidity, wind direction, wind speed, rainfall, air pressure, temperature, and visibility) collected from Envicloud (http://www.envicloud.cn/).
 The average accuracy of the models used in the two papers is compared in Tab.
@@
 The smallest errors are underlined.
-Numbers from the papers.
+Numbers from the original papers.
 Model Metrics 1h 2h 3h 4h 5h 6h
@@
 RMSE 14.25 19.80 24.72 28.14 30.63 33.44
-The hybrid interpretable neural network and graph neural network model (INNGNN) [art_5] is the most accurate model for all time horizons except 1 hour, where it is beaten by the meta-graph convolutional recurrent network model (MegaCRN) [MegaCRN].
-The authors attribute this result to the fact that in the short term the AQI is strongly influenced by nearby cities, and the MegaCRN model is more focused on local features.
+As from Tab.
+LABEL:t:risultati_chinese_aqi, the hybrid interpretable neural network and graph neural network model (INNGNN) [art_5] is the most accurate model for all time horizons except the 1 hour horizon, where it is beaten by the meta-graph convolutional recurrent network model (MegaCRN) [MegaCRN].
+The authors attribute this result to the fact that in the short term the AQI is strongly influenced by nearby cities, whereas the MegaCRN model is more focused on local features.
+The INNGNN model, proposed by Ding et al.
+in 2023, consists of a temporal and a spatial dependency module.
+The temporal module employs self-attention and interpretable neural networks (INN) to extract local and global features of time series data, effectively identifying periodicity and trend of the AQI time series.
+After that, the spatial module has an encoder-decoder architecture where GNN modules extract spatial dependencies across geographic locations, represented by an adjacency matrix based on spatial distance.
+The MegaCRN model, introduced by Jiang et al.
+in 2023 for traffic forecasting, includes a meta-graph learner module and an encoder-decoder spatial module built using a series of convolutional GNNs.
+The source code for the MegaCRN model can be accessed at https://github.com/deepkashiwa20/MegaCRN.
 5.3 Finance
-”Finance” is another emerging field of study, though not yet widely explored, with 9 out of 156 selected papers.
+“Finance” is another emerging theme of study, though not yet widely explored, with just 16 out of 281 selected papers.
 The difficulty of predicting financial data is a well-established fact.
@@
 This complexity makes financial time series highly volatile, and accurate forecasts and data modeling are essential to develop effective trading strategies.
-GNNs are taken into consideration to model the inter-dependencies between variables together with series dynamics.
+Spatio-temporal GNNs are taken into consideration to model the inter-dependencies between variables together with series dynamics.
 5.3.1 Overview
-The ”Finance” group is relatively narrow, in part due to the limited number of published papers.
-The majority of studies are concentrated on stock price prediction ([art_59], [art_60], [art_62], [art_63], [art_65]), with the aim of forecasting the future trend of the price of a company.
+The “Finance” thematic group is relatively narrow in terms of sub-themes.
+This is in part due to the limited number of published papers.
+The majority of studies are concentrated on stock price forecasting [art_59, art_60, art_62, art_63, art_65, art_377, art_382], with the aim of forecasting the future trend of the price of a company.
 Sometimes the stock prediction problem is seen as a classification task.
-Indeed, due to the stochastic nature and volatility of financial markets, predicting price movements (”up”, ”neutral”, ”down”) can be more feasible than accurately forecasting specific prices.
-For example, assuming a daily granularity, a rise (”up”) can be considered to occur when the next trading day’s closing price ($p_{t}$) is at least 0.55% higher than the previous day’s closing price ($p_{t-1}$), whereas if $p_{t}$ is more than 0.50% lower than $p_{t-1}$, the price movement can be classified as a ”down”.
-Otherwise, it is classified as ”neutral”.
-All the selected papers addressing stock prices approach the problem from a classification perspective.
-Another distinctive aspect in this group of papers is that researchers usually also evaluate how their predictions affect the stability of investment returns over a specified period.
-This is commonly done through trading simulations in the test set, evaluated with metrics such as the Sharpe ratio and overall returns.
+This is because, due to the stochastic nature and volatility of financial markets, predicting just the direction of price movements (“up”, “neutral”, “down”) can be sometimes more feasible than accurately forecasting price values.
+In this case, assuming a daily granularity, a rise (“up”) can be considered to occur when the next trading day’s closing price ($p_{t}$) is at least 0.55% higher than the previous day’s closing price ($p_{t-1}$), whereas if $p_{t}$ is more than 0.50% lower than $p_{t-1}$, the price movement can be classified as a “down”.
+Otherwise, it is classified as “neutral”.
+All the selected GNN papers addressing stock prices indeed approach the price forecasting problem from a classification perspective.
+Another distinctive aspect in this group of papers is that researchers usually evaluate, among others, how their predictions affect the stability of investment returns over a specified period.
+This is commonly done through trading simulations in the test set, evaluated with metrics such as the Sharpe ratio and overall returns [art_59, art_60, art_65, art_66, art_377].
 These simulations require modeling rather than forecasting.
-An examination of the selected papers indicates an interest in the application of GNNs in Finance from 2022.
+An examination of the selected papers indicates an interest in the application of spatio-temporal GNNs in Finance which starts from 2022.
 Two of the papers were published in IEEE Access by IEEE, while the remaining ones appeared in various other journals.
@@
 The number of datasets in this group is limited. The links to the public datasets referenced in the selected papers are listed in Tab. LABEL:t:link_dataset_finance.
-Table 10: List of public datasets in the ”Finance” group and their corresponding links.
+Table 14: List of public datasets in the “Finance” group and their corresponding links.
 Dataset Used by Link
-China A-Shares [art_63], [art_65], [art_66] https://finance.sina.com.cn/stock
+China A-Shares [art_63], [art_65], [art_66], [art_377], [art_378] https://finance.sina.com.cn/stock
+Criptocurriences [art_381] https://coinmarketcap.com/
 CSI 100 index [art_62] https://tushare.pro/
-CSI 300 index [art_59], [art_62] https://global.csmar.com/
+CSI 300 index [art_59], [art_62], [art_382] https://global.csmar.com/
+Gold and silver market [art_374] https://www.kitco.com/
+HAL agriculture [art_379] www.hal.gov.tr
 S&P 100 index [art_62] https://finance.yahoo.com/
@@
 Russell 3K index [art_62] https://finance.yahoo.com/
+US crude oil market [art_374] http://www.eia.gov/
 All the listed datasets represent collections of stock indices.
@@
 Many studies also include exogenous data, such as other market data, events, and news.
-As for the news, they can be processed by some language model like BERT (as in Refs.
-[art_59] and [art_66]).
-Most of the data have a granularity of 1 day (which is typical of financial data), and the forecasting horizon is also generally 1 day.
-Only one of the selected papers (Ref.
-[art_62]) analyses also longer time horizons for stock forecasting.
-As for pre-processing, half of the selected papers use normalization techniques, including min-max normalization, Z-score normalization, and logarithmic normalization.
-5.3.3 Proposed models
-In terms of the taxonomy of the proposed GNN models, 5 out of 9 papers utilize attentional GNNs, while the other 4 papers employ convolutional GNNs.
-As for the graph structure, some papers use models which learn the graph structure themselves ([art_62], [art_64]), while some others defines it a priori based on the relationships between companies and stocks ([art_63], [art_67], [art_68]).
-Since the majority of papers deal with a stock movement classification problem, the most common loss function is the cross entropy.
-The majority of the papers explicitly indicate that Python is the programming language utilized, with nearly half employing PyTorch and a few utilizing TensorFlow.
-Only Ref.
-[art_63] provides a link to the source code of the proposed knowledge graph and graph convolution neural network (KG-GCN) model, available at https://github.com/Gjl12321/KG_GCN_Stock_Price_Trend_Prediction_System.
+News can be processed by some language model like BERT (as in Refs.
+[art_59, art_66, art_378]).
+Most of the data have a granularity of 1 day, and the forecasting horizon is also generally 1 day.
+Two of the selected papers (Refs.
+[art_62, art_377]) analyses longer time horizons for stock forecasting, and others do the same for agricultural price (Ref.
+[art_379]) or cryptocurrencies’ price (Ref.
+[art_381]).
+As for pre-processing techniques, half of the selected papers use normalization techniques, including min-max normalization, Z-score normalization, and logarithmic normalization.
+5.3.3 Types of models
+In terms of the taxonomy of the proposed GNN models, 6 out of 16 papers utilize attentional GNNs, 9 papers employ convolutional GNNs, and one paper takes into account both convolutional and attentional GNNs (Ref.
+[art_374]).
+As for the graph structure, some papers use models which learn the graph structure themselves (e.g., Refs.
+[art_62, art_64, art_378, art_379, art_382]), whereas some others define it a priori, based on the relationships between companies and stocks (e.g., Refs.
+[art_63, art_67, art_68, art_381]).
+Half of the papers deal with a stock movement classification problem, so the most commonly used loss function is cross entropy.
+The majority of the papers explicitly indicate that Python is the programming language utilized, with nearly half employing PyTorch, and a few utilizing TensorFlow.
+Only Refs.
+[art_63] and [art_379] provide a link to the source code of the proposed models, which are, respectively, the knowledge graph with graph convolution neural network (KG-GCN) model, available at https://github.com/Gjl12321/KG_GCN_Stock_Price_Trend_Prediction_System, and the spectral temporal graph neural network (StemGNN), available at https://github.com/cevher/Time-Series-Analysis.
 5.3.4 Benchmark models
-Tab. LABEL:t:benchmark_models_finance displays the benchmark models utilized in the ”Finance” group by a minimum of two of the selected papers.
-Table 11: List of benchmark models in the ”Finance” group divided per category.
+Tab. LABEL:t:benchmark_models_finance displays the benchmark models utilized in the “Finance” group by a minimum of two of the selected papers.
+Table 15: List of benchmark models in the “Finance” group divided per category.
 Category Model Used by
-Mathematical and statistical methods Mean reversion (MR) [art_63], [art_65]
-Classical machine learning methods Convolutional neural network (CNN) [art_59], [art_64]
-Dual-stage attention-based recurrent neural network (DARNN) [art_63], [art_65]
-Feed-forward neural network (FNN) [art_59], [art_68]
-Gated recurrent unit (GRU) [art_60], [art_68]
-Long-short term memory network (LSTM) [art_59], [art_62], [art_63], [art_64], [art_65], [art_67], [art_68]
-Support vector machine (SVM) [art_64], [art_68]
-GNN methods Graph convolutional network (GCN) [art_59], [art_60], [art_62], [art_64], [art_68]
+Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_68], [art_375]
+Mean reversion (MR) [art_63], [art_65], [art_377]
+Non-GNN machine learning methods Convolutional neural network (CNN) [art_59], [art_64], [art_374], [art_379]
+Dual-stage attention-based recurrent neural network (DARNN) [art_63], [art_65], [art_377]
+Feed-forward neural network (FNN) [art_59], [art_68], [art_375], [art_378]
+Gated recurrent unit (GRU) [art_60], [art_68], [art_382]
+Long-short term memory network (LSTM) [art_59], [art_62], [art_63], [art_64], [art_65], [art_67], [art_68], [art_374], [art_375], [art_377], [art_379], [art_381]
+Support vector machine (SVM) [art_64], [art_68], [art_375], [art_378]
+GNN methods Graph convolutional network (GCN) [art_59], [art_60], [art_62], [art_64], [art_68], [art_382]
+Graph convolutional network with long short term memory network (GCN+LSTM) [art_65], [art_377]
 Financial Graph Attention Network (FinGAT) [FinGAT] [art_60], [art_62]
-Hierarchical graph attention network (HATS) [HATS] [art_62], [art_63], [art_65]
-Spatiotemporal hypergraph convolution network (STHGCN) [STHGCN] [art_63], [art_65]
-It is evident that there is no consensus regarding the most appropriate benchmark models, especially in the context of GNNs.
-The most common benchmark is the LSTM, which is considered an effective approach for processing long term information.
+Hierarchical graph attention network (HATS) [HATS] [art_62], [art_63], [art_65], [art_377]
+Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_62], [art_382]
+Spatiotemporal hypergraph convolution network (STHGCN) [STHGCN] [art_63], [art_65], [art_377]
+It is evident that there is no consensus regarding the most appropriate benchmark model to use, especially of GNN type.
+The limited number of GNN-based benchmarks can be attributed to the fact that GNN models have not been extensively explored in the financial domain.
+The most common benchmark is the LSTM network, which is considered an effective approach for processing long term information in time series.
 The most commonly used GNN model is a vanilla graph convolutional network (GCN), followed by a more complex hierarchical graph attention network (HATS) proposed by Raehyun et al.
 in 2019 [HATS].
-The HATS model has been proposed for predicting stock movements using relational data.
+The HATS model was proposed for predicting stock movements using relational data.
 It aggregates information from different types of relationships among data and incorporates this information into each stock representation.
+From an architectural point of view, the model starts with employing LSTM and GRU networks as feature extraction modules.
+Next, a hierarchical GNN-based architecture is designed to capture the significance of neighboring nodes and relation type across different layers.
+Finally, the node representations are fed into a task-specific module, which can be implemented as a simple linear transformation layer.
 The source code for the HATS model is available at https://github.com/dmis-lab/hats.
@@
 The accuracy of the proposed classification models is primarily evaluated in terms of accuracy, precision and F1 score.
-The only two papers with comparable results are Refs.
+The only two papers with results which can be compared are Refs.
 [art_63] and [art_65], which study 758 frequently traded stocks collected from the A-share market in China from January 2013 to December 2019 (available at https://finance.sina.com.cn/stock).
@@
 The two best models are the knowledge graph and graph convolution neural network (KG-GCN) [art_63] and the stock price movement prediction via convolutional network with adversarial hypergraph model (MONEY) [art_65], depending on the metrics considered.
+The KG-GCN model, proposed by Wang et al.
+in 2023, first constructs a knowledge graph to represent the semantic relationships and quantify correlations between stocks.
+Then a community detection module based on GCN node embeddings and K-means clustering is used to group similar stocks together.
+Finally, the historical prices of similar stocks are fed into LSTM and GRU networks to predict stock price trends.
+The MONEY model, introduced by Sun et al.
+in 2023, is composed of a convolutional GNN and an hypergraph convolution network to capture complex relations among similar stocks.
 5.4 Health
-The fourth group of papers is ”Health”, which includes 8 out of 156 publications.
+The fourth thematic group of papers is “Health”, which includes 15 out of 281 publications.
 These papers focus on critical aspects of health monitoring, disease modeling, and diagnostic tools, with particular emphasis on the spread and diagnosis of disease.
@@
 5.4.1 Overview
-Applications of GNNs in the ”Health” group have been studied starting from 2021.
-The reason of this spread may be attributed to two key factors.
-The first reason is the growing presence of sensors in hospitals, which enable the collection of multivariate time series signals which can be used for the timely diagnosis of chronic diseases.
-The second reason is the spread of the COVID-19 pandemic, which has motivated researchers to investigate and forecast the spread of the virus.
-Epidemic forecasting is the most common topic among the selected papers.
+Applications of GNNs to the “Health” thematic group have been studied starting from 2021.
+The reason of this expansion may be attributed to two key factors.
+The first factor is the growing presence of sensors in hospitals, which enable the collection of multivariate time series signals, which can be used for the timely diagnosis of chronic diseases.
+The second factor is the spread of the COVID-19 pandemic, which motivated researchers to investigate and to forecast the spread of the virus.
+Epidemic forecasting is indeed the most common topic among the selected papers.
 The prediction of epidemics presents two significant challenges.
-First, their seasonality can vary greatly over time, and they can manifest with different intensities and durations.
+First, epidemics’ seasonality can vary greatly over time, and they can manifest with different intensities and durations.
 Second, occasional pandemics can disrupt established patterns even for years.
-These complexities make the forecasting task more difficult, and machine learning and GNN models have been studied to address this issue.
-A common reference model for epidemic prediction is the SIR model [modello_SIR], an acronym referring to individuals which stands for susceptible (S), infectious (I), and recovered (R).
-The SIR model divides the population into individuals who are susceptible to the disease, those who are infected and can spread it, and those who have recovered and are now immune, and uses differential equations to describe how the disease spreads.
+These complexities make the forecasting task more difficult, and non-GNN machine learning, and GNN models themselves, have been studied to address this issue.
+Moreover, the spread of large-scale human diseases is closely related to both the mobility patterns of populations and the demographic characteristics of residents in different regions [art_413].
+Consequently, the application of GNN models to epidemic dynamics appears to be a suitable approach for analyzing and addressing such data.
+A common reference model for epidemic prediction is the SIR model [modello_SIR], an acronym referring to the state of the health of individuals.
+The three letters of the acronym stand for susceptible (S), infectious (I), and recovered (R).
+Namely, the SIR model divides the population into individuals who are susceptible to the disease, those who are infected and can spread it, and those who have recovered and are now immune.
+It uses a system of differential equations to describe how the disease spreads spatially and changes in time.
 In general, some papers use it as benchmark (e.g., Ref.
-[art_49]).
-All the selected papers are published in different journals, but primarily in the field of computer science.
+[art_49, art_413, art_415]).
+Other papers in this section focus on different types of time series, such as electroencephalogram (EEG) data for epilepsy detection (Ref.
+[art_14]) or emotion recognition (Refs.
+[art_411, art_412]).
+While emotion recognition may not seem directly related to the “Health” topic, it is included here because it relies on EEG data.
+The use of EEG in emotion recognition helps to identify more accurately true emotional states, as there physiological signals cannot be consciously controlled in the same way as facial expressions, tone of voice, or body posture [art_411].
+The selected papers were each published in different journals, with one paper per journal, primarily in the field of computer science.
 The majority of the corresponding authors are affiliated with academic institutions based in China.
@@
 Tab. LABEL:t:link_dataset_health provides a list of the public datasets utilized in the selected papers, along with links for accessing them.
-Table 13: List of public datasets in the ”Health” group and their corresponding links.
+Table 17: List of public datasets in the “Health” group and their corresponding links.
 Dataset Used by Link
+African COVID-19 dataset [art_410] https://www.worldometers.info/coronavirus/
 Bonn dataset [art_14] http://dx.doi.org/10.1103/PhysRevE.64.061907 (source paper)
 Brazilian COVID-19 data [art_52] https://github.com/lcaldeira/GrafoBrasilCovid
-COVID-19 data [art_49] https://www.2019ncov.chinacdc.cn/
-COVID-19 vaccinations [art_50] https://ourworldindata.org/us-states-vaccinations
+CDC US-State [art_414] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
+Chickenpox [art_410] https://www.kaggle.com/datasets/die9origephit/chickenpox-cases-hungary
+EEG DEAP dataset [art_412] http://www.eecs.qmul.ac.uk/mmv/datasets/deap/download.html
 eICU [art_189] https://doi.org/10.1038/sdata.2018.178 (source paper)
+European COVID-19 dataset [art_410] https://www.worldometers.info/coronavirus/
+ILI IDWR Japan-Prefecture [art_414] https://tinyurl.com/y5dt7stm
+ILI US-HHS US-Regional [art_414] https://tinyurl.com/y39tog3h
 MIMIC-IV [art_189] https://physionet.org/content/mimiciv/2.0/
-Nytimes COVID-19 data [art_50] https://github.com/nytimes/covid-19-data
+Nytimes COVID-19 data [art_50], [art_413] https://github.com/nytimes/covid-19-data
+OxCGRT dataset [art_416] https://www.bsg.ox.ac.uk/research/covid-19-government-response-tracker
 Parkinson’s Progression Markers Initiative (PPMI) [art_92] http://www.ppmi-info.org/
+Japan COVID-19 [art_415] https://github.com/swsoyee/2019-ncov-japan
 Parkinson Speech data set (PS) [art_92] https://doi.org/10.1109/jbhi.2013.2245674 (source paper)
+SJTU Emotion EEG Dataset (SEED) [art_411], [art_412] https://bcmi.sjtu.edu.cn/home/seed/
 Spanish COVID-19 data [art_51] https://cnecovid.isciii.es/
+Spanish COVID-19 data [art_414] https://dataforgood.fb.com/tools/disease-prevention-maps/
 Spikes and slow waves (SSW) dataset [art_14] http://dx.doi.org/10.1109/ICDM50108.2020.00067 (source paper)
+US COVID-19  [^2] This dataset includes multiple sources. [art_415] https://github.com/CSSEGISandData/COVID-19
+US COVID-19 [art_416] https://covidtracking.com/data/national
+US COVID-19 vaccinations [art_50] https://ourworldindata.org/us-states-vaccinations
 US influenza [art_48] https://github.com/aI-area/DVGSN/tree/main/dataset
-None of the papers in this section use the same dataset, so it is not possible to compare the results of the proposed models.
-Most of the datasets are related to the COVID-19 effects and have a daily granularity.
-In addition, some papers ([art_49], [art_52], [art_92], [art_189]) include exogenous variables that can be either context-related (e.g., geography, population, health conditions, vaccinations, economic factors), or specific to a single patient (including personal information such as age, gender, and medical conditions).
-5.4.3 Proposed models
-As for the taxonomy of the proposed models, 5 out of 8 papers use a convolutional GNN architecture.
-Two papers use a pure attentional model, while another paper presents a hybrid convolutional-attentional architecture.
+Only two pairs of papers in this thematic group use the same dataset, but none of the proposed models yield comparable results.
+Most of the datasets are related to COVID-19 and have a daily time granularity.
+In addition, some papers (Refs.
+[art_49, art_52, art_92, art_189, art_413]) include exogenous variables that can be either context-related (e.g., geography, population, health conditions, vaccinations, economic factors), or specific to a single patient (including personal information such as age, gender, and medical conditions).
+A paper that undertakes an in-depth study of exogenous variables is Ref.
+[art_416].
+The authors categorize these variables as holiday variables, time categorical variables (e.g., categorical week and month, and the number of days and weeks since the start of the dataset), and policy variables (e.g., school and workplace closing, movement restrictions, face coverings, and vaccination policies).
+They carry on a significance analysis of these variables, and explore the correlation among the significant ones.
+The COVID-19 policy dataset (OxCGRT) which includes three years of 23 policy indicators across more than 180 countries is available at https://www.bsg.ox.ac.uk/research/covid-19-government-response-tracker.
+5.4.3 Types of models
+As for the taxonomy of the proposed models, 11 out of 15 papers use a convolutional GNN architecture.
+Three papers use a pure attentional model, while another paper presents a hybrid convolutional-attentional architecture.
 Different approaches are used to define the graph structure.
 Some papers base their graphs on the geographical distribution of the data (as in Refs.
-[art_49], [art_52] and [art_92]), and others use models that learn the graph structure directly, as in Refs.
-[art_48] and [art_50].
+[art_49, art_52, art_92]), and others use models that learn the graph structure directly, as in Refs.
+[art_48, art_50, art_415, art_416].
 There is also a paper that uses a variant of the visibility algorithm [art_14].
-In almost all the cases, the graph is static.
-In terms of programming languages, this community uses only Python with PyTorch.
-As for loss functions, MSE is typically used for prediction tasks, and cross entropy is used for classification tasks.
+In half of the cases, the graph is static, while the dynamic graph approach appears more frequently in the papers from 2024.
+In terms of programming languages, the community of the “Health” thematic group uses only Python with PyTorch.
+As for the coice of the loss functions, MSE is typically used for prediction tasks, and cross entropy is used for classification tasks.
 Tab.
 LABEL:t:link_codici_health shows the links to the related source codes.
-Table 14: List of source codes of the ”Health” models in the review.
+Table 18: List of source codes of the “Health” models in the review.
 Model Link
+Backbone-based Dynamic Spatio-Temporal Graph Neural Network (BDSTGNN) [art_415] https://github.com/JkMaoya/BDSTGNN
 Dynamic virtual graph significance networks (DVGSN) [art_48] https://github.com/aI-area/DVGSN
 Deep learning of contagion dynamics on complex networks [art_51] https://doi.org/10.5281/zenodo.4974521
+Ordinary Differential Equation in Graph Neural Networks (GNNODENet) [art_413] https://github.com/xiongzhangxyq/GNNODENet
+Recurrent Graph Gate Fusion Transformers (ReGraFT) [art_416] https://github.com/mfriendly/ReGraFT
 Sparse spectra graph convolutional network (SSGCNet) [art_14] https://github.com/anonymous2020-source-code/WNFG-SSGCNet-ADMM
@@
 5.4.4 Benchmark models
-Tab. LABEL:t:benchmark_models_health lists the benchmark models in the ”Health” group that are discussed in at least two papers.
-Table 15: List of benchmark models in the ”Health” group divided per category.
+Tab. LABEL:t:benchmark_models_health lists the benchmark models in the “Health” group that are discussed in at least two papers.
+Table 19: List of benchmark models in the “Health” group divided per category.
 Category Model Used by
-Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_48], [art_49], [art_52]
-Classical machine learning methods Feed-forward neural network (FNN) [art_14], [art_48]
-Long-short term memory network (LSTM) [art_49], [art_50]
-There are no particularly relevant benchmark models, and none fall within the category of GNNs. This is because GNNs have not been widely studied in the context of health.
+Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_48], [art_49], [art_52], [art_410], [art_414], [art_415]
+SIR model [modello_SIR] [art_49], [art_413], [art_415]
+Non-GNN machine learning methods Feed-forward neural network (FNN) [art_14], [art_48]
+Bi-Hemisphere Domain Adversarial Neural Network (Bi-DANN) [Bi-DANN] [art_411], [art_412]
+A Novel Bi-Hemispheric Discrepancy Model for EEG Emotion Recognition (BiHDM) [BiHDM] [art_411], [art_412]
+Gated recurrent unit (GRU) [art_410], [art_413], [art_415], [art_416]
+Long-short term memory network (LSTM) [art_49], [art_50], [art_410], [art_413]
+Recurrent neural network (RNN) [art_189], [art_413], [art_414]
+Support vector machine (SVM) [art_189], [art_411]
+GNN methods Cross-location attention based graph neural network (ColaGNN) [ColaGNN] [art_414], [art_415], [art_416]
+Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_49], [art_413], [art_414], [art_416]
+Dynamical graph convolutional neural networks (DGCNN) [DGCNN] [art_411], [art_412]
+Graph attention network (GAT) [GAT] [art_48], [art_413]
+Graph convolutional network (GCN) [art_49], [art_413]
+Regularized Graph Neural Networks (RGNN) [RGNN] [art_411], [art_412]
+Spatio-temporal attention network for pandemic prediction (STAN) [STAN] [art_413], [art_415]
+Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_413], [art_415]
+There are no particularly outstanding benchmark models, and the ones categorized as GNNs appear only in papers from 2024.
+This is due to the fact that in the past years GNNs have not been extensively explored in the context of health.
+The most widely used GNN-based benchmark is the diffusion convolutional recurrent neural network (DCRNN), proposed by Li et al.
+in 2017 for traffic forecasting.
+The model aims to capture spatial dependencies using bidirectional random walks on the graph (whose adjacency matrix is based on node proximity), and temporal dependencies using an encoder-decoder architecture with scheduled sampling.
+The source code for the DCRNN model is available at https://github.com/liyaguang/DCRNN.
 5.4.5 Results
@@
 The most common error metrics are the MAE and the MSE for forecasting problems and accuracy, sensitivity and specificity for classification problems.
+The forecasting horizon spans from 1 day to more than 10 weeks ahead.
 A comparison of the accuracy of the different models is not possible due to the lack of common benchmark datasets.
 5.5 Mobility
-The topic that includes the largest number of papers identified by the query (39 out of 156) is ”Mobility”, which regards the movement of people.
+The thematic group that includes the largest number of papers identified by our query (66 out of 281) is “Mobility”, which regards the movement of people.
 It includes urban traffic, air travel, and bicycle demand, among other applications.
@@
 5.5.1 Overview
-Early traffic prediction is essential for improving the efficiency of transportation systems, helping drivers plan their trips more effectively, and preventing urban congestion.
-The advent of smart cities infrastructures and transportation systems has contributed to the collection of rich data from road sensors that can be used for this purpose.
-However, making accurate traffic forecasts is challenging due to constantly changing traffic patterns over time and space, as well as the influence of external factors such as weather conditions and special events.
-Unlike classical statistical and machine learning models, GNNs can be particularly effective for traffic forecasting due to their ability to model complex relationships in spatial and temporal domain.
-The increasing number of papers published each year in this group reflects both the growing research interest in this field and the growing confidence in GNN models.
-The most commonly investigated areas regard urban traffic ([art_93], [art_95], [art_97], [art_98], [art_99], [art_100], [art_103], [art_105], [art_106], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_122], [art_123], [art_124], [art_125], [art_126], [3], [art_128], [art_130], [art_131], [art_132], [art_135]), especially urban traffic flow and urban traffic speed.
-Among the journals where the selected papers were published, Applied Intelligence by Springer is the most common, with 5 out of 39 papers.
-This is followed by ACM Transactions on Intelligent Systems and Technology and ACM Transactions on Knowledge Discovery from Data, both published by the Association for Computing Machinery, each with 4 out of 39 papers.
+Early traffic prediction is essential for improving the efficiency of transportation systems, helping drivers plan their trips more effectively, and, in the context of cities, preventing urban congestion.
+The advent of smart cities infrastructures and transportation systems has contributed to the collection of rich data from road sensors.
+Making accurate traffic forecasts is challenging, due to constantly changing traffic patterns over time and space, as well as the influence of external factors such as weather conditions and special events.
+Unlike statistical and non-GNN machine learning models, GNNs can be particularly effective for traffic forecasting due to their ability to model complex relationships in spatial and temporal domain.
+The increasing number of papers published each year in this thematic group reflects both the growing research interest in this field and the growing confidence in GNN models.
+The most commonly investigated areas regard urban traffic (approximately 70% of the papers), especially urban traffic flow and urban traffic speed.
+Among the journals where the selected papers were published the most common are Applied Intelligence by Springer, ACM Transactions on Knowledge Discovery from Data, both by the Association for Computing Machinery, and IEEE Transactions on Intelligent Transportation Systems by the Institute of Electrical and Electronics Engineers, with 6 out of 66 papers each.
 Most of the corresponding authors are affiliated with Chinese institutions.
@@
 Tab. LABEL:t:link_dataset_traffico provides a list of the public datasets mentioned in the selected papers, along with the links to access them.
-Table 16: List of public datasets in the ”Mobility” group and their corresponding links.
+Table 20: List of public datasets in the “Mobility” group and their corresponding links.
 Dataset Used by Link
+Bureau of Transportation Statistics (BTS) [art_434], [art_443], [art_448] https://www.transtats.bts.gov/
 CAAC flights [art_129] http://www.caac.gov.cn/index.html
+Capital Bike Share [art_446] https://ride.capitalbikeshare.com/system-data
 CD-HW [art_126] https://doi.org/10.1109/ACCESS.2020.3027375 (source paper)
-City Bike NYC [art_94], [art_114] https://citibikenyc.com/system-data
-DiDi Beijing [art_102] https://github.com/didi/TrafficIndex
+DiDi Beijing [art_102], [art_447] https://github.com/didi/TrafficIndex
 DiDi Chengdu [art_93], [art_98] https://gaia.didichuxing.com/
-Huai’an Bus Transaction [art_101] https://www.pandabus.cn/panda/panda-minibus.html
+Divvy Bike Chicago [art_446] https://ride.divvybikes.com/system-data
+Huai’an Bus Transaction [art_452] https://www.pandabus.cn/panda/panda-minibus.html
 Los-loop [art_103] https://github.com/lehaifeng/T-GCN/tree/master/data
-METR-LA [art_93], [art_99], [art_100], [art_109], [art_110], [art_113], [art_116], [art_120], [art_122], [art_126], [art_128] https://paperswithcode.com/dataset/metr-la
+Lyft bike dataset [art_446] https://www.lyft.com/bikes/bay-wheels/system-data
+METR-LA [art_93], [art_99], [art_100], [art_109], [art_110], [art_113], [art_116], [art_120], [art_122], [art_126], [art_128], [art_408], [art_427], [art_430], [art_433], [art_435], [art_438], [art_441], [art_442], [art_444] https://paperswithcode.com/dataset/metr-la
 Milan network [art_123] https://github.com/arunasubbiah/milan-telecom-data-modeling
 NE-BJ [art_100] https://github.com/tsinghua-fib-lab/Traffic-Benchmark
-NYC-Taxi [art_96] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
-PeMS03 [art_108], [art_113], [art_115], [art_117], [art_125], [art_130] https://www.kaggle.com/datasets/elmahy/pems-dataset
-PeMS04 (PeMSD4) [art_95], [art_97], [art_99], [art_104], [art_108], [art_113], [art_115], [art_117], [art_122], [art_125], [3], [art_130], [art_131], [art_132] https://paperswithcode.com/dataset/pems04
-PeMS07 (PeMSD7) [art_108], [art_110], [art_113], [art_117], [art_125], [3], [art_130] https://paperswithcode.com/dataset/pems07
-PeMS08 (PeMSD8) [art_95], [art_97], [art_99], [art_108], [art_115], [art_117], [art_122], [art_125], [art_130], [art_131], [art_132] https://paperswithcode.com/dataset/pems08
-PEMS-BAY [art_100], [art_109], [art_110], [art_116], [art_120], [art_122] https://paperswithcode.com/dataset/pems-bay
+NYC-Bike [art_94], [art_96], [art_114], [art_446], [art_449], [art_451] https://citibikenyc.com/system-data
+NYC SafeGraph [art_449] https://www.safegraph.com/academics
+NYC-Taxi [art_96], [art_449] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
+PeMS03 [art_108], [art_113], [art_115], [art_117], [art_125], [art_130], [art_419], [art_435], [art_445], [art_453] https://www.kaggle.com/datasets/elmahy/pems-dataset
+PeMS04 (PeMSD4) [art_97], [art_99], [art_104], [art_108], [art_113], [art_115], [art_117], [art_122], [art_125], [3], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_424], [art_433], [art_435], [art_441], [art_445], [art_448], [art_449], [art_452], [art_453] https://paperswithcode.com/dataset/pems04
+PeMS07 (PeMSD7) [art_108], [art_110], [art_113], [art_117], [art_125], [3], [art_130], [art_406], [art_422], [art_424], [art_427], [art_432], [art_435], [art_438], [art_441], [art_444], [art_445], [art_453] https://paperswithcode.com/dataset/pems07
+PeMS08 (PeMSD8) [art_97], [art_99], [art_108], [art_115], [art_117], [art_122], [art_125], [art_130], [art_131], [art_132], [art_406], [art_408], [art_424], [art_432], [art_433], [art_435], [art_441], [art_444], [art_445], [art_448], [art_449], [art_452], [art_453] https://paperswithcode.com/dataset/pems08
+PEMS-BAY [art_100], [art_109], [art_110], [art_116], [art_120], [art_122], [art_406], [art_408], [art_419], [art_420], [art_422], [art_430], [art_433], [art_435], [art_441], [art_442], [art_444] https://paperswithcode.com/dataset/pems-bay
 Q-Traffic [art_124] https://github.com/JingqingZ/BaiduTraffic#Dataset
-RCOTP [art_95] https://developer.ibm.com/exchanges/data/all/airline/
+Seattle loop [art_441] https://github.com/zhiyongc/Seattle-Loop-Data
 UVDS [art_128] https://doi.org/10.1007/978-3-030-73280-6_6 (source paper)
-TaxiBJ [art_126] https://www.microsoft.com/en-us/research/publication/forecasting-citywide-crowd-flows-based-big-data/
+TaxiBJ [art_126], [art_451] https://www.microsoft.com/en-us/research/publication/forecasting-citywide-crowd-flows-based-big-data/
 Taxi Shenzhen [art_103], [art_135] https://github.com/lehaifeng/T-GCN/tree/master/data
 TLC Taxi [art_119] https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page
-The most popular datasets are PeMS ones (http://pems.dot.ca.gov/), provided by the California Transportation Agency Performance Measurement System, which include various versions differing in time range, spatial coverage, and the number of sensors, and METR-LA, collected from loop detectors in Los Angeles County highway.
-The number of sensors is in the hundreds, and so is the number of nodes in the graph.
+Urban-core [art_441] https://topis.seoul.go.kr/
+Xiecheng delay dataset [art_443] https://pan.baidu.com/s/1dEPyMGh#list/path\$=\$%2F
+The most popular datasets used in this group are those provided by the California Transportation Agency Performance Measurement System (PeMS, available at http://pems.dot.ca.gov/), which include various versions differing in time range, spatial coverage, and number of sensors, and the one called METR-LA, collected from loop detectors along the Los Angeles County highway.
+The number of sensors is in the hundreds, and such is also the number of nodes in the graph.
 For a more detailed overview of the benchmark datasets in the traffic domain, see Ref.
@@
 Other common public traffic datasets can be found at https://paperswithcode.com/task/traffic-prediction.
-In terms of external data, the most commonly used are weather conditions, time information (time of day and day of week) and events (e.g., holidays), as well as information about the road network (e.g., points of interest).
+As to external data, the most commonly used are data on weather conditions, time information (time of day and day of week) and events (e.g., holidays), as well as information about the road network, like points of interest.
 The majority of the papers utilize data with a 5-minute granularity, and provide forecasts for horizons ranging from 5 minutes to 1 hour ahead.
-In addition to imputation of missing values, in almost half of the cases the pre-processing of the data includes data normalization (min-max or Z-score normalization) and the selection of time windows regarding recent, daily and weekly data (as in 8 out of 39 papers).
-5.5.3 Proposed models
-As for the proposed models, 18 out of 39 belong to the pure convolutional GNN paradigm, followed by 11 attentional GNNs, 9 hybrid convolutional-attention architectures, and 1 recurrent GNN.
+In addition to imputation of missing values, in almost half of the cases the pre-processing of the data includes data normalization (min-max or Z-score normalization) and the selection of time windows regarding recent, daily and weekly data (as in 10 out of 66 papers).
+5.5.3 Types of models
+As for the proposed models, 35 out of 66 belong to the pure convolutional GNN paradigm, 12 to hybrid convolutional-attention architectures, followed by 14 attentional GNNs, and 1 recurrent GNN.
+Two papers use both convolutional and attentional models, and other two (Refs.
+[art_438, art_451]) do not fall into a specific class.
 Regarding the definition of the graph structure, most of the papers discuss it in detail.
-Half of the papers are based on a pre-defined adjacency matrix, and among the remaining papers that propose models that learn the adjacency matrix on their own, a further half is based on a pre-initialization of the matrix, as for example in Ref.
-[art_108].
-Here, more than in any other group, there are multiple ways to define the adjacency matrix, taking into account either the structure of the road network, the actual flow of people, or a combination of different information.
-More in detail, some models use the road connectivity ([art_104], [art_120], [art_123], [art_135]), others their spatial distance ([art_93], [art_96], [art_99], [art_110], [art_128], [art_132]), and some others incorporate extra information about the so-called points of interest (such as Refs.
-[art_94], [art_96]).
-Other models instead consider some similarity measures between the time series ([art_97], [art_117], [art_122], [art_123]), such as the correlation coefficient or cosine similarity.
-Still others directly use the number of people traveling on the road at a given time step ([art_94], [art_102], [art_129], [art_133]).
-The two most common loss functions used to train the models are the MSE and the MAE.
+Half of the papers published before 2024 are based on a pre-defined adjacency matrix, while the majority of papers of 2024 propose models that learn the adjacency matrix on their own, either from scratch or based on a pre-initialization (e.g., Refs.
+[art_108, art_444]).
+Here, more than in any other thematic group, there are multiple ways to define the adjacency matrix.
+This can be made by taking into account the structure of the road network, the actual flow of people, or a combination of different sources of information.
+More in detail, a first group of papers uses road connectivity (e.g., Refs.
+[art_104, art_120, art_123, art_135, art_431]); second group uses spatial distance (e.g., Refs.
+[art_93, art_96, art_99, art_110, art_128, art_132, art_434, art_454]); a third group incorporates extra information about the so-called points of interest (such as Refs.
+[art_94, art_96]); a fourth group considers some similarity measures between the time series (e.g., Refs.
+[art_97, art_117, art_122, art_123, art_435, art_445]), such as the correlation coefficient or cosine similarity; a fifth group directly uses the number of people traveling on the road at a given time step (e.g., Refs.
+[art_94, art_102, art_129, art_133]), or a combination of different aspects (e.g., Refs.
+[art_420, art_444, art_448, art_449]).
+The two most common loss functions used to train the models are the MAE and the MSE.
 Few papers consider a combination of loss functions, especially when dealing with a model that has to learn the graph structure by itself (Refs.
-[art_95], [art_126], [art_130]).
-Not all papers specify the language or library used for the code.
-Among those that do, Python is the most common language, with the majority using PyTorch, while only a few researchers use TensorFlow.
+[art_126, art_130, art_446, art_448]).
+Not all papers specify the language or library used for writing the code.
+Among those that do that, Python is the most common language, with the majority using PyTorch, whereas only a few researchers use TensorFlow.
 In addition, a few papers provide a link to the source code for the proposed model.
@@
 LABEL:t:link_codici_traffico.
-Table 17: List of source codes of the ”Mobility” models in the review.
+Table 21: List of source codes of the “Mobility” models in the review.
 Model Link
@@
 Adaptive generalized PageRank graph neural network (AGP-GNN) [art_116] https://github.com/guoxiaoyuatbjtu/agp-gnn
-Automated dilated spatio-temporal synchronous graph network (Auto-DSTSGN) [art_108] https://github.com/jinguangyi-uto-DSTSGN
+Adaptive multigraph convolutional networks (AMGCN) [art_444] https://github.com/hfimmortal/AMGCN
+Automated dilated spatio-temporal synchronous graph network (Auto-DSTSGN) [art_108] https://github.com/jinguangyin/Auto-DSTSGN
+BigST [art_447] https://github.com/usail-hkust/BigST
 Dynamic multi-view graph neural network for citywide traffic inference (CTVI+) [art_97] https://github.com/dsj96/TKDD
 Dynamic Graph Convolutional Recurrent Network (DGCRN) [art_100] https://github.com/tsinghua-fib-lab/Traffic-Benchmark
+Progressive Graph Convolutional Network (PGCN) [art_441] https://github.com/yuyolshin/PGCN
+Scale-aware neural architecture search framework for MTS forecasting (SNAS4MTF) [art_408] https://github.com/shangzongjiang/SNAS4MTF
 Spatio-temporal Causal Graph Attention Network (STCGAT) [art_125] https://github.com/zsqZZU/STCGAT/tree/v1.0.0
@@
 Spatio-Temporal Graph Mixformer (STGM) [art_110] https://github.com/mouradost/stgm
+Space-Time Separable Graph Convolutional Network (STSGCN) [art_443] https://github.com/FraLuca/STSGCN
+Two-level Resolution Neural Network (TwoResNet) [art_430] https://github.com/semink/TwoResNet
+Wavelet dynamic spatiotemporal aware Graph Neural Network (W-DSTAGNN) [art_419] https://github.com/yash-jakhmola/w-dstagnn
 5.5.4 Benchmark models
 Tab.
-LABEL:t:benchmark_models_traffico lists the benchmark models used in the papers in the ”Mobility” group.
-The benchmarks are categorized into mathematical and statistical methods, classical machine learning models, and GNN models, and only those appearing in more than two papers are included.
-Table 18: List of benchmark models in the ”Mobility” group divided per category.
+LABEL:t:benchmark_models_traffico lists the benchmark models used within the papers in the “Mobility” thematic group.
+These benchmarks are categorized into mathematical and statistical methods, non-GNN machine learning models, and GNN models, and only those appearing in more than two papers are included in the table.
+Table 22: List of benchmark models in the “Mobility” group divided per category.
 Category Model Used by
-Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_93], [art_94], [art_95], [art_96], [art_98], [art_100], [art_101], [art_103], [art_106], [art_107], [art_109], [art_116], [art_123], [art_126], [art_129], [art_132], [art_135]
-Historical average (HA) [art_93], [art_94], [art_96], [art_98], [art_100], [art_101], [art_102], [art_103], [art_109], [art_110], [art_120], [art_123], [art_125], [art_131], [art_132], [art_135]
-Vector autoregression (VAR) [art_99], [art_100], [art_126], [art_131], [art_132]
-Classical machine learning methods Feed-forward neural network (FNN) [art_93], [art_97], [art_100], [art_105], [art_114], [art_120], [art_121], [art_124], [art_126]
-Gated recurrent unit (GRU) [art_98], [art_99], [art_101], [art_106], [art_121], [art_123], [art_124], [3], [art_131], [art_132], [art_135]
-Long-short term memory network (LSTM) [art_93], [art_96], [art_98], [art_99], [art_100], [art_106], [art_107], [art_108], [art_109], [art_113], [art_115], [art_116], [art_120], [art_121], [art_122], [art_125], [art_126], [3], [art_130], [art_131], [art_132]
+Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_93], [art_94], [art_96], [art_98], [art_100], [art_103], [art_106], [art_109], [art_116], [art_123], [art_126], [art_129], [art_132], [art_135], [art_406], [art_408], [art_419], [art_421], [art_424], [art_430], [art_433], [art_438], [art_444], [art_446], [art_448], [art_449], [art_451], [art_452], [art_453]
+Historical average (HA) [art_93], [art_94], [art_96], [art_98], [art_100], [art_102], [art_103], [art_109], [art_110], [art_120], [art_123], [art_125], [art_131], [art_132], [art_135], [art_406], [art_408], [art_420], [art_424], [art_433], [art_435], [art_438], [art_441], [art_443], [art_446], [art_447], [art_449], [art_451], [art_452], [art_453]
+Vector autoregression (VAR) [art_99], [art_100], [art_126], [art_131], [art_132], [art_406], [art_419], [art_424], [art_435], [art_443], [art_445], [art_447], [art_453]
+Non-GNN machine learning methods Convolutional long-short term memory network (ConvLSTM) [ConvLSTM] [art_98], [art_446]
+Deep multi-view spatial-temporal network (DMVST-Net) [DMVST-Net] [art_96], [art_451]
+Feed-forward neural network (FNN) [art_93], [art_97], [art_100], [art_105], [art_114], [art_120], [art_121], [art_124], [art_126], [art_408], [art_420], [art_431], [art_438]
+Gated recurrent unit (GRU) [art_98], [art_99], [art_106], [art_121], [art_123], [art_124], [3], [art_131], [art_132], [art_135], [art_425], [art_452]
+Long-short term memory network (LSTM) [art_93], [art_96], [art_98], [art_99], [art_100], [art_106], [art_108], [art_109], [art_113], [art_115], [art_116], [art_120], [art_121], [art_122], [art_125], [art_126], [3], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_420], [art_421], [art_424], [art_425], [art_430], [art_433], [art_435], [art_438], [art_441], [art_442], [art_443], [art_444], [art_445], [art_449], [art_453]
 Long-short term time series network (LSTNet) [LSTNet] [art_102], [art_110]
-Seq2Seq [Seq2seq] [art_105], [art_124]
-Spatial and temporal normalization for multivariate time series forecasting (ST-Norm) [ST-Norm] [art_116], [art_117]
-Support vector regression (SVR) [art_99], [art_100], [art_103], [art_107], [art_109], [art_119], [art_123], [art_129], [art_130], [art_135]
+Seq2Seq [Seq2seq] [art_105], [art_124], [art_451]
+Spatial and temporal normalization for multivariate time series forecasting (ST-Norm) [ST-Norm] [art_116], [art_117], [art_406], [art_422]
+Spatial-temporal identity (STID) [STID] [art_116], [art_422], [art_433], [art_434]
+Support vector regression (SVR) [art_99], [art_100], [art_103], [art_109], [art_119], [art_123], [art_129], [art_130], [art_135], [art_408], [art_419], [art_421], [art_435], [art_444], [art_445], [art_453]
+Temporal convolutional network (TCN) [TCN] [art_446], [art_453]
+Temporal 2D-variation modeling for general time series analysis (TimesNet) [TimesNet] [art_434], [art_435]
+Transformer [attention] [art_424], [art_454]
 Extreme gradient boosting (XGBoost) [XGBoost] [art_97], [art_114]
-GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_95], [art_98], [art_100], [art_101], [art_110], [art_117], [art_124], [art_125], [art_130], [art_131]
-Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_93], [art_95], [art_98], [art_99], [art_100], [art_107], [art_108], [art_110], [art_113], [art_115], [art_117], [art_121], [art_125], [art_126], [3], [art_130], [art_131], [art_132]
+GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_98], [art_100], [art_110], [art_117], [art_124], [art_125], [art_130], [art_131], [art_406], [art_419], [art_435], [art_441], [art_447], [art_448], [art_452], [art_453], [art_454]
+Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_93], [art_98], [art_99], [art_100], [art_108], [art_110], [art_113], [art_115], [art_117], [art_121], [art_125], [art_126], [3], [art_130], [art_131], [art_132], [art_406], [art_421], [art_424], [art_427], [art_432], [art_433], [art_435], [art_442], [art_445], [art_446], [art_447], [art_448], [art_449], [art_453], [art_454]
+Attention based Spatial-Temporal Graph Neural Network (ASTGNN) [ASTGNN] [art_422], [art_435], [art_446]
+Automated spatio-temporal graph (AutoSTG) [AutoSTG] [art_108], [art_408]
 Dual-stage attention-based recurrent neural network (DA-RNN) [DA-RNN] [art_96], [art_105]
-Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_93], [art_94], [art_95], [art_96], [art_100], [art_101], [art_103], [art_107], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_122], [art_124], [art_125], [art_126], [art_128], [art_129], [art_130], [art_131], [art_132]
-Dynamic spatial-temporal aware graph neural network (DSTAGNN) [DSTAGNN] [art_95], [art_125]
+Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_93], [art_94], [art_96], [art_100], [art_103], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_122], [art_124], [art_125], [art_126], [art_128], [art_129], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_420], [art_421], [art_422], [art_424], [art_425], [art_427], [art_430], [art_432], [art_433], [art_435], [art_438], [art_441], [art_442], [art_444], [art_445], [art_447], [art_448], [art_451], [art_452], [art_453]
+Dynamic graph convolutional recurrent network (DGCRN) [art_100] [art_110], [art_406], [art_444]
+Dynamic and Multi-faceted SpatioTemporal Graph Convolution Network (DMSTGCN) [DMSTGCN] [art_441], [art_444]
+Dynamic spatial-temporal aware graph neural network (DSTAGNN) [DSTAGNN] [art_125], [art_419], [art_424], [art_447], [art_448]
+Fully connected gated graph architecture (FC-GAGA) [FC-GAGA] [art_116], [art_408], [art_427]
 GeoMAN [GeoMAN] [art_96], [art_132]
-Graph convolutional network (GCN) [art_104], [art_135]
-Graph multi-attention network (GMAN) [GMAN] [art_100], [art_109], [art_110], [art_116], [art_122], [art_128]
-Graph WaveNet (GWN) [GWN] [art_93], [art_95], [art_100], [art_101], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_122], [art_124], [art_128]
-Multi-range attentive bicomponent graph convolutional network (MRA-BGCN) [MRA-BGCN] [art_109], [art_116]
-Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_100], [art_109], [art_110], [art_113], [art_116], [art_117], [art_122], [art_124], [art_128]
+Graph convolutional network (GCN) [art_104], [art_135], [art_431]
+Graph multi-attention network (GMAN) [GMAN] [art_100], [art_109], [art_110], [art_116], [art_122], [art_128], [art_408], [art_419], [art_422], [art_427], [art_430], [art_444], [art_451]
+Graph WaveNet (GWN) [GWN] [art_93], [art_100], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_122], [art_124], [art_128], [art_406], [art_408], [art_419], [art_420], [art_422], [art_424], [art_430], [art_432], [art_433], [art_434], [art_435], [art_441], [art_444], [art_447], [art_448], [art_449], [art_452], [art_453], [art_454]
+Discrete graph structure learning for time series (GTS) [GTS] [art_420], [art_433]
+Long short-term graph convolutional networks (LSGCN) [LSGCN] [art_99], [art_432], [art_446], [art_453]
+Multi-range attentive bicomponent graph convolutional network (MRA-BGCN) [MRA-BGCN] [art_109], [art_116], [art_408]
+Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_100], [art_109], [art_110], [art_113], [art_116], [art_117], [art_122], [art_124], [art_128], [art_406], [art_408], [art_420], [art_422], [art_444], [art_445]
 Multi-view graph convolutional network (MVGCN) [MVGCN] [art_94], [art_96]
-Spatial-temporal fusion graph neural network (STFGNN) [STFGNN] [art_98], [art_101], [art_108], [art_115], [art_117], [art_125], [art_130]
-Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_93], [art_94], [art_95], [art_96], [art_99], [art_100], [art_101], [art_106], [art_107], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_121], [art_122], [art_124], [art_126], [3], [art_128], [art_129], [art_130], [art_131], [art_132]
+Spectral temporal graph neural network for multivariate time-series forecasting (StemGNN) [StemGNN] [art_406], [art_422]
+Spatial-temporal fusion graph neural network (STFGNN) [STFGNN] [art_98], [art_108], [art_115], [art_117], [art_125], [art_130], [art_406], [art_435], [art_445], [art_452], [art_453]
+Spatial-temporal graph to sequence model (STG2Seq) [STG2Seq] [art_99], [art_424]
+Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_93], [art_94], [art_96], [art_99], [art_100], [art_106], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_121], [art_122], [art_124], [art_126], [3], [art_128], [art_129], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_420], [art_421], [art_422], [art_424], [art_425], [art_433], [art_435], [art_438], [art_442], [art_443], [art_444], [art_445], [art_448], [art_449], [art_451], [art_452], [art_453], [art_454]
 Spatial temporal graph neural network (STGNN) [STGNN] [art_94], [art_103]
-Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_99], [art_108], [art_117], [art_125]
-Deep-meta-learning based model (ST-MetaNet) [ST-MetaNet] [art_96], [art_100], [art_109], [art_116]
-Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_98], [art_99], [art_100], [art_108], [art_115], [art_117], [art_125], [art_130], [art_131]
-Temporal graph convolutional network (T-GCN) [T-GCN] [art_103], [art_107], [art_113], [art_121], [art_123]
-As for the basic mathematical and statistical methods, two widely used benchmarks are the ARIMA model and historical average model (HA), which uses the average of past observations as the forecast for future values.
-Despite their apparent simplicity, these models are sometimes shown to be very powerful benchmarks, with an accuracy comparable to that of more complex models.
-As for the traditional machine learning models, LSTMs are widely recognized as a type of recurrent neural network capable of handling and long-term dependencies in traffic data.
-As for GNN-based models instead, the three most used models are, in order, the spatio-temporal graph convolutional network (STGCN) [STGCN], the diffusion convolutional recurrent neural network (DCRNN) [DCRNN], and the attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN], all developed specifically for the traffic forecasting problem.
-The STGCN model was proposed by Yu et al.
-in 2018, and it integrates graph convolution and gated temporal convolution through spatio-temporal convolutional blocks.
-The source code is available at https://github.com/VeritasYin/STGCN_IJCAI-18.
-The DCRNN model was proposed by Li et al.
-in 2017, and it aims to capture the spatial dependency using bidirectional random walks on the graph, and the temporal dependency using the encoder-decoder architecture with scheduled sampling.
+Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_99], [art_108], [art_117], [art_125], [art_406], [art_424], [art_447], [art_453]
+Spatio-temporal graph attention network (ST-GRAT) [ST-GRAT] [art_116], [art_444]
+Deep-meta-learning based model (ST-MetaNet) [ST-MetaNet] [art_96], [art_100], [art_109], [art_116], [art_408], [art_451]
+Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_98], [art_99], [art_100], [art_108], [art_115], [art_117], [art_125], [art_130], [art_131], [art_406], [art_419], [art_424], [art_435], [art_441], [art_442], [art_443], [art_453], [art_454]
+Temporal graph convolutional network (T-GCN) [T-GCN] [art_103], [art_113], [art_121], [art_123], [art_421], [art_433]
+Time zigzags at graph convolutional networks for time series forecasting (Z-GCNETs) [Z-GCNETs] [art_125], [art_424], [art_441]
+As for the basic mathematical and statistical methods, two widely used benchmarks are the ARIMA model and the historical average model (HA), which uses the average of past observations as the forecast for future values.
+Despite their apparent simplicity, these benchmarks are sometimes shown to be very powerful, with an accuracy comparable to that of more complex models.
+As for the non-GNN machine learning models, LSTM networks are widely recognized as a type of recurrent neural network capable of handling long-term dependencies in traffic data.
+As for GNN-based models instead, the three most used models are the spatio-temporal graph convolutional network (STGCN) [STGCN], the diffusion convolutional recurrent neural network (DCRNN) [DCRNN], and the attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN], all developed specifically for the traffic forecasting problem.
+The STGCN model proposed by Yu et al.
+in 2018 consists of two spatio-temporal convolutional blocks (with two temporal gated convolutional layers, one spatial graph convolutional layer in the middle, and residual connections) and a fully-connected output layer in the end.
+In its original implementation, it processes graphs a with weighted adjacency matrix based on spatial distance between sensors in the road network.
+The source code for the STGCN model is available at https://github.com/VeritasYin/STGCN_IJCAI-18.
+The DCRNN model proposed by Li et al.
+in 2017 aims to capture the spatial dependency by using bidirectional random walks on the graph (whose adjacency matrix is based on node proximity), and to capture the temporal dependencies by using an encoder-decoder architecture with scheduled sampling.
 The code can be accessed at https://github.com/liyaguang/DCRNN.
-Finally, the ASTGCN model was proposed by Guo et al.
-in 2019, and it models in parallel recent, daily-periodic and weekly-periodic dependencies as three independent components with a spatio-temporal attention mechanism to capture the dynamic spatio-temporal correlations, and a spatio-temporal convolution which simultaneously employs graph convolutions to capture the spatial patterns.
-The code is available at https://github.com/Davidham3/ASTGCN-2019-mxnet.
+The ASTGCN model proposed by Guo et al.
+in 2019 models at once recent, daily-periodic and weekly-periodic dependencies as three independent components, with a spatio-temporal attention mechanism for capturing the dynamic spatio-temporal correlations and a spatio-temporal convolution which simultaneously employs graph convolutions to capture the spatial patterns.
+The source code for the ASTGCN model is available at https://github.com/Davidham3/ASTGCN-2019-mxnet.
 For a more comprehensive list of source codes for the most dated benchmark GNN models in the traffic domain see Ref.
@@
 5.5.5 Results
-In all papers, the authors claim that the proposed GNN-based models are able to outperform the other benchmarks.
-However, there are a limited number of situations in which some simple statistical and mathematical models demonstrate comparable accuracy, as it can be observed in Tabs.
-LABEL:t:risultati_metr-la and LABEL:t:risultati_pems-bay.
-For the and comparison and evaluation of the models, the most commonly used error metrics are MAE, RMSE, and MAPE.
-Few papers also use different metrics, such as the coefficient of determination $R^{2}$.
-However, despite the use of the same metrics and the availability of widely used datasets, comparing models is not trivial because different papers often focus on different time windows of a dataset.
+In all papers mentioned, the authors claim that the proposed GNN-based models are able to outperform the benchmarks.
+However, there are a limited but interesting number of situations in which some simple statistical and mathematical models demonstrate comparable accuracy, as it can be observed in Tabs.
+LABEL:t:risultati_metr-la and LABEL:t:risultati_pems-bay [art_452].
+For the comparison and evaluation of these models, the most commonly used error metrics are MAE, RMSE, and MAPE.
+A few papers in addition use further metrics such as the coefficient of determination $R^{2}$.
+However, despite the use of the same metrics and the availability of widely used datasets, comparison among these models is not trivial, because different papers often focus on different time windows of a dataset.
 […]
 Table 23: Comparison of the average accuracy of the different models on the METR-LA dataset for 15, 30 and 60 minute time horizons, expressed in terms of MAE, MAPE and RMSE.
-The results of the models marked with are taken from papers in the ”Generic” group.
+The results of the models marked with are taken from papers in the “Generic” group.
 The smallest errors are underlined.
-Numbers from the papers.
+Numbers from the original papers.
 Time horizon 15 min 30 min 60 min
@@
 AGP-GNN [art_116] 2.71 7.21 5.25 3.04 8.15 6.20 3.41 9.70 7.20
+AMGCN [art_444] 2.52 6.13 4.48 2.97 7.70 5.35 3.38 8.93 6.21
 ARIMA 3.99 9.60 8.12 5.15 12.70 10.45 6.90 17.40 13.23
 ASTGCN 4.86 9.21 9.27 5.43 10.13 10.61 6.51 11.64 12.52
+ASTGNN 2.66 6.96 5.09 3.06 8.47 6.26 3.52 9.56 7.21
+AutoSTG 2.70 - 5.16 3.06 - 6.17 3.47 - 7.27
+CASTGNN [art_442] 2.88 6.67 6.44 3.25 7.91 7.56 3.70 9.76 8.92
 DCRNN 2.77 7.30 5.38 3.15 8.80 6.45 3.60 10.50 7.60
 DGCRN [art_100] 2.62 6.63 5.01 2.99 8.02 6.05 3.44 9.73 7.19
+DMSTGCN 2.85 7.54 5.54 3.26 9.19 6.56 3.72 10.96 7.55
 DST-GCNN [art_126] 2.68 7.20 5.35 3.01 8.50 6.23 3.41 10.30 7.47
+DualGraph [art_438] 2.62 7.10 5.36 2.83 7.90 5.89 3.15 9.10 6.67
 DyGCN-LSTM [art_113] 2.64 6.35 4.64 3.24 8.04 5.24 3.53 9.30 6.53
+DyHSL 2.58 6.47 4.98 2.95 8.05 6.13 3.41 9.41 7.03
 FC-GAGA 2.75 7.25 5.34 3.10 8.57 6.30 3.51 10.14 7.31
@@
 GGRU 2.71 6.99 5.24 3.12 8.56 6.36 3.64 10.62 7.65
+GL-STGTN [art_433] 2.61 6.94 5.08 2.96 8.23 5.97 3.37 9.81 6.99
 GMAN 2.77 7.25 5.48 3.07 8.35 6.34 3.40 9.72 7.22
+GTS 2.67 7.21 5.27 3.04 8.41 6.25 3.46 9.98 7.31
 GWN 2.69 6.90 5.15 3.07 8.37 6.22 3.53 10.01 7.37
 HA 4.16 13.00 7.80 4.16 13.00 7.80 4.16 13.00 7.80
+HighResNet 2.68 6.88 5.10 3.03 8.34 6.11 3.47 10.24 7.27
 LSTM 3.44 9.60 6.30 3.77 10.90 7.23 4.37 13.20 8.69
@@
 MRA-BGCN 2.67 6.80 5.12 3.06 8.30 6.17 3.49 10.00 7.30
-MS-GAT 2.65 6.43 2.99 4.83 8.31 5.91 6.48 9.48 8.11
 MTGNN 2.69 6.86 5.18 3.05 8.19 6.17 3.49 9.87 7.23
@@
 MVST-GNN [art_109] 2.70 5.19 6.97 3.06 8.12 6.08 3.40 9.60 6.98
+PAHNN [art_435] 2.51 6.34 4.82 2.87 7.63 5.73 3.25 9.32 6.67
+PDFormer 2.83 7.77 5.45 3.20 9.19 6.46 3.62 10.91 7.47
+PGCN [art_441] 2.70 6.98 5.16 3.08 8.38 6.22 3.54 9.94 7.36
 SDGCN 2.76 7.16 5.38 3.15 8.53 6.46 3.61 9.87 7.43
+SLCNN 2.53 6.70 5.18 2.88 8.00 6.15 3.30 9.70 7.20
+SNAS4MTF [art_408] 2.68 6.90 5.16 3.03 8.25 6.12 3.43 9.82 7.15
 STAG-GCN [art_93] 2.67 7.00 5.23 3.07 8.26 6.15 3.50 9.93 7.24
+STAEFormer 2.65 6.85 5.11 2.97 8.13 6.00 3.34 9.70 7.02
+STID 2.82 7.75 5.53 3.19 9.39 6.57 3.55 10.95 7.55
 STFGNN 2.57 6.51 4.73 2.83 7.46 5.46 3.18 8.81 6.40
@@
 T-GCN 3.03 7.81 5.26 3.52 9.45 6.12 4.30 11.80 7.31
+TGC-GRU 5.25 12.45 8.56 5.99 14.18 10.37 7.32 17.11 13.47
 VAR 4.42 10.20 7.89 5.41 12.70 9.13 6.52 15.80 10.11
+TimesNet 2.78 7.52 5.49 3.36 9.43 6.41 4.52 11.37 9.66
+TwoResNet [art_430] 2.65 6.78 5.08 3.01 8.14 6.07 3.39 9.71 7.08
 WaveNet 2.99 8.04 5.89 3.59 10.25 7.28 4.45 13.62 8.93
-The results on the METR-LA dataset are not uniform, and the best performing models vary depending on the metric and the time horizon.
-However, especially for longer time horizons, the Dynamic Graph Convolution LSTM Network (DyGCN-LSTM) [art_113] appears to be the most accurate one (excluding the ”Generic” models, among which the best model is the multichannel fusion graph neural network (MCFGNN) [art_86]).
+Z-GCNETs 3.23 7.87 7.48 3.93 9.75 9.40 4.83 12.04 11.57
+Clearly, the results on the METR-LA dataset are not uniform, and the best performing models vary depending on the metric and the time horizon.
+However, the multichannel fusion graph neural network (MCFGNN) [art_86] from the “Generic” group consistently exhibits low errors across different time horizons.
+The MCFGNN model, proposed by Chen et al.
+in 2022, is composed of three main components: a graph construction layer, a graph convolution layer, and a temporal convolution layer.
+The input data are first processed by the graph construction layer, which incorporates three channels to capture meta-features, time-varying patterns, and globally stabilized graphs.
+Spatial and temporal convolutions are then applied in pairs to capture spatio-temporal dependencies in the data.
+To mitigate the issue of vanishing gradients, residual connections are added to each spatio-temporal block.
 […]
@@
 The smallest errors are underlined.
-Numbers from the papers.
+Numbers from the original papers.
 Time horizon 15 min 30 min 60 min
@@
 AGP-GNN [art_116] 1.33 2.79 2.86 1.61 3.66 3.67 1.86 4.30 4.34
+AMGCN [art_444] 1.22 2.63 2.34 1.53 3.55 3.13 1.93 4.66 3.95
 ARIMA 1.62 3.50 3.30 2.33 5.40 4.76 3.38 8.30 6.50
 ASTGCN 1.52 3.22 3.13 2.01 4.48 4.27 2.61 6.00 5.42
-DCRNN 1.38 2.90 2.95 1.74 3.9 3.97 2.07 4.90 4.74
+ASTGNN 1.34 3.84 2.75 1.65 3.72 3.74 1.94 4.59 4.51
+DCRNN 1.38 2.90 2.95 1.74 3.90 3.97 2.07 4.90 4.74
 DGCRN [art_100] 1.28 2.66 2.69 1.59 3.55 3.63 1.89 4.43 4.42
+DMSTG - - - - - - 1.91 4.51 4.43
+DMSTGCN 1.33 2.80 2.83 1.67 3.81 3.79 1.99 4.78 4.54
+DSTCGCN - - - - - - 1.95 4.66 4.51
+DyHSL 1.30 2.73 2.71 1.60 3.57 3.67 1.95 4.62 4.53
 FC-GAGA 1.36 2.87 2.86 1.68 3.80 3.80 1.97 4.67 4.52
 FNN 2.20 5.19 4.42 2.30 5.43 4.63 2.46 5.89 4.98
+GL-STGTN [art_433] 1.26 2.79 2.73 1.55 3.63 3.58 1.80 4.38 4.30
 GMAN 1.34 2.81 2.82 1.62 3.62 3.72 1.86 4.31 4.32
+GSLNN [art_420] 1.29 2.71 2.66 1.64 3.67 3.74 1.93 4.51 4.46
+GTS 1.35 2.85 2.84 1.67 3.80 3.82 1.96 4.62 4.53
 GWN 1.30 2.73 2.74 1.63 3.67 3.70 1.95 4.63 4.52
-HA 2.88 6.80 5.59 2.88 6.80 5.59 2.88 6.80 5.59
+HA 1.89 4.16 4.30 2.50 5.62 5.82 3.31 7.65 7.54
 LSTM 2.05 4.80 4.19 2.20 5.20 4.55 2.37 5.70 4.96
+LSVR 1.85 3.80 3.59 2.48 5.50 5.18 3.28 8.00 7.08
+HighResNet 1.31 2.74 2.75 1.64 3.68 3.75 1.95 4.61 4.56
+MegaCRN 1.28 2.67 2.72 1.60 3.57 3.68 1.88 4.41 4.42
 MRA-BGCN 1.29 2.90 2.72 1.61 3.80 3.67 1.91 4.60 4.46
@@
 MVST-GNN [art_109] 1.30 2.73 2.74 1.57 3.56 3.54 1.80 4.19 4.26
+PAHNN [art_435] 1.27 2.60 2.65 1.52 3.48 3.54 1.83 4.35 4.27
+PDFormer 1.32 2.78 2.83 1.64 3.71 3.79 1.91 4.51 4.43
+PGCN [art_441] 1.30 2.72 2.73 1.62 3.63 3.67 1.92 4.45 4.45
+RGSL - - - - - - 2.03 4.77 4.64
+RST-LTG [art_406] - - - - - - 1.88 4.32 4.31
 SDGCN 1.35 2.89 2.88 1.69 3.87 3.91 1.99 4.69 4.59
+SLCNN 1.44 3.00 2.90 1.72 3.90 3.81 2.03 4.80 4.53
+SNAS4MTF [art_408] 1.30 2.72 2.77 1.61 3.63 3.68 1.89 4.44 4.36
+STAEFormer 1.31 2.76 2.78 1.62 3.62 3.68 1.88 4.41 4.34
+StemGNN 1.44 3.08 3.12 1.93 4.54 4.38 2.57 6.55 5.88
+STID 1.30 2.73 2.81 1.62 3.68 3.72 1.89 4.47 4.40
+STFGNN 1.42 2.92 2.97 1.67 3.92 3.94 1.98 4.83 4.74
 STGCN 1.36 2.90 2.96 1.81 4.17 4.27 2.49 5.79 5.69
@@
 STGRAT 1.29 2.67 2.71 1.61 3.63 3.69 1.95 4.64 4.54
-STID 1.30 2.73 2.81 1.62 3.68 3.72 1.89 4.47 4.4
 ST-MetaNet 1.36 2.82 2.90 1.76 4.00 4.02 2.20 5.45 5.06
@@
 SVR 1.85 3.80 3.59 2.48 5.50 5.18 3.28 8.00 7.08
+TD2MG2NN [art_422] 1.25 2.62 2.64 1.55 3.47 3.50 1.85 4.28 4.22
+TGC-GRU 2.34 5.08 4.30 2.44 5.40 4.64 2.57 5.78 5.01
+T-GCN 1.50 3.14 2.83 1.73 3.76 3.40 2.18 4.94 4.35
+TimesNet 1.32 2.94 2.72 1.76 3.74 4.16 2.18 4.84 5.20
+TwoResNet [art_430] 1.30 2.70 2.73 1.61 3.59 3.69 1.89 4.40 4.41
 VAR 1.74 3.60 3.16 2.32 5.00 4.25 2.93 6.50 5.44
 WaveNet 1.39 2.91 3.01 1.83 4.16 4.21 2.35 5.87 5.43
-The results on the the PEMS-BAY dataset indicate that the dynamic graph convolutional recurrent network (DGCRN) [art_100] is the most accurate model for the 15-minute forecasting horizon, while the multi-view spatial-temporal graph neural network (MVST-GNN) [art_109] is the best model for the longer horizons.
+Z-GCNETs 1.36 2.88 2.86 1.68 3.79 3.78 1.98 4.60 4.53
+The results on the the PEMS-BAY dataset vary depending on the metrics and the time horizon.
+The most accurate models are the adaptive multigraph convolutional networks (AMGCN) [art_444], the multi view spatial-temporal graph neural network (MVST-GNN) [art_109], and the temporal decomposition based dynamic multi granularity graph neural network (TD2MG2NN) [art_422].
+The AMGCN model was proposed by Li et al.
+in 2024, and includes a graph fusion block, an adaptive graph generator block, a stack of spatio-temporal convolution blocks, and an output module.
+The graph fusion block integrates information from five pre-defined graph structures, based on spatial distance, reachability, correlation, and distributions.
+Then, an independent adaptive graph generator learns an evolving graph structure to complement the pre-defined graph.
+Spatio-temporal features are finally extracted through a stack of graph convolutional modules with residual connections, and a time-mixing strategy.
+The MVST-GNN model, proposed by Li et al.
+in 2022, includes a multi-view temporal transformer module and a multi-view graph convolution module.
+The multi-view temporal transformer module is designed to capture dynamic temporal features from recent, daily, and weekly patterns using self-attention and a fusion module.
+Then, the graph convolution module operates on a self-learned adjacency matrix and passes its output through two fully connected linear layers to generate the final forecast.
+The TD2MG2NN model, instead, was introduced by Zang et al.
+in 2024 for evaluating the resilience of road networks in real time during heavy rain events.
+The process starts by decomposing the time series into seasonal and trend components.
+These components are processed individually through a season-temporal encoder layer with temporal convolutions, and a trend-temporal encoder layer with regression, respectively.
+The model then learns two feature matrices, one for daily and one for weekly patterns, which are fed into a dynamic multi-granularity graph convolution module.
+This module processes the feature matrices in parallel, and then passes them through a spatial decoder layer to produce the final output.
 […]
 Table 25: Comparison of the average accuracy of the different models on the PeMS03 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE.
-The results of the models marked with are taken from papers in the ”Generic” group.
+The results of the models marked with are taken from papers in the “Generic” group.
 The smallest errors are underlined.
-Numbers from the papers.
+Numbers from the original papers.
 Time horizon 60 min
 Metrics MAE MAPE RMSE
-AGCRN 17.69 19.40 29.66
+AGCRN 15.98 15.23 28.25
+ARIMA 35.31 33.78 47.59
 ASTGCN 17.69 19.40 29.66
+ASTGNN 14.78 14.79 25.00
 Auto-DSTSGN [art_108] 14.59 14.22 25.17
@@
 DCRNN 18.18 18.91 30.31
+DSTAGNN 15.57 14.68 27.21
+DyHSL 15.49 14.38 27.06
 Graph WaveNet 19.85 19.31 32.94
+HA 31.58 33.78 52.39
 LSGCN 17.94 - 29.85
 LSTM 21.33 23.33 35.11
+PAHNN [art_435] 14.02 13.19 23.67
 STAGCN [art_130] 15.40 14.48 26.23
@@
 STSGCN 17.48 16.78 29.21
+STWave+ [art_453] 14.71 14.88 26.31
 SVR 21.97 21.51 35.29
+TCN 19.32 19.93 33.55
+TimesNet 16.41 15.17 26.72
+VAR 23.65 24.51 38.26
+W-DSTAGNN [art_419] 15.31 14.49 26.59
 […]
 Table 26: Comparison of the average accuracy of the different models on the PeMS04 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE.
-The results of the models marked with are taken from papers in the ”Generic” group.
+The results of the models marked with are taken from papers in the “Generic” group.
 The smallest errors are underlined.
-Numbers from the papers.
+Numbers from the original papers.
 Time horizon 60 min
@@
 ASTGCN 22.93 16.56 35.22
+ASTGNN 18.60 12.36 30.91
 Auto-DSTSGN [art_108] 18.85 13.21 30.48
@@
 DCGCN [art_117] 20.28 - 31.65
-DCRNN 19.71 13.54 31.43
+DCRNN 24.70 17.12 38.12
 DGCRN 19.04 12.80 30.82
@@
 DSA-NET 22.79 16.03 35.77
+DSTAGNN 19.30 12.70 31.46
 DSTGN [art_73] 18.61 12.31 30.79
 DSTIGNN [art_74] 18.41 12.45 29.97
+DyHSL 17.66 12.42 29.46
+EnhanceNet 20.44 13.58 32.37
 FEDFormer 22.86 16.04 35.07
+GL-STGTN [art_433] 18.20 12.48 29.71
 GMAN 20.93 14.06 33.34
@@
 MTGODE 19.55 13.08 32.99
+PAHNN [art_435] 16.28 11.98 27.62
 SCINet 19.29 11.89 31.28
@@
 STG2Seq 25.20 18.77 38.48
-STGODE 20.84 13.77 32.82
+STGCN 21.16 13.83 34.89
+STGODE 20.84 13.76 32.84
+ST-GTNN 20.08 13.31 32.18
 STIDGCN 18.47 12.42 29.90
-STGCN 21.16 13.83 34.89
-STSGCN 21.08 13.88 33.83
-STGODE 20.84 13.76 32.84
+STNorm 19.21 13.05 32.30
 STSGCN 21.19 13.90 33.65
+STWawe+ [art_453] 18.25 12.21 30.14
 SVR 28.70 19.20 44.56
 TCN 23.22 15.59 37.26
-VAR 23.51 17.85 36.39
+TimesNet 21.63 13.15 34.90
+Transformer 21.10 15.13 31.46
+TSGDC [art_424] 18.80 12.67 31.08
+VAR 23.75 18.09 36.66
+W-DSTAGNN [art_419] 19.30 12.70 31.28
+Z-GCNETs 20.84 13.77 32.82
 […]
@@
 The smallest errors are underlined.
-Numbers from the papers.
+Numbers from the original papers.
 Time horizon 60 min
@@
 AGCRN 22.37 9.12 36.55
-ASTGCN 28.05 13.92 42.57
+AMGCN [art_444] 19.83 8.19 31.03
+ARIMA 38.17 19.46 59.27
+ASTGCN 24.01 10.73 37.87
 Auto-DSTSGN [art_108] 20.08 8.57 33.02
@@
 DCGCN [art_117] 22.06 - 34.66
-DCRNN 25.30 11.66 38.58
+DCRNN 25.22 11.82 38.61
+DGCRN 19.91 8.32 31.47
+DMSTG 19.98 8.56 33.08
+DMSTGCN 20.77 8.94 33.67
+DSTAGNN 21.42 9.01 34.51
+DSTCGCN 20.37 8.64 33.47
+DyHSL 18.84 8.11 31.65
+GMAN 21.56 9.51 34.97
 GWN 26.85 12.12 42.78
-LSGCN 27.31 - 41.46
+HA 45.12 24.51 65.64
+LSGCN 27.31 11.98 41.46
 LSTM 29.98 13.20 45.94
+MTGNN 20.57 9.12 33.54
+PAHNN [art_435] 17.35 7.64 29.78
+PGCN 20.03 8.78 32.54
+RGSL 20.73 8.71 34.48
+RST-LTG [art_406] 19.49 8.26 32.91
+SLCNN 20.38 8.85 33.27
 STAGCN [art_130] 21.10 8.92 34.10
+StemGNN 22.23 9.20 36.46
 STFGCN 22.07 9.21 35.80
+STFGNN 23.46 9.21 36.60
+STG2Seq 32.77 20.16 47.16
 STGCN 25.38 11.08 38.78
 STGODE 22.59 - 37.54
+ST-GRAT 20.69 9.06 33.01
+STNorm 20.59 8.61 34.86
 STSGCN 24.26 10.21 39.03
+STWave [art_453] 19.59 8.17 33.53
 SVR 32.49 14.26 50.22
+TCN 32.72 14.26 42.23
+TimesNet 25.12 10.60 40.71
+Transformer 22.07 10.12 34.21
+TSGDC [art_424] 19.96 8.54 33.25
+VAR 50.22 32.22 75.63
+Z-GCNETs 21.77 9.25 35.17
 […]
 Table 28: Comparison of the average accuracy of the different models on the PeMS08 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE.
-The results of the models marked with are taken from papers in the ”Generic” group.
+The results of the models marked with are taken from papers in the “Generic” group.
 The smallest errors are underlined.
-Numbers from the papers.
+Numbers from the original papers.
 Time horizon 60 min
@@
 AGCRN 15.95 10.09 25.22
+AJSTGL [art_448] 13.79 - 22.48
 ARIMA 31.09 22.73 44.32
-ASTGCN 18.61 13.08 28.16
+ASTGCN 18.25 11.64 28.06
+ASTGNN 15.00 9.50 24.70
 Auto-DSTSGN [art_108] 14.74 9.45 23.76
@@
 DCGCN [art_117] 15.68 - 24.39
-DCRNN 15.26 9.96 24.28
+DCRNN 16.82 10.92 26.36
+DHSL [art_400] 15.19 9.86 24.61
+DGCRN 16.22 12.06 26.10
 DMGF-Net [art_99] 16.48 10.56 25.73
+DMSTG 13.85 9.26 23.30
 DSA-NET 17.14 11.32 26.96
+DSTAGNN 15.67 - 25.37
+DSTCGCN 15.18 9.68 24.49
 DSTIGNN [art_74] 13.88 9.04 23.64
+DyHSL 14.01 8.87 22.91
+EnhanceNet 16.33 10.39 25.56
+FC-GAGA 19.41 13.53 30.22
 FEDformer 19.55 13.58 29.30
-GMAN 16.97 11.32 26.70
+FNN 24.24 16.21 37.57
+GL-STGTN [art_433] 14.00 9.50 23.41
+GMAN 15.31 10.13 24.92
 GODERN-FS [art_88] 15.59 9.95 24.79
@@
 HA 34.86 24.07 52.04
+iDCGCN 15.56 - 25.02
 LSGCN 17.73 11.20 26.76
 LSTM 22.20 14.20 34.06
-MTGNN 18.28 12.15 30.05
+MRA-BGCN 18.87 12.96 28.56
+MTGNN 15.71 10.03 24.62
 MTGODE 15.61 10.15 25.96
+PAHNN [art_435] 13.58 8.35 22.11
+RGSL 15.49 9.96 24.80
+RST-LTG [art_406] 13.61 9.05 23.15
 SCINet 15.78 9.97 24.60
 SDGL [art_78] 14.93 9.61 24.13
+SNAS4MTF [art_408] 15.07 9.67 24.00
 STAGCN [art_130] 15.36 9.80 24.32
 STCGNN [art_115] 15.55 10.02 24.61
-STFGCN 16.64 10.60 26.22
+StemGNN 15.91 10.90 25.44
+STFGNN 16.64 10.60 26.25
 STIDGCN 14.10 9.15 23.72
@@
 STGCN 17.50 11.29 27.09
+STGMN 16.04 - 24.18
 STGODE 16.79 10.58 26.01
+ST-GTNN 16.51 10.45 25.19
+STMetaNet 18.75 14.49 27.50
+STNorm 15.39 9.91 24.80
 STSGCN 17.13 10.96 26.80
+STWave+ [art_453] 13.21 8.63 23.04
 SVR 23.25 14.64 36.16
 TCN 22.72 14.03 35.79
-VAR 22.07 14.04 31.02
+TimesNet 19.01 11.83 30.65
+VAR 19.19 13.10 29.81
+As stated above, the results consistently indicate that the periodicity aware spatial-temporal adaptive hypergraph neural network (PAHNN) [art_435] proposed by Zhao et al.
+in 2024 is the most accurate model.
+The PAHNN model includes a temporal multi-periodic to time dependencies and periodic features of traffic time series, and a spatial adaptive hypergraph neural network to capture higher-order correlations among nodes.
 5.6 Predictive monitoring
-In modern industry, sensor-based monitoring of processes has become essential.
-This pairs with the rapid growth of the Internet of Things (IoT), which aims at a smarter management across various applications [art_9].
-In this framework, multi-sensor systems can be used to detect anomalies in complex scenarios and to monitor the overall status of a system.
-Predictive monitoring is a task that involves the continuous observation and analysis of a system, in order to forecast its future states and verify whether the predicted outcomes meet certain standards.
-This predictive monitoring task is important because it significantly improves efficiency, reliability, and cost of industrial operations.
-Related approaches include several techniques, such as anomaly detection, fault diagnosis and remaining useful life estimation.
-Despite the specificity of the field, 16 out of 156 papers address the topic of predictive monitoring.
-This includes anomaly detection ([art_9], [art_10]), fault diagnosis ([art_53], [art_54], [art_55], [art_56], [art_57], [art_58], [art_192]) and RUL estimation ([art_152], [art_153], [art_154], [art_155], [art_156], [art_157], [art_158]).
+The theme of predictive monitoring is related to sensor-based monitoring of processes, which has become essential in modern industry.
+The evolution of monitoring systems is concomitant with the rapid growth of the Internet of Things (IoT), which aims at a smarter management across various applications [art_9].
+Hence, in this context, multi-sensor systems can be used to detect anomalies in complex scenarios and to monitor the overall status of a system.
+In the specific, predictive monitoring is a task that involves the continuous observation and analysis of a system in order to forecast its future states and verify whether the predicted outcomes meet certain standards.
+The predictive monitoring task is important because it can significantly improve efficiency, reliability, and cost of industrial operations.
+Related data-based approaches include techniques such as anomaly detection, fault diagnosis, and estimation of the equipment’s remaining useful life.
+Despite the specificity of the field, 35 out of 281 papers address the topic of predictive monitoring.
+This includes anomaly detection [art_9, art_10, art_477, art_479, art_481, art_483, art_484, art_489], fault diagnosis [art_53, art_54, art_55, art_56, art_57, art_58, art_192, art_469, art_471, art_474, art_486, art_487, art_491] and RUL estimation [art_152, art_153, art_154, art_155, art_156, art_157, art_158, art_478, art_488].
 The use of GNNs in these analyses is justified by the necessity of using a model capable of capturing the spatio-temporal correlations among the many sensors present in the system.
 5.6.1 Overview
-Predictive monitoring includes several techniques, such as anomaly detection, fault diagnosis and remaining useful life estimation.
-Anomaly detection is a technique that aims to identify abnormal data that are not due to random deviations, and are rather generated by a different underlying mechanism [ref_anomaly_detection_1].
+As said, predictive monitoring is a research field which focuses on several techniques, such as anomaly detection, fault diagnosis and remaining useful life estimation.
+Anomaly detection is a technique that aims to identify abnormal data that are not due to random deviations from some regular pattern, but are generated by a different underlying mechanism [ref_anomaly_detection_1].
 It is a prerequisite part of fault diagnosis [ref_anomaly_detection_2].
-Fault diagnosis is the process of determining whether a fault has occurred in a system, including the identification of the time, location, type, and severity of the fault.
+Fault diagnosis is the process of determining whether a fault has occurred in a machine, including the identification of the time, location, type, and severity of the fault.
 It is usually considered to be a classification problem.
 As for remaining useful life (RUL) estimation, it aims to determine how long a machine will operate before it needs to be repaired or replaced, and it is useful for scheduling maintenance interventions.
-The selected papers implement these three techniques, with anomaly detection being less prevalent than the other two.
-The main applications studied are industrial (e.g., bearings fault diagnosis [art_54], [art_55], [art_57], [art_155], [art_156]) and mechanical (e.g, engines RUL estimation [art_152], [art_153], [art_154], [art_157]).
-There are also two papers dealing with energy related tasks, specifically fault diagnosis of photovoltaic systems [art_53] and energy networks [art_56].
-They are included in this section due to their specific focus on fault classification rather than on energy-related time series forecasting.
-It shoud also be underlined that the majority of the papers aim at solving a classification task, which is typical of anomaly detection and fault diagnosis.
+The selected papers implement these three techniques, with anomaly detection being less studied than the other two.
+The main applications developed are industrial (e.g., bearings fault diagnosis [art_54, art_55, art_57, art_155, art_156]) and mechanical (e.g, engines RUL estimation [art_152, art_153, art_154, art_157, art_478, art_488]).
+There are also some papers dealing with energy related tasks, specifically fault diagnosis of photovoltaic systems [art_53] and of energy networks [art_56, art_487].
+They are included in this thematic group because of their specific focus on fault classification rather than on energy-related time series forecasting.
+It should also be underlined that the majority of these papers aim at solving a classification task, which is typical of anomaly detection and fault diagnosis.
 As for RUL estimation, the quantity to be predicted is the RUL itself.
-For this reason, we do not refer to forecasting time horizons as in the other sections, since the goal is to estimate the RUL rather than to forecast data over future time intervals.
-The majority of the collected papers were published in 2023 by authors affiliated with Chinese universities.
-The most popular journal for these topics is IEEE Transactions on Instrumentation and Measurement by the Institute of Electrical and Electronics Engineers by IEEE, with 5 out of 15 papers.
-All other journals contain only one of the selected articles.
+For this reason, we do not make specific reference to forecasting time horizons, unlike in the other subsections, since the goal here is to estimate the RUL rather than to forecast data over future time intervals.
+The majority of the collected papers related to this theme were published in 2024 by authors affiliated with Chinese universities.
+The most popular journal for these topics is IEEE Transactions on Instrumentation and Measurement by IEEE, with 6 out of 35 papers.
 It can be observed that the majority of the articles are published in IEEE journals.
@@
 Tab. LABEL:t:link_dataset_predictive_monitoring contains a list of the public datasets used in the selected papers, with the links to access them.
-Table 25: List of public datasets in the ”Predictive monitoring” group and their corresponding links.
+Table 29: List of public datasets in the “Predictive monitoring” group and their corresponding links.
 Dataset Used by Link
@@
 Air Quality - TPS [art_9] https://www.kaggle.com/amritpal333/tps-july-2021-original-dataset-clean
-C-MAPSS [art_152], [art_153], [art_154], [art_157] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
-CWRU bearing dataset [art_55], [art_192] http://csegroups.case.edu/bearingdatacenter/home
+C-MAPSS [art_152], [art_153], [art_154], [art_157], [art_478], [art_488] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
+CWRU bearing dataset [art_55], [art_192], [art_491] http://csegroups.case.edu/bearingdatacenter/home
+Herbert river water level [art_483] https://zenodo.org/records/8053359
+Intelligent maintenance system (IMS) [art_475] https://www.nasa.gov/intelligent-systems-division
 JNU bearing dataset [art_55] http://mad-net.org:8765/explore.html?id=9&t=0.9355271549540183
@@
 MFPT [art_192] https://api.semanticscholar.org/CorpusID:26870683 (source paper)
+Mars science laboratory rover (MSL) [art_481], [art_489] https://doi.org/10.1145/3219819.3219845 (source paper)
 NREL PV [art_53] https://www.nrel.gov/
 PRONOSTIA [art_155], [art_156] https://hal.science/hal-00719503/document (source paper)
+Pooled server metrics (PSM) [art_479], [art_481] https://doi.org/10.1145/3447548.3467174 (source paper)
+Server Machine Dataset (SMD) [art_479], [art_484] https://github.com/NetManAIOps/OmniAnomaly
 SEU bearing dataset [art_55] https://github.com/cathysiyu/Mechanical-datasets
-XJTUGearbox and XJTUSpurgear [art_192] http://dx.doi.org/10.1016/j.ymssp.2021.108653 (source paper)
-The two most popular datasets are C-MAPSS and N-CMAPSS, two well-known datasets for RUL prediction provided by NASA.
+Secure water treatment (SWaT) [art_477], [art_479], [art_481], [art_484], [art_489] https://itrust.sutd.edu.sg/itrust-labs_datasets/dataset_info/
+Soil moisture active passive satellite (SMAP) [art_479], [art_481], [art_484], [art_489] https://doi.org/10.1145/3219819.3219845 (source paper)
+Water distribution (WaDi) [art_477], [art_479], [art_484], [art_489] https://itrust.sutd.edu.sg/itrust-labs_datasets/dataset_info/
+Wireless capsule endoscopy images [art_475] https://doi.org/10.1055/s-0043-105488 (source paper)
+XJTUGearbox and XJTUSpurgear [art_192], [art_474] http://dx.doi.org/10.1016/j.ymssp.2021.108653 (source paper)
+The two popular datasets are C-MAPSS and N-CMAPSS, two well-known datasets for RUL prediction provided by NASA.
 C-MAPSS describes the deterioration of aircraft engines by recording temperature, pressure, and fan speed during the cruise phase (https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf).
 N-CMAPSS is an extended version of C-MAPSS, whose records cover climbing, cruising, and descending flight conditions (https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/).
-Practically no papers use exogenous variables, with the exception of Ref.
+Almost no papers use exogenous variables, with the exception of Ref.
 [art_53], which evaluates the impact of meteorological variables on fault diagnosis of photovoltaic systems.
-The granularity of the studied time series is highly variable, ranging from 1 hour to fractions of seconds.
-Many papers report the frequency of the sensors, with values ranging from 1 Hz to 50 kHz.
+The time granularity of the studied time series is highly variable, ranging from 1 hour to fractions of seconds.
+Many papers report the frequency of the studied sensors, with values ranging from 1 Hz to 50 kHz.
 Most datasets are normalized, by either using min-max normalization or Z-score normalization.
 In some papers, a data augmentation procedure is employed to mitigate the natural imbalance between the number of normal and anomalous conditions.
-5.6.3 Proposed models
-Half of the proposed GNN models (9 out of 16 papers) fall under the category of pure convolutional models, followed by 4 attentional GNNs, 2 hybrid models, and one whose description does not clarify which class of the taxonomy it belongs to (Ref.
-[art_155]).
+5.6.3 Types of models
+More than half of the proposed spatio-temporal GNN models (19 out of 35 papers) fall under the category of pure convolutional GNNs, followed by 10 attentional GNNs, 2 hybrid models, and one recurrent model (Ref.
+[art_470]).
+One paper (Ref.
+[art_57]) uses both convolutional and attentional GNNs, one paper (Ref.
+[art_475]) uses a variant of the convolutional GNN, and the description of another one (Ref.
+[art_155]) does not clarify which class of the taxonomy it belongs to.
 Almost all the models have a graph learning module, so that the graph structure is rarely defined a priori by the researchers.
-Three papers ([art_55], [art_57], [art_154]) use the visibility graph algorithm proposed by Lacasa et al.
+Three papers [art_55, art_57, art_154] use the visibility graph algorithm proposed by Lacasa et al.
 in Ref.
 [visibility_graph] for converting the time series into a graph.
-The most common loss functions in this group are the cross entropy function for classification problems, and MSE or RMSE for RUL forecasting problems.
-Most of the papers mention the Python language and the PyTorch library, but only two papers provide a link to the source code of their model, and in both cases, it is for RUL prediction.
-The source code for the convolution-graph attention network (ConvGAT) [art_153] is available at https://github.com/CUG-FDGroup, while the one for the graph network with causal connectivity and temporal convolutional neural network feature extractor proposed in Ref.
-[art_155] is available at https://github.com/mylonasc/gnn-tcnn.
+The most common loss functions used within this thematic group are the cross entropy function for classification problems, and MSE or RMSE for RUL forecasting problems.
+Most of the papers mention the Python language and the PyTorch library, but only the papers listed in Tab.
+LABEL:t:link_codici_predictive_monitoring provide the links to the related source codes.
+Table 30: List of source codes of the “Health” models in the review.
+Model Link
+Convolution-graph attention network (ConvGAT) [art_153] https://github.com/CUG-FDGroup
+Graph Network with Causal Connectivity and Temporal Convolutional Neural Network feature extractor [art_155] https://github.com/mylonasc/gnn-tcnn
+Graph spatiotemporal process and anomaly scorer (GST-Pro) [art_477] https://github.com/huankoh/GST-Pro
+Causality Enhanced Global-Local Graph Neural Network (CEGLo-GNN) [art_482] https://github.com/YueYueXia/CEGLo-GNN
+Graph Deviation Network+ (GDN+) [art_483] https://github.com/KatieBuc/gnnad/releases
 5.6.4 Benchmark models
-Tab. LABEL:t:benchmark_models_predictive_monitoring presents the benchmark models employed in the ”Predictive monitoring” group by at least two articles examined in this review.
-Table 26: List of benchmark models in the ”Predictive monitoring” group divided per category.
+Tab. LABEL:t:benchmark_models_predictive_monitoring presents the benchmark models employed in the “Predictive monitoring” thematic group by at least two articles examined.
+Table 31: List of benchmark models in the “Predictive monitoring” group divided per category.
 Category Model Used by
-Classical machine learning methods Attention-based long-short term memory network [art_153], [art_157]
-Feature-attention based bidirectional GRU and CNN model (AGCNN) [AGCNN] [art_153], [art_154], [art_157]
-Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_152], [art_153], [art_154], [art_157]
-Convolutional neural network (CNN) [art_9], [art_56], [art_152], [art_153], [art_154], [art_156], [art_157]
-Decision tree (DT) [art_9], [art_56]
-Feed-forward neural network (FNN) [art_55], [art_155]
-Gradient boosting (GB) [art_56], [art_154]
-Gated recurrent unit (GRU) [art_53], [art_56], [art_156]
+Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_483], [art_487]
+Non-GNN machine learning methods Attention-based long-short term memory network [art_153], [art_157], [art_488]
+Feature-attention based bidirectional GRU and CNN model (AGCNN) [AGCNN] [art_153], [art_154], [art_157], [art_478], [art_488]
+Anomalous Rhythm Detection using Adversarially Generated Time Series (BeatGAN) [BeatGAN] [art_481], [art_484]
+Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_152], [art_153], [art_154], [art_157], [art_470], [art_488]
+Convolutional neural network (CNN) [art_9], [art_56], [art_152], [art_153], [art_154], [art_156], [art_157], [art_461], [art_469], [art_471], [art_486], [art_487], [art_488]
+Convolutional neural network with long-short term memory network (CNN-LSTM) [art_154], [art_470]
+Crossformer [Crossformer] [art_152], [art_482]
+Deep autoencoding gaussian model (DAGMM) [DAGMM] [art_481], [art_484]
+Deep belief network (DBN) [DBN] [art_157], [art_488]
+Decision tree (DT) [art_9], [art_56], [art_487]
+Feed-forward neural network (FNN) [art_55], [art_155], [art_469], [art_478]
+Genetic algorithm approach on restricted Boltzmann-based long-short term memory network (GA+RBM+LSTM) [GA+RBM+LSTM] [art_157], [art_488]
+Gradient boosting (GB) [art_56], [art_154], [art_487]
+Gated recurrent unit (GRU) [art_53], [art_56], [art_156], [art_469], [art_487]
+Integrated multi-head dual sparse self-attention network (IMDSSN) [IMDSSN] [art_154], [art_478], [art_488]
+Isolation forest [art_10], [art_481]
+$k$-nearest neighbors (KNN) [art_56], [art_475], [art_487]
+Local outlier factor (LOF) [art_475], [art_481]
 Logistic regression (LR) [art_9], [art_56]
-Long-short term memory network (LSTM) [art_55], [art_56], [art_154], [art_157], [art_158]
-Random forest (RF) [art_56], [art_154]
-Support vector machine (SVM) [art_9], [art_56]
-Transformer [attention] [art_152], [art_154]
-GNN methods Graph attention network (GAT) [GAT] [art_54], [art_56], [art_192]
-Graph convolutional network (GCN) [art_54], [art_55], [art_156], [art_192]
-Hierarchical attention graph convolutional network (HAGCN) [HAGCN] [art_152], [art_153]
-Spatio-temporal fusion attention (STFA) [art_157] [art_153], [art_154]
-Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_58], [art_152]
-In this group of papers there is a high variety of benchmarks used, and only few papers share a benchmark model, especially in the GNN category.
-The most popular benchmarks are two classic well-known machine learning models, the CNN and the LSTM network.
-Not surprisingly, there are no classical statistical or mathematical benchmark models, since they are more frequently used in forecasting problems compared to classification problems.
-Among the GNN benchmarks, there are two models that have been specifically developed for remaining useful life prediction, namely the hierarchical attention graph convolutional network (HAGCN) [HAGCN] and the spatio-temporal fusion attention (STFA) [art_157].
+Long-short term memory network (LSTM) [art_55], [art_56], [art_154], [art_157], [art_158], [art_461], [art_470], [art_487], [art_488]
+Long-short term memory-based variational auto-encoder (LSTM-VAE) [art_10], [art_477], [art_479], [art_484], [art_489]
+OmniAnomaly [OmniAnomaly] [art_477], [art_479], [art_481], [art_484], [art_489]
+Residual network (ResNet) [ResNet] [art_471], [art_491]
+Random forest (RF) [art_56], [art_154], [art_478], [art_487]
+Support vector machine (SVM) [art_9], [art_56], [art_471], [art_478], [art_486], [art_487]
+Deep transformer networks for anomaly detection (TranAD) [TranAD] [art_484], [art_489]
+Transformer [attention] [art_152], [art_154], [art_470], [art_478], [art_479], [art_488], [art_489]
+Unsupervised anomaly detection on multivariate time series (USAD) [5] [art_479], [art_484], [art_489]
+GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_58], [art_482]
+Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_480], [art_482]
+Fused sparse autoencoder and graph network (FuSAGNet) [FuSAGNet] [art_10], [art_479]
+Graph attention network (GAT) [GAT] [art_54], [art_56], [art_192], [art_470], [art_487]
+Graph convolutional network (GCN) [art_54], [art_55], [art_156], [art_192], [art_461], [art_474], [art_486]
+Graph deviation network (GDN) [GDN] [art_10], [art_477], [art_479], [art_481], [art_484]
+Gated graph convolutional network (GGCN) [GGCN] [art_153], [art_478]
+Graph WaveNet (GWN) [GWN] [art_58], [art_480], [art_482]
+Hierarchical attention graph convolutional network (HAGCN) [HAGCN] [art_152], [art_153], [art_478]
+Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_470], [art_483]
+Spatio-temporal fusion attention (STFA) [art_157] [art_153], [art_154], [art_478], [art_488]
+Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_58], [art_152], [art_470], [art_478], [art_480]
+In this thematic group there is a high variety of benchmarks used, and before 2024 only few papers share a benchmark model in the GNN category.
+The most popular benchmarks are two classic machine learning models, the CNN and the LSTM network.
+Not surprisingly, almost no statistical or mathematical benchmark models were used, since they are more frequently used in forecasting problems than in classification problems.
+Among the GNN benchmarks, some of them have been specifically developed for remaining useful life prediction or predictive monitoring tasks, such as the hierarchical attention graph convolutional network (HAGCN) [HAGCN] and the spatio-temporal fusion attention (STFA) [art_157].
 In the HAGCN model, proposed by Li et al.
@@
 As for the STFA model, proposed by Kong et al.
-in 2022, it integrates a priori knowledge about the equipment’s structure with a spatio-temporal deep learning architecture which employs LSTM cells and an attention mechanism.
+in 2022, it integrates some a priori knowledge about the equipment’s structure with a spatio-temporal deep learning architecture which employs LSTM cells and an attention mechanism.
 Neither of the two original papers includes a link to the source code.
 5.6.5 Results
-The results demonstrate that the proposed GNN models outperform the benchmarks.
+The reported results show that the proposed GNN models clearly outperform the benchmarks.
 The accuracy of the models is evaluated using different metrics, depending on whether the context is classification or forecasting.
@@
 The smallest errors are underlined.
-Numbers from the papers.
+Numbers from the original papers.
 Sub-dataset FD001 FD002 FD003 FD004
@@
 AGCNN 12.42 225.51 19.43 1492.76 13.39 227.09 21.50 3392.60
-Attention-based LSTM 14.53 322.44 - - - - 27.08 5649.14
 ATT-LSTM 13.95 320.00 17.65 2102.00 12.72 223.00 20.21 3100.00
@@
 CNN-LSTM 14.40 290.00 27.23 9869.00 14.32 316.00 26.69 6594.00
+CNN-Transformer 12.25 198.00 17.08 1575.00 13.39 290.00 19.86 1741.00
 ConvGAT [art_153] 11.34 197.43 14.12 771.61 10.97 235.26 15.51 1231.17
@@
 DBN 15.21 417.59 27.12 9031.64 14.71 442.43 29.88 7954.51
-DCNN 12.61 273.70 22.36 10412.00 12.64 284.1 23.31 12466.00
+DCNN 12.61 273.70 22.36 10412.00 12.64 284.10 23.31 12466.00
+DSAN 13.40 242.00 22.06 2869.00 15.12 497.00 21.03 2677.00
+DS-STFN [art_478] 10.92 161.35 13.77 946.34 10.01 150.42 15.53 1079.91
 Earlier CNN 18.45 1287.00 30.29 13570.00 19.82 1596.00 29.16 7886.00
@@
 GA+RBM+LSTM 12.56 231.00 22.73 3366.00 12.10 251.00 22.66 2840.00
+GA-Transformer 11.63 215.00 15.99 1133.00 11.35 228.00 20.15 2672.00
+GAT-DAT 13.83 318.60 14.81 1163.80 14.85 438.50 16.80 1928.60
 GAT-EdgePool 13.53 309.59 15.07 1380.41 14.66 470.74 17.60 1726.53
@@
 HDNN 13.02 245.00 15.24 1282.42 12.22 287.72 18.16 1527.42
+HGNN-ACGF 12.58 218.04 21.67 4584.97 12.40 248.47 22.43 2737.86
 IMDSSN 12.14 206.11 17.40 1775.15 12.35 229.54 19.78 2852.81
@@
 STFA [art_157] 11.35 194.44 19.17 2493.09 11.64 224.53 21.41 2760.13
-Transformer 13.52 287.07 19.32 1436.74 13.44 263.64 20.38 2784.62
-In general, the best results on the different sub-datasets are achieved by the GNN-based comprehensive dynamic structure GNN (CDSG) [art_154] and convolution-graph attention network (ConvGAT) [art_153] models.
+STP-GNN [art_488] 11.38 192.17 15.81 1094.81 11.02 199.58 17.77 1692.19
+TATFA-Transformer 12.21 261.50 15.07 1359.70 11.23 210.21 18.81 2506.35
+Transformer 12.92 276.27 22.64 1960.25 13.04 284.41 24.81 3892.24
+In general, the best results on the different sub-datasets are achieved by the convolution-graph attention network (ConvGAT) [art_153] and the the dual-stream spatio-temporal fusion network (DS-STFN) [art_478] models.
+The ConvGAT model, proposed by Chen et al.
+in 2023, applies independent convolutional kernels to each feature of the input time series, and passes the resulting values to a graph attention network, with graph structure defined using cosine similarity.
+Finally, a fully connected layer outputs the RUL prediction.
+The DS-STFN model, proposed by Zhang et al.
+in 2024, organizes the data into grid-structured and graph-structured formats, with the graph structure defined by cosine similarity.
+A temporal convolutional network with self-attention extracts temporal features from the grid data, and, in parallel, a convolutional graph transformer module captures spatial features from the graph data.
+The spatial and temporal features are then fused to predict the RUL.
 5.7 Generic
-This subsection includes all the papers that do not explicitly address a specific problem.
-They still present the results on some benchmark case studies, but the related datasets may cover different groups.
-In particular, the focus is on papers that do not discuss or mention any specific case study prior to the experimental section.
+This subsection includes all the papers that do not specifically address a problem in the specific themes discussed so far.
+These papers present results on benchmark case studies, for which the related datasets span different thematic groups at once.
+In particular, the focus is on papers that do not discuss or mention any specific case study prior to the experimental section, or analyze datasets from multiple fields.
 The reason for this choice is that they are considered to explore broader methodologies or frameworks that are applied to diverse datasets without a declared focus on a particular case.
 5.7.1 Overview
-A significant number of papers (24 out of 156) have been categorized under the ”Generic” group.
-The majority of these works commit to address a multivariate time series forecasting problem which is not necessarily limited to a specific field.
-This ”Generic” subset includes one paper from 2021, 4 from 2022, and 19 from 2023, highlighting the recent interest in applying GNNs in more general contexts.
-These papers are distributed across various journals, with the exception of Expert Systems with Applications by Elsevier, IEEE Transactions on Knowledge and Data Engineering by IEEE, and Information Sciences by Elsevier, each publishing 2 of the selected papers.
+A significant number of papers (44 out of 281) have been categorized under the “Generic” thematic group.
+The majority of these works commit themselves to address a multivariate time series forecasting problem not necessarily limited to a specific field.
+This “Generic” set includes one paper from 2021, 4 from 2022, 19 from 2023, and 20 from 2024, highlighting the recent interest in applying GNNs in cross-disciplinary contexts.
+These papers are scattered across various journals, with the exception of Expert Systems with Applications by Elsevier, IEEE Transactions on Knowledge and Data Engineering by IEEE, and Information Sciences by Elsevier, which published 4 and 3 of the selected papers, respectively.
 The majority of authors are affiliated with institutions in China.
@@
 The selected papers make use of the datasets listed in Tab. LABEL:t:link_dataset_generic in their experiments.
-Table 28: List of public datasets in the ”Generic” papers and their corresponding links.
+Table 33: List of public datasets in the “Generic” papers and their corresponding links.
 Dataset Used by Link
@@
 CCMP wind data [art_73] http://data.remss.com
+CDC US States [art_401] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
 CI earthquakes [art_160] https://doi.org/10.5194/adgeo-43-31-2016 (source paper)
+CityBike NYC [art_464] https://www.citibikenyc.com/system-data
 Chickenpox cases [art_160] https://doi.org/10.48550/arXiv.2102.08100 (source paper)
+C-MAPSS [art_403] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
+CORAv2.0 [art_396] https://mds.nmdis.org.cn/pages/home.html
 CW earthquakes [art_89] https://doi.org/10.5194/adgeo-43-31-2016 (source paper)
-Electricity consumption [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_88], [art_90] https://github.com/laiguokun/multivariate-time-series-data
+ECG5000 [art_397] https://www.cs.ucr.edu/ẽamonn/time_series_data/
+Electricity consumption [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_88], [art_90], [art_384], [art_405] https://github.com/laiguokun/multivariate-time-series-data
 Electricity Consuming Load (ECL) [art_82], [art_85] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014
-Electricity transformer Temperature (ETT) [art_76], [art_82], [art_85] https://github.com/zhouhaoyi/ETDataset
-Energy [art_77] http://dx.doi.org/10.1016/j.enbuild.2017.01.083 (source paper)
+Electricity transformer Temperature (ETT) [art_76], [art_82], [art_85], [art_393], [art_417] https://github.com/zhouhaoyi/ETDataset
+Energy [art_77], [art_388] http://dx.doi.org/10.1016/j.enbuild.2017.01.083 (source paper)
 Enron [art_202] https://doi.org/10.1073/pnas.1800683115 (source paper)
-Eu-Core [art_202] https://doi.org/10.1145/ (source paper)
-Exchange-rate [art_73], [art_74], [art_76], [art_77], [art_78], [art_79], [art_80], [art_81], [art_85], [art_87], [art_89], [art_90] https://github.com/laiguokun/multivariate-time-series-data
+Eu-Core [art_202] https://doi.org/10.1145/3018661.3018731 (source paper)
+Exchange-rate [art_73], [art_74], [art_76], [art_77], [art_78], [art_79], [art_80], [art_81], [art_85], [art_87], [art_89], [art_90], [art_384], [art_393], [art_405] https://github.com/laiguokun/multivariate-time-series-data
 Facebook [art_202] http://networkrepository.com/socfb
@@
 Hypertext [art_202] http://networkrepository.com/ia-infect-hyper (source paper)
+IDWR Japan Prefectures [art_401] https://www.niid.go.jp/niid/en/idwr-e.html
 ILI [art_76], [art_85], [art_89] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
+ISRUC-SLEEP Dataset [art_403] https://sleeptight.isr.uc.pt/ISRUC_Sleep/
 Jena Weather [art_76] https://www.bgc-jena.mpg.de/wetter/
+KnowAir [art_400] https://github.com/shuowang-ai/PM2.5-GNN
+LA Traffic [art_404] https://ladot.lacity.gov/residents/traffic-counts
 Loan [art_202] https://ai.ppdai.com/mirror/showCompetitionRisk
-METR-LA [art_72], [art_75], [art_79], [art_80], [art_86], [art_160] https://paperswithcode.com/dataset/metr-la
+Los-loop [art_383] https://github.com/lehaifeng/T-GCN/tree/master/data
+Large-Scale Traffic and Weather Events (LSTW) [art_404] https://smoosavi.org/datasets/lstw
+METR-LA [art_72], [art_75], [art_79], [art_80], [art_86], [art_160], [art_387], [art_397], [art_401], [art_417], [art_464] https://paperswithcode.com/dataset/metr-la
+Metro [art_400] https://lbs.amap.com/
 MIMIC-III Heart Failure [art_71] https://doi.org/10.1038/sdata.2016.35 (source paper)
@@
 Nasdaq [art_77], [art_80] https://doi.org/10.48550/arXiv.1704.02971 (source paper)
-NOAA Weather [art_82] https://www.ncei.noaa.gov/data/local-climatological-data/
+NOAA Weather [art_82], [art_393] https://www.ncei.noaa.gov/data/local-climatological-data/
+NYC-Taxi [art_398] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
 PeMS03 [art_74] https://www.kaggle.com/datasets/elmahy/pems-dataset
-PeMS04 (PeMSD4) [art_73], [art_74], [art_78], [art_79], [art_88] https://paperswithcode.com/dataset/pems04
-PeMS08 (PeMSD8) [art_73], [art_74], [art_78], [art_79], [art_88] https://paperswithcode.com/dataset/pems08
-PEMS-BAY [art_72], [art_75], [art_160] https://paperswithcode.com/dataset/pems-bay
+PeMS04 (PeMSD4) [art_73], [art_74], [art_78], [art_79], [art_88], [art_384], [art_396], [art_397], [art_400], [art_405] https://paperswithcode.com/dataset/pems04
+PeMS08 (PeMSD8) [art_73], [art_74], [art_78], [art_79], [art_88], [art_383], [art_384], [art_396], [art_397], [art_400], [art_405] https://paperswithcode.com/dataset/pems08
+PEMS-BAY [art_72], [art_75], [art_160], [art_401] https://paperswithcode.com/dataset/pems-bay
 PhysioNet [art_71] https://archive.physionet.org/challenge/2012/papers/
 Reddit network [art_70] http://snap.stanford.edu/jodie/reddit.csv
-Solar-Energy [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_88], [art_90] https://github.com/laiguokun/multivariate-time-series-data
+Seattle-loop [art_401] https://github.com/zhiyongc/Seattle-Loop-Data
+Sleep-EDF-20 [art_395] https://gist.github.com/emadeldeen24/a22691e36759934e53984289a94cb09b
+Solar-Energy [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_88], [art_90], [art_384], [art_387], [art_405] https://github.com/laiguokun/multivariate-time-series-data
+Secure Water Treatment (SWaT) [art_387] https://itrust.sutd.edu.sg/itrust-labs_datasets/
+Stocktwits [art_398] https://stocktwits.com/
 TAIEX [art_84] https://finance.yahoo.com/
 TaoBao [art_202] https://tianchi.aliyun.com/competition/entrance/
-Traffic [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_85], [art_87], [art_90] https://github.com/laiguokun/multivariate-time-series-data
+Taiwan PM2.5 [art_464] https://taqm.epa.gov.tw/taqm/en/default.aspx
+T-drive [art_383] https://www.kaggle.com/datasets/arashnic/tdriver
+Tiingo [art_398] https://www.tiingo.com/
+Traffic [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_85], [art_87], [art_90], [art_388] https://github.com/laiguokun/multivariate-time-series-data
+UCI Human Activity Recognition [art_403] https://archive.ics.uci.edu/dataset/240/human+activity+recognition+using+smartphones
+University of East Anglia (UEA) datasets [art_385], [art_386], [art_389], [art_407] https://timeseriesclassification.com/
 US stock market price [art_89] https://doi.org/10.48550/arXiv.1810.09936 (source paper)
+Water Distribution (WaDi) [art_387] https://itrust.sutd.edu.sg/itrust-labs_datasets/
+WESAD dataset [art_404] https://ubi29.informatik.uni-siegen.de/usi/data_wesad.html
 Wikipedia network [art_70] http://snap.stanford.edu/jodie/wikipedia.csv
 Wind-Speed [art_76] https://www.kaggle.com/datasets/fedesoriano/wind-speed-prediction-dataset
-It should be noted that there are multiple recurring datasets that can be considered as benchmarks for the community.
-Some of these have already been discussed in the previous sections.
-If the results for these datasets have already been shown in previous sections, they will not be repeated again here.
-Instead, they will be just marked with a star together with the benchmark models in the tables of the previous sections.
+Wind Generation [art_405] https://www.kaggle.com/datasets/sohier/30-years-of-european-wind-generation
+WISDM [art_395] https://archive.ics.uci.edu/dataset/507/wisdm+smartphone+and+smartwatch+activity+and+biometrics+dataset)
+It should be noted that multiple recurring datasets can be considered as benchmarks for this community.
+Some of these were discussed in the previous subsections (e.g., METR-LA, PeMS datasets).
+Hence, if the results for these datasets were already shown in previous subsections, they will not be discussed again here.
+Instead, they will be just marked with the symbol in the tables of the previous subsections.
 The four most common datasets are Electricity consumption, Exchange-rate, Solar-Energy and Traffic, all available at https://github.com/laiguokun/multivariate-time-series-data.
-The Electricity consumption dataset contains the hourly electricity consumption in kWh for 321 customers recorded from 2012 to 2014.
+The Electricity consumption dataset contains the hourly electricity consumption in kWh for 321 customers, recorded from 2012 to 2014.
 The Traffic dataset describes the road occupancy rates measured by different sensors on the San Francisco Bay area road network from 2015 to 2016, with an hourly granularity.
@@
 Finally, Exchange rate contains daily exchange rates of 8 countries including Australia, British, Canada, Switzerland, China, Japan, New Zealand and Singapore, from 1990 to 2016.
-In general, the data granularity varies significantly, ranging from fractions of seconds to one week, indicating that there is no particular focus on a specific granularity.
+The University of East Anglia (UEA) dataset includes 30 multivariate time series from many real-world applications, including heart records, electroencephalogram signals, motion data, audio spectra, written characters.
+Since a large portion of the dataset contains data not related to the focus of this review, the details of the results and models used will not be discussed here.
+Interested readers are referred to Refs.
+[art_385, art_386, art_389, art_407].
+In general, the data granularity varies significantly, ranging from fractions of seconds to one month, indicating that there is no particular focus on a specific granularity.
 Individual papers often cover multiple granularities, and many papers take into account datasets with granularity from 5 minutes to 1 day.
@@
 Only a few papers mention data pre-processing, and usually utilize min-max or Z-score normalization.
-5.7.3 Proposed models
-Among the 24 ”Generic” papers, 19 use a purely convolutional GNN approach, and two employ a purely attentional model.
-One paper lacks sufficient detail to determine its classification within the taxonomy [art_85], and another declares it is using an aggregating approach that is simpler than either convolutional or attentional approaches [art_81].
-One more paper proposes a multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) model, where the continuous dynamics of simplified graph propagation is described by an ordinary differential equation (ODE) [art_75].
+5.7.3 Types of models
+Among the 43 papers in the “Generic” thematic group, 33 use a purely convolutional GNN approach, and four employ a purely attentional model.
+Five paper lack sufficient detail to be classified within the taxonomy chosen (Refs.
+[art_85, art_385, art_386, art_389, art_407]), and another one declares that it uses an aggregating approach that is simpler than either convolutional or attentional approaches [art_81].
+Another paper proposes a so-called multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) model, where the continuous dynamics of simplified graph propagation is described by an ordinary differential equation (ODE) [art_75].
+This is not the only paper in this section using differential equations.
+In Ref.
+[art_401], two domain specific ODEs (one for traffic speed evolution, and the other for disease propagation) are integrated into the convolutional GNN model.
+The authors argue that this dynamics-informed model approach enhances robustness, allow easier generalization when the equation of the dynamics is known, and reduces the number of model parameters.Additionally, during inference, the model only requires a single data point, as the ODEs explicitly incorporate the immediate dynamics.
 As for the definition of the graph structure, the majority of these models use self-learning techniques.
-This is a straightforward choice, since all the papers in this section want to ensure the applicability of the model in different contexts, and the use of pre-defined rules for the definition of the graph limits the model’s universal applicability.
-Half of the papers employ a dynamic graph structure, which adapts to the varying temporal patterns observed in the data.
+This is a straightforward choice, since all the papers in this group want to ensure the applicability of the model in different contexts, and the use of pre-defined rules for the definition of the graph limits the model’s universal applicability.
+More than half of the papers employ a dynamic graph structure, which adapts to the varying temporal patterns observed in the data.
 The two most commonly utilized loss functions are the MSE and the MAE.
 Since the graph structure is typically learned by the models themselves, many papers include a regularization term in the loss function for the optimization of the graph structure (e.g., Refs.
-[art_73], [art_78], [art_79], [art_86]).
+[art_73, art_78, art_79, art_86]).
 The majority of the papers specify the language and libraries used for the code, namely Python with PyTorch, PyTorch Geometric [PyTorch_Geometric], and Torch Spatiotemporal [Cini_Torch_Spatiotemporal_2022].
@@
 LABEL:t:link_codici_generic.
-Table 29: List of source codes of the ”Generic” models in the review.
+Table 34: List of source codes of the “Generic” models in the review.
 Model Link
 Adaptive dependency learning neural network (ADLNN) [art_83] https://github.com/AbishekSriramulu/ADLGNN.git
+Attention-adjusted Graph Spatio-Temporal Network (AGSTN) [art_464] https://github.com/l852888/AGSTN
+Cross-attention temporal dynamic graph neural network (CATodyNet) [art_386] https://github.com/HQYWY/CATodyNet
 Graph neural network with neural Granger causality (CauGNN) [art_77] https://github.com/RRRussell/CauGNN
+Dynamic diffusion-variational GNN (DVGNN) [art_383] https://github.com/gorgen2020/DVGNN
 Gformer [art_82] https://github.com/wxh453751461/Gformer
+GraphSensor [art_395] https://github.com/jldx-gjq/Graphnet/tree/main/graphsensor-code-main
 Multi-scale adaptive graph neural network (MAGNN) [art_80] https://github.com/shangzongjiang/MAGNN
+Multi-task Dynamic Graph generation framework for multivariate time series forecasting (MDG) [art_397] https://github.com/gortwwh/MDG
 Multivariate time series deep spatiotemporal forecasting model with a graph neural network (MDST-GNN) [art_90] https://github.com/yiminghzc/MDST-GNN
+Multivariate Time Series Classification based on Fusion Features (MSTC_FF) [art_385] https://github.com/dumingsen/MTSC_FF
 Multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) [art_75] https://github.com/GRAND-Lab/MTGODE
@@
 Temporal graph convolution and attention (T-GAN) [art_202] https://github.com/malei666666/T_GAN
+Temporal Dynamic Graph Neural Network (TodyNet) [art_389] https://github.com/liuxz1011/TodyNet
 Two GNN models with different graph structures [art_160] https://github.com/StefanBloemheuvel/graph_comparison
 5.7.4 Benchmark models
-Tab. LABEL:t:benchmark_models_generic displays the benchmark models utilized by at least two of the selected ”Generic” papers.
-Table 30: List of benchmark models in the ”Generic” group divided per category.
+Tab. LABEL:t:benchmark_models_generic displays the benchmark models utilized by at least two of the selected “Generic” papers.
+Table 35: List of benchmark models in the “Generic” group divided per category.
 Category Model Used by
-Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_73], [art_74], [art_78], [art_79], [art_80], [art_81], [art_83], [art_84], [art_88], [art_89]
+Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_73], [art_74], [art_78], [art_79], [art_80], [art_81], [art_83], [art_84], [art_88], [art_89], [art_384], [art_393], [art_397], [art_400], [art_405], [art_464]
 Gaussian process (GP) [art_73], [art_74], [art_78], [art_80], [art_81], [art_83]
-Historical average (HA) [art_78], [art_88]
-Vector autoregression (VAR) [art_73], [art_74], [art_77], [art_78], [art_79], [art_81], [art_84], [art_88]
-Classical machine learning methods Autoformer [Autoformer] [art_74], [art_76], [art_85]
-Dual self-attention network for multivariate time series forecasting (DSANet) [DSANet] [art_73], [art_78]
+Historical average (HA) [art_78], [art_88], [art_400]
+Vector autoregression (VAR) [art_73], [art_74], [art_77], [art_78], [art_79], [art_81], [art_84], [art_88], [art_384], [art_400], [art_405]
+Non-GNN machine learning methods Autoformer [Autoformer] [art_74], [art_76], [art_85]
+DeepAR [DeepAR] [art_69], [art_393]
+DeepSleepNet [DeepSleepNet] [art_395], [art_403]
+Dual self-attention network for multivariate time series forecasting (DSANet) [DSANet] [art_73], [art_78], [art_405]
 Feed-forward neural network (FNN) [art_84], [art_89]
-Gated recurrent unit (GRU) [art_73], [art_78], [art_82], [art_88], [art_90]
+Gated recurrent unit (GRU) [art_73], [art_78], [art_82], [art_88], [art_90], [art_384], [art_398], [art_404], [art_405]
 Hybrid framework based on fully Dilated CNN (HyDCNN) [HyDCNN] [art_75], [art_83]
-Informer [Informer] [art_74], [art_76], [art_82], [art_85]
+Informer [Informer] [art_74], [art_76], [art_82], [art_85], [art_397]
+$k$-nearest neighbors (KNN) [art_84], [art_387]
 LogTrans [LogTrans] [art_82], [art_85]
-Long-short term memory network (LSTM) [art_71], [art_74], [art_79], [art_82], [art_84], [art_85], [art_86], [art_88], [art_89]
-Long-short term time series network (LSTNet) [LSTNet] [art_73], [art_74], [art_75], [art_77], [art_78], [art_80], [art_81], [art_82], [art_83], [art_85], [art_87], [art_89], [art_90]
+Long-short term memory network (LSTM) [art_71], [art_74], [art_79], [art_82], [art_84], [art_85], [art_86], [art_88], [art_89], [art_396], [art_397], [art_398], [art_400], [art_464]
+Long-short term time series network (LSTNet) [LSTNet] [art_73], [art_74], [art_75], [art_77], [art_78], [art_80], [art_81], [art_82], [art_83], [art_85], [art_87], [art_89], [art_90], [art_387], [art_393]
 Multi-Level Construal Neural Network (MLCNN) [MLCNN] [art_77], [art_87]
 Neural basis expansion analysis for interpretable time series forecasting (N-BEATS) [N-BEATS] [art_69], [art_85]
-Recurrent neural network with fully connected gated recurrent units (RNN-GRU) [art_73], [art_77], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87]
+Recurrent neural network with fully connected gated recurrent units (RNN-GRU) [art_73], [art_77], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_405]
 Reformer [Reformer] [art_76], [art_82], [art_85]
 SCINet [SCINet] [art_74], [art_90]
-Temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM] [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_90]
-Transformer [attention] [art_71], [art_77]
-Hybrid of the multilayer perception and autoregressive model (VAR-MLP) [VAR-MLP] [art_73], [art_74], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_90]
-GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_73], [art_74], [art_78], [art_79], [art_80], [art_88]
-Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_78], [art_79]
-Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_73], [art_74], [art_75], [art_78], [art_79], [art_84], [art_86], [art_87], [art_88]
+Support vector machine (SVM) [art_84], [art_396], [art_464]
+Temporal convolutional network (TCN) [TCN] [art_88], [art_387], [art_396], [art_397]
+Temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM] [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_90], [art_405]
+Transformer [attention] [art_71], [art_77], [art_393]
+Hybrid of the multilayer perception and autoregressive model (VAR-MLP) [VAR-MLP] [art_73], [art_74], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_90], [art_405]
+GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_73], [art_74], [art_78], [art_79], [art_80], [art_88], [art_384], [art_396], [art_400], [art_464]
+Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_78], [art_79], [art_383], [art_384], [art_396], [art_397], [art_400], [art_464]
+Attention based Spatial-Temporal Graph Neural Network (ASTGNN) [ASTGNN] [art_74], [art_396]
+Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_73], [art_74], [art_75], [art_78], [art_79], [art_84], [art_86], [art_87], [art_88], [art_383], [art_384], [art_396], [art_397], [art_400], [art_405]
+Dynamic Graph Convolutional Recurrent Network (DGCRN) [art_100] [art_397], [art_398]
+Dynamic and Multi-faceted SpatioTemporal Graph Convolution Network (DMSTGCN) [DMSTGCN] [art_396], [art_417]
+Dynamic Spatial-Temporal Aware Graph Neural Network (DSTAGNN) [DSTAGNN] [art_383], [art_384]
 Graph auto encoder (GAE) [GAE] [art_70], [art_202]
 Graph multi-attention network (GMAN) [GMAN] [art_73], [art_75], [art_86], [art_88]
-Graph for time series (GTS) [GTS] [art_72], [art_78]
-Graph WaveNet (GWN) [GWN] [art_74], [art_75], [art_79], [art_80], [art_86], [art_87]
-Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_72], [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_86], [art_87], [art_88], [art_90]
+Discrete graph structure learning for time series (GTS) [GTS] [art_72], [art_78], [art_384], [art_397], [art_401]
+Graph WaveNet (GWN) [GWN] [art_74], [art_75], [art_79], [art_80], [art_86], [art_87], [art_396], [art_397], [art_417]
+Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_72], [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_86], [art_87], [art_88], [art_90], [art_393], [art_396], [art_397], [art_401], [art_405], [art_417]
+Multivariate Time series with dynamic Graph neural Ordinary Differential Equations (MTGODE) [art_75] [art_88], [art_401], [art_417]
 Node2vec [node2vec] [art_70], [art_202]
-Spatial-temporal fusion graph neural network (STFGNN) [STFGNN] [art_73], [art_86]
-Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_73], [art_75], [art_78], [art_79], [art_84], [art_86], [art_88]
-Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_75], [art_88]
-Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_73], [art_74], [art_78]
-There is a wide variety of benchmark models in this ”Generic” group.
+Spectral temporal graph neural network for multivariate time-series forecasting (StemGNN) [StemGNN] [art_397], [art_405]
+Spatial-temporal fusion graph neural network (STFGNN) [STFGNN] [art_73], [art_86], [art_405]
+Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_73], [art_75], [art_78], [art_79], [art_84], [art_86], [art_88], [art_384], [art_397], [art_400], [art_401]
+Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_75], [art_88], [art_383]
+Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_73], [art_74], [art_78], [art_387], [art_397], [art_400], [art_405]
+Temporal graph convolutional network (T-GCN) [T-GCN] [art_383], [art_384], [art_393]
+Time zigzags at graph convolutional networks for time series forecasting (Z-GCNETs) [Z-GCNETs] [art_400], [art_405]
+This “Generic” thematic group includes a wide variety of benchmark models, also due to the large number of papers within this group.
 As for mathematical and statistical models, autoregressions and ARIMA models are very popular for their simplicity and interpretability.
-Among classical machine learning models, the most widely used models are LSTMs, temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM], and long-short term time series network (LSTNet) [LSTNet].
+Among non-GNN machine learning models, the most widely used models are LSTM networks, the so-called temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM], and the long-short term time series network (LSTNet) [LSTNet].
 The TPA-LSTM model, introduced by Shih et al.
@@
 The source code for the model is accessible at https://github.com/laiguokun/LSTNet.
-Among GNN models, the most common ones are DCRNN [DCRNN] and MTGNN [MTGNN], already discussed in the previous sections.
+Among GNN models, the most common ones are DCRNN [DCRNN] and MTGNN [MTGNN] models, already discussed in the previous subsections.
 5.7.5 Results
-The most common error metrics used in the evaluation of the models in the ”Generic” group are, in order, the mean absolute error (MAE), correlation coefficient (CORR), root mean square error (RMSE), root relative squared error (RRSE), mean absolute percentage error (MAPE), and mean squared error (MSE).
+The most common error metrics used in the evaluation of the models in the “Generic” group are, in order, the mean absolute error (MAE), correlation coefficient (CORR), root mean squared error (RMSE), root relative squared error (RRSE), mean absolute percentage error (MAPE), and mean squared error (MSE).
 Differently from other metrics, where lower values indicate higher accuracy, the correlation coefficient (CORR) quantifies the strength of the the linear relationship between predicted and actual values, with higher values indicating a better performance.
-This Results subsection compares the results obtained for the Electricity consumption, Exchange rate, Solar energy and Traffic datasets.
+This subsection about results compares the results obtained for the Electricity consumption, Exchange rate, Solar energy and Traffic datasets.
 The results of the proposed models and benchmarks on METR-LA and PEMS datasets were already reported with a in Subsec.
@@
 AGLG-GRU [art_79] 0.0738 0.9434 0.0864 0.9302 0.0912 0.9283 0.0947 0.9274
-AR 0.0995 0.8845 0.1035 0.8632 0.105 0.8591 0.1054 0.8595
+AR 0.0995 0.8845 0.1035 0.8632 0.1050 0.8591 0.1054 0.8595
 ARIMA 0.0917 0.8902 0.1002 0.8834 0.1028 0.8604 0.1042 0.8487
+ASTCRF [art_384] 0.0742 0.9593 0.0872 0.9511 0.0913 0.9388 - -
 Autoformer 0.1258 0.9147 0.1344 0.9001 0.1357 0.8921 0.1554 0.8704
 Crossformer 0.0742 0.9452 0.0855 0.9351 0.0901 0.9280 0.0967 0.9205
+DPGNN [art_405] 0.0720 0.9527 0.0810 0.9422 0.0893 0.9357 0.0935 0.9291
 DSTGN [art_73] 0.0713 0.9518 0.0821 0.9424 0.0887 0.9357 - -
-DSTIGNN [art_74] 0.0733 0.9515 0.082 0.9412 0.0899 0.9338 0.0954 0.9289
+DSTIGNN [art_74] 0.0733 0.9515 0.0820 0.9412 0.0899 0.9338 0.0954 0.9289
 ESG 0.0718 0.9494 0.0844 0.9372 0.0898 0.9321 0.0962 0.9279
-FEDformer 0.0889 0.9321 0.1006 0.9191 0.1154 0.908 0.1202 0.9012
+FEDformer 0.0889 0.9321 0.1006 0.9191 0.1154 0.9080 0.1202 0.9012
 GP 0.1500 0.8670 0.1907 0.8334 0.1621 0.8394 0.1273 0.8818
+GRU 0.1102 0.8597 0.1144 0.8623 0.1183 0.8472 - -
 GTS 0.0790 0.9291 0.0884 0.9187 0.0957 0.9135 0.0951 0.9098
@@
 SDGL [art_78] 0.0698 0.9534 0.0805 0.9445 0.0889 0.9351 0.0935 0.9301
-SDLGNN 0.0726 0.9502 0.082 0.9384 0.0896 0.9304 0.0947 0.9257
+SDLGNN 0.0726 0.9502 0.0820 0.9384 0.0896 0.9304 0.0947 0.9257
 SDLGNN-Corr 0.0737 0.9475 0.0841 0.9346 0.0923 0.9263 0.0971 0.9227
 STG-NCDE 0.6152 0.8739 0.6584 0.8663 0.7302 0.8728 - -
+StemGNN 0.0799 0.9490 0.0909 0.9397 0.0989 0.9342 0.1019 0.9209
 TDG4-MSF [art_81] 0.0731 0.9499 0.0828 0.9371 0.0894 0.9306 0.0969 0.9246
@@
 The accuracy of the various models varies depending on the studied forecasting horizon but, overall, one can notice that the GNN models perform better than the others, with the static and dynamic graph learning network (SDGL) [art_78] being the most accurate.
+The SDGL model, proposed by Li et al.
+in 2023, first employs a static and dynamic graph learning module that incorporates graph regularization to control the smoothness, connectivity, and sparsity of the learned graph.
+The graph input then passes through gated temporal convolutional networks and graph convolution modules, which are equipped with residual and skip connections that deliver the information to the output layer.
 Tab.
@@
 ARIMA 0.0198 0.9754 0.0261 0.9721 0.0344 0.9548 0.0445 0.9301
+ASTCRF [art_384] 0.0165 0.9792 0.0235 0.9709 0.0344 0.9572 - -
 Autoformer 0.0291 0.9631 0.0379 0.9397 0.0441 0.9201 0.0501 0.9012
 Crossformer 0.0226 0.9759 0.0269 0.9656 0.0358 0.9436 0.0489 0.9291
+DPGNN [art_405] 0.0184 0.9797 0.0246 0.9715 0.0346 0.9564 0.0456 0.9386
 DSTGN [art_73] 0.0179 0.9782 0.0250 0.9715 0.0346 0.9562 - -
@@
 GP 0.0239 0.8713 0.0272 0.8193 0.0394 0.8484 0.0580 0.8278
+GRU 0.0192 0.9786 0.0264 0.9712 0.0408 0.9351 - -
 GTS 0.0180 0.9898 0.0260 0.9824 0.0333 0.9701 0.0442 0.9518
-GWN 0.0251 0.9740 0.0300 0.9640 0.0381 0.951 0.0486 0.9294
+GWN 0.0251 0.9740 0.0300 0.9640 0.0381 0.9510 0.0486 0.9294
 Informer 0.0882 0.9563 0.1081 0.9321 0.1301 0.8911 0.1521 0.8021
@@
 SDGL [art_78] 0.0180 0.9808 0.0249 0.9730 0.0342 0.9583 0.0455 0.9402
+StemGNN 0.0506 0.8871 0.0674 0.8703 0.0676 0.8499 0.0685 0.8738
 TDG4-MSF [art_81] 0.0172 0.9825 0.0244 0.9718 0.0330 0.9569 0.0433 0.9386
@@
 VAR-MLP 0.0265 0.8609 0.0394 0.8725 0.0407 0.8280 0.0578 0.7675
-Also in this case, GNN models are the best overall.
-The discrete graph structure learning for time series model (GTS), proposed in 2021 by Shang et al.
+Also in this case, GNN models are the best.
+The so-called discrete graph structure learning for time series model (GTS), proposed in 2021 by Shang et al.
 in Ref.
@@
 It is therefore a suitable benchmark for comparison.
+The GTS model was introduced by Shang et al.
+in 2021 for traffic and sensor network forecasting.
+It includes a probabilistic graph learning module that optimizes the mean performance over the graph distribution, and a recurrent graph convolutional module to generate the forecasts.
 Tab.
@@
 Metrics RRSE CORR RRSE CORR RRSE CORR RRSE CORR
-ADLGNN [art_83] 0.1708 0.9866 0.2188 0.9768 0.2897 0.9551 0.4128 0.906
+ADLGNN [art_83] 0.1708 0.9866 0.2188 0.9768 0.2897 0.9551 0.4128 0.9060
 AGCRN 0.1840 0.9841 0.2432 0.9708 0.3185 0.9487 0.4141 0.9087
@@
 ARIMA 0.2328 0.9739 0.3413 0.9402 0.4531 0.8886 0.5810 0.8133
+ASTCRF [art_384] 0.1776 0.9855 0.2347 0.9743 0.3107 0.9503 - -
 Autoformer 0.1935 0.9784 0.2604 0.9651 0.3959 0.9111 0.6064 0.818
 Crossformer 0.1758 0.9801 0.2304 0.9679 0.3101 0.9437 0.4115 0.9001
+DPGNN [art_405] 0.1794 0.9850 0.2364 0.9724 0.3003 0.9543 0.4091 0.9117
 DSTGN [art_73] 0.1787 0.9867 0.2358 0.9721 0.3079 0.9513 - -
@@
 GP 0.2259 0.9751 0.3286 0.9448 0.5200 0.8518 0.7973 0.5971
+GRU 0.1932 0.9823 0.2628 0.9675 0.4163 0.9150 - -
 GTS 0.1842 0.9842 0.2691 0.9645 0.3259 0.9481 0.4796 0.8678
@@
 Informer 0.2134 0.9715 0.2701 0.9549 0.4331 0.8985 0.7017 0.7921
-LSTNet 0.1843 0.9843 0.2559 0.969 0.3254 0.9467 0.4643 0.887
+LSTNet 0.1843 0.9843 0.2559 0.9690 0.3254 0.9467 0.4643 0.887
 MAGNN [art_80] 0.1771 0.9853 0.2361 0.9724 0.3015 0.9539 0.4108 0.9097
@@
 SDLGNN-Corr 0.1806 0.9848 0.2378 0.9722 0.3042 0.9534 0.4173 0.9067
+StemGNN 0.1839 0.9841 0.2612 0.9679 0.3564 0.9395 0.4768 0.8740
 STG-NCDE 0.2346 0.9748 0.2908 0.9605 0.5149 0.8639 - -
@@
 The results indicate that the dynamic spatiotemporal interactive graph neural network (DSTIGNN) model [art_74] and the static and dynamic graph learning network (SDGL) model [art_78] are the most accurate on the solar energy dataset, for shorter and longer time horizons, respectively.
+The DSTIGNN model, introduced by Gao et al.
+in 2021, employs a hierarchical structure consisting of a spatio-temporal interactive learning module, a dynamic graph inference module, multiple spatio-temporal layers, and an output module.
+The spatio-temporal interactive learning module splits the time series into two subsequences to capture both temporal and spatial correlations.
+Then, the graph inference module learns a dynamic adjacency matrix for the graph.
+Once the graph structure is determined, a stack of graph convolutional layers, enhanced with residual connections, along with a convolutional output module, generate the final forecast.
+The description of the SDGL instead has already been provided in this subsection, in the discussion of the results on the Electricity consumption dataset.
 Finally, Tab.
@@
 TDG4-MSF [art_81] 0.4029 0.9014 0.4196 0.8925 0.4294 0.8864 0.4366 0.8834
-Theta 0.6071 0.7768 0.6241 0.7606 0.6271 0.7591 0.6012 0.783
+Theta 0.6071 0.7768 0.6241 0.7606 0.6271 0.7591 0.6012 0.7830
 TPA-LSTM 0.4487 0.8812 0.4658 0.8717 0.4641 0.8717 0.4765 0.8629
@@
 The accuracy of the models varies by horizon and metric.
-However, the best one in terms of RRSE over all horizons is the temporal decomposition enhanced graph neural network for multivariate time series forecasting (TDG4MSF) [art_81].
+However, the best one in terms of RRSE over all horizons is the temporal decomposition enhanced graph neural network for multivariate time series forecasting (TDG4MSF) [art_81], proposed by Miao et al.
+in 2021.
+The TDG4MSF model is composed of four components: temporal decomposition enhanced representation learning for extracting short and long term periodic patterns, graph structure learning, gated graph neural network depending on linear transformation for representation learning, and MLP-based output layer.
 5.8 Other topics
-This final subsection includes 20 papers that fall outside the previously discussed groups.
+This final subsection includes 29 papers that fall outside the previously discussed groups.
 The case studies include cellular traffic prediction, recommendation systems for user preferences, and other forecasting or classification problems.
@@
 5.8.1 Overview
-Since GNNs have recently gained considerable popularity, many recent papers focus on applications that do not relate to the categories discussed so far.
-A frequently discussed topic in this ”Other topics” category is cellular traffic prediction ([art_138], [art_144], [art_148]), in terms of SMS messages, calls, internet connections, phone’s traffic statistics or locations, which are essential for the cellular network resource management system.
-Another topic is modulation classification ([art_12], [art_13], [4]), which is the process of determining the modulation used at the transmitter based on observations of the received signal [modulation_classification_definition].
-A further area of investigation is the study of user preferences ([art_147], [art_151]) to enable personalized experiences and recommendations.
-Other applications discussed in the papers of this section are encrypted traffic classification [art_118], urban spatio-temporal event prediction [art_136], prediction of the quality of operational processes [art_137], seismic intensity forecasting [art_139] (with source code available at https://github.com/StefanBloemheuvel/GCNTimeseriesRegression), prediction of spatio-temporal dynamics with complex structures [art_140], tailing dam monitoring [art_141], forecasting of educational video engagement [art_142], subsurface production forecasting [art_143], time series prediction for data centers maintenance [art_145], hydraulic runoff [art_146], water demand forecasting [art_149], and human activity recognition [art_159] (with source code available at https://github.com/riktimmondal/HAR-Sensor).
+The interest in GNNs has grown so extensively that the selected themes are actually not sufficient to cover all relevant application domains.
+However, since only a few papers would belong to each potential additional theme, they are grouped under the label “Other Topics”.
+A frequently discussed topic in this “Other topics” category is cellular traffic prediction [art_138, art_144, art_148, art_468, art_457], in terms of SMS messages, calls, internet connections, phone’s traffic statistics or locations, which are essential for the cellular network resource management system.
+Another topic included is modulation classification [art_12, art_13, art_303, art_459, art_463], which is the process of determining the modulation used at the transmitter based on observations of the received signal [modulation_classification_definition].
+Two further areas of investigation to be added are the study of user preferences [art_147, art_151] to enable personalized experiences and recommendations, and anomalous IoT intrusion detection [4, art_460].
+Other applications discussed in the papers of this subsection are encrypted traffic classification [art_118], urban spatio-temporal event prediction [art_136], prediction of the quality of operational processes [art_137], seismic intensity forecasting [art_139] (with source code available at https://github.com/StefanBloemheuvel/GCNTimeseriesRegression), prediction of spatio-temporal dynamics with complex structures [art_140], tailing dam monitoring [art_141], forecasting of educational video engagement [art_142], subsurface production forecasting [art_143], time series prediction for data centers maintenance [art_145], hydraulic runoff [art_146], water demand forecasting [art_149], human activity recognition [art_159] (with source code available at https://github.com/riktimmondal/HAR-Sensor ), crime prediction [art_465] (with source code available at https://github.com/ZJUDataIntelligence/HDM-GNN), polypropylene production data [art_455], and tungsten flotation process [art_458] (with source code available at https://github.com/ywmcsu/SASGNN).
 5.8.2 Datasets
-Among the papers selected here, there are only three shared datasets, two for modulation classification and one for cellular traffic prediction.
+Among the papers selected here, there are only four shared datasets, two for modulation classification, one for cellular traffic prediction, and one for botnet detection in IoT network traffic.
 However, the datasets are handled in different ways, either in terms of pre-processing techniques, selected time windows, or train/test split percentages.
-For this reason, the results of these papers are not directly comparable among each other.
-The two datasets for modulation classification used by Refs.
-[art_12], [art_13] and [4] are RML2016.10a (https://www.kaggle.com/datasets/raindrops12/rml201610a, source paper: http://dx.doi.org/10.1007/978-3-319-44188-7_16) and RML2016.10b (https://www.kaggle.com/datasets/marwanabudeeb/rml201610b, source paper: https://pubs.gnuradio.org/index.php/grcon/article/view/11).
+For this reason, the results of these papers are not directly comparable to each other.
+The two datasets for modulation classification used in Refs.
+[art_12, art_13, art_303, art_463] are RML2016.10a (https://www.kaggle.com/datasets/raindrops12/rml201610a, source paper: http://dx.doi.org/10.1007/978-3-319-44188-7_16) and RML2016.10b (https://www.kaggle.com/datasets/marwanabudeeb/rml201610b, source paper: https://pubs.gnuradio.org/index.php/grcon/article/view/11).
 They include 220,000 and 1.2 million samples respectively, in many modulation types.
@@
 As for the cellular traffic prediction dataset used in Refs.
-[art_138] and [art_144], it was proposed by Telecom Italia and MIT Media Lab during the ”Telecom Italia Big Data Challenge”, and it is available at https://dataverse.harvard.edu/dataverse/bigdatachallenge.
+[art_138, art_144, art_468], it was proposed by Telecom Italia and MIT Media Lab during the “Telecom Italia Big Data Challenge”, and it is available at https://dataverse.harvard.edu/dataverse/bigdatachallenge.
 The investigated geographical area is divided in a grid, and each cell records the number of SMS messages, calls, and wireless network traffic data in a 10 minutes interval.
+Finally, the Bot-IoT dataset used in Refs.
+[4, art_460] and available at https://www.unsw.adfa.edu.au/unsw-canberra-cyber/cybersecurity/ADFA-NB15-Datasets/bot_iot.php, was created by designing a realistic network environment in the Cyber Range Lab of UNSW Canberra, and incorporates a combination of normal and botnet traffic.
 6 Discussion
-The aim of this SLR is to provide a comprehensive and detailed overview of the use of spatio-temporal GNN models for time series classification and forecasting in various fields, and to assess their performance in different application domains.
-Existing literature has been collected and synthesized in order to present an overview of datasets, models, and tables of results to support researchers in their future work.
-In the Introduction two sets of research questions were defined: a generic set and a specific set.
-Now, the following discussion will answer those research questions.
-Actually, most of the general research questions were already addressed in the overview presented in Sec.
+As stated in the Introduction, the aim of this SLR is to provide a comprehensive and detailed overview of spatio-temporal GNN models for time series classification and forecasting in different fields, and to assess their performance.
+Existing literature was collected and synthesized to provide an overview of datasets, models, and tables of results that will hopefully assist researchers in their future work.
+In the Introduction two sets of research questions were stated: a generic set and a specific set.
+This section will collect the answers to those research questions.
+Most of the general overview questions (GQs) were addressed in the overview presented in Sec.
 3.
 As for the tools, the GNN community mostly uses Python and PyTorch, although some researchers use TensorFlow.
-Unfortunately, only few papers provide a link to the source code of the proposed model.
+Unfortunately only few papers provide a link to the source code of the proposed model.
 Almost all papers are funded by public or private entities.
-As for the specific questions, the answers are the following.
+As for the specific questions (SQs), the answers are the following.
 SQ1 (Applications).
-The three most investigated fields are, in order, ”Mobility”, ”Environment”, and ”Generic”.
-Regarding the first two groups, their datasets can be naturally translated into graphs, which explains why they have been studied more extensively over time.
-In contrast, the ”Generic” group has recently gained a significant interest for the applicability to broader contexts.
-As for the differences between the various fields, the main ones are related to the definition of the graph, which is not always explicit, and the benchmarks used, often related to the mindset of specific communities.
+The three most investigated fields are “Mobility”, “Environment”, and “Generic”.
+Regarding the first two thematic groups, their datasets can be naturally translated into graphs, which explains why they have been studied more extensively over time.
+In contrast, the “Generic” thematic group has only recently gained a significant interest.
+As for the differences among the thematic groups, the main ones are related to the definition of the graph, which is not always explicit, and the benchmarks used, often related to the mindset of specific communities.
 It is not easy to compare results across different applications and to determine the most promising fields.
-The studies in the ”Generic” group appear promising across a wide range of domains.
+The studies in the “Generic” thematic group appear promising across a wide range of domains.
 SQ2 (Graph construction).
@@
 As observed in Ref.
-[survey_spatiotemporal], there are two main approaches to the design of spatio-temporal GNNs: one that treats the spatial and temporal components in separate modules, and one that integrates and processes them together.
+[survey_spatiotemporal], there are two main approaches for the design of spatio-temporal GNNs: one that treats the spatial and temporal components in separate modules, and another that integrates and processes them together.
 The analysis of the collected papers reveals that the most common approach in literature is that of separate modules.
-Specifically, researchers often address the spatial and temporal aspects of the problem independently from each other, and focus on each module separately.
-The proposed taxonomy of GNN models refers only to the spatial component, and the review indicates that the convolutional and the attentional approaches are the prevalent ones.
-Approximately 62% of the models are purely convolutional, 25% are purely attentional, and 8% are hybrid convolutional-attentional.
+Specifically, researchers very often address the spatial and temporal aspects of the problem independently from each other, and focus on each module separately.
+Hence, the proposed taxonomy of GNN models takes its main perspective from the point of view of the spatial component.
+The review indicates that the convolutional and the attentional approaches are the prevalent ones.
+Approximately 62% of the models are purely convolutional, 21% are purely attentional, and 9% are hybrid convolutional-attentional.
 With regard to the temporal component, the recurrent structures of GRUs and the attention mechanism are widely used.
@@
 As for the benchmark models, there are many options available, and their choice depends on the specific application.
-In fields such as energy, Finance, health, and predictive monitoring, the focus tends to be on simpler statistical and classical machine learning benchmarks, with only a limited use of GNN benchmarks.
-In the ”Generic” group, many recent machine learning benchmark models based on the Transformer architecture have emerged.
-Notably, in ”Mobility”, ”Environment”, and ”Generic” groups (which are also the most investigated fields), there are many reference GNN benchmark models.
-The most prevalent benchmarks are the graph convolutional neural network with long-short term memory (GC-LSTM) [GC-LSTM], spatio-temporal graph convolutional network (STGCN) [STGCN], attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN], diffusion convolutional recurrent neural network (DCRNN) [DCRNN], spatio-temporal graph convolutional network (STGCN) [STGCN], and multivariate time series forecasting with graph neural network (MTGNN) [MTGNN].
+In thematic groups such as “Energy”, “Finance”, “Health”, and “Predictive monitoring”, the focus tends to be on simpler statistical and non-GNN machine learning benchmarks, with only a limited use of GNN benchmarks.
+In the “Generic” group, many recent machine learning benchmark models based on the Transformer architecture have emerged.
+Notably, in “Mobility”, “Environment”, and “Generic” groups (which are also the most investigated fields), there are many reference GNN benchmark models.
+Among the most prevalent benchmarks there are: the graph convolutional neural network with long-short term memory (GC-LSTM) [GC-LSTM], spatio-temporal graph convolutional network (STGCN) [STGCN], attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN], diffusion convolutional recurrent neural network (DCRNN) [DCRNN], spatio-temporal graph convolutional network (STGCN) [STGCN], and multivariate time series forecasting with graph neural network (MTGNN) [MTGNN].
 SQ5 (Benchmark datasets).
 The datasets mentioned in the selected papers are closely related to the specific case of study.
-Even though some benchmark datasets are listed in the ”Generic” group, there is no common standard dataset for the entire GNN research community.
-It would be beneficial to agreeing on, and start adopting some of the most commonly used datasets such as the traffic ones, with the goal of developing a shared benchmark dataset for the whole research community.
+Even though some benchmark datasets are listed in the “Generic” thematic group, there is no common standard dataset for the entire spatio-temporal GNN research community.
+Hence, it would be beneficial for the community to agree on, and start adopting, some of the most commonly used datasets, with the goal of developing a shared benchmark dataset.
+A possible common benchmark could be the traffic one.
 Hopefully, this review could serve this purpose by providing comprehensive tables of the results of the models on different datasets to facilitate comparison between spatio-temporal GNN models.
 SQ6 (Modeling paradigms).
-As for the modeling paradigm, most of the selected papers work with a homogeneous graph, which models the relationships among multiple entities of the same nature.
+As for the modeling paradigm used in each paper, most of the selected papers work with a homogeneous graph, which models the relationships among multiple entities of the same nature.
 This is due to the fact that many basic GNN algorithms were originally developed for homogeneous graphs, and because it is easier to identify relationships between quantities of the same type.
@@
 SQ7 (Metrics).
-Regarding the error metrics, their choice is highly dependent on the specific case study.
+Regarding the error metrics used in the papers, their choice is highly dependent on the specific case study.
 However, the most common ones are mean absolute error (MAE), mean squared error (MSE), and mean absolute percentage error (MAPE) for forecasting problems, and accuracy for classification problems.
+A meta-analysis of papers from different areas indicates that some thematic groups have seen more intense research on GNN models than others.
+For instance, the papers in the “Mobility” and the “Generic” groups present more complex models than others, this being likely due to the more extensive research conducted in these fields, as also evidenced by the larger number of published papers.
+In contrast, the “Finance” and “Health” thematic groups have less papers, and often have only a limited number of benchmarks available for comparison.
+Another notable example is the “Energy” group, which has a relatively large number of papers, particularly in 2024.
+However, the energy community tends to align more closely with traditional econometrics or non-GNN machine learning models, and the proposed GNN models in this field are generally less complex than those found in other groups.
+This suggests that some fields are currently more prone to the development of innovative and complex models, whereas others require a greater focus on adapting existing models to real-world problems.
+Furthermore, papers in the “Energy” group often test their models on proprietary dataset, which further complicates the development of models across the entire group.
+Consequently, we expect some form of cross-pollination between the communities, which is likely to reduce these differences over time.
+A final aspect to highlight is the importance of reproducibility and comparability of these studies, which is essential for the advancement of scientific knowledge.
+In some fields, a greater attention to reproducibility can be observed, also thanks to the availability of datasets, a larger number of benchmarks, and more detailed descriptions of the models.
+Tab.
+40 shows the average number of datasets and benchmark models used, as well as the percentage of papers reporting the source code for each group.
+Table 40: Average number of datasets and benchmark models used, and percentage of papers with the link to the source code for each group.
+Group Datasets (avg.) Benchmarks (avg.) Code availability
+Energy 1.6 5.8 3%
+Environment 1.3 6.5 26%
+Finance 1.8 5.6 12%
+Generic 4.2 10.0 41%
+Health 1.7 7.7 53%
+Mobility 2.6 8.4 27%
+Other 2.1 5.6 14%
+Predictive monitoring 2.1 8.1 14%
 7 Limits, challenges, and future research directions
@@
 To alleviate this problem, this review presented all the information gathered from the selected papers, including datasets, benchmarks, codes, and tables of results.
-It is intended to serve as a detailed overview and a foundation for further exploration, in the hope that researchers will begin to examine the data collected in the presented tables and, over time, identify relevant datasets and benchmark models.
+All that is intended to serve as a detailed overview and a foundation for further exploration, in the hope that researchers will begin to examine the data collected in the presented tables and, over time, identify relevant datasets and benchmark models.
+Moreover, we encourage the authors to discuss the relationship between complexity and performance of the proposed models to facilitate a more comprehensive comparison.
 Reproducibility.
@@
 Moreover, in many papers the authors do not give detailed information about the model and the graph structure, like definition of the graph, number of nodes, calculation of edge weights.
-This issues complicate the verification of results and the assessment of the reproducibility of experiments, which are essential to drive further research in this field.
-Without access to these fundamental resources, it is extremely difficult to assess the accuracy and significance of the proposed models.
-Explainability.
-A crucial aspect of GNN models is explainability, intended as the ability to interpret and understand their decision-making processes.
-Despite their attempt to explicitly model spatial relationships between series, GNNs are often considered as ”black box” models.
-None of the selected papers adequately tackle the concept of explainability.
-However, recent literature is beginning to investigate this important topic.
-For instance, few papers studied perturbation-based explanation methods, which however have proven to be somewhat ineffective as reported in Ref.
-[interpret_3].
-Other approaches, such as GNNExplainer [interpret_2], aim to identify a small subset of node features that have a crucial role in GNN’s predictions.
-Recently, some researchers have started to explore graph counterfactuals as a means of generating explanations, as reviewed in Ref.
-[interpret_1].
+This issue complicates the verification of results and the assessment of the reproducibility of experiments, which are essential to drive further research in this field.
+Without access to this fundamental information, it is extremely difficult to assess the accuracy and significance of the proposed models.
 Poor information capacity.
-Another limitation of GNNs lies in the poor information behind the design of the graph structure and the model definition.
+Another limitation of GNNs lies in the limited amount of information used in the design of the graph structure and the model definition.
 A poorly constructed graph with unrepresentative or overly sparse connections, as well as a lack of physical constraints in the equations of the model, can significantly affect the performance.
@@
 Therefore, it is essential to develop new techniques to overcome these limitations.
-Heterogeneity.
-The majority of current GNN models are designed to deal with homogeneous graphs, where nodes and edges are all of the same type.
-As a result, it is challenging to use these GNNs with heterogeneous graphs, which have different types of nodes and edges or different inputs.
-Therefore, further research is necessary to develop GNN models that can effectively capture the interactions between various types of nodes and edges.
 Scalability.
 Spatio-temporal GNNs are widely used for modeling and analyzing large and complex time series-based network structures.
-However, GNN models often require a significant amount of memory to compute the adjacency matrix and node embeddings, especially in the case of dynamic graph structures.
-This scalability challenge results in high computational costs, and in many cases the necessity to use significant GPU resources.
+However, GNN models often scale badly, and require a significant amount of memory and computing resources to calculate the adjacency matrix and node embeddings, especially in the case of dynamic graph structures.
+This challenge can result in high computational costs, and in many cases in the necessity of using intensively GPU devices.
 Therefore, the development of more scalable GNN models is crucial to facilitate time series analysis even in environments with limited computing power.
+Heterogeneity.
+The majority of current GNN models are designed to deal with homogeneous graphs, where nodes and edges are all of the same nature.
+As a result, it would be challenging to use these GNNs with heterogeneous graphs which have different types of nodes and edges or different inputs.
+Therefore, further research is necessary to develop GNN models that can effectively capture the interactions between various types of nodes and edges.
+Explainability.
+A crucial aspect for GNN models is explainability, intended as the ability to interpret and understand the decision-making processes implemented in them.
+Despite their attempt to explicitly model spatial relationships between series, GNNs are often considered as “black box” models.
+Remarkably, none of the selected papers adequately tackle the concept of explainability.
+Some models are claimed to be explainable because of the integration of non-linear physical laws (e.g., Ref.
+[art_401]), while other papers claim explainability on the basis of the possibility to visualize a graph (e.g., Ref.
+[art_360]) or because they present a feature aggregation form that can be intuitively understood.
+In addition, other approaches focus on determining feature importance through tools such as Captum [Captum], which enables the interpretation and understanding of PyTorch model predictions by exploring the features that contribute to a specific outcome (e.g., Ref.
+[art_361]).
+This highlights a research gap in this area.
+However, recent literature is beginning to investigate this important topic.
+For instance, few papers studied perturbation-based explanation methods, which however have proven to be somewhat ineffective, as reported in Ref.
+[interpret_3].
+Other approaches, such as GNNExplainer [interpret_2], aim to identify a small subset of node features that have a crucial role in GNN’s predictions, as in Ref.
+[art_349, art_357].
+Recently, some researchers have started to explore graph counterfactuals as a means of generating explanations, as reviewed in Ref.
+[interpret_1].
+Quasi-equilibrium learning.
+As pointed out in Subsec.
+4.2, GNNs were originally introduced in Ref.
+[Scarselli2009] by promoting an extension of the backpropagation algorithm which exploits a fixed-point principle.
+The subsequent introduction of the notion of convolutional graph has fundamentally changed such computational scheme.
+It is worth mentioning that in contemporary GNNs there is no need to wait for the system to reach equilibrium.
+For example, the deep neural network scheme no longer requires weight-sharing.
+These improvements have led to successful applications in many real-world problems, where modern features contribute significantly to both accuracy and computational efficiency.
+However, when analyzing the processing of temporal information, GNNs are often implemented within a single graph domain where a temporal function is defined.
+This is a case where the original computational approach proposed in Ref.
+[Scarselli2009] seems to be more adequate, especially when considering that the rate of weight updates is generally much slower than the rate of external temporal signals.
+This suggests the necessity of revisiting the process of convergence to the fixed point, and of exploring computational schemes which take place at quasi-equilibrium points.
+We expect that this approach can significantly alleviate the computational burden that is typical of the convergence to fixed-points required for spatio-temporal graph collections.
 8 Conclusions
 This paper presented the results of a SLR on the application of spatio-temporal GNN models to time series classification and forecasting problems in different fields.
-Lately, GNNs have gained significant popularity due to their ability to process graph-structured data.
-This has led in more recent years to the development of spatio-temporal GNNs in the field of time series analysis, due to their ability to model dependencies between variables and across time points.
-This SLR has brought forward on two set of questions: generic questions which can be answered at the level of bibliographic overview, and more specific questions regarding particular aspects of the proposed spatio-temporal GNN models.
-By answering these questions, it has emerged that the majority of the models in the selected papers is characterized by a convolutional spatial aggregation.
-However, several graph attentional models are also emerging.
-In addition, while the majority of the selected papers focus on models with a pre-defined graph structure, an increasing number of studies, particularly in the ”Generic” group, are beginning to develop models that learn the graph structure autonomously.
-A first key point which results from the presented overview is that the current literature on spatio-temporal GNNs is very fragmented, as it was pictorially reported in Fig.
-4.
-This can be attributed to the fact that the involved researchers come from different communities with focus on specific application domains.
-As a result, there seems to be a lack of standardised datasets or benchmarks.
-The objective of this review was also to collect information on datasets, proposed models, links to source codes, benchmarks, and results, in order to provide a foundation for future studies.
-A second major key point resulting from the proposed SLR seems hence to be that there is a need for the GNN research community to work on common datasets and develop comparable and reproducible models.
-This would enhance transparency and make it easier to evaluate advancements in the field and to share them more easily.
+The interest on this subject comes from the fact that lately, GNNs have gained significant popularity due to their ability to process graph-structured data.
+This has led in more recent years to the development of spatio-temporal GNNs and their application to time series analysis, due to their ability to model dependencies between variables and across time points at once.
+This SLR has brought forward two sets of research questions: generic questions which can be answered at the level of a bibliographic overview, and more specific questions regarding particular aspects of the proposed spatio-temporal GNN models.
+The answers to these questions provide a list of interesting considerations.
+It has emerged that the majority of the models in the selected papers is characterized by a convolutional spatial aggregation.
+However, graph attentional models are also emerging as important tools.
+In addition, while the majority of the selected papers focus on models with a pre-defined graph structure, an increasing number of studies, particularly in the “Generic” thematic group, are beginning to develop models that learn the graph structure autonomously.
+Another key point which results from the presented overview is that the current literature on spatio-temporal GNNs is very fragmented, as it was pictorially reported in Fig.
+3.
+This can be attributed to the fact that the involved researchers come from different communities which focus on distinct application domains.
+As a result, there seems to be a lack of standardized datasets or benchmarks in the studies.
+The objective of this review was also to collect information on datasets, proposed models, links to source codes, benchmarks, and results, in order to provide common background information for future studies.
+A further major key point resulting from our review, which we believe should be emphasized in these conclusions, is the need for the spatio-temporal GNN research community to work on common datasets and develop comparable and reproducible models.
+That would enhance transparency, and it would make it easier for all researchers to reciprocally evaluate advancements in the field and to share them more easily.
+Supplementary material
+Link to the GitHub repository with the links to the 281 reviewed papers, organized by groups: https://github.com/FlaGer99/SLR-Spatio-Temporal-GNN.git.
+CRediT author statement
+Flavio Corradini: Project administration, Funding acquisition.
+Flavio Gerosa: Investigation, Writing - Review & Editing, Visualization.
+Marco Gori: Writing - Review & Editing, Supervision.
+Carlo Lucheroni: Writing - Review & Editing, Supervision.
+Marco Piangerelli: Conceptualization, Methodology, Writing - Review & Editing, Supervision.
+Martina Zannotti: Conceptualization, Methodology, Validation, Investigation, Writing - Original Draft, Writing - Review & Editing, Visualization.
+Acknowledgments
+This work has been funded by the European Union - NextGenerationEU, Mission 4, Component 2, under the Italian Ministry of University and Research (MUR) National Innovation Ecosystem grant ECS00000041 - VITALITY - CUPJ13C22000430001; and the NextGenerationEU - National Recovery and Resilience Plan (PNRR) grant CUP J11J23001450006, with the contribution of Syeew srl.
 Appendix A List of selected journal papers
 Tab.
-LABEL:t:selected_papers lists all the journal papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (e.g., classification or forecasting).
-Table 35: List of selected journal papers, with year, group they belong to, case study, and nature of the task.
+LABEL:t:selected_papers lists all the journal papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (classification or forecasting).
+Table 41: List of the 281 selected journal papers, with year, group they belong to, case study, and nature of the task.
 Ref. Year Group Case study Task
@@
 [art_24] 2020 Energy Wind power Forecasting
-[art_25] 2022 Energy Urban buildings energy consumption Forecasting
+[art_25] 2022 Energy Urban energy consumption Forecasting
 [art_26] 2023 Energy Solar power Forecasting
 [art_27] 2022 Energy Wind power Forecasting
+[art_332] 2024 Energy Wind power Forecasting
+[art_333] 2024 Energy Wind power Forecasting
+[art_334] 2024 Energy Air conditioning control Forecasting
+[art_335] 2024 Energy Air conditioning control Forecasting
+[art_337] 2024 Energy Electrical demand and prices Forecasting
+[art_338] 2024 Energy Electric load Forecasting
+[art_339] 2024 Energy Electric load Forecasting
+[art_340] 2024 Energy Chiller energy consumption Forecasting
+[art_341] 2024 Energy Electricity demand Forecasting
+[art_342] 2024 Energy Electric power anomaly detection Classification
+[art_343] 2024 Energy Electricity demand Forecasting
+[art_344] 2024 Energy Electricity consumption Forecasting
+[art_345] 2024 Energy District heating networks Forecasting
+[art_347] 2024 Energy Power quality disturbances Classification
+[art_348] 2024 Energy Residential load Forecasting
+[art_349] 2024 Energy Photovoltaic energy Forecasting
+[art_350] 2024 Energy Electric vehicle charging stations Forecasting
+[art_351] 2024 Energy Natural gas consumption Forecasting
+[art_353] 2024 Energy Networked microgrids Forecasting
+[art_390] 2024 Energy Multivariate time series Forecasting
 [art_1] 2023 Environment Air quality index Forecasting
@@
 [art_36] 2023 Environment Sea temperature Forecasting
-[art_38] 2022 Environment Frost Classification/forecasting
+[art_38] 2022 Environment Frost Both
 [art_39] 2022 Environment Sea temperature Forecasting
@@
 [art_47] 2023 Environment Sea temperature Forecasting
+[art_354] 2024 Environment PM2.5 prediction Forecasting
+[art_355] 2024 Environment Sea clutter Forecasting
+[art_356] 2024 Environment Earthquakes Forecasting
+[art_357] 2024 Environment Plankton dynamics Forecasting
+[art_359] 2024 Environment Groundwater data Forecasting
+[art_360] 2024 Environment Carbon emission Forecasting
+[art_361] 2024 Environment PM2.5 prediction Forecasting
+[art_363] 2024 Environment Water demand Forecasting
+[art_364] 2024 Environment Water quality and flow rate Forecasting
+[art_365] 2024 Environment Weather Forecasting
+[art_366] 2024 Environment Wind speed scenario Forecasting
+[art_367] 2024 Environment Carbon emission Forecasting
+[art_368] 2024 Environment Air quality Forecasting
+[art_370] 2024 Environment Temperature Forecasting
+[art_371] 2024 Environment Air quality Forecasting
+[art_373] 2024 Environment Greenhouse gas concentrations Forecasting
 [art_59] 2022 Finance Stock prediction Classification
@@
 [art_68] 2023 Finance Sale prediction Forecasting
+[art_374] 2024 Finance Oil and metal market Classification
+[art_375] 2024 Finance Supply chain demand Forecasting
+[art_377] 2024 Finance Stock price Classification
+[art_378] 2024 Finance Stock market volatility Forecasting
+[art_379] 2024 Finance Agricultural price Forecasting
+[art_381] 2024 Finance Cryptocurrencies’ price Forecasting
+[art_382] 2024 Finance Stock trend Classification
 [art_69] 2023 Generic Batch workloads Forecasting
 [art_70] 2023 Generic Node classification for websites Classification
-[art_71] 2023 Generic Clinical risk Classification
+[art_71] 2023 Generic Clinical risk classification Classification
 [art_72] 2023 Generic Sensor network Forecasting
@@
 [art_75] 2023 Generic Multivariate time series Forecasting
-[art_76] 2023 Generic Long-term series forecasting Forecasting
+[art_76] 2023 Generic Long term series forecasting Forecasting
 [art_77] 2023 Generic Multivariate time series Forecasting
@@
 [art_81] 2023 Generic Multivariate time series Forecasting
-[art_82] 2023 Generic Multivariate long sequence time series Forecasting
+[art_82] 2023 Generic Multivariate long sequences Forecasting
 [art_83] 2023 Generic Multivariate time series Forecasting
-[art_84] 2022 Generic Multivariate time series Classification/forecasting
+[art_84] 2022 Generic Multivariate time series Both
 [art_85] 2023 Generic Multivariate time series Forecasting
@@
 [art_87] 2022 Generic Multivariate time series Forecasting
-[art_88] 2023 Generic Multi-channel time series Forecasting
+[art_88] 2023 Generic Multivariate time series Forecasting
 [art_89] 2023 Generic Multivariate time series Forecasting
@@
 [art_160] 2023 Generic Impact of graph construction Forecasting
-[art_202] 2021 Generic Generic Classification/forecasting
+[art_202] 2021 Generic Generic Both
+[art_383] 2024 Generic Multivariate time series Forecasting
+[art_384] 2024 Generic Multivariate time series Forecasting
+[art_385] 2024 Generic Multivariate time series Classification
+[art_386] 2024 Generic Multivariate time series Classification
+[art_387] 2024 Generic Multivariate time series Both
+[art_388] 2024 Generic Multivariate time series Forecasting
+[art_389] 2024 Generic Multivariate time series Classification
+[art_393] 2024 Generic Multivariate time series Forecasting
+[art_395] 2024 Generic Multivariate time series Classification
+[art_396] 2024 Generic Multivariate time series Forecasting
+[art_397] 2024 Generic Multivariate time series Forecasting
+[art_398] 2024 Generic Multivariate time series Forecasting
+[art_400] 2024 Generic Multivariate time series Forecasting
+[art_401] 2024 Generic Multivariate time series Forecasting
+[art_403] 2024 Generic Multivariate time series Both
+[art_404] 2024 Generic Multivariate time series Forecasting
+[art_405] 2024 Generic Multivariate time series Forecasting
+[art_407] 2024 Generic Multivariate time series Classification
+[art_417] 2024 Generic Multivariate time series Forecasting
+[art_464] 2024 Generic Multivariate time series Forecasting
 [art_14] 2023 Health Epilepsy diagnosis Classification
@@
 [art_189] 2023 Health Clinical risk classification Classification
+[art_410] 2024 Health Infectious disease Forecasting
+[art_411] 2024 Health EEG emotion recognition Classification
+[art_412] 2024 Health EEG emotion recognition Classification
+[art_413] 2024 Health Epidemic Forecasting
+[art_414] 2024 Health Epidemic Forecasting
+[art_415] 2024 Health Epidemic Forecasting
+[art_416] 2024 Health Epidemic Forecasting
 [art_93] 2022 Mobility Urban traffic Forecasting
 [art_94] 2021 Mobility Crowd flow Forecasting
-[art_95] 2023 Mobility Flight delays, urban traffic Forecasting
 [art_96] 2022 Mobility Crowd flow Forecasting
@@
 [art_100] 2023 Mobility Urban traffic Forecasting
-[art_101] 2023 Mobility Bus station profiling Profiling
 [art_102] 2021 Mobility Passenger demand Forecasting
 [art_103] 2023 Mobility Urban traffic Forecasting
-[art_104] 2022 Mobility Internet of vehicles Forecasting
+[art_104] 2022 Mobility Internet of vehicles traffic Forecasting
 [art_105] 2020 Mobility Urban traffic Forecasting
 [art_106] 2023 Mobility Urban traffic Forecasting
-[art_107] 2023 Mobility Metro flow Forecasting
 [art_108] 2023 Mobility Urban traffic Forecasting
@@
 [art_135] 2022 Mobility Urban traffic Forecasting
-[art_12] 2023 Other Modulation classification Classification
-[art_13] 2022 Other Modulation classification Classification
-[art_118] 2023 Other Encrypted traffic classification Classification
-[art_136] 2022 Other Spatial temporal events Forecasting
-[art_137] 2023 Other Quality indicators Forecasting
-[art_138] 2023 Other Cellular traffic Forecasting
-[art_139] 2023 Other Seismic data Forecasting
-[art_140] 2023 Other Structural dynamics on irregular domains Forecasting
-[art_141] 2023 Other Tailing dam monitoring Forecasting
-[art_142] 2023 Other Video engagement Forecasting
-[art_143] 2023 Other Subsurface production Forecasting
-[art_144] 2023 Other Cellular traffic Forecasting
-[art_145] 2023 Other Data centers maintenance Forecasting
-[art_146] 2023 Other Hydraulic runoff Forecasting
-[art_147] 2022 Other App popularity Forecasting
-[art_148] 2021 Other Cellular traffic Forecasting
-[art_149] 2022 Other Water demand Forecasting
-[art_151] 2023 Other User preference Edge prediction
-[art_159] 2021 Other Human activity recognition Classification
-[4] 2023 Other Modulation classification Classification
+[art_406] 2024 Mobility Multivariate time series Forecasting
+[art_408] 2024 Mobility Multivariate time series Forecasting
+[art_419] 2024 Mobility Urban traffic Forecasting
+[art_420] 2024 Mobility Urban traffic Forecasting
+[art_421] 2024 Mobility Metro flow traffic Forecasting
+[art_422] 2024 Mobility Urban traffic Forecasting
+[art_424] 2024 Mobility Urban traffic Forecasting
+[art_425] 2024 Mobility Network traffic prediction Forecasting
+[art_427] 2024 Mobility Urban traffic Forecasting
+[art_430] 2024 Mobility Urban traffic Forecasting
+[art_431] 2024 Mobility Urban traffic Forecasting
+[art_432] 2024 Mobility Urban traffic Forecasting
+[art_433] 2024 Mobility Urban traffic Forecasting
+[art_434] 2024 Mobility Airport performances Forecasting
+[art_435] 2024 Mobility Urban traffic Forecasting
+[art_438] 2024 Mobility Urban traffic Forecasting
+[art_439] 2024 Mobility Urban traffic Forecasting
+[art_441] 2024 Mobility Urban traffic Forecasting
+[art_442] 2024 Mobility Urban traffic Forecasting
+[art_443] 2024 Mobility Flight delay prediction Forecasting
+[art_444] 2024 Mobility Urban traffic Forecasting
+[art_445] 2024 Mobility Urban traffic Forecasting
+[art_446] 2024 Mobility Bike sharing demand Forecasting
+[art_447] 2024 Mobility Urban traffic Forecasting
+[art_448] 2024 Mobility Urban traffic Forecasting
+[art_449] 2024 Mobility Human mobility Forecasting
+[art_451] 2024 Mobility Urban traffic Forecasting
+[art_452] 2024 Mobility Bus station Profiling
+[art_453] 2024 Mobility Urban traffic Forecasting
+[art_454] 2024 Mobility Bus stop mobility Forecasting
+[art_12] 2023 Other topics Modulation classification Classification
+[art_13] 2022 Other topics Modulation classification Classification
+[art_118] 2023 Other topics Encrypted traffic classification Classification
+[art_136] 2022 Other topics Spatio-temporal events Forecasting
+[art_137] 2023 Other topics Quality indicators prediction Forecasting
+[art_138] 2023 Other topics Cellular traffic prediction Forecasting
+[art_139] 2023 Other topics Seismic data Forecasting
+[art_140] 2023 Other topics Structural dynamics Forecasting
+[art_141] 2023 Other topics Tailing dam monitoring Forecasting
+[art_142] 2023 Other topics Educational videos engagement Forecasting
+[art_143] 2023 Other topics Subsurface production Forecasting
+[art_144] 2023 Other topics Cellular Network Traffic Forecasting
+[art_145] 2023 Other topics Data centers maintenance Forecasting
+[art_146] 2023 Other topics Hydraulic runoff Forecasting
+[art_147] 2022 Other topics App popularity Forecasting
+[art_148] 2021 Other topics Cellular Network Traffic Forecasting
+[art_149] 2022 Other topics Water demand Forecasting
+[art_151] 2023 Other topics User preference Edge prediction
+[art_159] 2021 Other topics Human activity recognition Classification
+[art_303] 2023 Other topics Modulation classification Classification
+[art_455] 2024 Other topics Polypropylene production data Forecasting
+[4] 2024 Other topics Botnet detection Classification
+[art_457] 2024 Other topics Cellular KPI prediction Forecasting
+[art_458] 2024 Other topics Tungsten floatation plant Forecasting
+[art_459] 2024 Other topics Modulation recognition Classification
+[art_460] 2024 Other topics IoT intrusion detection Classification
+[art_463] 2024 Other topics Modulation recognition Classification
+[art_465] 2024 Other topics Crime prediction Forecasting
+[art_468] 2024 Other topics Mobile traffic prediction Forecasting
 [art_9] 2022 Predictive monitoring Anomaly detection Classification
@@
 [art_192] 2023 Predictive monitoring Fault diagnosis Classification
+[art_461] 2024 Predictive monitoring Motor temperature Forecasting
+[art_469] 2024 Predictive monitoring Fault diagnosis Classification
+[art_470] 2024 Predictive monitoring Root cause localization Forecasting
+[art_471] 2024 Predictive monitoring Mechanical fault diagnosis Classification
+[art_474] 2024 Predictive monitoring Mechanical fault diagnosis Classification
+[art_475] 2024 Predictive monitoring Outlier detection Classification
+[art_477] 2024 Predictive monitoring Anomaly detection Forecasting
+[art_478] 2024 Predictive monitoring Remaining useful life Forecasting
+[art_479] 2024 Predictive monitoring Anomaly detection Classification
+[art_480] 2024 Predictive monitoring Sensor data Forecasting
+[art_481] 2024 Predictive monitoring Anomaly detection Classification
+[art_482] 2024 Predictive monitoring Process factors Forecasting
+[art_483] 2024 Predictive monitoring Anomaly detection Classification
+[art_484] 2024 Predictive monitoring Anomaly detection Classification
+[art_486] 2024 Predictive monitoring Fault diagnosis Classification
+[art_487] 2024 Predictive monitoring Fault diagnosis Classification
+[art_488] 2024 Predictive monitoring Remaining useful life Forecasting
+[art_489] 2024 Predictive monitoring Anomaly detection Classification
+[art_491] 2024 Predictive monitoring Fault diagnosis Classification

### v4

@@
 The objective of this systematic literature review is hence to provide a comprehensive overview of the various modeling approaches and application domains of GNNs for time series classification and forecasting.
-A database search was conducted, and 281 journal papers were selected for a detailed examination of the current state-of-the-art in the field.
-This examination is intended to offer to the reader a comprehensive collection of proposed models, links to related source code, available datasets, benchmark models, and fitting results.
-All this information is hoped to assist researchers in future studies.
-To the best of our knowledge, this is the first systematic literature review presenting a detailed comparison of results of current spatio-temporal GNN models applied to different domains.
+A database search was conducted, and 366 papers were selected for a detailed examination of the current state-of-the-art in the field.
+This examination is intended to offer to the reader a comprehensive review of proposed models, links to related source code, available datasets, benchmark models, and fitting results.
+All this information is hoped to assist researchers in their studies.
+To the best of our knowledge, this is the first and broadest systematic literature review presenting a detailed comparison of results from current spatio-temporal GNN models applied to different domains.
 In its final part, this review discusses current limitations and challenges in the application of spatio-temporal GNNs, such as comparability, reproducibility, explainability, poor information capacity, and scalability.
+This paper is complemented by a GitHub repository at https://github.com/FlaGer99/SLR-Spatio-Temporal-GNN.git providing additional interactive tools to further explore the presented findings.
+footnotetext: Accepted for publication in Neural Networks
 1 Introduction
-In recent years, graph neural networks (GNNs) have emerged as a powerful class of artificial neural network models aimed at processing data that can be represented as graphs [Zheng2024].
-GNNs are in fact particularly well-suited for a wide range of practical and engineering applications where data naturally lend themselves to be represented as graph structures, such as applications to transportation networks [review_traffico_jiang], image analysis [GNN_image], classification of molecular structure [GNN_molecule], and natural language processing [GNN_NLP].
-The intuitive concept behind the GNN approach is that the nodes of a graph can be put in correspondence to objects or concepts, whereas the edges can be assumed to represent their mutual relationships [Scarselli2009].
+In recent years, graph neural networks (GNNs) have emerged as a powerful class of artificial neural network models aimed at processing data that can be represented as graphs [516].
+GNNs are in fact particularly well-suited for a wide range of practical and engineering applications where data naturally lend themselves to be represented as graph structures, such as applications to transportation networks [168], image analysis [313], classification of molecular structure [280], and natural language processing [284].
+The intuitive concept behind the GNN approach is that the nodes of a graph can be put in correspondence to objects or concepts, whereas the edges can be assumed to represent their mutual relationships [332].
 Data for these quantities can then be processed at once by a GNN model.
-GNNs can be used at three different graph levels, related to three distinct classes of problems: at graph level, at edge level, and at node level [review_traffico_con_codici].
+GNNs can be used at three different levels, related to three distinct classes of problems: at graph-level, at edge-level, and at node-level [521].
 In graph-level problems, the goal is to predict a property of a graph based on its entire structure, rather than on individual nodes or edges.
@@
 As an example, this task can be used in recommendation systems to predict potential connections between users and items on the basis of past interactions.
-In node-level problems, the goal is to predict the role or value of each node within a graph, in order to solve either a classification or a regression task.
+In node-level problems, the goal is to predict the role or value of each node within a graph, in order to solve either a step-wise classification or a regression task.
 Noticeably, regression tasks can sometimes be put in the form of temporal “autoregressions”, that is, they can include time.
+When introduced, GNNs were at first applied to static datasets.
 The inclusion of time in GNN modeling has proven to have far-reaching consequences.
-For this reason, this review examines the available journals’ literature on the specific use of GNNs on time series related tasks, with a particular focus on time series classification and forecasting.
-The idea behind the use of GNNs for time series problems is that GNNs can be made capable of capturing complex relationships, both inter-variable (connections between different variables within a multivariate series) but also inter-temporal (dependencies between different points in time) at once [review_Alippi].
-The inclusion of this capability in GNN models results in a spatio-temporal GNN approach.
+Spatio-temporal GNNs have been applied to a wide range of time series tasks, including forecasting, classification, anomaly detection, and data imputation.
+This review examines the available literature on the use of spatio-temporal GNNs for time series step-wise classification and future data forecasting, as these are the tasks most commonly addressed in the literature and best illustrate the current developments in spatio-temporal GNNs.
+The idea behind the use of GNNs for time series problems is that GNNs can be made capable of capturing complex relationships, both inter-variable (connections between different variables within a multivariate series) but also inter-temporal (dependencies between different points in time) at once [176].
+The inclusion of this capability in GNN models results in a characteristic spatio-temporal approach.
 Within this approach, the spatial dimension is related to the multivariate framework, where nodes typically correspond to different features of a multivariate time series at a given time step $t$, and edges to their relationships, and the temporal dimension is related to the temporal nature of the data.
-Namely, spatio-temporal GNNs are a class of GNN models designed to handle data in both spatial and temporal dimensions at once.
+Spatio-temporal GNNs can thus be seen as a class of GNN models especially designed to handle data in both spatial and temporal dimensions at once.
 In a spatio-temporal graph, data in each node or edge can evolve over time, and the challenge is then to model their related dynamic interactions over time.
+In the broader landscape of time series modeling, a variety of other approaches have been developed alongside spatio-temporal GNNs.
+Early statistical frameworks, such as ARIMA, and, more prominently, neural networks designed for sequential data, including RNNs, LSTMs, and GRUs, have been for a long time the standard solution for modeling temporal data.
+These architectures effectively capture short and medium term time dependencies, but struggle with very long sequences and with explicitly modeling interactions among multiple variables.
+The attention mechanism and the Transformer architecture, originally developed for handling long, static word sequences in Large Language Models (LLMs), were later adopted to better address long term dependencies in time series [494].
+Building on these architectural advances, and supported by the availability of modern GPU hardware, many experiments have been conducted.
+However, it remains unclear whether these highly parallelizable architectures are truly suitable for time series tasks [367].
+The development of LLMs has also stimulated a further line of research in time series modeling.
+LLMs, trained on massive text datasets for capturing general patterns in language, can then be adapted to a variety of downstream task, that is, as foundational models.
+Inspired by this approach, researchers have begun developing time series foundation models (TSFMs), designed to learn general representations from diverse temporal datasets [247].
+These models, which do not necessarily use LLMs, are trained on large collections of time series to enable them to generalize effectively to new sequences.
+All these approaches share the ability to model the dynamics of temporal data, and are inherently sequential models.
+What differentiates spatio-temporal GNNs from conventional machine learning, LLMs applied to dynamics, and TSFMs is how they handle the mutual relationships among features in multivariate time series.
+In TSFMs, these relationships are supposed to be learned end-to-end.
+In contrast, spatio-temporal GNNs explicitly prioritize these relationships.
+In the case of GNN models based on pre-defined static adjacency matrices, the inter-variable relationships are effectively “forced” into the model, like handcrafted features (which TSFMs could not be able to learn).
+As for GNNs using self-learned adjacency matrices, these more recent models are naturally biased toward explicitly identifying these relationships, offering a clearly interpretable representation of inter-dependencies.
 Notwithstanding all the potential interest for such important features, to the best of our knowledge there is a lack of systematic literature reviews (SLRs) on spatio-temporal GNN models for time series applications.
 Most of the existing surveys are not SLRs, or they are not recent, or only focus on a limited number of application domains and specific issues not necessarily related to time series.
-For example, recent reviews focus on algorithmic features [cit_rev2, cit_rev3, survey_Lucheroni, survey_Chikwendu], GNN characterization [cit_rev8, review_Alippi], specific fields [cit_rev4], and distributed training [cit_rev5].
-Only one SLR on spatio-temporal GNNs was actually recently published, by [survey_spatiotemporal].
-However, due to the restrictive search query used in that paper, that SLR covers 52 publications only.
-In addition, [survey_kumar_spatiotemporal_data] proposed a recent survey on spatio-temporal predictive modeling techniques for different domains.
+A much broader view on the subject is nowadays necessary.
+For example, recent reviews focus on algorithmic features [273, 423, 35, 55], GNN characterization [415, 176], specific fields [174, 78], and distributed training [253].
+Only two SLR on spatio-temporal GNNs were actually recently published, in [467] and [78].
+However, Ref.
+[467] uses a restrictive search query that limits the SLR to only 52 publications, whereas Ref.
+[78] examines a broader collection of papers, but it is only focused on traffic flow prediction.
+In addition, [197] is a recent survey on spatio-temporal predictive modeling techniques for different domains.
 However, the authors approach the topic from a more general perspective, analyzing different machine learning models without placing too much emphasis on spatio-temporal GNNs.
@@
 LABEL:t:tabella_altre_review.
-Table 1: Short description of main reviews on spatio-temporal GNN models.
+Table 1: Short description of main reviews on spatio-temporal GNN models. At the time of writing this SLR, articles marked with a $\diamond$ were only available as preprints.
 Review Description
 Wu et al.
-(2020) [cit_rev3] One of the first reviews on the use of GNNs in machine learning and data mining.
+(2020) [423] One of the first reviews on the use of GNNs in machine learning and data mining.
 It introduces a taxonomy that classifies GNN architectures into four categories: recurrent, convolutional, graph autoencoders, and spatial-temporal GNNs.
@@
 Zhou et al.
-(2020) [review_traffico_con_codici] The authors outline a general design pipeline for GNN models, covering key aspects such as the graph structure, pooling techniques, loss function design, and components of the propagation module (including recurrent and convolutional operators, and skip connections).
+(2020) [521] The authors outline a general design pipeline for GNN models, covering key aspects such as the graph structure, pooling techniques, loss function design, and components of the propagation module (including recurrent and convolutional operators, and skip connections).
 They present some applications of GNNs, including physical systems modeling, chemical analysis, text processing, image classification, recommendation systems, and analysis of other types of data.
 Jin et al.
-(2023) [cit_rev4] Review on the applications of spatio-temporal GNNs to predictive learning in urban computing.
+(2023) [174] Review on the applications of spatio-temporal GNNs to predictive learning in urban computing.
 In the specific, the authors describe four methods for the construction of pre-defined graphs: topology-based, distance-based, similarity-based and interaction-based approaches.
@@
 Li et al.
-(2023) [li2023graphneuralnetworkspatiotemporal] The authors aim to provide a systematic review of spatio-temporal mining using GNNs.
+(2023) [236] The authors aim to provide a systematic review of spatio-temporal mining using GNNs.
 After providing an overview of the graph structure and the spatial and temporal components of GNNs, the authors describe some applications in fields such as transportation, disaster management, environment, power systems, and neuroscience.
@@
 Lin et al.
-(2023) [cit_rev5] This review is specialized on distributed training of GNNs.
+(2023) [253] This review is specialized on distributed training of GNNs.
 It investigates optimization techniques used in this context, including parallel mini-batch generation, dynamic mini-batch allocation, mini-batch transmission pipelining, and parallel aggregation with edge partitioning.
 Longa et al.
-(2023) [cit_rev2] The authors offer an overview of GNNs for temporal graphs.
+(2023) [273] The authors offer an overview of GNNs for temporal graphs.
 They describe supervised and unsupervised learning tasks on temporal graphs, and present a taxonomy of temporal GNNs that groups the model in snapshot-based and event-based approaches.
 Chen et al.
-(2024) [survey_Lucheroni] One of the first reviews on the use of GNNs for time series modeling.
+(2024) [35] One of the first reviews on the use of GNNs for time series modeling.
 The authors thoroughly describe a wide range of models for time series classification, forecasting, and anomaly detection, based either on recurrent/convolutional GNNs and attentional GNNs.
@@
 Chikwendu et al.
-(2024) [survey_Chikwendu] This survey provides an overview of graph neural architecture search, and explores methods for representing nodes, edges, and sub-graphs, which add context and semantics to graphs.
+(2024) [55] This survey provides an overview of graph neural architecture search, and explores methods for representing nodes, edges, and sub-graphs, which add context and semantics to graphs.
 The authors also describe some applications, including computer vision, brain networks, recommendation system, biomedical application, natural language processing, and traffic.
 Feng et al.
-(2024) [feng2024comprehensivesurveydynamicgraph] This survey examines 81 dynamic GNN models.
+(2024) [90] This survey examines 81 dynamic GNN models.
 It introduces a novel taxonomy that categorizes GNNs into discrete-time dynamic graphs and continuous-time dynamic graphs based on their structures, features, and dynamic modeling methods.
@@
 Jin et al.
-(2024) [review_Alippi] The authors present a comprehensive review of GNNs for time series analysis, covering forecasting, classification, anomaly detection, and imputation.
+(2024) [176] The authors present a comprehensive review of GNNs for time series analysis, covering forecasting, classification, anomaly detection, and imputation.
 They categorize spatio-temporal GNNs according to their three main components: the spatial module, the temporal module, and the overall model architecture.
@@
 Khemani et al.
-(2024) [Khemani2024] The review discusses general concepts of GNNs, including architectures, techniques, datasets, applications, and challenges, offering a broad analysis that emphasizes these aspects without focusing specifically on time series analysis.
+(2024) [182] The review discusses general concepts of GNNs, including architectures, techniques, datasets, applications, and challenges, offering a broad analysis that emphasizes these aspects without focusing specifically on time series analysis.
 It provides a brief description of papers in the literature, detailing application areas, datasets used, models applied, and performance evaluation.
 Kumar et al.
-(2024) [survey_kumar_spatiotemporal_data] This SLR includes 200 papers from high-impact journals and prominent conferences on spatio-temporal predictive modeling techniques across different domains.
+(2024) [197] This SLR includes 200 papers from high-impact journals and prominent conferences on spatio-temporal predictive modeling techniques across different domains.
 Although its structure is similar to that of our review, its focus is broader and not specifically on spatio-temporal GNNs, as it also covers and discusses statistical and non-GNN machine learning models.
+Zeghina et al.
+(2024) [467] This SLR includes 52 papers on spatio-temporal GNNs across various application domains.
+Compared to our SLR, it covers a limited set of papers due to a restricted search query and it does not provide detailed and comprehensive tables of datasets, benchmarks, and model results.
+Capone et al.
+(2025) [28] This review provides an overview of recent spatio-temporal GNN models and classifies them as time-then-space (TTS), space-then-time (STT), and combined space-and-time (S&T) models.
+Compared to our SLR, this review is not systematic, it includes a limited number of papers, and it does not provide detailed quantitative comparisons of model performance across datasets.
+Fan et al.
+(2025) [78] This SLR covers 86 studies on the use of GNNs for traffic flow prediction.
+It examines data processing, model deployment, graph construction methods, and network architectures, discussing their respective strengths and weaknesses.
+However, unlike our review, its focus is on a single application domain.
 Wu et al.
-(2024) [cit_rev8] This review offers a comprehensive survey of GNN models from the perspective of computation theory.
+(2025) [415] This review offers a comprehensive survey of GNN models from the perspective of computation theory.
 In particular, the authors compare existing GNN models through various perspectives: time, memory, communication, parallelism, and execution.
-Zeghina et al.
-(2024) [survey_spatiotemporal] This SLR includes 52 papers on spatio-temporal GNNs across various application domains.
-Compared to our SLR, it covers a limited set of papers due to a restricted search query and it does not provide detailed and comprehensive tables of datasets, benchmarks, and model results.
-The intention of the SLR presented here is thus to provide a broad and comprehensive overview of the applications of spatio-temporal GNNs for time series classification and forecasting in different areas.
-Necessarily, this SLR also aims to assess if the current popularity of spatio-temporal GNN models is indeed due to their effectiveness, and to evaluate their efficacy and accuracy in different fields.
-Additionally, it collects results and benchmarks used in the discussed papers in order to assist researchers in their work by synthesizing the available literature on GNN applications.
-To the best of our knowledge, this is the first review presenting comprehensive tables with all the results from the various models and benchmarks proposed in a highly fragmented literature.
+While most existing reviews focus on specific subfields, the present work takes a unique broader perspective.
+Its aim is to systematically collect and synthesize a wide range of publications on spatio-temporal GNNs for time series tasks, covering multiple application domains and methodological approaches.
+As a result, the number of articles included in this survey is larger than in more narrowly focused reviews.
+Although this approach results in a longer text, we hope it is able to provide a unique overview of the field, capturing both the diversity of applications and recent developments that cannot be fully appreciated in more specialized surveys.
+The intention of the SLR presented here is therefore to: (i) offer a wide-ranging overview of spatio-temporal GNNs for time series classification and forecasting in different domains, (ii) assess whether their popularity stems from actual effectiveness, (iii) evaluate their efficacy and accuracy in different fields, (iv) collect results and benchmarks from the literature on GNN applications to support researchers in their work.
+In addition, to the best of our knowledge, this is the first review presenting comprehensive tables with all the results from the various models and benchmarks proposed in a highly fragmented literature.
+The tables presented in the paper are also available in the accompanying GitHub repository in an interactive format, enabling users to sort, filter, and explore the results of the various models.
 This synthesis and comparison of the results is performed by gathering data from the existing literature, without conducting any numerical experiment on data.
@@
 During the development of this review, we formulated two sets of questions, one general and one specific, which are outlined below.
-The answers to the first set want to provide a general overview of the publications in the field, whereas the answers to the second set focus on specific aspects of the proposed GNN models.
+The answers to the first set want to provide a general overview of the publications in the field, whereas the answers to the second set focus on specific aspects of the proposed spatio-temporal GNN models.
 The general overview questions (GQs) are:
   1. GQ1)
-    Trend: What is the temporal trend of the number of publications on GNN models? Is there a growing interest in GNNs?
+    Trend: What is the temporal trend of the number of publications on spatio-temporal GNN models? Is there a growing interest in GNNs?
   2. GQ2)
@@
   4. GQ4)
+    Conferences: What are the most prominent conferences where GNN research is presented?
+  5. GQ5)
     Research groups: Which are the most active research groups on GNNs?
-  5. GQ5)
+  6. GQ6)
     Tools: What tools (programming languages, libraries, frameworks) are used to implement the GNN models?
-  6. GQ6)
+  7. GQ7)
     Fundings: Did the authors of the papers receive public or private funding for their research?
@@
 At the same time, a general overview of such GNN models will be assembled by highlighting similarities and differences between the approaches across different fields.
-In case of need of an even more detailed analysis of a specific discipline, or other aspects of GNN models, the reader is invited to consult the specialized reviews mentioned before [cit_rev1, cit_rev2, cit_rev3, survey_Lucheroni, cit_rev8, cit_rev4, cit_rev5, cit_rev6].
+In case of need of an even more detailed analysis of a specific discipline, or other aspects of GNN models, the reader is invited to consult the specialized reviews mentioned before [14, 273, 423, 35, 415, 174, 253, 339, 78, 521, 236, 55, 90, 176, 182, 197, 467, 28].
 The paper is organized as follows.
@@
 8 concludes the review.
-A provides a list of all journal papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (e.g., classification or forecasting).
+Appendix A provides a list of all papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (e.g., classification or forecasting).
 All links mentioned in the papers are included in this review.
 For the reader’s convenience, links that no longer work at the time of submission are marked with a .
+The overall structure of the paper is visually summarized in Fig.
+1.
+Figure 1: Overview of the structure of the SLR.
+To maximize the accessibility and usability of the collected information, the paper is supported by a GitHub repository, available at https://github.com/FlaGer99/SLR-Spatio-Temporal-GNN.git, which complements the review with interactive resources.
+Beyond listing the 366 reviewed papers organized by groups, the repository offers interactive tables for sorting and filtering models by any metric, dynamic plots to visualize performance and trends, and tools to explore top-performing models and most common benchmarks.
+These features make it easy for readers to analyze, compare, and gain deeper insights from the collected data.
+We strongly encourage exploring the repository as an interactive complement to this paper, especially when examining the extensive and complex tables comparing model results, as it allows readers to easily sort and visualize the data.
 2 Methodology for the systematic literature review
@@
 Prior to commencing the literature search, some exclusion criteria were defined.
-First, papers from the current year 2025 were excluded because it has not yet ended, hence the review will only cover the literature up to the end of year 2024.
-Second, only journal articles were considered, whereas both book chapters and conference papers were excluded.
-Such a selection is made for two reasons.
-Whereas it is true that new research directions often emerge from conference papers, which can contain outstanding results, most often the results presented in this form are considered preliminary even by the authors themselves, frequently because felt as lacking the stability required for standard journal publication.
-In addition, the exclusion of conference papers is made for readability reasons, specifically to ensure that a manageable number of articles can be discussed in full within a limited number of pages.
-Nevertheless, conference papers that provide useful benchmarks and which are frequently referenced as accurate benchmark models in other journal articles, are anyway referenced throughout the text.
+First, papers from the current year 2025 were excluded because it has only just begun, hence the review will only cover the literature up to the end of year 2024.
+Second, only journal articles published in Q1 and Q2 journals, as ranked by the Scimago Journal Rank, and conference papers from A and A venues, according to the ICORE conference rankings, were included.
+This selection was made to ensure that the analysis focused exclusively on high-quality, peer-reviewed research published in top-tier venues.
+Additionally, book chapters, workshops, and tutorials were excluded.
 An advanced search query to be submitted to the databases was designed to capture the breadth of the topic while maintaining specificity.
-Its general form is (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting)  [^1] Specific search queries for the different databases: Scopus: TITLE-ABS-KEY((("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting)) AND PUBYEAR < 2025 AND (EXCLUDE(DOCTYPE,"cr") OR EXCLUDE (DOCTYPE,"cp")).
+Its general form is (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting) Specific search queries for the different databases: Scopus: TITLE-ABS-KEY((("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting)) AND PUBYEAR < 2025.
 IEEE Xplore: (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting).
-Filters applied: Journals, 2000 - 2024.
+Filters applied: 2000 - 2024.
 Web of Science: (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting).
-Filters applied: Article, 2000 - 2024.
+Filters applied: 2000 - 2024.
 ACM: [[All: "graph neural network"] OR [All: gnn]] AND [All: "time series"] AND [[All: classification] OR [All: forecasting]] AND [E-Publication Date: (01/01/2000 TO 12/31/2024)].
@@
   2. 2.
-    Duplicate removal: duplicate records were identified and removed.
+    Duplicate removal and venue filtering: duplicate records were removed, along with publications from journals outside Q1–Q2 and conferences not ranked A or A.
   3. 3.
@@
 A graphical representation of this selection process is provided in Fig.
-1.
-A total of 879 records were identified, 355 from Scopus, 67 from IEEE Xplore, 194 from Web of Science, and 263 from ACM.
-After that, 253 duplicates were removed.
-Then, the titles and abstracts of the remaining 626 records were manually evaluated to select the pertinent papers.
-Among these, 333 were discarded because they were considered outside the scope of this review, and 12 were discarded because they were not in English.
-Hence, 281 records were included in the review.
-Figure 1: PRISMA flowchart summarizing the identification and selection process, adapted from [PRISMA_statement].
-The flowchart is structured vertically into three key phases: identification (collection of records from databases and removal of duplicates), screening (assessment for eligibility and exclusion of out-of-scope studies), and inclusion (inclusion of papers in the review).
+2.
+A total of 2663 records were identified, 766 from Scopus, 258 from IEEE Xplore, 292 from Web of Science, and 1347 from ACM.
+After that, 1618 duplicates and publications from non top-tier venues were removed.
+Then, the titles and abstracts of the remaining 1045 records were manually evaluated to select the pertinent papers.
+Among these, 667 were discarded because they were considered outside the scope of this review, and 12 were discarded because they were not in English.
+As a result, 366 records were included in the review, comprising 262 journal papers and 104 conference papers.
+The complete list of selected paper can be found in Appendix A and in the accompanying GitHub repository.
+Next, Sec.
+3 provides an overview of the selected publications.
+Figure 2: PRISMA flowchart summarizing the identification and selection process, adapted from [302].
+The flowchart is structured vertically into three key phases: identification (collection of records from databases, and removal of duplicates and non-top-tier venues), screening (assessment for eligibility and exclusion of out-of-scope studies), and inclusion (inclusion of papers in the review).
 The boxes on the right indicate the number of studies excluded at each stage.
-As for conference papers, a total of 1,765 records were identified across the four included databases.
-Assuming the same proportion of duplicates and articles to be excluded as those found for journal papers, about 600 conference papers would remain, making the total number of papers excessive.
-As previously stated, these conference papers have been excluded from this review.
-Next, Sec.
-3 provides an overview of the selected journal publications.
 3 Overview of the publications
-This section provides a comprehensive overview of the 281 selected journal papers.
+This section provides a comprehensive overview of the 366 selected papers.
 After their selection, the papers were grouped according to their domain of application, as better explained in the thematic analysis of Sec.
@@
 Fig.
-2 displays the number of publications over time across different groups.
-A cumulative positive trend can be observed over time, reflecting an increasing interest in spatio-temporal GNN models across various fields, since 2020.
-The four most common groups that have seen the greatest increase in publications are “Environment”, “Generic”, “Mobility”, and “Predictive monitoring”.
+3 presents a stacked column chart showing the number of journal and conference publications over time, broken down by research group.
+The height of each column represents the total number of publications per year, while the colored segments represent contributions from each group.
+A clear upward trend can be observed over time, especially in journal papers, reflecting an increasing interest in spatio-temporal GNN models across various fields, since 2020.
+The groups that experienced the most significant increase are “Environment”, “Generic”, “Mobility”, and “Predictive monitoring”, among which “Generic” and “Mobility” stand out in terms of both number of publications and the novelty of their contributions.
 For more details, see also the group numerosity reported in Tab.
 2.
-Figure 2: Number of journal publications over time across the different groups.
-Table 2: Pivot table illustrating the number of publications by year and group.
+Figure 3: Number of journal and conference papers over time across the different groups.
+Table 2: Pivot table illustrating the number of publications by year and group (journal and conference papers combined).
 Field / Year 2020 2021 2022 2023 2024 Total
-Energy 1 - 6 6 20 33
-Environment - 2 4 21 16 43
-Finance - - 3 6 7 16
-Generic - 1 4 19 20 44
-Health - 1 2 5 7 15
-Mobility 2 4 15 15 30 66
-Other - 2 4 14 9 29
-Predictive monitoring - 2 3 11 19 35
-Total 3 12 41 97 128 281
-The catalog of the publication sources indicates that, in the studies period, 145 different journals have published articles on GNN models for time series forecasting or classification.
+Energy 1 1 6 5 20 33
+Environment 0 2 4 22 15 43
+Finance 0 0 5 6 3 14
+Health 1 3 5 7 7 23
+Mobility 11 6 24 27 34 102
+Predictive monitoring 0 2 4 12 22 40
+Generic 4 7 11 28 27 77
+Other 0 4 5 14 11 34
+Total 17 25 64 121 139 366
+The catalog of the publication sources indicates that, in the studies period, 131 different Q1 and Q2 journals have published articles on GNN models for time series forecasting or classification.
 Tab.
@@
 IEEE Access 14
-ACM Trans. Knowl. Discov. Data 10
+ACM Transactions on Knowledge Discovery from Data 9
 Knowledge-Based Systems 9
+Applied Intelligence 8
 IEEE Transactions on Instrumentation and Measurement 8
-Applied Intelligence 8
+ACM Transactions on Intelligent Systems and Technology 7
+Expert Systems with Applications 7
 IEEE Transactions on Knowledge and Data Engineering 7
-Expert Systems with Applications 7
-ACM Trans. Intell. Syst. Technol. 7
+Applied Energy 6
+IEEE Transactions on Industrial Informatics 6
+IEEE Transactions on Intelligent Transportation Systems 6
 Sensors 6
-IEEE Transactions on Intelligent Transportation Systems 6
-IEEE Transactions on Industrial Informatics 6
-Applied Energy 6
 The diversity of the journals indicates a vast range of potential applications for spatio-temporal GNN models in different fields.
@@
 Such diversity also underlines the interdisciplinary nature of GNNs research, clearly spanning multiple domains.
-As for the publishers, there are around 40 different publishers, and the ones with more than 5 papers are listed in Tab.
+As for ranked A* and A conference venues, 18 different venues have hosted publications on GNN models for time series forecasting or classification.
+The conference venues with 5 or more papers are listed in Tab.
 4.
-Table 4: Publishers with more than 5 published paper.
-Journal Papers
-Elsevier 82
-Institute of Electrical and Electronics Engineers Inc. 42
-Association for Computing Machinery 27
-Multidisciplinary Digital Publishing Institute (MDPI) 27
-IEEE 18
-IEEE Computer Society 13
-Springer 10
-Springer Science and Business Media 10
-John Wiley and Sons Inc 5
-In order to examine the degree of collaboration among researchers, a co-authorship network was generated by using the VOSviewer software [VOSviewer].
+Table 4: Conference venues with more than 5 published paper.
+Conference venue Papers
+ACM Int. Conf. on Information and Knowledge Management (CIKM) 30
+ACM Int. Conf. on Knowledge Discovery and Data Mining (KDD) 18
+ACM Int. Conf. on Advances in Geographic Information Systems (SIGSPATIAL) 11
+Advances in Neural Information Processing Systems (NeurIPS) 7
+Int. Conf. on Machine Learning (ICML) 5
 Fig.
-3 depicts authors as nodes of the network, and collaborations as connections, revealing some research clusters.
+4 shows a pie chart of how corresponding authors are distributed among different countries.
+This visualization provides an overview of the global distribution of research activities.
+Nearly 70% of the authors are affiliated with institutions based in China.
+Figure 4: Pie chart of the distribution of corresponding authors in different countries.
+In order to examine the degree of collaboration among researchers, a co-authorship network was generated by using the VOSviewer software [373].
+In this network, authors are represented as nodes, and co-authorships as connections, revealing clusters of researchers who frequently collaborate.
 Fig.
-3 shows two of these clusters enlarged.
+5 shows the complete network, while Fig.
+5 provides an enlarged view of two selected clusters.
 (a) Map of the complete collaboration network.
@@
 Figure 5: Network of collaboration between researchers, created using VOSviewer software.
-In Fig.
-3, the color of the nodes is related to the average publication year of an author, so that the evolution of research groups over time can be evaluated.
-Two key observations can be made.
-First, the number of clusters suggests that there are many distinct research groups, indicating a high degree of fragmentation within the community.
-Second, the color of the nodes indicates that some groups were active only at the beginning and have subsequently disappeared, while others have recently emerged, as indicated by their yellow color.
-Some authors have had various collaborations over time, as also shown in Fig.
-3.
-Fig.
-4 shows a pie chart of how corresponding authors are distributed among different countries.
-This visualization provides an overview of the global distribution of research activities.
-More than the 70% of the authors are affiliated with institutions in China.
-Figure 4: Pie chart of the distribution of corresponding authors in different countries.
+In the figures, the color of each node corresponds to the author’s average publication year, allowing the temporal evolution of research groups to be visualized.
+From the complete network in Fig.
+5, two main insights emerge.
+First, the large number of clusters indicates a highly fragmented research community, with many small, relatively independent groups.
+Second, the color of the nodes indicates that some groups were active primarily in earlier years and have since faded, while others have emerged more recently, as indicated by the yellow color.
+Additionally, several authors appear in multiple clusters, reflecting collaborations that span different groups over time, as also shown in Fig.
+5.
+Additional figures and analyses of trends in the collected papers are available on our GitHub repository.
 4 Graph neural networks
@@
 in 2005 in Ref.
-[gori_aggiunto] and by Scarselli et al.
+[107] and by Scarselli et al.
 in 2009 in Ref.
-[Scarselli2009] as an architecture designed to implement a function that maps a graph $\mathcal{G}$ with $n$ nodes into an $m$-dimensional Euclidean space.
+[332] as an architecture designed to implement a function that maps a graph $\mathcal{G}$ with $n$ nodes into an $m$-dimensional Euclidean space.
 Noticeably, this definition was intended in general for graph data, not specifically time-dependent data, which is the focus of this SLR.
 In Ref.
-[Scarselli2009] the target node’s embedding was learned in two steps.
+[332] the target node’s embedding was learned in two steps.
 First, neighbor information is iteratively propagated with fixed weights within the graph until a stable fixed point is reached (whose existence is guaranteed under certain assumptions by Banach’s fixed point theorem).
@@
 in Ref.
-[survey_Lucheroni].
+[35].
 More details about that will be given in Subsec.
@@
 A large part of the research on spatio-temporal GNNs is focused on multivariate time series, as they can be naturally abstracted into spatio-temporal graphs, as depicted in Fig.
-5.
+6.
 In this context, nodes usually correspond to different variables at a given time step, and edges represent relationships between variables.
 Figure 6: Representation of a spatio-temporal graph. Here, each node has a dynamic feature, given by the points of the time series.
-This specific form of spatio-temporal modeling assumes that the feature information of a node depends on both its own historical values and on the historical data from its neighboring nodes [survey_spatiotemporal].
+This specific form of spatio-temporal modeling assumes that the feature information of a node depends on both its own historical values and on the historical data from its neighboring nodes [467].
 Notice that once the the graph structure is given, information is propagated (in “internal time”) within it until the output is obtained, as better explained in the following.
 Our discussion will build up from this perspective.
-When it comes to the development of a spatio-temporal GNN model, there are two primary approaches: one that handles spatial and temporal substructures in separate blocks, and another that integrates and processes these two substructures at once [survey_spatiotemporal].
+When it comes to the development of a spatio-temporal GNN model, there are two primary approaches: one that handles spatial and temporal substructures in separate blocks, and another that integrates and processes these two substructures at once [467].
 These approaches are implemented by means of modules that are purely spatial, purely temporal, or hybrid combinations of the two.
@@
 In the taxonomy of spatio-temporal GNNs used in this review, the focus is exclusively on the description of the spatial module that propagates the information within the nodes at fixed time $t$ (not explicitly included in the following equations for simplicity), which, in the reviewed literature, is a fundamental aspect that differentiates one GNN model from another.
-Based on the mechanism at the basis of the information propagation in the graph at a given time step, spatio-temporal GNNs can be further classified as recurrent GNNs, convolutional GNNs, and attentional GNNs, as discussed in the following.
+Based on the mechanism at the basis of the information propagation in the graph at a given time step, spatio-temporal GNNs can be further classified as recurrent GNNs, convolutional GNNs, and attentional GNNs, as represented in Fig.
+7 and discussed in the following.
 These classes are not mutually exclusive, as some models may fit into hybrid categories.
@@
 However, for the purpose of this review, the taxonomy discussed above will be considered sufficient.
+Figure 7: Taxonomy of the spatial component of spatio-temporal GNNs, categorizing models into recurrent, convolutional, and attentional.
 Recurrent GNNs.
@@
 As detailed in the seminal Ref.
-[Scarselli2009], which introduced recurrent GNNs, in order to ensure the convergence of Eq.
+[332], which introduced recurrent GNNs, in order to ensure the convergence of Eq.
 (2) it is necessary that the iterative function $f$ is a contraction mapping.
@@
 They were introduced by Kipf and Welling in 2017 in Ref.
-[graph_convolution_kipf].
-Unlike recurrent GNNs, which apply the same iterative contraction mapping $f$ until an equilibrium is reached, convolutional GNNs can use different parameters at each updating step, by stacking multiple graph convolutional layers to extract the node embeddings [cit_rev3].
-Noticeably, as the number of convolutional layers increases, the convolutional perceptual field expands, and a more abstract representation of information is thus obtained [art_461].
+[190].
+Unlike recurrent GNNs, which apply the same iterative contraction mapping $f$ until an equilibrium is reached, convolutional GNNs can use different parameters at each updating step, by stacking multiple graph convolutional layers to extract the node embeddings [423].
+Noticeably, as the number of convolutional layers increases, the convolutional perceptual field expands, and a more abstract representation of information is thus obtained [31].
 A basic form for the equation for the $k^{\text{th}}$ layer of a convolutional GNN with $n$ nodes is
 $\begin{gathered}H^{(k)}=f_{k}\left(\tilde{\mathcal{A}}H^{(k-1)}\Theta_{k-1}\right)\>,\quad\tilde{\mathcal{A}}=\hat{\mathcal{D}}^{-\frac{1}{2}}\hat{\mathcal{A}}\hat{\mathcal{D}}^{-\frac{1}{2}}\>,\quad\hat{\mathcal{A}}=\mathcal{A}+I\>,\quad\\ \hat{\mathcal{D}}_{ii}=\sum_{j}\hat{\mathcal{A}}_{ij}\>,\quad\quad H^{(0)}=X\>,\end{gathered}$ (3)
-where $f_{k}$ represents the convolutional operation performed at layer $k\in\{1,\dots,K\}$ (with $K$ typically kept small to prevent issues like over-smoothing [oversmoothing]), $\mathcal{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix, $I\in\mathbb{R}^{n\times n}$ is the identity matrix (used to add self-loops in the adjacency matrix $\mathcal{A}$), $\hat{\mathcal{D}}\in\mathbb{R}^{n\times n}$ is the diagonal degree matrix associated to $\hat{\mathcal{A}}$, $H^{(k)}\in\mathbb{R}^{n\times d_{k}}$ are the $d_{k}$-dimensional node embeddings $h^{(k)}_{v}$ of the entire graph produced by the $k^{\text{th}}$ layer, $X\in\mathbb{R}^{n\times d_{0}}$ is the collection of all node attributes $x_{v}$ of the entire graph, and $\Theta_{k}\in\mathbb{R}^{d_{k-1}\times d_{k}}$ is a trainable weight matrix for the $k^{\text{th}}$ layer [survey_Lucheroni].
+where $f_{k}$ represents the convolutional operation performed at layer $k\in\{1,\dots,K\}$ (with $K$ typically kept small to prevent issues like over-smoothing [105]), $\mathcal{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix, $I\in\mathbb{R}^{n\times n}$ is the identity matrix (used to add self-loops in the adjacency matrix $\mathcal{A}$), $\hat{\mathcal{D}}\in\mathbb{R}^{n\times n}$ is the diagonal degree matrix associated to $\hat{\mathcal{A}}$, $H^{(k)}\in\mathbb{R}^{n\times d_{k}}$ are the $d_{k}$-dimensional node embeddings $h^{(k)}_{v}$ of the entire graph produced by the $k^{\text{th}}$ layer, $X\in\mathbb{R}^{n\times d_{0}}$ is the collection of all node attributes $x_{v}$ of the entire graph, and $\Theta_{k}\in\mathbb{R}^{d_{k-1}\times d_{k}}$ is a trainable weight matrix for the $k^{\text{th}}$ layer [35].
 The mechanism of node embedding generation from Eq.
 (3) is illustrated in Fig.
-6(b), where it is put in contrast with the mechanism of a recurrent GNNs, shown in Fig.
-6(a).
+8(b), where it is put in contrast with the mechanism of a recurrent GNNs, shown in Fig.
+8(a).
 (a) Recurrent GNNs use the same graph recurrent layer in the iterative generation of the node embeddings. Here $\Theta$ represents the weights of the parametric trainable function $f$.
@@
 Attentional GNNs.
-In attentional GNNs, the aggregation process uses the attention mechanism [attention] to combine node features.
+In attentional GNNs, the aggregation process uses the attention mechanism [375] to combine node features.
 The equations to compute the node embedding $h_{u}$ of layer $k+1$ from the embeddings of layer $k$ for each node $u$ are
@@
 (4b) computes a pairwise masked attention score between nodes $u$ and $v$.
-This is done by applying the concatenation operation ||||| | to the vectors $z_{u}^{(k)}$ and $z_{v}^{(k)}$, then calculating a dot product of the concatenation with a learnable weight vector $\vec{a}^{\,(k)}$, and finally applying the LeakyReLU nonlinearity.
+This is done by applying the concatenation operation $||$ to the vectors $z_{u}^{(k)}$ and $z_{v}^{(k)}$, then calculating a dot product of the concatenation with a learnable weight vector $\vec{a}^{\,(k)}$, and finally applying the LeakyReLU nonlinearity.
 The quantity $\epsilon_{uv}$ indicates the importance of the features of node $v$ to node $u$.
@@
 Finally, Eq.
-(4d) computes the embedding $h_{u}^{(k+1)}$ of the next layer by aggregating the information in neighboring nodes, weighted by the attention scores [GAT].
+(4d) computes the embedding $h_{u}^{(k+1)}$ of the next layer by aggregating the information in neighboring nodes, weighted by the attention scores [376].
 Graph attention networks can in addition use a multi-head attention mechanism to stabilize the learning process of self-attention.
-In practice, the node updating operation in graph attention networks is a generalization of the traditional averaging or max-pooling of neighboring nodes, which allows each node to compute a weighted average of its neighbors and identify the most relevant ones [Brody2021HowAA].
+In practice, the node updating operation in graph attention networks is a generalization of the traditional averaging or max-pooling of neighboring nodes, which allows each node to compute a weighted average of its neighbors and identify the most relevant ones [20].
 As anticipated, the taxonomy discussed so far focuses only on the propagation mechanisms of the spatial component.
@@
 This spatial operation cannot directly account for temporal information.
-In contrast, the temporal module, often based on architectures such as LSTM, GRU, or self-attention, focuses on the evolution of data over time, independently from cross-sectional node interactions [art_384].
+In contrast, the temporal module, often based on architectures such as LSTM, GRU, or self-attention, focuses on the evolution of data over time, independently from cross-sectional node interactions.
 When these different modules are stacked together, spatial and temporal information end up to be used simultaneously, although they are processed in separate modules.
@@
 4.4 Determination of the graph structure
-A crucial issue when dealing with spatio-temporal GNNs for time series problems is the determination of the graph structure, i.e., the connectivity of the nodes.
-Namely, some time series datasets come by their own nature with a pre-defined graph structure (e.g., a road network), whereas some others do not.
-When a natural pre-defined graph structure is available, it helps the model to better capture the underlying dynamics of the system.
-When it is not directly available, it must be defined by the user in some way (based on domain knowledge or some metrics), or determined by some algorithm (such as visibility graphs [visibility_Gori, visibility_graph, art_14]), or learned by the model itself.
+A crucial issue when dealing with spatio-temporal GNNs for time series tasks is the determination of the graph structure, i.e., the connectivity of the nodes.
+The type of spatial information captured by the graph can vary widely across domains.
+For instance, in traffic networks, nodes often represent sensors or intersections, and edges correspond to actual roads or physical connections.
+In contrast, in domains such as healthcare or finance, nodes may represent patients, brain regions, or financial instruments, and edges encode correlations, interactions, or other relational information.
+This variability across datasets highlights the importance of carefully designing the graph structure to reflect the spatial relationships relevant to each domain.
+The specific choices made in different application areas are discussed in the corresponding subsections of this review.
+Namely, some time series datasets come by their own nature with a pre-defined graph structure (e.g., a road network for traffic datasets), whereas some others do not.
+When a natural graph structure is available, it helps the model capture the underlying dynamics more effectively.
+When it is not directly available, it must be defined by the user in some way (based on domain knowledge or some metrics), or determined by some algorithm (such as visibility graphs [468, 201, 383]), or learned by the model itself.
 Once the graph structure is determined, it is necessary to define the weighted graph adjacency matrix, which is a generalization of the graph adjacency matrix $\mathcal{A}$ whose entries are given by the edge weights of the graph.
@@
 For references and a more detailed explanation of these and other techniques for constructing pre-defined structures, see Ref.
-[art_160].
-As for automatic graph structure learning techniques, a distinction can be made between models that rely on predefined graph structures (sometimes referred to as supplementary graph structure learning models), and those that learn the graph structure entirely from scratch.
-In the first category, there are models such as Graph WaveNet (GWN) [GWN], which constructs an adaptive matrix as a complement to a predefined graph structure [art_405].
-These techniques still rely on some form of prior assumptions.
-To address this limitation, models that learn the graph structure from scratch have been proposed.
-One of them is the MTGNN model [MTGNN], which extracts uni-directional relationships from node embeddings as it learns, and achieves state-of-the-art results across different datasets.
-It is one of the first models to use self-learned graph structures [art_417].
+[15].
+As for automatic graph structure learning techniques, a distinction can be made between models that rely on pre-defined graph structures (sometimes referred to as supplementary graph structure learning models), and those that learn the graph structure entirely from scratch.
+Two notable examples from Wu et al.
+illustrate these approaches.
+The first example is Graph WaveNet (GWN) [425], which constructs a self-adaptive adjacency matrix $A$ using two randomly initialized learnable node embedding matrices $E_{1}$ and $E_{2}$ as:
+$A=\text{softmax}\left(\text{ReLU}\left(E_{1}\,E_{2}^{\top}\right)\right)\>,$ (5)
+where $E_{1}$ and $E_{2}$ correspond to the source and target node embeddings, respectively.
+Their multiplication represents the spatial dependency weights between source nodes and target nodes.
+In the original work, these learned hidden graph dependencies are combined with a pre-defined graph structure.
+The authors note that when such a structure is unavailable, the self-adaptive adjacency matrix alone can be used to model hidden spatial dependencies.
+As a second example, which overcomes the reliance on a pre-defined graph, the same authors proposed the MTGNN model [424], which extracts uni-directional relationships from node embeddings as follows:
+$M_{1}=\tanh(\eta\,E_{1}\Phi_{1})$ (6)
+$M_{2}=\tanh(\eta\,E_{2}\Phi_{2})$
+$A=\text{ReLU}\left(\tanh\left(\eta\left(M_{1}M_{2}^{\top}-M_{2}M_{1}^{\top}\right)\right)\right)$
+$\text{idx}=\text{argtopk}(A[i,:])\quad\forall i=1,\ldots,n$
+$A[i,-\text{idx}]=0\>,$
+where $E_{1}$ and $E_{2}$ are randomly initialized learnable node embedding matrices, $\Phi_{1}$ and $\Phi_{2}$ are model parameters, $\eta$ is a hyperparameter controlling the saturation rate of the activation function, and $\text{argtopk}(\cdot)$ is a function that returns the index of the top-k largest values of a vector.
+Unlike GWN, which can be used either standalone or to augment a pre-defined graph, MTGNN learns the graph structure fully end-to-end.
+These two models are among the first to explore self-learned graph structures.
+Although other methods have since been proposed, GWN and MTGNN continue to serve as prototypical models for self-learned graph structures.
 When the graph structure is learned, some authors introduce a graph regularization term to the training loss function to optimize the learned graph adjacency matrix.
-This regularization term is typically designed to control the sparsity of the matrix, and to promote graph proximity, encouraging larger entries when the series in the corresponding nodes are “close enough” [art_448].
-Some researchers argue that models which generate graphs without prior knowledge may overlook local conditions when constructing adjacency matrices [art_397].
+This regularization term is typically designed to control the sparsity of the matrix, and to promote graph proximity, encouraging larger entries when the series in the corresponding nodes are “close enough”.
+Some researchers argue that models which generate graphs without prior knowledge may overlook local conditions when constructing adjacency matrices [152].
 Therefore, it is important to conduct further research in this area to identify the most effective techniques for generating informative graphs.
 For a more detailed explanation of graph learning models, see Refs.
-[art_396, art_405].
+[137, 240].
 Another key aspect in the determination of the graph structure is the choice between using a static or dynamic adjacency matrix, i.e.
@@
 In contrast, when graphs are learned by the model, dynamic adjacency matrices are typically employed.
-The use of a dynamic adjacency matrix has the advantage of capturing the dynamic nature of the problem [art_447].
+The use of a dynamic adjacency matrix has the advantage of capturing the dynamic nature of the problem [126].
 However, the learned dynamic graphs often show significant differences between adjacent time steps, whereas real-world graph structures typically evolve smoothly over time.
@@
 An example of this intermediate approach can be found in Ref.
-[art_415], which also provides a comparison between this intermediate approach and the classical ones.
+[285], which also provides a comparison between this intermediate approach and the classical ones.
 In general, there is no rule that can determine in advance the best approach to follow.
@@
 For this reason, we do not explore this aspect further.
+As an example, for a detailed discussion of graph construction techniques in traffic flow prediction models, the reader is referred to Ref.
+[78].
 Each subsection of the thematic analysis in Sec.
@@
 Evaluating the performance of a time series forecasting or classification model is essential to assess its reliability.
-Moreover, experimental results often show that even when datasets involve the same forecasting task, the transferability of model architectures is limited [art_408].
+Moreover, experimental results often show that even when datasets involve the same forecasting task, the transferability of model architectures is limited [32].
 Therefore, a proper evaluation of the model is a crucial aspect.
@@
 Each metric provides its own insight into different aspects of prediction quality, so multiple metrics are typically used for a more comprehensive evaluation.
-In addition to presenting a table of the error metrics computed for the entire graph, some authors provide a spatial visualization of these errors across the various nodes, as for example in Refs.
-[art_28, art_449].
+In addition to presenting a table of the error metrics computed for the entire graph, some authors provide a spatial visualization of these errors across the various nodes, as for example in Ref.
+[437].
 As for forecasting problems, commonly used evaluation metrics include the mean absolute error (MAE), mean squared error (MSE), root mean squared error (RMSE), mean absolute percentage error (MAPE), root relative squared error (RRSE), the coefficient of determination (R2), and the correlation coefficient (CORR).
@@
 As for classification problems, common evaluation metrics include accuracy, precision, recall, and F1 score.
-Given $TP$, $TN$, $FP$, and $FN$ the number of true positives, true negatives, false positives, and false negatives, respectively, the formulas for the binary classification problem are reported in Tab.
+Given $TP$, TNTN, FPFP, and $FN$ the number of true positives, true negatives, false positives, and false negatives, respectively, the formulas for the binary classification problem are reported in Tab.
 6.
+Similar formulas can be applied for the multi-class classification problem.
 Table 6: Evaluation metrics for binary classification tasks and their formulas.
@@
 F1 score $\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
-Similar formulas can be applied for the multi-class classification problem.
 Besides accuracy, another important aspect to consider is the complexity of the model.
@@
 Consequently, it was not possible to provide an accurate and detailed comparison of the relationship between complexity and performance across the different models in this review.
-Although creating a plot to compare accuracy and computational times of the different models would be very useful (as in Ref.
-[art_407]), it is not possible to provide such a comparison here.
+Although creating a plot to compare accuracy and computational times of the different models would be very useful (as in Refs.
+[140] and [81]), it is not possible to provide such a comparison here.
 The authors of the various papers should be encouraged to incorporate such comparisons in their work.
@@
 The goal is to provide an overview of available empirical studies, the most common approaches, and the results obtained concerning the discussed themes.
-In each following subsection, one per theme, a list of the benchmark models that appear in the selected journal papers will be provided.
+In each following subsection, one per theme, a list of the benchmark models that appear in the selected papers will be provided.
 Within each theme, they will be classified as: mathematical and statistical models coming from Econometrics, non-GNN machine learning models, and GNN-based models.
@@
 All results presented in the tables of this SLR are taken directly from the tables of the selected papers, as we did not conduct any further experiments to obtain additional results.
+When examining the results tables, the reader is invited to consult the GitHub repository associated with this review, which provides interactive functionality for sorting and filtering models by any metric and generating plots, making it easy to analyze and compare the results.
 Note that in the tables which list benchmark models we have included references to any GNN model found.
@@
 Further information about the models used as benchmarks can be found either in the benchmark tables or in the papers where they are used as benchmarks.
+The most common benchmarks used across all groups of papers are also listed in the GitHub repository.
 The thematic groups of papers are presented in alphabetical order, with the exception of the “Generic” applications group (which cannot be directly attributed to a specific case study) and the “Other topics” group (focusing on specific problems of other disciplines).
@@
 5.1 Energy
-The first thematic group is “Energy” (with 33 out of 281 papers), which includes all the studies related to energy systems.
+The first thematic group is “Energy” (with 33 out of 366 papers), which includes all the studies related to energy systems.
 In particular, it includes research on electricity load, energy consumption, and power generation forecasting.
@@
 5.2 will refer to the study of environmental data that may or may not be related to energy applications.
-Wind speed forecasting, for instance, is connected to wind power generation due to the cubic relationship between wind speed and wind power [art_29].
+Wind speed forecasting, for instance, is connected to wind power generation due to the cubic relationship between wind speed and wind power [11].
 However, wind speed forecasting may also be of interest for other reasons, and thus it is included in the more general Subsec.
@@
 An overview of the selected papers reveals a relevant interest in the application of GNNs in the energy field starting from 2022.
-The most frequently studied fields include forecasting of wind [art_20, art_24, art_27, art_332, art_333] and solar power [art_19, art_21, art_22, art_26, art_349], which are renewable energy sources characterized by their intermittency.
-The most popular journal among the selected papers is Applied Energy by Elsevier, with 5 published papers.
-The majority of the corresponding authors are affiliated with institutions in China.
+The most frequently studied fields include forecasting of wind [132, 460, 216, 439, 452] and solar power [350, 351, 317, 377], which are renewable energy sources characterized by their intermittency.
+The most popular journal among the selected papers is Applied Energy by Elsevier, with 5 published papers, and the only “Energy” conference paper was presented at IJCAI.
 5.1.2 Datasets
@@
 Dataset Used by Link
-Commercial and Residential Load [art_348] https://www.osti.gov/biblio/1788456
-Danish smart heat meters [art_15] https://doi.org/10.5281/zenodo.6563114
-DKA Solar Centre [art_19] https://dkasolarcentre.com.au/
-IEEE bus systems [art_23] https://doi.org/10.1109/UPEC.2015.7339813 (source paper)
-Individual Household Electric Power Consumption [art_344] https://archive.ics.uci.edu/ml/datasets/individual+household+electric+power+consumption
-JERICHO-E-usage [art_15] https://doi.org/10.6084/m9.figshare.c.5245457.v1
-NREL [art_20], [art_22], [art_24] https://www.nrel.gov/
-OPSD [art_16] https://open-power-system-data.org/
-PJM Hourly Energy Consumption [art_390] https://www.pjm.com/markets-and-operations/data-dictionary
-PVGIS [art_349] http://re.jrc.ec.europa.eu/pvgis/apps4/pvest.php
-PV Switzerland [art_21], [art_22] https://doi.org/10.1109/IJCNN48605.2020.9207573 (source paper)
-Residential heat and electricity demand [art_341] https://ieee-dataport.org/open-access/8-years-hourly-heat-and-electricity-demand-residential-building
-SGCC dataset [art_342] http://www.sgcc.com.cn/
-Southern Power Grid [art_344] https://www.csg.cn/
-UMass Smart* Dataset [art_17] https://traces.cs.umass.edu/docs/traces/smartstar/
-Three of the selected papers [art_20, art_22, art_24] use data from the National Renewable Energy Laboratory (NREL), a national laboratory of the U.S.
+Commercial and residential load [87] https://www.osti.gov/biblio/1788456
+Danish smart heat meters [405] https://doi.org/10.5281/zenodo.6563114
+IEEE bus systems [509] https://doi.org/10.1109/UPEC.2015.7339813 (source paper)
+Individual household electric power consumption [156] https://archive.ics.uci.edu/ml/datasets/individual+household+electric+power+consumption
+JERICHO-E-usage [405] https://doi.org/10.6084/m9.figshare.c.5245457.v1
+NREL [132], [351], [460] https://www.nrel.gov/
+OpenEI residential load [336] https://openei.org/datasets/files/961/pub/RESIDENTIAL_LOAD_DATA_E_PLUS_OUTPUT/HIGH/
+OPSD [486] https://open-power-system-data.org/
+PJM hourly energy consumption [319] https://www.pjm.com/markets-and-operations/data-dictionary
+PVGIS [377] http://re.jrc.ec.europa.eu/pvgis/apps4/pvest.php
+PV Switzerland [350], [351] https://doi.org/10.1109/IJCNN48605.2020.9207573 (source paper)
+Residential heat and electricity demand [288] https://ieee-dataport.org/open-access/8-years-hourly-heat-and-electricity-demand-residential-building
+SGCC dataset [161] http://www.sgcc.com.cn/
+Southern power grid [156] https://www.csg.cn/
+UMass smart* [398] https://traces.cs.umass.edu/docs/traces/smartstar/
+Three of the selected papers [132, 351, 460] use data from the National Renewable Energy Laboratory (NREL), a national laboratory of the U.S.
 Department of Energy (https://www.nrel.gov/).
-Additional datasets were used for exogenous weather variables and satellite imagery (such as Himawari-8, available at https://www.data.jma.go.jp/mscweb/en/himawari89/space_segment/spsg_sample.html in Ref.
-[art_19]), but they are of course highly dependent on the location of the study.
 The data time granularity ranges from 1 minute to 1 day, and the forecasting horizon from tens of minutes to days.
@@
 5.1.3 Types of models
-As for the taxonomy of the proposed GNN models, 17 out of 33 papers use a convolutional GNN, and the other 9 use an attentional GNN.
-Six papers from 2024 use both convolutional and attentional approaches, and one paper [art_338] uses the original recurrent approach on a graph defined via the visibility algorithm.
+As for the taxonomy of the proposed GNN models, 17 out of 33 papers use a convolutional GNN, and 8 use an attentional GNN.
+Seven papers use both convolutional and attentional approaches, and one paper [104] uses the original recurrent approach on a graph defined via the visibility algorithm.
 As to the graph structure used, the papers typically provide only a brief description.
@@
 In some papers, the graph structure reflects the geographic location of the elements (e.g., wind or solar power plants, as in Refs.
-[art_24, art_26, art_349]), whereas in others it is based on the correlation or similarity between time series (as in Ref.
-[art_19]).
+[460, 317, 377]), whereas in others it is based on the correlation or similarity between time series (as in Ref.
+[323]).
 In some studies, a pre-defined structure is imposed and the weighted adjacency matrix is learned directly from the model, often using the attention mechanism as in Ref.
-[art_22].
+[351].
 The majority of papers in this group rely on a static graph.
 The most common loss function used to train these models is the MSE.
-Many papers mention the use of the programming languages Python and Matlab, and few of them mention the use of the Python libraries TensorFlow [2], PyTorch [PyTorch], and PyTorch Geometric [PyTorch_Geometric] for the implementation of the GNN model.
+Many papers mention the use of the programming languages Python and Matlab, and few of them mention the use of the Python libraries TensorFlow [2], PyTorch [309], and PyTorch Geometric [91] for the implementation of the GNN model.
 In this group, only Ref.
-[art_17] provides the link to the source code of the proposed residential load forecasting with multiple correlation temporal graph neural networks (RLF-MGNN) model (available at https://codeocean.com/capsule/9294192/tree/v1).
+[398] provides the link to the source code of the proposed residential load forecasting with multiple correlation temporal graph neural networks (RLF-MGNN) model (available at https://codeocean.com/capsule/9294192/tree/v1).
 5.1.4 Benchmark models
@@
 Category Model Used by
-Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_15], [art_16], [art_22], [art_25], [art_333], [art_337], [art_338], [art_341], [art_343], [art_349]
-Historical average (HA) [art_350], [art_390]
-Vector ARIMA [art_332], [art_333], [art_343], [art_350], [art_390]
-Non-GNN machine learning methods Autoformer [Autoformer] [art_333], [art_341]
-Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_15], [art_19], [art_339]
-Convolutional neural network (CNN) [art_15], [art_18], [art_21], [art_22], [art_343], [art_347]
-Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [art_15], [art_17], [art_351]
-Encoder-decorer architectures (ED) [art_15], [art_19], [art_21], [art_22]
-Feed-forward neural network (FNN) [art_15], [art_18], [art_19], [art_25], [art_27], [art_351]
-Gated recurrent unit (GRU) [art_15], [art_25], [art_334], [art_338], [art_343], [art_347]
-Informer [Informer] [art_15], [art_16]
-$k$-nearest neighbors (KNN) [art_18], [art_24], [art_342], [art_343]
-Long-short term memory network (LSTM) [art_15], [art_16], [art_19], [art_20], [art_23], [art_24], [art_333], [art_334], [art_335], [art_340], [art_341], [art_343], [art_344], [art_347], [art_348], [art_349], [art_353]
-Long-short term time series network (LSTNet) [LSTNet] [art_15], [art_332], [art_334]
-Recurrent neural network (RNN) [art_16], [art_17], [art_19], [art_20]
-Seq2Seq [Seq2seq] [art_335], [art_343]
-Support vector regression (SVR) [art_21], [art_22], [art_24], [art_27], [art_338], [art_342], [art_343], [art_344], [art_348]
-Transformer [attention] [art_16], [art_19], [art_335], [art_343]
-Extreme gradient boosting (XGBoost) [art_25], [art_334], [art_335], [art_341]
-GNN methods Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_349], [art_350], [art_390]
-Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_332], [art_340], [art_341]
-Spatio-temporal autoregressive model (ST-AR) [ST-AR] [art_21], [art_22]
-Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_339], [art_341], [art_348], [art_351], [art_353]
+Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [405], [486], [351], [147], [452], [323], [104], [288], [324], [377], [256]
+Historical average (HA) [500], [319]
+Vector ARIMA [439], [452], [324], [500], [319]
+Non-GNN machine learning methods Autoformer [414] [452], [288]
+Bidirectional long-short term memory network (BiLSTM) [109] [405], [477]
+Convolutional neural network (CNN) [405], [93], [350], [351], [324], [24]
+Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [405], [398], [315]
+Encoder-decoder architectures (ED) [405], [350], [351]
+Feed-forward neural network (FNN) [405], [93], [147], [216], [315], [256]
+Gated recurrent unit (GRU) [405], [147], [531], [104], [324], [24]
+Informer [520] [405], [486]
+$k$-nearest neighbors (KNN) [93], [460], [161], [324]
+Long-short term memory network (LSTM) [405], [486], [132], [509], [460], [452], [531], [148], [265], [288], [324], [156], [24], [87], [377], [463], [256]
+Long-short term time series network (LSTNet) [202] [405], [439], [531]
+Recurrent neural network (RNN) [486], [398], [132]
+Seq2Seq [365] [148], [324]
+Support vector regression (SVR) [350], [351], [460], [216], [104], [161], [324], [156], [87], [256]
+Transformer [375] [486], [148], [324]
+Extreme gradient boosting (XGBoost) [147], [531], [148], [288]
+GNN methods Diffusion convolutional recurrent neural network (DCRNN) [237] [377], [500], [319]
+Multivariate time series forecasting with graph neural network (MTGNN) [424] [439], [265], [288]
+Spatio-temporal autoregressive model (ST-AR) [29] [350], [351]
+Spatio-temporal graph convolutional network (STGCN) [457] [477], [288], [87], [315], [463]
 The majority of the papers only use non-GNN machine learning benchmarks.
@@
 In general, the authors of the papers in this group claim that their proposed GNN model outperforms the benchmarks.
-However, in some cases, it turns out that basic statistical methods may achieve good results as well [art_390].
+However, in some cases, it turns out that basic statistical methods may achieve good results as well [319].
 The most common error metrics used are the MAE, the RMSE, and the Normalized Root mean squared Error (NRMSE).
@@
 An example is provided in Ref.
-[art_349], where the authors also use the GNNExplainer [interpret_2] to assess the influence of different variables.
+[377], where the authors also use the GNNExplainer [455] to assess the influence of different variables.
 It is not possible to compare the accuracy of the models included in this group, because no common benchmark datasets were used.
 The only two papers that study the same dataset are written by the same authors (Refs.
-[art_21, art_22]), and their most recent paper includes a comparison with the models proposed in the first one.
+[350, 351]), and their most recent paper includes a comparison with the models proposed in the first one.
 5.2 Environment
-The thematic group “Environment” includes a relevant number of selected papers (43 out of 281).
+The thematic group “Environment” includes a relevant number of selected papers (43 out of 366).
 This suggests a considerable research interest on the application of GNNs in this field.
@@
 For example, in Refs.
-[art_29] and [art_34] the authors estimate the wind power forecasting error obtained by using wind speed forecasting models.
-As another example, predicting sea surface temperature can be useful for weather forecasting, fishing directions, and disaster warnings [art_28].
+[11] and [12] the authors estimate the wind power forecasting error obtained by using wind speed forecasting models.
+As another example, predicting sea surface temperature can be useful for weather forecasting, fishing directions, and disaster warnings [437].
 Moreover, monitoring and forecasting environmental data are crucial for evaluating the impact of human activities on the environment and tracking the progression of climate change.
@@
 The most investigated data are related to sea temperature (Refs.
-[art_28, art_33, art_35, art_36, art_39, art_41, art_47]), wind speed (Refs.
-[art_29, art_31, art_34, art_40, art_44, art_46, art_366]), and PM2.5 concentration forecasting (Refs.
-[art_2, art_3, art_4, art_6, art_7, art_8, art_354, art_361, art_371]).
+[437, 185, 304, 98, 392, 361, 345, 42]), wind speed (Refs.
+[11, 326, 12, 102, 97, 73, 269]), and PM2.5 concentration forecasting (Refs.
+[267, 357, 223, 293, 183, 310, 248, 325, 426]).
 Almost every selected paper is published in a different journal, with the exception of Energy, Engineering Applications of Artificial Intelligence, Journal of Cleaner Production, Knowledge-Based Systems, Science of the Total Environment by Elsevier, each with two published papers.
-The majority of the corresponding authors are affiliated with institutions in China.
+As for conference papers, the three “Environment” papers were presented at KDD, IJCAI and CIKM.
 5.2.2 Datasets
@@
 Dataset Used by Link
-AgroData-10 [art_28] https://www.aoml.noaa.gov/phod/argo
-Carbon Monitor [art_360] https://carbonmonitor.org/
-CCMP wind data [art_31], [art_40], [art_44] http://data.remss.com
-China Environmental Monitoring Station PM2.5 [art_2], [art_8], [art_371] https://www.cnemc.cn/
-China National Urban Air Quality [art_1], [art_5] https://github.com/Friger/GAGNN
-Copernicus 3D thermohaline data [art_35] https://resources.marine.copernicus.eu/products
-ECMWF East China Sea [art_45] https://github.com/Boltzxuann/GIPMN/tree/dataset
-ECMWF South and East China Sea [art_32] http://apps.ecmwf.int/datasets/data/interim-full-daily
-HadCRUT5 [art_370] https://doi.org/10.1029/2019JD032361 (source paper)
-Jinan PM2.5 [art_354] https://github.com/liaohaibing/MGST_GNN/tree/master
-KnowAir [art_6] https://github.com/shuowang-ai/PM2.5-GNN
-Korean Peninsula sea data [art_33] https://www.nifs.go.kr/risa/main.risa
-NOAA sea datasets [art_28], [art_36], [art_39], [art_46], [art_47], [art_365] https://www.ndbc.noaa.gov/
-Norwegian offshore wind [art_29], [art_34] https://frost.met.no/index.html
-Overbetuwe groundwater data [art_359] https://github.com/mlttac/GroundwaterFlowGNN
-Taiwan air quality dataset [art_3] https://data.gov.tw/en
-UrbanAir [art_6] https://doi.org/10.1016/j.chemosphere.2018.12.128 (source paper)
-US EPA PM2.5 [art_1], [art_4] https://aqs.epa.gov/aqsweb/airdata/download_files.html
-US Geological Survey [art_355] https://earthquake.usgs.gov/earthquakes/search/
-Yangtze River PM2.5 [art_354] https://github.com/liaohaibing/MGST_GNN/tree/master
-Six of the selected papers [art_28, art_36, art_39, art_46, art_47, art_365] use National Oceanic and Atmospheric Administration (NOAA) sea datasets (https://www.noaa.gov/) from the U.S.
+AgroData-10 [437] https://www.aoml.noaa.gov/phod/argo
+Carbon monitor [106] https://carbonmonitor.org/
+CCMP wind data [326], [102], [97] http://data.remss.com
+China environmental monitoring station PM2.5 [267], [310], [426] https://www.cnemc.cn/
+China national urban air quality [41], [72] https://github.com/Friger/GAGNN
+Copernicus 3D thermohaline data [304] https://resources.marine.copernicus.eu/products
+Delaware river basin [42] https://waterdata.usgs.gov/nwis
+ECMWF east China sea [311] https://github.com/Boltzxuann/GIPMN/tree/dataset
+ECMWF south and east China sea [502] http://apps.ecmwf.int/datasets/data/interim-full-daily
+HadCRUT5 [403] https://crudata.uea.ac.uk/cru/data/temperature/
+HUC12 [5] https://hawqs.tamu.edu/
+Jinan PM2.5 [248] https://github.com/liaohaibing/MGST_GNN/tree/master
+KnowAir [293] https://github.com/shuowang-ai/PM2.5-GNN
+Korean peninsula sea data [185] https://www.nifs.go.kr/risa/main.risa
+NARR weather conditions [133] https://psl.noaa.gov/data/gridded/data.narr.html
+NOAA sea datasets [437], [98], [392], [73], [345], [336] https://www.ndbc.noaa.gov/
+Norwegian offshore wind [11], [12] https://frost.met.no/index.html
+Overbetuwe groundwater data [366] https://github.com/mlttac/GroundwaterFlowGNN
+Taiwan air quality dataset [357] https://data.gov.tw/en
+UrbanAir [293] https://doi.org/10.1016/j.chemosphere.2018.12.128 (source paper)
+US EPA PM2.5 [41], [223] https://aqs.epa.gov/aqsweb/airdata/download_files.html
+WeatherBench [133] https://github.com/pangeo-data/WeatherBench
+Yangtze river PM2.5 [248] https://github.com/liaohaibing/MGST_GNN/tree/master
+Six of the selected papers [437, 98, 392, 73, 345, 336] use National Oceanic and Atmospheric Administration (NOAA) sea datasets (https://www.noaa.gov/) from the U.S.
 Department of Commerce (US DOC).
@@
 5.2.3 Types of models
-Regarding the taxonomy of the proposed GNN models, the majority of papers present convolutional GNNs (28 out of 43 papers), followed by 9 attentional GNNs, and by 3 hybrid architectures.
-The description in 3 of the papers (Refs.
-[art_5, art_354, art_361]), however, does not allow us to specify the classification of the model within the proposed taxonomy.
+Regarding the taxonomy of the proposed GNN models, the majority of papers present convolutional GNNs (28 out of 43 papers), followed by 8 attentional GNNs, and by 3 hybrid architectures.
+The description in 4 of the papers (Refs.
+[72, 248, 325, 42]), however, does not allow us to specify the classification of the model within the proposed taxonomy.
 As for the graph structure, not all of the papers describe it with sufficient precision, and some of them are vague about the graph construction and its static or dynamic nature.
@@
 Finally, there are a few papers where more than one graph structure is fused together in order to capture more complex dynamics, as in Ref.
-[art_2].
+[267].
 Almost all the papers in this thematic group deal with multivariate time series, with the exception of Ref.
-[art_370], which studies a univariate time series of global average temperature.
+[403], which studies a univariate time series of global average temperature.
 In this work, the authors treat each time step $t$ as a node $v_{t}$ in the graph, and the adjacency matrix is constructed in such a way that $v_{t}$ is influenced by $v_{t-1}$ and $v_{t+1}$.
@@
 Model Link
-Dynamic Spatial-Temporal Graph Convolutional Recurrent Network (DSTGCRN) [art_360] https://github.com/PalaceTony/DSTGCRN
-Dynamic spatiotemporal graph neural network (DST-GNN) [art_354] https://github.com/liaohaibing/MGST_GNN/tree/master
-Group-aware graph neural network (GAGNN) [art_1] https://github.com/Friger/GAGNN
-Gridded information propagation and mixing network (GIPMN) [art_45] https://github.com/Boltzxuann/GIPMN/tree/dataset
-GNNCoder [art_356] https://doi.org/10.5281/zenodo.13358007
-GroundwaterFlowGNN [art_359] https://github.com/mlttac/GroundwaterFlowGNN
-Hierarchical graph recurrent network (HiGRN) [art_28] https://github.com/Neoyanghc/HiGRN
-Physics-aware graph networks (PaGN) [art_365] https://github.com/eunviho/novel_pagn
-Spatio-temporal FFTransformer (ST-FFTransformer) [art_29] https://github.com/LarsBentsen/FFTransformer
+Domain inspired temporal graph convolution neural network [5] https://github.com/IBM/tgcn-soil-moisture
+Dynamic spatial-temporal graph convolutional recurrent network (DSTGCRN) [106] https://github.com/PalaceTony/DSTGCRN
+Dynamic spatiotemporal graph neural network (DST-GNN) [248] https://github.com/liaohaibing/MGST_GNN/tree/master
+Group-aware graph neural network (GAGNN) [41] https://github.com/Friger/GAGNN
+Gridded information propagation and mixing network (GIPMN) [311] https://github.com/Boltzxuann/GIPMN/tree/dataset
+GroundwaterFlowGNN [366] https://github.com/mlttac/GroundwaterFlowGNN
+Hierarchical graph recurrent network (HiGRN) [437] https://github.com/Neoyanghc/HiGRN
+Physics-aware graph networks (PaGN) [336] https://github.com/eunviho/novel_pagn
+Spatio-temporal FFTransformer (ST-FFTransformer) [11] https://github.com/LarsBentsen/FFTransformer
 5.2.4 Benchmark models
@@
 Category Model Used by
-Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_28], [art_33], [art_34], [art_38], [art_363], [art_370], [art_371]
-Historical average (HA) [art_3], [art_6], [art_28], [art_38]
-Naive [art_29], [art_34], [art_366]
-Vector autoregression (VAR) [art_28], [art_34], [art_38], [art_360]
-Non-GNN machine learning methods Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_33], [art_46]
-Combined FC-LSTM and convolution neural network (CFCC-LSTM) [CFCC-LSTM] [art_35], [art_39], [art_370]
-DLinear [DLinear] [art_31], [art_364]
-Hybrid convolutional neural network and gated recurrent unit (CGRU) [CGRU] [art_31], [art_40], [art_44]
-Convolutional neural network (CNN) [art_30], [art_46]
-Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [art_4], [art_41], [art_368], [art_371]
-Convolutional long-short term memory network (ConvLSTM) [ConvLSTM] [art_28], [art_41], [art_355]
-Feed-forward neural network (FNN) [art_2], [art_7], [art_38], [art_360], [art_365], [art_370]
-Gated recurrent unit encoder–decoder (GED) [GED] [art_36], [art_39], [art_47]
-Gated recurrent unit (GRU) [art_2], [art_6], [art_7], [art_8], [art_38], [art_43], [art_46], [art_361], [art_364], [art_365], [art_367], [art_370]
-Informer [Informer] [art_30], [art_47], [art_354], [art_364], [art_371]
-Long-short term time series network (LSTNet) [LSTNet] [art_355], [art_373]
-Predictive deep convolutional neural network (PDCNN) [PDCNN] [art_31], [art_40]
-Random forest (RF) [art_6], [art_364], [art_370]
-Support vector regression (SVR) [art_3], [art_6], [art_35], [art_36], [art_38], [art_360], [art_370], [art_371]
-Recurrent neural network (RNN) [art_46], [art_365], [art_367]
-Temporal convolutional network (TCN) [TCN] [art_6], [art_30], [art_355], [art_356], [art_368], [art_371], [art_373]
-Temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM] [art_6], [art_31], [art_32], [art_35], [art_40], [art_44]
-Transformer [attention] [art_30], [art_33]
-Hybrid wavelet random forest and deep belief network (Wavelet-DBN-RF) [Wavelet-DBN-RF] [art_31], [art_40], [art_44]
-Extreme gradient boosting (XGBoost) [art_1], [art_2], [art_5], [art_46], [art_364]
-GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_6], [art_360], [art_365], [art_368]
-Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_363], [art_368]
-Deeper graph convolutional network (DeeperGCN) [DeeperGCN] [art_1], [art_5]
-Graph convolutional neural network and long-short term memory (GC-LSTM) [GC-LSTM] [art_1], [art_2], [art_5], [art_6], [art_7], [art_354], [art_361]
-Graph convolutional network (GCN) [art_38], [art_41], [art_357], [art_367]
-Graph WaveNet (GWN) [GWN] [art_1], [art_6], [art_44], [art_47], [art_363], [art_368]
-Hierarchical graph convolution network (HGCN) [HGCN] [art_1], [art_28]
-Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_6], [art_32], [art_44], [art_355], [art_360], [art_373]
-Multi long-short term memory network (Multi-LSTM) [Multi-LSTM] [art_31], [art_40], [art_44]
-Superposition graph neural network (SGNN) [art_24] [art_31], [art_40]
-Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_3], [art_6], [art_28], [art_35], [art_47], [art_363], [art_368], [art_371]
-Spatio-temporal correlation graph neural network (STGN) [art_40] [art_31], [art_44]
-Spatio-Temporal U-Network (ST-UNet) [ST-UNet] [art_1], [art_5]
-Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_368], [art_371]
-Temporal graph convolutional network (T-GCN) [T-GCN] [art_6], [art_30], [art_35]
+Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [437], [185], [12], [257], [416], [403], [426], [133]
+Historical average (HA) [357], [293], [437], [257], [133]
+Naive [11], [12], [269]
+Vector autoregression (VAR) [437], [12], [257], [106]
+Non-GNN machine learning methods Bidirectional long-short term memory network (BiLSTM) [109] [185], [73]
+Combined FC-LSTM and convolution neural network (CFCC-LSTM) [442] [304], [392], [403]
+DLinear [469] [326], [417]
+Hybrid convolutional neural network and gated recurrent unit (CGRU) [433] [326], [102], [97]
+Convolutional neural network (CNN) [210], [73]
+Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [223], [361], [45], [426]
+Convolutional long-short term memory network (ConvLSTM) [346] [437], [361], [229]
+Feed-forward neural network (FNN) [267], [183], [257], [106], [336], [403]
+Gated recurrent unit encoder–decoder (GED) [430] [98], [392], [345]
+Gated recurrent unit (GRU) [267], [293], [183], [310], [257], [10], [73], [325], [417], [336], [403], [133]
+Informer [520] [210], [345], [248], [417], [426], [133]
+Long-short term memory network (LSTM) [41], [267], [357], [223], [72], [293], [183], [310], [437], [210], [502], [185], [304], [98], [257], [392], [10], [73], [345], [184], [325], [417], [336], [45], [403], [426], [17], [42], [5], [133]
+Long-short term time series network (LSTNet) [202] [229], [17]
+Predictive deep convolutional neural network (PDCNN) [525] [326], [102]
+Random forest (RF) [293], [417], [403]
+Recurrent neural network (RNN) [73], [336]
+Support vector regression (SVR) [357], [293], [304], [98], [257], [106], [403], [426]
+Temporal convolutional network (TCN) [9] [293], [210], [229], [45], [426], [17]
+Temporal pattern attention long-short term memory network (TPA-LSTM) [347] [293], [326], [502], [304], [102], [97]
+Transformer [375] [210], [185]
+Hybrid wavelet random forest and deep belief network (Wavelet-DBN-RF) [164] [326], [102], [97]
+Extreme gradient boosting (XGBoost) [41], [267], [72], [73], [417]
+GNN methods Adaptive graph convolutional recurrent network (AGCRN) [8] [293], [106], [336], [45]
+Attention-based spatial-temporal graph convolutional network (ASTGCN) [116] [416], [45], [133]
+Diffusion convolutional recurrent neural network (DCRNN) [237] [416], [133]
+Graph convolutional neural network and long-short term memory (GC-LSTM) [316] [41], [267], [72], [293], [183], [248], [325]
+Graph convolutional network (GCN) [257], [361], [184]
+Graph WaveNet (GWN) [425] [41], [293], [97], [345], [416], [45]
+Hierarchical graph convolution network (HGCN) [115] [41], [437]
+Multivariate time series forecasting with graph neural network (MTGNN) [424] [293], [502], [97], [229], [106], [17], [133]
+Multi long-short term memory network (Multi-LSTM) [103] [326], [102], [97]
+Superposition graph neural network (SGNN) [460] [326], [102]
+Spatio-temporal graph convolutional network (STGCN) [457] [357], [293], [437], [304], [345], [416], [45], [426], [133]
+Spatio-temporal correlation graph neural network (STGN) [102] [326], [97]
+Spatio-temporal synchronous graph convolutional network (STSGCN) [353] [45], [426]
+Temporal graph convolutional network (T-GCN) [508] [293], [210], [304], [5], [133]
 The majority of the benchmarks used in this thematic group fall within the category of non-GNN machine learning methods, although there are several GNN benchmarks as well.
@@
 These models are widely used by environmental data forecasters due to their ability to capture temporal dependencies, and are considered a reliable benchmark in the field.
-As for GNN models, the most common benckmarks are the graph convolutional neural network and long-short term memory newtork (GC-LSTM) [GC-LSTM] (which appears only in quality air forecasting problems), and the spatio-temporal graph convolutional network (STGCN) [STGCN].
+As for GNN models, the most common benchmarks are the graph convolutional neural network and long-short term memory network (GC-LSTM) [316] (which appears only in quality air forecasting problems), and the spatio-temporal graph convolutional network (STGCN) [457].
 The GC-LSTM network is a model that integrates graph convolutional networks and long-short-term memory networks.
@@
 in 2020 in Ref.
-[MTGNN] is often reported in the papers using it as the second-best performing model, thus it probably deserves special attention and further investigation in future research.
+[424] is often reported in the papers using it as the second-best performing model, thus it probably deserves special attention and further investigation in future research.
 5.2.5 Results
@@
 The three papers using it (Refs.
-[art_31, art_40, art_44]) select 120 points from the area and use data records from January 2010 to April 2019 at 6 hours intervals, with a total of 13,624 samples at each node, for 10 different forecasting horizons.
+[326, 102, 97]) select 120 points from the area and use data records from January 2010 to April 2019 at 6 hours intervals, with a total of 13,624 samples at each node, for 10 different forecasting horizons.
 All papers present results for the complete set of 120 nodes and a subset of nodes.
@@
 RMSE 2.100 2.277 2.461 2.594 2.911 3.042 3.141 3.198 3.245 3.285
-DAGLN [art_31] MAE 1.068 1.219 1.346 1.443 1.716 1.852 1.935 1.984 2.019 2.067
+DAGLN [326] MAE 1.068 1.219 1.346 1.443 1.716 1.852 1.935 1.984 2.019 2.067
 RMSE 1.423 1.625 1.793 1.917 2.255 2.428 2.512 2.570 2.607 2.656
-DASTGN [art_44] MAE 1.039 1.184 1.297 1.387 1.662 1.806 1.900 1.945 1.995 2.028
+DASTGN [97] MAE 1.039 1.184 1.297 1.387 1.662 1.806 1.900 1.945 1.995 2.028
 RMSE 1.384 1.572 1.721 1.829 2.170 2.338 2.449 2.498 2.554 2.600
@@
 RMSE 3.222 3.246 3.264 3.270 3.275 3.328 3.392 3.431 3.453 3.458
-STGN [art_40] MAE 1.084 1.281 1.407 1.523 1.809 1.942 2.046 2.073 2.137 2.189
+STGN [102] MAE 1.084 1.281 1.407 1.523 1.809 1.942 2.046 2.073 2.137 2.189
 RMSE 1.430 1.697 1.863 2.013 2.358 2.514 2.631 2.649 2.717 2.771
@@
 RMSE 1.855 2.038 2.226 2.328 2.617 2.703 2.744 2.766 2.788 2.798
-For all the time horizons considered in the papers, the dynamic adaptive spatio-temporal graph neural network (DASTGN) model [art_44] proposed by Gao et al.
+For all the time horizons considered in the papers, the dynamic adaptive spatio-temporal graph neural network (DASTGN) model [97] proposed by Gao et al.
 in 2023 has the highest accuracy.
@@
 The spatio-temporal convolution module consists of an input module, a stack of spatio-temporal layers connected with residual and skip connections, and an output module that generates wind speed forecasts.
-Other two well performing models are the data-based adaptive graph learning network (DAGLN) [art_31], and the multivariate time series forecasting with graph neural network (MTGNN) [MTGNN].
+Other two well performing models are the data-based adaptive graph learning network (DAGLN) [326], and the multivariate time series forecasting with graph neural network (MTGNN) [424].
 The DAGLN model was proposed by Ren at al.
@@
 The MTGNN model, proposed by Wu et al.
-in 2020, is a well-known state-of-the-art model for multivariate time series forecasting.
+in 2020, is a well-known state-of-the-art model for multivariate time series forecasting, here included in the “Generic” group of papers.
 It incorporates a graph learning layer that learns a graph adjacency matrix from scratch, followed by a stack of graph convolution and temporal convolution modules to capture spatial and temporal dependencies, and an output module.
@@
 The second dataset shared by two papers (Refs.
-[art_1, art_5]) is called China National Urban Air Quality dataset (available at https://github.com/Friger/GAGNN).
+[41, 72]) is called China National Urban Air Quality dataset (available at https://github.com/Friger/GAGNN).
 It contains air quality index data (AQI) of 209 cities, collected from January 2017 to April 2019 at 1 hour granularity, for a total of 20,370 samples.
@@
 RMSE 11.49 17.23 21.32 24.31 26.72 28.71
-DeeperGCN MAE 6.54 9.74 11.77 13.40 15.29 16.41
-RMSE 13.67 18.93 21.14 23.83 26.25 28.02
 FGA MAE 5.87 9.14 11.71 13.75 15.42 16.80
 RMSE 11.36 17.01 21.05 24.09 26.55 28.52
-GAGNN [art_1] MAE 5.56 8.59 10.80 12.52 13.91 15.10
+GAGNN [41] MAE 5.56 8.59 10.80 12.52 13.91 15.10
 RMSE 10.81 16.17 19.84 22.51 24.63 26.37
@@
 RMSE 10.80 16.10 19.85 22.70 24.91 26.40
-INNGNN [art_5] MAE 5.48 8.49 10.67 12.34 13.72 14.91
+INNGNN [72] MAE 5.48 8.49 10.67 12.34 13.72 14.91
 RMSE 10.70 16.03 19.66 22.29 24.37 26.11
@@
 RMSE 11.27 16.84 20.77 23.60 25.80 27.38
-ST-UNet MAE 5.95 9.30 11.58 13.38 14.82 16.02
-RMSE 11.74 18.01 21.34 23.90 25.94 27.64
 XGBoost MAE 6.85 10.89 13.99 16.27 18.14 19.56
@@
 As from Tab.
-LABEL:t:risultati_chinese_aqi, the hybrid interpretable neural network and graph neural network model (INNGNN) [art_5] is the most accurate model for all time horizons except the 1 hour horizon, where it is beaten by the meta-graph convolutional recurrent network model (MegaCRN) [MegaCRN].
+LABEL:t:risultati_chinese_aqi, the hybrid interpretable neural network and graph neural network model (INNGNN) [72] is the most accurate model for all time horizons except the 1 hour horizon, where it is beaten by the meta-graph convolutional recurrent network model (MegaCRN) [166].
 The authors attribute this result to the fact that in the short term the AQI is strongly influenced by nearby cities, whereas the MegaCRN model is more focused on local features.
@@
 5.3 Finance
-“Finance” is another emerging theme of study, though not yet widely explored, with just 16 out of 281 selected papers.
+“Finance” is another emerging theme of study, though not yet widely explored, with just 14 out of 366 selected papers.
 The difficulty of predicting financial data is a well-established fact.
@@
 This is in part due to the limited number of published papers.
-The majority of studies are concentrated on stock price forecasting [art_59, art_60, art_62, art_63, art_65, art_377, art_382], with the aim of forecasting the future trend of the price of a company.
+Half of the studies are concentrated on stock price forecasting [151, 307, 371, 391, 362, 37, 429], with the aim of forecasting the future trend of the price of a company.
 Sometimes the stock prediction problem is seen as a classification task.
@@
 Another distinctive aspect in this group of papers is that researchers usually evaluate, among others, how their predictions affect the stability of investment returns over a specified period.
-This is commonly done through trading simulations in the test set, evaluated with metrics such as the Sharpe ratio and overall returns [art_59, art_60, art_65, art_66, art_377].
+This is commonly done through trading simulations in the test set, evaluated with metrics such as the Sharpe ratio and overall returns [151, 307, 362, 53, 429].
 These simulations require modeling rather than forecasting.
@@
 Two of the papers were published in IEEE Access by IEEE, while the remaining ones appeared in various other journals.
-Most of the corresponding authors are affiliated with academic institutions in China.
+The only two conference papers of this group were presented at CIKM.
 5.3.2 Datasets
@@
 Dataset Used by Link
-China A-Shares [art_63], [art_65], [art_66], [art_377], [art_378] https://finance.sina.com.cn/stock
-Criptocurriences [art_381] https://coinmarketcap.com/
-CSI 100 index [art_62] https://tushare.pro/
-CSI 300 index [art_59], [art_62], [art_382] https://global.csmar.com/
-Gold and silver market [art_374] https://www.kitco.com/
-HAL agriculture [art_379] www.hal.gov.tr
-S&P 100 index [art_62] https://finance.yahoo.com/
-S&P 500 index [art_59] https://finance.yahoo.com/
-Russell 1K index [art_62] https://finance.yahoo.com/
-Russell 3K index [art_62] https://finance.yahoo.com/
-US crude oil market [art_374] http://www.eia.gov/
+China A-Shares [391], [362], [53], [205] https://finance.sina.com.cn/stock
+CSI 100 index [371] https://tushare.pro/
+CSI 300 index [151], [371], [37], [429] https://global.csmar.com/
+Nordpool dataset [343] https://www.nordpoolgroup.com/
+S&P 100 index [371] https://finance.yahoo.com/
+S&P 500 index [151], [429] https://finance.yahoo.com/
+Russell 1K index [371] https://finance.yahoo.com/
+Russell 3K index [371] https://finance.yahoo.com/
 All the listed datasets represent collections of stock indices.
@@
 News can be processed by some language model like BERT (as in Refs.
-[art_59, art_66, art_378]).
+[151, 53, 205]).
 Most of the data have a granularity of 1 day, and the forecasting horizon is also generally 1 day.
-Two of the selected papers (Refs.
-[art_62, art_377]) analyses longer time horizons for stock forecasting, and others do the same for agricultural price (Ref.
-[art_379]) or cryptocurrencies’ price (Ref.
-[art_381]).
+One of the selected papers (Ref.
+[371]) analyses longer time horizons for stock forecasting.
 As for pre-processing techniques, half of the selected papers use normalization techniques, including min-max normalization, Z-score normalization, and logarithmic normalization.
 5.3.3 Types of models
-In terms of the taxonomy of the proposed GNN models, 6 out of 16 papers utilize attentional GNNs, 9 papers employ convolutional GNNs, and one paper takes into account both convolutional and attentional GNNs (Ref.
-[art_374]).
+In terms of the taxonomy of the proposed GNN models, 7 out of 14 papers utilize attentional GNNs, 6 papers employ convolutional GNNs, and another one (Ref.
+[343]) does not allow us to specify the classification of the model within the proposed taxonomy.
 As for the graph structure, some papers use models which learn the graph structure themselves (e.g., Refs.
-[art_62, art_64, art_378, art_379, art_382]), whereas some others define it a priori, based on the relationships between companies and stocks (e.g., Refs.
-[art_63, art_67, art_68, art_381]).
+[371, 225, 205, 37]), whereas some others define it a priori, based on the relationships between companies and stocks (e.g., Refs.
+[391, 158, 444]).
+One paper (Ref.
+[448]) examines how graph structures change across different observation scales, based on the idea that variable correlations observed over short timescales may differ significantly from those seen over longer periods.
 Half of the papers deal with a stock movement classification problem, so the most commonly used loss function is cross entropy.
 The majority of the papers explicitly indicate that Python is the programming language utilized, with nearly half employing PyTorch, and a few utilizing TensorFlow.
-Only Refs.
-[art_63] and [art_379] provide a link to the source code of the proposed models, which are, respectively, the knowledge graph with graph convolution neural network (KG-GCN) model, available at https://github.com/Gjl12321/KG_GCN_Stock_Price_Trend_Prediction_System, and the spectral temporal graph neural network (StemGNN), available at https://github.com/cevher/Time-Series-Analysis.
+Only Ref.
+[391] provides a link to the source code of the proposed models, which is the knowledge graph with graph convolution neural network (KG-GCN) model, available at https://github.com/Gjl12321/KG_GCN_Stock_Price_Trend_Prediction_System.
 5.3.4 Benchmark models
@@
 Category Model Used by
-Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_68], [art_375]
-Mean reversion (MR) [art_63], [art_65], [art_377]
-Non-GNN machine learning methods Convolutional neural network (CNN) [art_59], [art_64], [art_374], [art_379]
-Dual-stage attention-based recurrent neural network (DARNN) [art_63], [art_65], [art_377]
-Feed-forward neural network (FNN) [art_59], [art_68], [art_375], [art_378]
-Gated recurrent unit (GRU) [art_60], [art_68], [art_382]
-Long-short term memory network (LSTM) [art_59], [art_62], [art_63], [art_64], [art_65], [art_67], [art_68], [art_374], [art_375], [art_377], [art_379], [art_381]
-Support vector machine (SVM) [art_64], [art_68], [art_375], [art_378]
-GNN methods Graph convolutional network (GCN) [art_59], [art_60], [art_62], [art_64], [art_68], [art_382]
-Graph convolutional network with long short term memory network (GCN+LSTM) [art_65], [art_377]
-Financial Graph Attention Network (FinGAT) [FinGAT] [art_60], [art_62]
-Hierarchical graph attention network (HATS) [HATS] [art_62], [art_63], [art_65], [art_377]
-Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_62], [art_382]
-Spatiotemporal hypergraph convolution network (STHGCN) [STHGCN] [art_63], [art_65], [art_377]
+Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [444], [297]
+Mean reversion (MR) [391], [362]
+Non-GNN machine learning methods Convolutional neural network (CNN) [151], [225]
+Dual-stage attention-based recurrent neural network (DARNN) [391], [362]
+Feed-forward neural network (FNN) [151], [444], [297], [205]
+Gated recurrent unit (GRU) [307], [444], [37], [429]
+Long-short term memory network (LSTM) [151], [371], [391], [225], [362], [158], [444], [297], [429]
+Support vector machine (SVM) [225], [444], [297], [205]
+Transformer [375] [37], [429]
+GNN methods Graph convolutional network (GCN) [151], [307], [371], [225], [444], [37]
+Financial graph attention network (FinGAT) [142] [307], [371]
+Hierarchical graph attention network (HATS) [188] [371], [391], [362], [429]
+Multivariate time series forecasting with graph neural network (MTGNN) [424] [371], [37]
+Spatiotemporal hypergraph convolution network (STHGCN) [331] [391], [362]
+Temporal graph convolution (TGC) [88] [391], [429]
 It is evident that there is no consensus regarding the most appropriate benchmark model to use, especially of GNN type.
@@
 The most commonly used GNN model is a vanilla graph convolutional network (GCN), followed by a more complex hierarchical graph attention network (HATS) proposed by Raehyun et al.
-in 2019 [HATS].
+in 2019 [188].
 The HATS model was proposed for predicting stock movements using relational data.
@@
 The only two papers with results which can be compared are Refs.
-[art_63] and [art_65], which study 758 frequently traded stocks collected from the A-share market in China from January 2013 to December 2019 (available at https://finance.sina.com.cn/stock).
+[391] and [362], which study 758 frequently traded stocks collected from the A-share market in China from January 2013 to December 2019 (available at https://finance.sina.com.cn/stock).
 These two papers compare the results obtained by using 5, 10 and 20 trading days as records.
@@
 HGTAN 40.02% 41.77% 39.03% 40.32%
-KG-GCN [art_63] 51.38% 65.72% 31.45% 39.27%
+KG-GCN [391] 51.38% 65.72% 31.45% 39.27%
 LSTM 35.03% 36.43% 34.23% 35.20%
 MOM 35.73% 35.19% 32.82% 33.96%
-MONEY [art_65] 39.90% 43.92% 40.61% 42.20%
+MONEY [362] 39.90% 43.92% 40.61% 42.20%
 MR 35.32% 38.03% 33.60% 35.68%
@@
 TGC 37.81% 36.96% 34.49% 35.67%
-The two best models are the knowledge graph and graph convolution neural network (KG-GCN) [art_63] and the stock price movement prediction via convolutional network with adversarial hypergraph model (MONEY) [art_65], depending on the metrics considered.
+The two best models are the knowledge graph and graph convolution neural network (KG-GCN) [391] and the stock price movement prediction via convolutional network with adversarial hypergraph model (MONEY) [362], depending on the metrics considered.
 The KG-GCN model, proposed by Wang et al.
@@
 5.4 Health
-The fourth thematic group of papers is “Health”, which includes 15 out of 281 publications.
+The fourth thematic group of papers is “Health”, which includes 23 out of 366 publications.
 These papers focus on critical aspects of health monitoring, disease modeling, and diagnostic tools, with particular emphasis on the spread and diagnosis of disease.
@@
 These complexities make the forecasting task more difficult, and non-GNN machine learning, and GNN models themselves, have been studied to address this issue.
-Moreover, the spread of large-scale human diseases is closely related to both the mobility patterns of populations and the demographic characteristics of residents in different regions [art_413].
+Moreover, the spread of large-scale human diseases is closely related to both the mobility patterns of populations and the demographic characteristics of residents in different regions [445].
 Consequently, the application of GNN models to epidemic dynamics appears to be a suitable approach for analyzing and addressing such data.
-A common reference model for epidemic prediction is the SIR model [modello_SIR], an acronym referring to the state of the health of individuals.
+A common reference model for epidemic prediction is the SIR model [180], an acronym referring to the state of the health of individuals.
 The three letters of the acronym stand for susceptible (S), infectious (I), and recovered (R).
@@
 It uses a system of differential equations to describe how the disease spreads spatially and changes in time.
+The main advantage of this model is its inherent interpretability; it produces coefficients that are easy to interpret.
+However, its strong and simplistic assumptions also restrict its forecasting capability [514].
 In general, some papers use it as benchmark (e.g., Ref.
-[art_49, art_413, art_415]).
-Other papers in this section focus on different types of time series, such as electroencephalogram (EEG) data for epilepsy detection (Ref.
-[art_14]) or emotion recognition (Refs.
-[art_411, art_412]).
+[462, 445, 285]).
+One study (Ref. [334]) also analyzes how air quality and satellite observations of temperature and humidity relate to COVID-19 mortality and severity.
+Other papers in this section focus on different types of time series, such as electroencephalogram (EEG) data for epilepsy detection (Refs.
+[383, 38, 23]) or emotion recognition (Refs.
+[527, 119, 384]).
 While emotion recognition may not seem directly related to the “Health” topic, it is included here because it relies on EEG data.
-The use of EEG in emotion recognition helps to identify more accurately true emotional states, as there physiological signals cannot be consciously controlled in the same way as facial expressions, tone of voice, or body posture [art_411].
+The use of EEG in emotion recognition helps to identify more accurately true emotional states, as there physiological signals cannot be consciously controlled in the same way as facial expressions, tone of voice, or body posture [527].
 The selected papers were each published in different journals, with one paper per journal, primarily in the field of computer science.
-The majority of the corresponding authors are affiliated with academic institutions based in China.
+In terms of conference venues, KDD and CIKM are the two most popular, with three papers each.
 5.4.2 Datasets
@@
 Dataset Used by Link
-African COVID-19 dataset [art_410] https://www.worldometers.info/coronavirus/
-Bonn dataset [art_14] http://dx.doi.org/10.1103/PhysRevE.64.061907 (source paper)
-Brazilian COVID-19 data [art_52] https://github.com/lcaldeira/GrafoBrasilCovid
-CDC US-State [art_414] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
-Chickenpox [art_410] https://www.kaggle.com/datasets/die9origephit/chickenpox-cases-hungary
-EEG DEAP dataset [art_412] http://www.eecs.qmul.ac.uk/mmv/datasets/deap/download.html
-eICU [art_189] https://doi.org/10.1038/sdata.2018.178 (source paper)
-European COVID-19 dataset [art_410] https://www.worldometers.info/coronavirus/
-ILI IDWR Japan-Prefecture [art_414] https://tinyurl.com/y5dt7stm
-ILI US-HHS US-Regional [art_414] https://tinyurl.com/y39tog3h
-MIMIC-IV [art_189] https://physionet.org/content/mimiciv/2.0/
-Nytimes COVID-19 data [art_50], [art_413] https://github.com/nytimes/covid-19-data
-OxCGRT dataset [art_416] https://www.bsg.ox.ac.uk/research/covid-19-government-response-tracker
-Parkinson’s Progression Markers Initiative (PPMI) [art_92] http://www.ppmi-info.org/
-Japan COVID-19 [art_415] https://github.com/swsoyee/2019-ncov-japan
-Parkinson Speech data set (PS) [art_92] https://doi.org/10.1109/jbhi.2013.2245674 (source paper)
-SJTU Emotion EEG Dataset (SEED) [art_411], [art_412] https://bcmi.sjtu.edu.cn/home/seed/
-Spanish COVID-19 data [art_51] https://cnecovid.isciii.es/
-Spanish COVID-19 data [art_414] https://dataforgood.fb.com/tools/disease-prevention-maps/
-Spikes and slow waves (SSW) dataset [art_14] http://dx.doi.org/10.1109/ICDM50108.2020.00067 (source paper)
-US COVID-19  [^2] This dataset includes multiple sources. [art_415] https://github.com/CSSEGISandData/COVID-19
-US COVID-19 [art_416] https://covidtracking.com/data/national
-US COVID-19 vaccinations [art_50] https://ourworldindata.org/us-states-vaccinations
-US influenza [art_48] https://github.com/aI-area/DVGSN/tree/main/dataset
-Only two pairs of papers in this thematic group use the same dataset, but none of the proposed models yield comparable results.
+African COVID-19 dataset [270] https://www.worldometers.info/coronavirus/
+Bonn dataset [383] http://dx.doi.org/10.1103/PhysRevE.64.061907 (source paper)
+Brazilian COVID-19 data [298] https://github.com/lcaldeira/GrafoBrasilCovid
+CDC US-State [528], [68] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
+Chickenpox [270], [160] https://www.kaggle.com/datasets/die9origephit/chickenpox-cases-hungary
+CSSE COVID-19 [334], [282], [514] https://github.com/CSSEGISandData/COVID-19
+EEG DEAP dataset [119] http://www.eecs.qmul.ac.uk/mmv/datasets/deap/download.html
+eICU [434] https://doi.org/10.1038/sdata.2018.178 (source paper)
+European COVID-19 dataset [270] https://www.worldometers.info/coronavirus/
+ILI IDWR Japan-Prefecture [528], [68] https://tinyurl.com/y5dt7stm
+ILI US-HHS US-Regional [528], [68] https://tinyurl.com/y39tog3h
+ISRUC-Sleep Subgroup III (ISRUC-S3) [384] https://sleeptight.isr.uc.pt/
+Japan COVID-19 [285] https://github.com/swsoyee/2019-ncov-japan
+Montreal archive of sleep studies (MASS-SS3) [384] http://www.ceams-carsm.ca/en/MASS
+Meta COVID-19 data [528], [160] https://dataforgood.fb.com/tools/disease-prevention-maps/
+MIMIC-IV [434] https://physionet.org/content/mimiciv/2.0/
+Nytimes COVID-19 data [235], [445] https://github.com/nytimes/covid-19-data
+OxCGRT dataset [187] https://www.bsg.ox.ac.uk/research/covid-19-government-response-tracker
+Parkinson’s progression markers initiative (PPMI) [498] http://www.ppmi-info.org/
+Parkinson speech dataset (PS) [498] https://doi.org/10.1109/jbhi.2013.2245674 (source paper)
+SJTU emotion EEG dataset (SEED) [527], [119], [384] https://bcmi.sjtu.edu.cn/home/seed/
+Spanish COVID-19 data [290] https://cnecovid.isciii.es/
+Spikes and slow waves (SSW) [383] http://dx.doi.org/10.1109/ICDM50108.2020.00067 (source paper)
+TUH EEG seizure corpus [23] https://isip.piconepress.com/projects/nedc/html/tuh_eeg/
+US COVID-19 Johns Hopkins This dataset includes multiple sources. [285] https://github.com/CSSEGISandData/COVID-19
+US COVID-19 tracking project [187] https://covidtracking.com/data/national
+US COVID-19 vaccinations [235] https://ourworldindata.org/us-states-vaccinations
+US influenza [483] https://github.com/aI-area/DVGSN/tree/main/dataset
+Only few subsets of papers in this thematic group use the same dataset, but none of the proposed models yield comparable results.
 Most of the datasets are related to COVID-19 and have a daily time granularity.
 In addition, some papers (Refs.
-[art_49, art_52, art_92, art_189, art_413]) include exogenous variables that can be either context-related (e.g., geography, population, health conditions, vaccinations, economic factors), or specific to a single patient (including personal information such as age, gender, and medical conditions).
+[462, 298, 498, 434, 445, 334, 282]) include exogenous variables that can be either context-related (e.g., geography, population, health conditions, vaccinations, economic factors, air quality), or specific to a single patient (including personal information such as age, gender, and medical conditions).
 A paper that undertakes an in-depth study of exogenous variables is Ref.
-[art_416].
+[187].
 The authors categorize these variables as holiday variables, time categorical variables (e.g., categorical week and month, and the number of days and weeks since the start of the dataset), and policy variables (e.g., school and workplace closing, movement restrictions, face coverings, and vaccination policies).
@@
 5.4.3 Types of models
-As for the taxonomy of the proposed models, 11 out of 15 papers use a convolutional GNN architecture.
-Three papers use a pure attentional model, while another paper presents a hybrid convolutional-attentional architecture.
+As for the taxonomy of the proposed models, 18 out of 23 papers use a convolutional GNN architecture.
+Three papers use a pure attentional model, while two other papers present a hybrid convolutional-attentional architecture.
 Different approaches are used to define the graph structure.
 Some papers base their graphs on the geographical distribution of the data (as in Refs.
-[art_49, art_52, art_92]), and others use models that learn the graph structure directly, as in Refs.
-[art_48, art_50, art_415, art_416].
-There is also a paper that uses a variant of the visibility algorithm [art_14].
+[462, 298, 498, 334, 160]), and others use models that learn the graph structure directly, as in Refs.
+[483, 235, 285, 187, 282].
+There is also a paper that uses a variant of the visibility algorithm [383].
 In half of the cases, the graph is static, while the dynamic graph approach appears more frequently in the papers from 2024.
 In terms of programming languages, the community of the “Health” thematic group uses only Python with PyTorch.
-As for the coice of the loss functions, MSE is typically used for prediction tasks, and cross entropy is used for classification tasks.
+As for the choice of the loss functions, MSE is typically used for prediction tasks, and cross entropy is used for classification tasks.
 Tab.
@@
 Model Link
-Backbone-based Dynamic Spatio-Temporal Graph Neural Network (BDSTGNN) [art_415] https://github.com/JkMaoya/BDSTGNN
-Dynamic virtual graph significance networks (DVGSN) [art_48] https://github.com/aI-area/DVGSN
-Deep learning of contagion dynamics on complex networks [art_51] https://doi.org/10.5281/zenodo.4974521
-Ordinary Differential Equation in Graph Neural Networks (GNNODENet) [art_413] https://github.com/xiongzhangxyq/GNNODENet
-Recurrent Graph Gate Fusion Transformers (ReGraFT) [art_416] https://github.com/mfriendly/ReGraFT
-Sparse spectra graph convolutional network (SSGCNet) [art_14] https://github.com/anonymous2020-source-code/WNFG-SSGCNet-ADMM
-Spatial-Temporal Graph Convolutional Network (STGCN) [art_52] https://github.com/lcaldeira/CovidPandemicForecasting
-Time-aware Context-Gated Graph Attention Network (T-ContextGGAN) [art_189] https://github.com/OwlCitizen/TContext-GGAN
+Backbone-based dynamic spatio-temporal graph neural network (BDSTGNN) [285] https://github.com/JkMaoya/BDSTGNN
+DCRNN for COVID-19 forecasting [334] https://github.com/Covid-19-papers/Recurrent-Graph-Neural-Network-on-COVID-19-and-NASA-Satellite-data
+Deep learning of contagion dynamics on complex networks [290] https://doi.org/10.5281/zenodo.4974521
+Dynamic virtual graph significance networks (DVGSN) [483] https://github.com/aI-area/DVGSN
+Ordinary differential equation in graph neural networks (GNNODENet) [445] https://github.com/xiongzhangxyq/GNNODENet
+Hierarchical spatial-temporal framework for COVID-19 trend forecasting (HierST) [514] https://github.com/shun-zheng/HierST
+Recurrent graph gate fusion transformers (ReGraFT) [187] https://github.com/mfriendly/ReGraFT
+Sparse spectra graph convolutional network (SSGCNet) [383] https://github.com/anonymous2020-source-code/WNFG-SSGCNet-ADMM
+Spatial-temporal graph convolutional network (STGCN) [298] https://github.com/lcaldeira/CovidPandemicForecasting
+Time-aware context-gated graph attention network (T-ContextGGAN) [434] https://github.com/OwlCitizen/TContext-GGAN
+Temporal multiresolution graph neural networks (TMGNN) [160] https://github.com/bachnguyenTE/temporal-mgn
 5.4.4 Benchmark models
@@
 Category Model Used by
-Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_48], [art_49], [art_52], [art_410], [art_414], [art_415]
-SIR model [modello_SIR] [art_49], [art_413], [art_415]
-Non-GNN machine learning methods Feed-forward neural network (FNN) [art_14], [art_48]
-Bi-Hemisphere Domain Adversarial Neural Network (Bi-DANN) [Bi-DANN] [art_411], [art_412]
-A Novel Bi-Hemispheric Discrepancy Model for EEG Emotion Recognition (BiHDM) [BiHDM] [art_411], [art_412]
-Gated recurrent unit (GRU) [art_410], [art_413], [art_415], [art_416]
-Long-short term memory network (LSTM) [art_49], [art_50], [art_410], [art_413]
-Recurrent neural network (RNN) [art_189], [art_413], [art_414]
-Support vector machine (SVM) [art_189], [art_411]
-GNN methods Cross-location attention based graph neural network (ColaGNN) [ColaGNN] [art_414], [art_415], [art_416]
-Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_49], [art_413], [art_414], [art_416]
-Dynamical graph convolutional neural networks (DGCNN) [DGCNN] [art_411], [art_412]
-Graph attention network (GAT) [GAT] [art_48], [art_413]
-Graph convolutional network (GCN) [art_49], [art_413]
-Regularized Graph Neural Networks (RGNN) [RGNN] [art_411], [art_412]
-Spatio-temporal attention network for pandemic prediction (STAN) [STAN] [art_413], [art_415]
-Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_413], [art_415]
+Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [483], [462], [298], [270], [528], [285], [160], [68], [282]
+Historical average (HA) [160], [282]
+Naive [298], [160]
+SIR model [180] [462], [445], [285]
+Non-GNN machine learning methods Feed-forward neural network (FNN) [383], [483]
+Bi-hemisphere domain adversarial neural network (Bi-DANN) [238] [527], [119]
+A novel bi-hemispheric discrepancy model for EEG emotion recognition (BiHDM) [234] [527], [119]
+Convolutional neural network and recurrent neural network with residual module (CNNRNN-Res) [421] [528], [68]
+Gated recurrent unit (GRU) [270], [445], [285], [187]
+Long-short term memory network (LSTM) [462], [235], [270], [445], [38], [384], [160], [68], [282]
+Long-short term time series network (LSTNet) [202] [528], [68]
+MiniRocket [65] [38], [23]
+Random forest (RF) [483], [384]
+Recurrent neural network (RNN) [434], [445], [528], [68]
+Recurrent neural network with attention (RNN+Att) [528], [68]
+Support vector machine (SVM) [434], [527], [384]
+Time-Series representation learning framework via Temporal and Contextual Contrasting (TS-TCC) [77] [38], [23]
+Time series transformer (TST) [471] [235], [23]
+GNN methods Cross-location attention based graph neural network (ColaGNN) [68] [528], [285], [187]
+CovidGNN [179] [285], [282]
+Diffusion convolutional recurrent neural network (DCRNN) [237] [462], [445], [528], [187], [68]
+Dynamical graph convolutional neural networks (DGCNN) [354] [527], [119], [384]
+Graph attention network (GAT) [376] [483], [445]
+Graph convolutional network (GCN) [462], [445]
+Regularized graph neural networks (RGNN) [517] [527], [119]
+Spatio-temporal attention network for pandemic prediction (STAN) [94] [445], [285], [282]
+Spatio-temporal graph convolutional network (STGCN) [457] [445], [68]
+Spatial-temporal graph ODE networks (STGODE) [84] [445], [285]
 There are no particularly outstanding benchmark models, and the ones categorized as GNNs appear only in papers from 2024.
@@
 5.5 Mobility
-The thematic group that includes the largest number of papers identified by our query (66 out of 281) is “Mobility”, which regards the movement of people.
+The thematic group that includes the largest number of papers identified by our query (102 out of 366) is “Mobility”, which regards the movement of people.
 It includes urban traffic, air travel, and bicycle demand, among other applications.
@@
 The increasing number of papers published each year in this thematic group reflects both the growing research interest in this field and the growing confidence in GNN models.
-The most commonly investigated areas regard urban traffic (approximately 70% of the papers), especially urban traffic flow and urban traffic speed.
+The most commonly investigated areas regard urban traffic (approximately $1/4$ of the papers), especially urban traffic flow and urban traffic speed.
 Among the journals where the selected papers were published the most common are Applied Intelligence by Springer, ACM Transactions on Knowledge Discovery from Data, both by the Association for Computing Machinery, and IEEE Transactions on Intelligent Transportation Systems by the Institute of Electrical and Electronics Engineers, with 6 out of 66 papers each.
-Most of the corresponding authors are affiliated with Chinese institutions.
+The three conference venues with the highest number of papers are CIKM with 13 papers, SIGSPATIAL with 9 papers, and KDD with 8 papers.
 5.5.2 Datasets
@@
 Dataset Used by Link
-Bureau of Transportation Statistics (BTS) [art_434], [art_443], [art_448] https://www.transtats.bts.gov/
-CAAC flights [art_129] http://www.caac.gov.cn/index.html
-Capital Bike Share [art_446] https://ride.capitalbikeshare.com/system-data
-CD-HW [art_126] https://doi.org/10.1109/ACCESS.2020.3027375 (source paper)
-DiDi Beijing [art_102], [art_447] https://github.com/didi/TrafficIndex
-DiDi Chengdu [art_93], [art_98] https://gaia.didichuxing.com/
-Divvy Bike Chicago [art_446] https://ride.divvybikes.com/system-data
-Huai’an Bus Transaction [art_452] https://www.pandabus.cn/panda/panda-minibus.html
-Los-loop [art_103] https://github.com/lehaifeng/T-GCN/tree/master/data
-Lyft bike dataset [art_446] https://www.lyft.com/bikes/bay-wheels/system-data
-METR-LA [art_93], [art_99], [art_100], [art_109], [art_110], [art_113], [art_116], [art_120], [art_122], [art_126], [art_128], [art_408], [art_427], [art_430], [art_433], [art_435], [art_438], [art_441], [art_442], [art_444] https://paperswithcode.com/dataset/metr-la
-Milan network [art_123] https://github.com/arunasubbiah/milan-telecom-data-modeling
-NE-BJ [art_100] https://github.com/tsinghua-fib-lab/Traffic-Benchmark
-NYC-Bike [art_94], [art_96], [art_114], [art_446], [art_449], [art_451] https://citibikenyc.com/system-data
-NYC SafeGraph [art_449] https://www.safegraph.com/academics
-NYC-Taxi [art_96], [art_449] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
-PeMS03 [art_108], [art_113], [art_115], [art_117], [art_125], [art_130], [art_419], [art_435], [art_445], [art_453] https://www.kaggle.com/datasets/elmahy/pems-dataset
-PeMS04 (PeMSD4) [art_97], [art_99], [art_104], [art_108], [art_113], [art_115], [art_117], [art_122], [art_125], [3], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_424], [art_433], [art_435], [art_441], [art_445], [art_448], [art_449], [art_452], [art_453] https://paperswithcode.com/dataset/pems04
-PeMS07 (PeMSD7) [art_108], [art_110], [art_113], [art_117], [art_125], [3], [art_130], [art_406], [art_422], [art_424], [art_427], [art_432], [art_435], [art_438], [art_441], [art_444], [art_445], [art_453] https://paperswithcode.com/dataset/pems07
-PeMS08 (PeMSD8) [art_97], [art_99], [art_108], [art_115], [art_117], [art_122], [art_125], [art_130], [art_131], [art_132], [art_406], [art_408], [art_424], [art_432], [art_433], [art_435], [art_441], [art_444], [art_445], [art_448], [art_449], [art_452], [art_453] https://paperswithcode.com/dataset/pems08
-PEMS-BAY [art_100], [art_109], [art_110], [art_116], [art_120], [art_122], [art_406], [art_408], [art_419], [art_420], [art_422], [art_430], [art_433], [art_435], [art_441], [art_442], [art_444] https://paperswithcode.com/dataset/pems-bay
-Q-Traffic [art_124] https://github.com/JingqingZ/BaiduTraffic#Dataset
-Seattle loop [art_441] https://github.com/zhiyongc/Seattle-Loop-Data
-UVDS [art_128] https://doi.org/10.1007/978-3-030-73280-6_6 (source paper)
-TaxiBJ [art_126], [art_451] https://www.microsoft.com/en-us/research/publication/forecasting-citywide-crowd-flows-based-big-data/
-Taxi Shenzhen [art_103], [art_135] https://github.com/lehaifeng/T-GCN/tree/master/data
-TLC Taxi [art_119] https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page
-Urban-core [art_441] https://topis.seoul.go.kr/
-Xiecheng delay dataset [art_443] https://pan.baidu.com/s/1dEPyMGh#list/path\$=\$%2F
+Baoding dataset [493] https://github.com/usail-hkust/ASeer
+BRN+ [63] https://www.microsoft.com/en-us/research/publication/t-drive-trajectory-data-sample/
+Bruxelles transport data [320] https://www.kaggle.com/datasets/giobbu/belgium-obu
+Bureau of transportation statistics (BTS) [420], [419] https://www.transtats.bts.gov/
+CAAC flights [25] http://www.caac.gov.cn/index.html
+Capital bike share [89] https://ride.capitalbikeshare.com/system-data
+CarPark1918 [171] https://github.com/JIANGYUE61610306/SAGDFN/tree/main/data/model
+CD-HW [39] https://doi.org/10.1109/ACCESS.2020.3027375 (source paper)
+DiDi Beijing [402], [126] https://github.com/didi/TrafficIndex
+DiDi Chengdu [274], [220], [276], [144], [219], [275] https://gaia.didichuxing.com/
+Didi Shenzhen [276], [144] https://outreach.didichuxing.com/app-outreach/CTIE
+Divvy bike Chicago [89], [135] https://ride.divvybikes.com/system-data
+Highways England [127] http://tris.highwaysengland.co.uk/detail/trafficflowdata
+Los-loop [263] https://github.com/lehaifeng/T-GCN/tree/master/data
+Lyft bike dataset [89] https://www.lyft.com/bikes/bay-wheels/system-data
+METR-LA [274], [218], [212], [217], [200], [198], [120], [490], [181], [39], [22], [32], [13], [199], [242], [510], [406], [348], [378], [208], [341], [171], [276], [393], [178], [342], [524], [144], [85], [213], [523], [222], [33], [281], [308], [411] https://www.kaggle.com/datasets/annnnguyen/metr-la-dataset
+NE-BJ [212] https://github.com/tsinghua-fib-lab/Traffic-Benchmark
+NYC-Bike [428], [211], [245], [89], [388], [259], [95], [422] https://citibikenyc.com/system-data
+NYC-Taxi [211], [259], [95], [422], [523], [496] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
+PeMS03 [173], [198], [438], [111], [511], [112], [163], [510], [250], [82], [81], [203], [84], [259], [80], [95], [213], [63], [461], [145] https://www.kaggle.com/datasets/elmahy/pems-dataset
+PeMS04 (PeMSD4) [62], [218], [489], [173], [198], [438], [111], [181], [511], [112], [69], [27], [227], [32], [163], [407], [242], [510], [348], [250], [82], [341], [127], [81], [342], [203], [84], [8], [80], [95], [530], [85], [213], [63], [370], [461], [145] https://www.kaggle.com/datasets/elmahy/pems-dataset
+PeMS07 (PeMSD7) [173], [200], [198], [111], [511], [112], [227], [466], [407], [13], [352], [510], [406], [348], [208], [250], [82], [178], [81], [203], [84], [80], [95], [85], [213], [370], [461], [145] https://www.kaggle.com/datasets/elmahy/pems-dataset
+PeMS08 (PeMSD8) [62], [218], [173], [438], [111], [181], [511], [112], [69], [27], [227], [32], [407], [352], [242], [510], [348], [208], [250], [82], [127], [81], [342], [203], [84], [259], [8], [80], [95], [530], [213], [63], [370], [219], [461], [281], [145] https://www.kaggle.com/datasets/elmahy/pems-dataset
+PEMS-BAY [212], [217], [200], [120], [490], [181], [227], [32], [163], [159], [466], [199], [242], [510], [348], [378], [208], [341], [276], [393], [178], [342], [524], [144], [85], [213], [222], [281], [308], [411] https://zenodo.org/records/4263971
+Q-Traffic [278] https://github.com/JingqingZ/BaiduTraffic#Dataset
+SafeGraph [123] https://www.safegraph.com/
+Seattle loop [348], [33] https://github.com/zhiyongc/Seattle-Loop-Data
+UVDS [22] https://doi.org/10.1007/978-3-030-73280-6_6 (source paper)
+TaxiBJ [39], [388], [49] https://www.microsoft.com/en-us/research/publication/forecasting-citywide-crowd-flows-based-big-data/
+Taxi Shenzhen [263], [169] https://github.com/lehaifeng/T-GCN/tree/master/data
+TLC Taxi [251] https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page
+Urban-core [348] https://topis.seoul.go.kr/
+Xiecheng delay [419] https://pan.baidu.com/s/1dEPyMGh#list/path\$=\$%2F
+Zhuzhou dataset [493] https://github.com/usail-hkust/ASeer
 The most popular datasets used in this group are those provided by the California Transportation Agency Performance Measurement System (PeMS, available at http://pems.dot.ca.gov/), which include various versions differing in time range, spatial coverage, and number of sensors, and the one called METR-LA, collected from loop detectors along the Los Angeles County highway.
@@
 For a more detailed overview of the benchmark datasets in the traffic domain, see Ref.
-[review_traffico_con_dataset].
-Other common public traffic datasets can be found at https://paperswithcode.com/task/traffic-prediction.
+[167].
+Other common public traffic datasets can be found at https://web.archive.org/web/20241212092408/https://paperswithcode.com/task/traffic-prediction.
 As to external data, the most commonly used are data on weather conditions, time information (time of day and day of week) and events (e.g., holidays), as well as information about the road network, like points of interest.
 The majority of the papers utilize data with a 5-minute granularity, and provide forecasts for horizons ranging from 5 minutes to 1 hour ahead.
-In addition to imputation of missing values, in almost half of the cases the pre-processing of the data includes data normalization (min-max or Z-score normalization) and the selection of time windows regarding recent, daily and weekly data (as in 10 out of 66 papers).
+In addition to imputation of missing values, in almost half of the cases the pre-processing of the data includes data normalization (min-max or Z-score normalization) and the selection of time windows regarding recent, daily and weekly data.
 5.5.3 Types of models
-As for the proposed models, 35 out of 66 belong to the pure convolutional GNN paradigm, 12 to hybrid convolutional-attention architectures, followed by 14 attentional GNNs, and 1 recurrent GNN.
-Two papers use both convolutional and attentional models, and other two (Refs.
-[art_438, art_451]) do not fall into a specific class.
+As for the proposed models, 53 out of 102 belong to the pure convolutional GNN paradigm, followed by 21 hybrid convolutional-attention architectures, 20 attentional GNNs, and 1 recurrent GNN.
+Two papers use both convolutional and attentional models, and other five (Refs.
+[406, 388, 84, 370, 411]) do not fall into a specific class.
 Regarding the definition of the graph structure, most of the papers discuss it in detail.
-Half of the papers published before 2024 are based on a pre-defined adjacency matrix, while the majority of papers of 2024 propose models that learn the adjacency matrix on their own, either from scratch or based on a pre-initialization (e.g., Refs.
-[art_108, art_444]).
+Almost half of the papers published before 2024 are based on a pre-defined adjacency matrix, while the majority of papers since 2024 propose models that learn the adjacency matrix on their own, either from scratch or based on a pre-initialization (e.g., Refs.
+[173, 208]).
 Here, more than in any other thematic group, there are multiple ways to define the adjacency matrix.
@@
 More in detail, a first group of papers uses road connectivity (e.g., Refs.
-[art_104, art_120, art_123, art_135, art_431]); second group uses spatial distance (e.g., Refs.
-[art_93, art_96, art_99, art_110, art_128, art_132, art_434, art_454]); a third group incorporates extra information about the so-called points of interest (such as Refs.
-[art_94, art_96]); a fourth group considers some similarity measures between the time series (e.g., Refs.
-[art_97, art_117, art_122, art_123, art_435, art_445]), such as the correlation coefficient or cosine similarity; a fifth group directly uses the number of people traveling on the road at a given time step (e.g., Refs.
-[art_94, art_102, art_129, art_133]), or a combination of different aspects (e.g., Refs.
-[art_420, art_444, art_448, art_449]).
+[489, 490, 169, 330, 134, 80]); second group uses spatial distance (e.g., Refs.
+[274, 211, 218, 200, 22, 27, 420, 193]); a third group incorporates extra information about the so-called points of interest (such as Refs.
+[428, 211]); a fourth group considers some similarity measures between the time series (e.g., Refs.
+[62, 111, 181, 510, 250]), such as the correlation coefficient or cosine similarity; a fifth group directly uses the number of people traveling on the road at a given time step (e.g., Refs.
+[428, 402, 25, 71]), or a combination of different aspects (e.g., Refs.
+[159, 208]).
+For a more detailed overview of graph construction techniques in traffic flow prediction models, the reader is referred to Ref.
+[78].
 The two most common loss functions used to train the models are the MAE and the MSE.
 Few papers consider a combination of loss functions, especially when dealing with a model that has to learn the graph structure by itself (Refs.
-[art_126, art_130, art_446, art_448]).
+[39, 112, 89, 341, 276]).
 Not all papers specify the language or library used for writing the code.
@@
 Model Link
-3-Dimensional Graph Convolution Network (3DGCN) [art_94] https://github.com/FIBLAB/3D-DGCN
-Adaptive generalized PageRank graph neural network (AGP-GNN) [art_116] https://github.com/guoxiaoyuatbjtu/agp-gnn
-Adaptive multigraph convolutional networks (AMGCN) [art_444] https://github.com/hfimmortal/AMGCN
-Automated dilated spatio-temporal synchronous graph network (Auto-DSTSGN) [art_108] https://github.com/jinguangyin/Auto-DSTSGN
-BigST [art_447] https://github.com/usail-hkust/BigST
-Dynamic multi-view graph neural network for citywide traffic inference (CTVI+) [art_97] https://github.com/dsj96/TKDD
-Dynamic Graph Convolutional Recurrent Network (DGCRN) [art_100] https://github.com/tsinghua-fib-lab/Traffic-Benchmark
-Progressive Graph Convolutional Network (PGCN) [art_441] https://github.com/yuyolshin/PGCN
-Scale-aware neural architecture search framework for MTS forecasting (SNAS4MTF) [art_408] https://github.com/shangzongjiang/SNAS4MTF
-Spatio-temporal Causal Graph Attention Network (STCGAT) [art_125] https://github.com/zsqZZU/STCGAT/tree/v1.0.0
-Spatial-Temporal Graph Convolutional Neural Network with LSTM layers (STGCN-LSTM) [3] https://github.com/ant-agafonov/stgcn-lstm
-Spatio-Temporal Graph Mixformer (STGM) [art_110] https://github.com/mouradost/stgm
-Space-Time Separable Graph Convolutional Network (STSGCN) [art_443] https://github.com/FraLuca/STSGCN
-Two-level Resolution Neural Network (TwoResNet) [art_430] https://github.com/semink/TwoResNet
-Wavelet dynamic spatiotemporal aware Graph Neural Network (W-DSTAGNN) [art_419] https://github.com/yash-jakhmola/w-dstagnn
+3-Dimensional graph convolution network (3DGCN) [428] https://github.com/FIBLAB/3D-DGCN
+Adaptive graph convolutional recurrent network (AGCRN) [8] https://github.com/LeiBAI/AGCRN
+Adaptive generalized PageRank graph neural network (AGP-GNN) [120] https://github.com/guoxiaoyuatbjtu/agp-gnn
+Adaptive multigraph convolutional networks (AMGCN) [208] https://github.com/hfimmortal/AMGCN
+Asynchronous spatio-temporal graph convolutional network (ASeer) [493] https://github.com/usail-hkust/ASeer
+Automated dilated spatio-temporal synchronous graph network (Auto-DSTSGN) [173] https://github.com/jinguangyin/Auto-DSTSGN
+BigST [126] https://github.com/usail-hkust/BigST
+Dynamic multi-view graph neural network for citywide traffic inference (CTVI+) [62] https://github.com/dsj96/TKDD
+Decoupled dynamic spatial-temporal graph neural network (D2STGNN) [342] https://github.com/zezhishao/D2STGNN
+Domain adversarial spatial-temporal network (DastNet) [370] https://github.com/YihongT/DASTNet
+Dynamic graph convolutional recurrent network (DGCRN) [212] https://github.com/tsinghua-fib-lab/Traffic-Benchmark
+Deep Kalman filtering network (DKFN) [33] https://github.com/Fanglanc/DKFN
+Dynamic and multi-faceted spatio-temporal graph convolution network (DMSTGCN) [127] https://github.com/liangzhehan/DMSTGCN
+Dynamic spatial-temporal aware graph neural network (DSTAGNN) [203] https://github.com/SYLan2019/DSTAGNN
+Graph bootstrap (Gboot) [95] https://github.com/wangzz-yyzz/Gboot
+Hierarchical information enhanced spatio-temporal (HIEST) [281] https://github.com/VAN-QIAN/CIKM23-HIEST
+Multi-step dependency relation (MSDR) [259] https://github.com/dcliu99/MSDR
+Multi-range spatial-temporal transformer network (MultiSPANS) [530] https://github.com/SELGroup/MultiSPANS
+Progressive graph convolutional network (PGCN) [348] https://github.com/yuyolshin/PGCN
+Scalable adaptive graph diffusion forecasting network (SAGDFN) [171] https://github.com/JIANGYUE61610306/SAGDFN
+Scale-aware neural architecture search framework for MTS forecasting (SNAS4MTF) [32] https://github.com/shangzongjiang/SNAS4MTF
+Spatio-temporal adaptive gated graph convolution network (STAG-GCN) [275] https://github.com/RobinLu1209/STAG-GCN
+Spatial-temporal convolutional graph attention network (ST-CGA) [496] https://github.com/jbdj-star/cga
+Spatio-temporal causal graph attention network (STCGAT) [511] https://github.com/zsqZZU/STCGAT/tree/v1.0.0
+STGNN enhanced by a scalable time series pre-training model (STEP) [341] https://github.com/zezhishao/STEP
+Spatial-temporal graph ordinary differential equation networks (STGODE) [84] https://github.com/square-coder/STGODE
+Spatio-temporal graph mixformer (STGM) [200] https://github.com/mouradost/stgm
+Spatio-temporal graph prompting (STGP) [144] https://github.com/hjf1997/STGP
+Space-time separable graph convolutional network (STSGCN) [419] https://github.com/FraLuca/STSGCN
+Spatio-temporal network with discrete wavelet transform (STWave) [81] https://github.com/LMissher/STWave
+Sparse unstructured spatio temporal reconstruction (SUSTeR) [411] https://github.com/ywoelker/SUSTeR
+Transferable graphs for traffics (TransGTR) [178] https://github.com/KL4805/TransGTR/
+Two-level resolution neural network (TwoResNet) [199] https://github.com/semink/TwoResNet
+Wavelet dynamic spatiotemporal aware graph neural network (W-DSTAGNN) [163] https://github.com/yash-jakhmola/w-dstagnn
 5.5.4 Benchmark models
@@
 Category Model Used by
-Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_93], [art_94], [art_96], [art_98], [art_100], [art_103], [art_106], [art_109], [art_116], [art_123], [art_126], [art_129], [art_132], [art_135], [art_406], [art_408], [art_419], [art_421], [art_424], [art_430], [art_433], [art_438], [art_444], [art_446], [art_448], [art_449], [art_451], [art_452], [art_453]
-Historical average (HA) [art_93], [art_94], [art_96], [art_98], [art_100], [art_102], [art_103], [art_109], [art_110], [art_120], [art_123], [art_125], [art_131], [art_132], [art_135], [art_406], [art_408], [art_420], [art_424], [art_433], [art_435], [art_438], [art_441], [art_443], [art_446], [art_447], [art_449], [art_451], [art_452], [art_453]
-Vector autoregression (VAR) [art_99], [art_100], [art_126], [art_131], [art_132], [art_406], [art_419], [art_424], [art_435], [art_443], [art_445], [art_447], [art_453]
-Non-GNN machine learning methods Convolutional long-short term memory network (ConvLSTM) [ConvLSTM] [art_98], [art_446]
-Deep multi-view spatial-temporal network (DMVST-Net) [DMVST-Net] [art_96], [art_451]
-Feed-forward neural network (FNN) [art_93], [art_97], [art_100], [art_105], [art_114], [art_120], [art_121], [art_124], [art_126], [art_408], [art_420], [art_431], [art_438]
-Gated recurrent unit (GRU) [art_98], [art_99], [art_106], [art_121], [art_123], [art_124], [3], [art_131], [art_132], [art_135], [art_425], [art_452]
-Long-short term memory network (LSTM) [art_93], [art_96], [art_98], [art_99], [art_100], [art_106], [art_108], [art_109], [art_113], [art_115], [art_116], [art_120], [art_121], [art_122], [art_125], [art_126], [3], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_420], [art_421], [art_424], [art_425], [art_430], [art_433], [art_435], [art_438], [art_441], [art_442], [art_443], [art_444], [art_445], [art_449], [art_453]
-Long-short term time series network (LSTNet) [LSTNet] [art_102], [art_110]
-Seq2Seq [Seq2seq] [art_105], [art_124], [art_451]
-Spatial and temporal normalization for multivariate time series forecasting (ST-Norm) [ST-Norm] [art_116], [art_117], [art_406], [art_422]
-Spatial-temporal identity (STID) [STID] [art_116], [art_422], [art_433], [art_434]
-Support vector regression (SVR) [art_99], [art_100], [art_103], [art_109], [art_119], [art_123], [art_129], [art_130], [art_135], [art_408], [art_419], [art_421], [art_435], [art_444], [art_445], [art_453]
-Temporal convolutional network (TCN) [TCN] [art_446], [art_453]
-Temporal 2D-variation modeling for general time series analysis (TimesNet) [TimesNet] [art_434], [art_435]
-Transformer [attention] [art_424], [art_454]
-Extreme gradient boosting (XGBoost) [XGBoost] [art_97], [art_114]
-GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_98], [art_100], [art_110], [art_117], [art_124], [art_125], [art_130], [art_131], [art_406], [art_419], [art_435], [art_441], [art_447], [art_448], [art_452], [art_453], [art_454]
-Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_93], [art_98], [art_99], [art_100], [art_108], [art_110], [art_113], [art_115], [art_117], [art_121], [art_125], [art_126], [3], [art_130], [art_131], [art_132], [art_406], [art_421], [art_424], [art_427], [art_432], [art_433], [art_435], [art_442], [art_445], [art_446], [art_447], [art_448], [art_449], [art_453], [art_454]
-Attention based Spatial-Temporal Graph Neural Network (ASTGNN) [ASTGNN] [art_422], [art_435], [art_446]
-Automated spatio-temporal graph (AutoSTG) [AutoSTG] [art_108], [art_408]
-Dual-stage attention-based recurrent neural network (DA-RNN) [DA-RNN] [art_96], [art_105]
-Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_93], [art_94], [art_96], [art_100], [art_103], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_122], [art_124], [art_125], [art_126], [art_128], [art_129], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_420], [art_421], [art_422], [art_424], [art_425], [art_427], [art_430], [art_432], [art_433], [art_435], [art_438], [art_441], [art_442], [art_444], [art_445], [art_447], [art_448], [art_451], [art_452], [art_453]
-Dynamic graph convolutional recurrent network (DGCRN) [art_100] [art_110], [art_406], [art_444]
-Dynamic and Multi-faceted SpatioTemporal Graph Convolution Network (DMSTGCN) [DMSTGCN] [art_441], [art_444]
-Dynamic spatial-temporal aware graph neural network (DSTAGNN) [DSTAGNN] [art_125], [art_419], [art_424], [art_447], [art_448]
-Fully connected gated graph architecture (FC-GAGA) [FC-GAGA] [art_116], [art_408], [art_427]
-GeoMAN [GeoMAN] [art_96], [art_132]
-Graph convolutional network (GCN) [art_104], [art_135], [art_431]
-Graph multi-attention network (GMAN) [GMAN] [art_100], [art_109], [art_110], [art_116], [art_122], [art_128], [art_408], [art_419], [art_422], [art_427], [art_430], [art_444], [art_451]
-Graph WaveNet (GWN) [GWN] [art_93], [art_100], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_122], [art_124], [art_128], [art_406], [art_408], [art_419], [art_420], [art_422], [art_424], [art_430], [art_432], [art_433], [art_434], [art_435], [art_441], [art_444], [art_447], [art_448], [art_449], [art_452], [art_453], [art_454]
-Discrete graph structure learning for time series (GTS) [GTS] [art_420], [art_433]
-Long short-term graph convolutional networks (LSGCN) [LSGCN] [art_99], [art_432], [art_446], [art_453]
-Multi-range attentive bicomponent graph convolutional network (MRA-BGCN) [MRA-BGCN] [art_109], [art_116], [art_408]
-Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_100], [art_109], [art_110], [art_113], [art_116], [art_117], [art_122], [art_124], [art_128], [art_406], [art_408], [art_420], [art_422], [art_444], [art_445]
-Multi-view graph convolutional network (MVGCN) [MVGCN] [art_94], [art_96]
-Spectral temporal graph neural network for multivariate time-series forecasting (StemGNN) [StemGNN] [art_406], [art_422]
-Spatial-temporal fusion graph neural network (STFGNN) [STFGNN] [art_98], [art_108], [art_115], [art_117], [art_125], [art_130], [art_406], [art_435], [art_445], [art_452], [art_453]
-Spatial-temporal graph to sequence model (STG2Seq) [STG2Seq] [art_99], [art_424]
-Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_93], [art_94], [art_96], [art_99], [art_100], [art_106], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_121], [art_122], [art_124], [art_126], [3], [art_128], [art_129], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_420], [art_421], [art_422], [art_424], [art_425], [art_433], [art_435], [art_438], [art_442], [art_443], [art_444], [art_445], [art_448], [art_449], [art_451], [art_452], [art_453], [art_454]
-Spatial temporal graph neural network (STGNN) [STGNN] [art_94], [art_103]
-Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_99], [art_108], [art_117], [art_125], [art_406], [art_424], [art_447], [art_453]
-Spatio-temporal graph attention network (ST-GRAT) [ST-GRAT] [art_116], [art_444]
-Deep-meta-learning based model (ST-MetaNet) [ST-MetaNet] [art_96], [art_100], [art_109], [art_116], [art_408], [art_451]
-Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_98], [art_99], [art_100], [art_108], [art_115], [art_117], [art_125], [art_130], [art_131], [art_406], [art_419], [art_424], [art_435], [art_441], [art_442], [art_443], [art_453], [art_454]
-Temporal graph convolutional network (T-GCN) [T-GCN] [art_103], [art_113], [art_121], [art_123], [art_421], [art_433]
-Time zigzags at graph convolutional networks for time series forecasting (Z-GCNETs) [Z-GCNETs] [art_125], [art_424], [art_441]
+Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [274], [428], [211], [220], [212], [263], [262], [217], [120], [39], [25], [27], [169], [227], [32], [163], [385], [407], [199], [242], [406], [208], [89], [388], [82], [171], [127], [276], [393], [178], [81], [524], [84], [144], [85], [213], [222], [33], [320], [219], [496], [145], [275]
+Historical average (HA) [274], [428], [211], [220], [212], [402], [263], [217], [200], [490], [511], [69], [27], [169], [227], [32], [159], [407], [242], [510], [406], [348], [419], [89], [126], [388], [82], [341], [127], [134], [493], [276], [135], [393], [81], [342], [524], [259], [8], [144], [222], [370], [33], [123], [320], [219], [275], [301]
+$k$-nearest neighbors (KNN) [62], [144], [275]
+Naive [493], [123], [301]
+Vector autoregression (VAR) [218], [212], [39], [69], [27], [227], [163], [407], [510], [419], [250], [126], [82], [341], [171], [127], [393], [81], [342], [8], [530], [213], [461], [145], [301]
+Non-GNN machine learning methods Adaptive recurrent neural network (AdaRNN) [75] [276], [144]
+Convolutional long-short term memory network (ConvLSTM) [346] [220], [89], [123], [301]
+Deep multi-view spatial-temporal network (DMVST-Net) [447] [211], [388], [496]
+Feed-forward neural network (FNN) [274], [62], [212], [431], [245], [490], [506], [278], [39], [32], [159], [330], [406], [213], [214], [275], [301]
+Gated recurrent unit (GRU) [220], [218], [262], [506], [278], [69], [27], [169], [312], [493], [8], [370], [219]
+Long-short term memory network (LSTM) [274], [211], [220], [218], [212], [262], [173], [217], [198], [438], [120], [490], [506], [181], [511], [39], [112], [69], [27], [227], [32], [163], [159], [385], [407], [312], [199], [242], [510], [406], [348], [378], [419], [208], [250], [82], [341], [171], [134], [493], [135], [393], [81], [342], [524], [203], [259], [80], [530], [85], [213], [387], [222], [33], [219], [214], [461], [496], [145], [275], [301]
+Long-short term time series network (LSTNet) [202] [402], [200]
+Multitask deep-learning (MDL) [482] [388], [523]
+Recurrent neural network (RNN) [388], [286], [135], [393]
+Seq2Seq [365] [431], [278], [388]
+Spatial-temporal dynamic network (STDN) [446] [388], [496]
+Spatial-temporal identity (STID) [340] [120], [466], [242], [420], [524], [422], [387]
+Spatial and temporal normalization for multivariate time series forecasting (ST-Norm) [67] [120], [111], [227], [466]
+Spatio-temporal residual networks (ST-ResNet) [480] [388], [496]
+Support vector regression (SVR) [218], [212], [263], [217], [251], [25], [112], [169], [32], [163], [385], [510], [208], [250], [82], [341], [171], [393], [81], [342], [530], [422], [213], [370], [33], [320], [461], [281], [496]
+Temporal convolutional network (TCN) [9] [89], [82], [493], [81], [203], [80], [63]
+Temporal 2D-variation modeling for general time series analysis (TimesNet) [413] [420], [510]
+Transformer [375] [407], [193], [523]
+Extreme gradient boosting (XGBoost) [43] [62], [245], [127]
+GNN methods Adaptive graph convolutional recurrent network (AGCRN) [8] [220], [212], [200], [111], [278], [511], [112], [69], [227], [163], [510], [348], [126], [82], [193], [171], [81], [524], [203], [80], [95], [213], [63], [370], [33], [214], [145]
+Attention-based spatial-temporal graph convolutional network (ASTGCN) [116] [274], [220], [218], [212], [173], [200], [198], [438], [111], [506], [511], [39], [112], [69], [27], [227], [385], [407], [13], [352], [242], [510], [378], [250], [89], [126], [82], [193], [341], [171], [81], [342], [524], [203], [84], [259], [8], [80], [95], [530], [422], [85], [213], [63], [33], [123], [320], [214], [461], [145], [275]
+Attention based spatial-temporal graph neural network (ASTGNN) [117] [466], [510], [89]
+Automated spatio-temporal graph (AutoSTG) [305] [173], [32], [213], [33], [214]
+Coupled layer-wise convolutional recurrent neural network (CCRNN) [449] [259], [422]
+Decoupled dynamic spatial-temporal graph neural network (D2STGNN) [342] [171], [411]
+Dual-stage attention-based recurrent neural network (DA-RNN) [318] [211], [431]
+Diffusion convolutional recurrent neural network (DCRNN) [237] [274], [428], [211], [212], [263], [173], [217], [200], [198], [438], [120], [111], [490], [181], [278], [511], [39], [22], [25], [112], [69], [27], [227], [32], [163], [159], [385], [466], [407], [312], [13], [199], [352], [242], [510], [406], [348], [378], [208], [250], [126], [388], [82], [341], [171], [127], [493], [393], [81], [342], [524], [203], [84], [122], [259], [8], [144], [80], [95], [530], [422], [85], [213], [387], [222], [370], [33], [123], [214], [461], [281], [145], [275], [308], [301]
+Deep spatial-temporal network (DeepST) [481] [388], [496]
+Dynamic graph convolutional recurrent network (DGCRN) [212] [200], [227], [208], [342]
+Dynamic and multi-faceted spatiotemporal graph convolution network (DMSTGCN) [127] [348], [208]
+Dynamic spatial-temporal aware graph neural network (DSTAGNN) [203] [511], [163], [407], [126], [144], [80], [95], [63], [461], [145]
+Fully connected gated graph architecture (FC-GAGA) [470] [120], [32], [13]
+Gated attention networks for learning on large and spatiotemporal graphs (GaAN) [479] [393], [308]
+Graph convolutional network (GCN) [489], [169], [330], [135], [370]
+GeoMAN [246] [211], [27]
+Graph multi-attention network (GMAN) [513] [212], [217], [200], [120], [181], [22], [32], [163], [466], [13], [199], [208], [388], [341], [171], [127], [342], [524], [530], [85], [213], [222], [33], [214], [461], [281], [145], [308]
+Discrete graph structure learning for time series (GTS) [338] [159], [242], [341], [171], [178], [524], [387], [281]
+Graph WaveNet (GWN) [425] [274], [212], [173], [217], [200], [198], [438], [120], [111], [181], [278], [22], [227], [32], [163], [159], [466], [407], [199], [352], [242], [420], [510], [348], [208], [126], [82], [193], [341], [171], [127], [493], [178], [81], [342], [524], [84], [122], [259], [144], [95], [530], [422], [85], [213], [387], [523], [222], [33], [214], [461], [281], [145], [308], [301]
+Hierarchical graph convolution network (HGCN) [115] [385], [281]
+Long short-term graph convolutional networks (LSGCN) [154] [218], [352], [89], [82], [81]
+Multi-graph convolutional networks (MGCN) [30] [245], [134]
+Multi-range attentive bicomponent graph convolutional network (MRA-BGCN) [44] [217], [120], [32], [222]
+Multivariate time series forecasting with graph neural network (MTGNN) [424] [212], [217], [200], [198], [120], [111], [181], [278], [22], [227], [32], [159], [466], [208], [250], [341], [171], [127], [342], [524], [530], [213], [222], [33], [214], [281]
+Multi-view graph convolutional network (MVGCN) [359] [428], [211]
+Propagation delay-aware dynamic long-range transformer (PDFormer) [165] [242], [493], [95], [461]
+a Spatio-temporal adaptive embedding transformer (STAEformer) [260] [493], [63]
+Spectral temporal graph neural network for multivariate time-series forecasting (StemGNN) [26] [227], [466], [387], [123]
+STGNN enhanced by a scalable time series Pre-training model (STEP) [341] [171], [144], [387]
+Spatial-temporal fusion graph neural network (STFGNN) [228] [220], [173], [438], [111], [511], [112], [227], [510], [250], [82], [81], [203], [259], [80], [95], [213], [33], [214], [145]
+Spatial-temporal graph to sequence model (STG2Seq) [7] [218], [407], [259], [80], [422], [523]
+Spatio-temporal graph convolutional network (STGCN) [457] [274], [428], [211], [218], [212], [262], [173], [217], [200], [198], [438], [120], [111], [490], [506], [181], [278], [39], [22], [25], [112], [69], [27], [227], [32], [163], [159], [385], [466], [407], [312], [242], [510], [406], [378], [419], [208], [250], [388], [82], [193], [171], [81], [203], [84], [122], [259], [8], [80], [95], [530], [85], [213], [387], [222], [370], [33], [214], [461], [281], [496], [145], [275], [308], [301], [411]
+Spatio-temporal graph few shot learning (ST-GFSL) [276] [178], [144]
+Spatial temporal graph neural network (STGNN) [393] [428], [263], [393]
+Spatial-temporal graph ODE networks (STGODE) [84] [218], [173], [111], [511], [227], [407], [126], [82], [81], [203], [122], [80], [95], [422], [213], [370], [33]
+Spatio-temporal graph attention network (ST-GRAT) [308] [120], [208]
+Deep-meta-learning based model (ST-MetaNet) [306] [211], [212], [217], [120], [32], [388], [524], [213], [222], [33], [214], [496]
+Spatiotemporal multi-graph convolution network (ST-MGCN) [101] [388], [496]
+Spatio-temporal synchronous graph convolutional network (STSGCN) [353] [220], [218], [212], [173], [438], [111], [511], [112], [69], [227], [163], [407], [510], [348], [378], [419], [82], [193], [341], [171], [81], [342], [524], [203], [84], [259], [8], [80], [95], [213], [33], [214], [461], [145]
+Temporal graph convolutional network (T-GCN) [508] [263], [198], [506], [385], [242], [393], [530], [422], [370], [320], [281]
+Time zigzags at graph convolutional networks for time series forecasting (Z-GCNETs) [50] [511], [407], [348], [203], [80], [145]
 As for the basic mathematical and statistical methods, two widely used benchmarks are the ARIMA model and the historical average model (HA), which uses the average of past observations as the forecast for future values.
@@
 As for the non-GNN machine learning models, LSTM networks are widely recognized as a type of recurrent neural network capable of handling long-term dependencies in traffic data.
-As for GNN-based models instead, the three most used models are the spatio-temporal graph convolutional network (STGCN) [STGCN], the diffusion convolutional recurrent neural network (DCRNN) [DCRNN], and the attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN], all developed specifically for the traffic forecasting problem.
+As for GNN-based models instead, the most used models are the spatio-temporal graph convolutional network (STGCN) [457], the diffusion convolutional recurrent neural network (DCRNN) [237], Graph WaveNet (GWN) [425], and the attention-based spatial-temporal graph convolutional network (ASTGCN) [116], all developed specifically for the traffic forecasting problem.
 The STGCN model proposed by Yu et al.
@@
 The code can be accessed at https://github.com/liyaguang/DCRNN.
+The GWN model was introduced by Wu et al.
+in 2021 as a novel approach to spatio-temporal forecasting, particularly in traffic tasks.
+It is known for its ability to learn an adaptive dependency matrix from node embeddings.
+The model combines graph convolutional and temporal convolutional layers to effectively capture spatio-temporal dependencies.
+The source code is available at https://github.com/nnzhan/Graph-WaveNet.
 The ASTGCN model proposed by Guo et al.
@@
 For a more comprehensive list of source codes for the most dated benchmark GNN models in the traffic domain see Ref.
-[review_traffico_jiang].
+[168].
 5.5.5 Results
@@
 However, there are a limited but interesting number of situations in which some simple statistical and mathematical models demonstrate comparable accuracy, as it can be observed in Tabs.
-LABEL:t:risultati_metr-la and LABEL:t:risultati_pems-bay [art_452].
+LABEL:t:risultati_metr-la and LABEL:t:risultati_pems-bay.
 For the comparison and evaluation of these models, the most commonly used error metrics are MAE, RMSE, and MAPE.
@@
 AGCRN 2.87 7.70 5.58 3.23 9.00 6.58 3.62 10.38 7.51
-AGP-GNN [art_116] 2.71 7.21 5.25 3.04 8.15 6.20 3.41 9.70 7.20
-AMGCN [art_444] 2.52 6.13 4.48 2.97 7.70 5.35 3.38 8.93 6.21
+AGP-GNN [120] 2.71 7.21 5.25 3.04 8.15 6.20 3.41 9.70 7.20
+AMGCN [208] 2.52 6.13 4.48 2.97 7.70 5.35 3.38 8.93 6.21
 ARIMA 3.99 9.60 8.12 5.15 12.70 10.45 6.90 17.40 13.23
@@
 AutoSTG 2.70 - 5.16 3.06 - 6.17 3.47 - 7.27
-CASTGNN [art_442] 2.88 6.67 6.44 3.25 7.91 7.56 3.70 9.76 8.92
+AutoSTS [213] 2.57 6.55 4.93 2.89 7.85 5.87 3.28 9.43 6.86
+CASTGNN [378] 2.88 6.67 6.44 3.25 7.91 7.56 3.70 9.76 8.92
+D2STGNN [342] 2.56 6.48 4.88 2.90 7.78 5.89 3.35 9.40 7.03
 DCRNN 2.77 7.30 5.38 3.15 8.80 6.45 3.60 10.50 7.60
-DGCRN [art_100] 2.62 6.63 5.01 2.99 8.02 6.05 3.44 9.73 7.19
+DetectorNet [222] - - - 3.06 8.12 6.08 3.40 9.60 6.98
+DGCRN [212] 2.62 6.63 5.01 2.99 8.02 6.05 3.44 9.73 7.19
 DMSTGCN 2.85 7.54 5.54 3.26 9.19 6.56 3.72 10.96 7.55
-DST-GCNN [art_126] 2.68 7.20 5.35 3.01 8.50 6.23 3.41 10.30 7.47
-DualGraph [art_438] 2.62 7.10 5.36 2.83 7.90 5.89 3.15 9.10 6.67
-DyGCN-LSTM [art_113] 2.64 6.35 4.64 3.24 8.04 5.24 3.53 9.30 6.53
+DST-GCNN [39] 2.68 7.20 5.35 3.01 8.50 6.23 3.41 10.30 7.47
+DualGraph [406] 2.62 7.10 5.36 2.83 7.90 5.89 3.15 9.10 6.67
+DyGCN-LSTM [198] 2.64 6.35 4.64 3.24 8.04 5.24 3.53 9.30 6.53
 DyHSL 2.58 6.47 4.98 2.95 8.05 6.13 3.41 9.41 7.03
@@
 FNN 3.99 9.90 7.94 4.23 12.90 8.17 4.49 14.00 8.69
-GA-LSTM [art_120] 2.30 7.01 4.97 2.98 8.50 6.00 3.55 9.97 7.23
+GA-LSTM [490] 2.30 7.01 4.97 2.98 8.50 6.00 3.55 9.97 7.23
+GaAN 2.71 6.99 5.24 3.12 8.56 6.36 3.64 10.62 7.65
 GGRU 2.71 6.99 5.24 3.12 8.56 6.36 3.64 10.62 7.65
-GL-STGTN [art_433] 2.61 6.94 5.08 2.96 8.23 5.97 3.37 9.81 6.99
+GL-STGTN [242] 2.61 6.94 5.08 2.96 8.23 5.97 3.37 9.81 6.99
 GMAN 2.77 7.25 5.48 3.07 8.35 6.34 3.40 9.72 7.22
-GTS 2.67 7.21 5.27 3.04 8.41 6.25 3.46 9.98 7.31
+GTS [338] 2.64 6.80 4.95 3.01 8.20 5.85 3.41 9.90 6.74
 GWN 2.69 6.90 5.15 3.07 8.37 6.22 3.53 10.01 7.37
@@
 LSTM 3.44 9.60 6.30 3.77 10.90 7.23 4.37 13.20 8.69
-MCFGNN [art_86] 2.31 5.86 5.29 2.66 6.89 6.27 3.05 7.65 7.20
+MCFGNN [52] 2.31 5.86 5.29 2.66 6.89 6.27 3.05 7.65 7.20
 MRA-BGCN 2.67 6.80 5.12 3.06 8.30 6.17 3.49 10.00 7.30
-MTGNN 2.69 6.86 5.18 3.05 8.19 6.17 3.49 9.87 7.23
-MTGODE [art_75] 2.66 6.87 5.10 3.00 8.19 6.05 3.39 9.80 7.05
-MVST-GNN [art_109] 2.70 5.19 6.97 3.06 8.12 6.08 3.40 9.60 6.98
-PAHNN [art_435] 2.51 6.34 4.82 2.87 7.63 5.73 3.25 9.32 6.67
+MTGNN [424] 2.69 6.86 5.18 3.05 8.19 6.17 3.49 9.87 7.23
+MTGODE [177] 2.66 6.87 5.10 3.00 8.19 6.05 3.39 9.80 7.05
+MVST-GNN [217] 2.70 5.19 6.97 3.06 8.12 6.08 3.40 9.60 6.98
+PAHNN [510] 2.51 6.34 4.82 2.87 7.63 5.73 3.25 9.32 6.67
 PDFormer 2.83 7.77 5.45 3.20 9.19 6.46 3.62 10.91 7.47
-PGCN [art_441] 2.70 6.98 5.16 3.08 8.38 6.22 3.54 9.94 7.36
+PGCN [348] 2.70 6.98 5.16 3.08 8.38 6.22 3.54 9.94 7.36
+SAGDFN [171] 2.56 6.50 5.00 2.94 7.90 6.05 3.37 9.50 7.17
 SDGCN 2.76 7.16 5.38 3.15 8.53 6.46 3.61 9.87 7.43
 SLCNN 2.53 6.70 5.18 2.88 8.00 6.15 3.30 9.70 7.20
-SNAS4MTF [art_408] 2.68 6.90 5.16 3.03 8.25 6.12 3.43 9.82 7.15
-STAG-GCN [art_93] 2.67 7.00 5.23 3.07 8.26 6.15 3.50 9.93 7.24
+SNAS4MTF [32] 2.68 6.90 5.16 3.03 8.25 6.12 3.43 9.82 7.15
+STAG-GCN [274] 2.67 7.00 5.23 3.07 8.26 6.15 3.50 9.93 7.24
 STAEFormer 2.65 6.85 5.11 2.97 8.13 6.00 3.34 9.70 7.02
+STEP [341] 2.61 6.60 4.98 2.96 7.96 5.97 3.37 9.61 6.99
 STID 2.82 7.75 5.53 3.19 9.39 6.57 3.55 10.95 7.55
@@
 ST-MetaNet 2.69 6.91 5.17 3.10 8.57 6.28 3.59 10.63 7.52
+STSD [524] 2.60 6.55 4.96 2.96 7.93 5.98 3.37 9.51 7.03
 STSGCN 3.31 8.06 7.62 4.13 10.29 9.77 5.06 12.91 11.66
 SVR 3.99 9.30 8.45 5.05 12.10 10.87 6.72 16.70 13.76
-TF-GAN [art_122] 2.63 6.55 4.94 3.06 8.36 6.20 3.32 9.48 7.11
+TF-GAN [181] 2.63 6.55 4.94 3.06 8.36 6.20 3.32 9.48 7.11
 T-GCN 3.03 7.81 5.26 3.52 9.45 6.12 4.30 11.80 7.31
@@
 TimesNet 2.78 7.52 5.49 3.36 9.43 6.41 4.52 11.37 9.66
-TwoResNet [art_430] 2.65 6.78 5.08 3.01 8.14 6.07 3.39 9.71 7.08
+TwoResNet [199] 2.65 6.78 5.08 3.01 8.14 6.07 3.39 9.71 7.08
 WaveNet 2.99 8.04 5.89 3.59 10.25 7.28 4.45 13.62 8.93
@@
 Clearly, the results on the METR-LA dataset are not uniform, and the best performing models vary depending on the metric and the time horizon.
-However, the multichannel fusion graph neural network (MCFGNN) [art_86] from the “Generic” group consistently exhibits low errors across different time horizons.
+However, the multichannel fusion graph neural network (MCFGNN) [52] from the “Generic” group consistently exhibits low errors across different time horizons.
 The MCFGNN model, proposed by Chen et al.
@@
 AGCRN 1.37 2.94 2.87 1.69 3.87 3.85 1.96 4.64 4.54
-AGP-GNN [art_116] 1.33 2.79 2.86 1.61 3.66 3.67 1.86 4.30 4.34
-AMGCN [art_444] 1.22 2.63 2.34 1.53 3.55 3.13 1.93 4.66 3.95
+AGP-GNN [120] 1.33 2.79 2.86 1.61 3.66 3.67 1.86 4.30 4.34
+AMGCN [208] 1.22 2.63 2.34 1.53 3.55 3.13 1.93 4.66 3.95
 ARIMA 1.62 3.50 3.30 2.33 5.40 4.76 3.38 8.30 6.50
@@
 ASTGNN 1.34 3.84 2.75 1.65 3.72 3.74 1.94 4.59 4.51
+AutoSTG 1.33 2.79 2.78 1.63 3.66 3.62 1.91 4.57 4.42
+AutoSTS [213] 1.28 2.66 2.63 1.57 3.54 3.50 1.83 4.41 4.27
+D2STGNN [342] 1.24 2.58 2.60 1.55 3.49 3.52 1.85 4.37 4.30
 DCRNN 1.38 2.90 2.95 1.74 3.90 3.97 2.07 4.90 4.74
-DGCRN [art_100] 1.28 2.66 2.69 1.59 3.55 3.63 1.89 4.43 4.42
+DetectorNet [222] - - - 1.57 3.56 3.54 1.80 4.19 4.26
+DGCRN [212] 1.28 2.66 2.69 1.59 3.55 3.63 1.89 4.43 4.42
 DMSTG - - - - - - 1.91 4.51 4.43
@@
 FNN 2.20 5.19 4.42 2.30 5.43 4.63 2.46 5.89 4.98
-GL-STGTN [art_433] 1.26 2.79 2.73 1.55 3.63 3.58 1.80 4.38 4.30
+GL-STGTN [242] 1.26 2.79 2.73 1.55 3.63 3.58 1.80 4.38 4.30
 GMAN 1.34 2.81 2.82 1.62 3.62 3.72 1.86 4.31 4.32
-GSLNN [art_420] 1.29 2.71 2.66 1.64 3.67 3.74 1.93 4.51 4.46
-GTS 1.35 2.85 2.84 1.67 3.80 3.82 1.96 4.62 4.53
+GSLNN [159] 1.29 2.71 2.66 1.64 3.67 3.74 1.93 4.51 4.46
+GTS [338] 1.32 2.80 2.62 1.64 3.60 3.41 1.91 4.40 3.97
 GWN 1.30 2.73 2.74 1.63 3.67 3.70 1.95 4.63 4.52
+Informer 2.30 5.02 4.21 2.40 5.32 4.49 2.55 5.73 4.85
+LDS 1.33 2.80 2.81 1.67 3.80 3.80 1.99 4.80 4.59
 HA 1.89 4.16 4.30 2.50 5.62 5.82 3.31 7.65 7.54
@@
 MRA-BGCN 1.29 2.90 2.72 1.61 3.80 3.67 1.91 4.60 4.46
-MTGNN 1.32 2.77 2.79 1.65 3.69 3.74 1.94 4.53 4.49
-MTGODE [art_75] 1.29 2.72 2.73 1.61 3.61 3.66 1.88 4.39 4.31
-MVST-GNN [art_109] 1.30 2.73 2.74 1.57 3.56 3.54 1.80 4.19 4.26
-PAHNN [art_435] 1.27 2.60 2.65 1.52 3.48 3.54 1.83 4.35 4.27
+MTGNN [424] 1.32 2.77 2.79 1.65 3.69 3.74 1.94 4.53 4.49
+MTGODE [177] 1.29 2.72 2.73 1.61 3.61 3.66 1.88 4.39 4.31
+MVST-GNN [217] 1.30 2.73 2.74 1.57 3.56 3.54 1.80 4.19 4.26
+PAHNN [510] 1.27 2.60 2.65 1.52 3.48 3.54 1.83 4.35 4.27
 PDFormer 1.32 2.78 2.83 1.64 3.71 3.79 1.91 4.51 4.43
-PGCN [art_441] 1.30 2.72 2.73 1.62 3.63 3.67 1.92 4.45 4.45
+PGCN [348] 1.30 2.72 2.73 1.62 3.63 3.67 1.92 4.45 4.45
 RGSL - - - - - - 2.03 4.77 4.64
-RST-LTG [art_406] - - - - - - 1.88 4.32 4.31
+RST-LTG [227] - - - - - - 1.88 4.32 4.31
 SDGCN 1.35 2.89 2.88 1.69 3.87 3.91 1.99 4.69 4.59
 SLCNN 1.44 3.00 2.90 1.72 3.90 3.81 2.03 4.80 4.53
-SNAS4MTF [art_408] 1.30 2.72 2.77 1.61 3.63 3.68 1.89 4.44 4.36
+SNAS4MTF [32] 1.30 2.72 2.77 1.61 3.63 3.68 1.89 4.44 4.36
 STAEFormer 1.31 2.76 2.78 1.62 3.62 3.68 1.88 4.41 4.34
 StemGNN 1.44 3.08 3.12 1.93 4.54 4.38 2.57 6.55 5.88
+STEP [341] 1.26 2.59 2.73 1.55 3.43 3.58 1.79 4.18 4.20
 STID 1.30 2.73 2.81 1.62 3.68 3.72 1.89 4.47 4.40
@@
 STGCN 1.36 2.90 2.96 1.81 4.17 4.27 2.49 5.79 5.69
+STG-NCDE 1.38 2.91 2.93 1.71 3.91 3.84 2.03 4.82 4.58
 STGODE 1.43 2.99 2.88 1.84 3.84 3.90 2.30 4.61 4.89
-STG-NCDE 1.38 2.91 2.93 1.71 3.91 3.84 2.03 4.82 4.58
-STGRAT 1.29 2.67 2.71 1.61 3.63 3.69 1.95 4.64 4.54
+STGRAT [308] 1.29 2.67 2.71 1.61 3.63 3.69 1.95 4.64 4.54
 ST-MetaNet 1.36 2.82 2.90 1.76 4.00 4.02 2.20 5.45 5.06
 ST-Norm 1.34 2.82 2.88 1.67 3.75 3.83 1.96 4.62 4.52
+STSD [524] 1.28 2.64 2.69 1.60 3.54 3.63 1.88 4.35 4.36
 STSGCN 1.44 3.04 3.01 1.83 4.17 4.18 2.26 5.40 5.21
 SVR 1.85 3.80 3.59 2.48 5.50 5.18 3.28 8.00 7.08
-TD2MG2NN [art_422] 1.25 2.62 2.64 1.55 3.47 3.50 1.85 4.28 4.22
+TD2MG2NN [466] 1.25 2.62 2.64 1.55 3.47 3.50 1.85 4.28 4.22
 TGC-GRU 2.34 5.08 4.30 2.44 5.40 4.64 2.57 5.78 5.01
@@
 TimesNet 1.32 2.94 2.72 1.76 3.74 4.16 2.18 4.84 5.20
-TwoResNet [art_430] 1.30 2.70 2.73 1.61 3.59 3.69 1.89 4.40 4.41
+TPGNN [271] 1.26 2.56 2.64 1.65 3.47 3.65 2.05 4.40 4.58
+TwoResNet [199] 1.30 2.70 2.73 1.61 3.59 3.69 1.89 4.40 4.41
 VAR 1.74 3.60 3.16 2.32 5.00 4.25 2.93 6.50 5.44
@@
 The results on the the PEMS-BAY dataset vary depending on the metrics and the time horizon.
-The most accurate models are the adaptive multigraph convolutional networks (AMGCN) [art_444], the multi view spatial-temporal graph neural network (MVST-GNN) [art_109], and the temporal decomposition based dynamic multi granularity graph neural network (TD2MG2NN) [art_422].
+The most accurate models are the adaptive multigraph convolutional networks (AMGCN) [208], and the spatio-temporal GNN enhanced by a scalable time series pre-training model (STEP) [341].
 The AMGCN model was proposed by Li et al.
@@
 Spatio-temporal features are finally extracted through a stack of graph convolutional modules with residual connections, and a time-mixing strategy.
+The STEP model was proposed by Shao et a.
+in 2021 as a framework to enhance spatio-temporal GNNs through a pre-training approach.
+Specifically, the pre-training model is designed to learn temporal patterns from long-term historical time series data and generate segment-level representations.
+These representations, which capture rich contextual information, are then integrated into a spatio-temporal GNN to improve its performance on short-term forecasting tasks.
+To demonstrate the effectiveness of the framework, the authors use Graph WaveNet as a representative backbone model within STEP.
+Other models that, overall, achieve good results are the multi view spatial-temporal graph neural network (MVST-GNN) [217], and the temporal decomposition based dynamic multi granularity graph neural network (TD2MG2NN) [466].
 The MVST-GNN model, proposed by Li et al.
@@
 Metrics MAE MAPE RMSE
+AdaSTNet 15.49 14.84 27.57
 AGCRN 15.98 15.23 28.25
-ARIMA 35.31 33.78 47.59
+ARIMA 35.41 33.78 47.59
 ASTGCN 17.69 19.40 29.66
 ASTGNN 14.78 14.79 25.00
-Auto-DSTSGN [art_108] 14.59 14.22 25.17
+Auto-DSTSGN [173] 14.59 14.22 25.17
 AutoSTG 16.27 16.10 27.63
-DCGCN [art_117] 15.29 - 25.98
+AutoSTS [213] 14.61 14.18 24.71
+DCGCN [111] 15.29 - 25.98
 DCRNN 18.18 18.91 30.31
-DSTAGNN 15.57 14.68 27.21
+DeepGLO 17.25 19.27 23.25
+DeepState 15.59 18.69 20.21
+DSTAGNN [203] 15.57 14.68 27.21
 DyHSL 15.49 14.38 27.06
-Graph WaveNet 19.85 19.31 32.94
+FourierGNN 17.27 15.88 27.20
+Gboot [95] 15.43 14.51 26.42
+GMAN 15.52 15.19 26.53
+GMSDR [259] 15.78 15.33 26.82
+GWN 19.85 19.31 32.94
 HA 31.58 33.78 52.39
@@
 LSTM 21.33 23.33 35.11
-PAHNN [art_435] 14.02 13.19 23.67
-STAGCN [art_130] 15.40 14.48 26.23
-STCGNN [art_115] 15.81 14.59 27.23
+LSTNet 19.07 17.73 29.67
+N-BEATS 18.45 18.35 31.23
+PAHNN [510] 14.02 13.19 23.67
+PDFormer 21.82 21.47 36.75
+PYNet [145] 14.94 14.94 25.27
+SFM 17.67 18.33 30.01
+SPGCL 23.31 21.88 37.37
+STAGCN [112] 15.40 14.48 26.23
+STCGNN [438] 15.81 14.59 27.23
+StemGNN 14.32 16.24 21.64
 STFGCN 16.77 16.30 28.34
+STGBN [80] 15.22 14.78 27.09
 STGCN 17.49 17.15 30.12
+STG2SEC 19.03 21.55 29.73
 STGODE 16.53 16.68 27.79
 STSGCN 17.48 16.78 29.21
-STWave+ [art_453] 14.71 14.88 26.31
+STWave [81] 14.93 15.05 26.50
+STWave+ [82] 14.71 14.88 26.31
 SVR 21.97 21.51 35.29
+TAMP-S2GCNets [51] 13.91 13.40 23.77
 TCN 19.32 19.93 33.55
@@
 VAR 23.65 24.51 38.26
-W-DSTAGNN [art_419] 15.31 14.49 26.59
+W-DSTAGNN [163] 15.31 14.49 26.59
+Z-GCNETs 14.20 13.88 25.29
 […]
@@
 Metrics MAE MAPE RMSE
-AGCRN 19.83 12.97 32.26
+AdaSTNet 19.25 12.77 31.26
+AGCRN [8] 19.83 12.97 32.26
 ARIMA 33.73 24.18 48.80
@@
 ASTGNN 18.60 12.36 30.91
-Auto-DSTSGN [art_108] 18.85 13.21 30.48
+Auto-DSTSGN [173] 18.85 13.21 30.48
 Autoformer 23.76 18.01 36.59
 AutoSTG 20.38 14.12 32.51
+AutoSTS [213] 18.76 12.84 30.31
+CL4ST [368] 18.49 12.00 30.17
 Crossformer 20.40 14.62 32.79
-DCGCN [art_117] 20.28 - 31.65
+DCGCN [111] 20.28 - 31.65
 DCRNN 24.70 17.12 38.12
 DGCRN 19.04 12.80 30.82
-DMGF-Net [art_99] 20.59 13.63 32.43
+DMGF-Net [218] 20.59 13.63 32.43
 DSA-NET 22.79 16.03 35.77
-DSTAGNN 19.30 12.70 31.46
-DSTGN [art_73] 18.61 12.31 30.79
-DSTIGNN [art_74] 18.41 12.45 29.97
+DSTAGNN [203] 19.30 12.70 31.46
+DSTGN [241] 18.61 12.31 30.79
+DSTIGNN [99] 18.41 12.45 29.97
 DyHSL 17.66 12.42 29.46
@@
 FEDFormer 22.86 16.04 35.07
-GL-STGTN [art_433] 18.20 12.48 29.71
+FOGS 19.74 13.05 31.66
+FourierGNN 22.98 15.14 36.23
+Gboot [95] 19.28 12.58 31.02
+GL-STGTN [242] 18.20 12.48 29.71
 GMAN 20.93 14.06 33.34
-GODERN-FS [art_88] 19.17 12.74 30.96
+GMSDR 20.49 14.15 32.13
+GODERN-FS [473] 19.17 12.74 30.96
 GRU 23.68 16.44 39.27
@@
 MTGODE 19.55 13.08 32.99
-PAHNN [art_435] 16.28 11.98 27.62
+PAHNN [510] 16.28 11.98 27.62
+PDFormer 25.75 17.55 42.09
 SCINet 19.29 11.89 31.28
-SDGL [art_78] 18.65 12.38 31.30
-STAGCN [art_130] 19.02 12.46 30.75
-STCGNN [art_115] 19.39 12.71 31.17
+SDGL [243] 18.65 12.38 31.30
+STAGCN [112] 19.02 12.46 30.75
+STCGNN [438] 19.39 12.71 31.17
+StemGNN 21.61 16.10 33.80
+STExplainer [369] 18.57 12.13 30.14
 STFGCN 19.83 13.02 31.88
@@
 STGCN 21.16 13.83 34.89
+STGDN [80] 19.10 12.60 30.83
+STGNCDE 19.21 12.76 31.09
 STGODE 20.84 13.76 32.84
@@
 STSGCN 21.19 13.90 33.65
-STWawe+ [art_453] 18.25 12.21 30.14
+STWave [81] 18.50 12.43 30.39
+STWawe+ [82] 18.25 12.21 30.14
 SVR 28.70 19.20 44.56
+TAMP-S2GCNets 19.74 13.22 31.74
 TCN 23.22 15.59 37.26
@@
 Transformer 21.10 15.13 31.46
-TSGDC [art_424] 18.80 12.67 31.08
+TSGDC [407] 18.80 12.67 31.08
 VAR 23.75 18.09 36.66
-W-DSTAGNN [art_419] 19.30 12.70 31.28
-Z-GCNETs 20.84 13.77 32.82
+W-DSTAGNN [163] 19.30 12.70 31.28
+Z-GCNETS [50] 19.50 12.78 31.61
 […]
@@
 Metrics MAE MAPE RMSE
+AdaSTNet 21.33 8.94 35.04
 AGCRN 22.37 9.12 36.55
-AMGCN [art_444] 19.83 8.19 31.03
+AMGCN [208] 19.83 8.19 31.03
 ARIMA 38.17 19.46 59.27
 ASTGCN 24.01 10.73 37.87
-Auto-DSTSGN [art_108] 20.08 8.57 33.02
+Auto-DSTSGN [173] 20.08 8.57 33.02
 AutoSTG 23.22 9.95 36.47
-DCGCN [art_117] 22.06 - 34.66
+AutoSTS [213] 20.26 8.54 33.09
+CL4ST [368] 20.20 8.53 34.06
+DCGCN [111] 22.06 - 34.66
 DCRNN 25.22 11.82 38.61
@@
 DMSTGCN 20.77 8.94 33.67
-DSTAGNN 21.42 9.01 34.51
+DSTAGNN [203] 21.42 9.01 34.51
 DSTCGCN 20.37 8.64 33.47
 DyHSL 18.84 8.11 31.65
+Gboot 21.35 9.02 34.43
 GMAN 21.56 9.51 34.97
+GMSDR 22.27 9.86 34.94
 GWN 26.85 12.12 42.78
+FOGS 21.28 8.95 34.88
+FourierGNN 25.47 10.76 39.69
 HA 45.12 24.51 65.64
@@
 MTGNN 20.57 9.12 33.54
-PAHNN [art_435] 17.35 7.64 29.78
+PAHNN [510] 17.35 7.64 29.78
+PDFormer 23.92 11.62 36.76
 PGCN 20.03 8.78 32.54
+PYNet [145] 19.61 8.36 32.85
 RGSL 20.73 8.71 34.48
-RST-LTG [art_406] 19.49 8.26 32.91
+RST-LTG [227] 19.49 8.26 32.91
 SLCNN 20.38 8.85 33.27
-STAGCN [art_130] 21.10 8.92 34.10
+SPGCL 31.35 18.32 46.34
+STAGCN [112] 21.10 8.92 34.10
 StemGNN 22.23 9.20 36.46
+STExplainer [369] 20.00 8.51 33.45
 STFGCN 22.07 9.21 35.80
@@
 STG2Seq 32.77 20.16 47.16
+STGBN [80] 21.33 8.94 35.04
 STGCN 25.38 11.08 38.78
-STGODE 22.59 - 37.54
+STGNCDE 20.53 8.80 33.84
+STGODE 22.59 10.14 37.54
 ST-GRAT 20.69 9.06 33.01
@@
 STSGCN 24.26 10.21 39.03
-STWave [art_453] 19.59 8.17 33.53
+STWave [81] 19.94 8.38 33.88
+STWave+ [82] 19.59 8.17 33.53
 SVR 32.49 14.26 50.22
+TAMP-S2GCNets 21.84 9.24 35.42
 TCN 32.72 14.26 42.23
@@
 Transformer 22.07 10.12 34.21
-TSGDC [art_424] 19.96 8.54 33.25
+TSGDC [407] 19.96 8.54 33.25
 VAR 50.22 32.22 75.63
@@
 Metrics MAE MAPE RMSE
-AGCRN 15.95 10.09 25.22
-AJSTGL [art_448] 13.79 - 22.48
+AdaSTNet 15.72 9.98 24.72
+AGCRN [8] 15.95 10.09 25.22
 ARIMA 31.09 22.73 44.32
@@
 ASTGNN 15.00 9.50 24.70
-Auto-DSTSGN [art_108] 14.74 9.45 23.76
+Auto-DSTSGN [173] 14.74 9.45 23.76
 Autoformer 21.04 15.98 31.33
 AutoSTG 16.37 10.36 25.46
+AutoSTS [213] 14.65 9.49 23.52
+CL4ST [368] 14.74 9.61 24.17
 Crossformer 16.25 11.14 26.14
-DCGCN [art_117] 15.68 - 24.39
+DCGCN [111] 15.68 - 24.39
 DCRNN 16.82 10.92 26.36
-DHSL [art_400] 15.19 9.86 24.61
+DHSL [390] 15.19 9.86 24.61
 DGCRN 16.22 12.06 26.10
-DMGF-Net [art_99] 16.48 10.56 25.73
+DMGF-Net [218] 16.48 10.56 25.73
 DMSTG 13.85 9.26 23.30
 DSA-NET 17.14 11.32 26.96
-DSTAGNN 15.67 - 25.37
+DSTAGNN [203] 15.67 - 25.37
 DSTCGCN 15.18 9.68 24.49
-DSTIGNN [art_74] 13.88 9.04 23.64
+DSTIGNN [99] 13.88 9.04 23.64
 DyHSL 14.01 8.87 22.91
@@
 FNN 24.24 16.21 37.57
-GL-STGTN [art_433] 14.00 9.50 23.41
+FOGS 15.73 9.88 24.92
+Gboot [95] 15.54 9.76 24.53
+GL-STGTN [242] 14.00 9.50 23.41
 GMAN 15.31 10.13 24.92
-GODERN-FS [art_88] 15.59 9.95 24.79
+GMSDR [259] 16.36 10.28 25.58
+GODERN-FS [473] 15.59 9.95 24.79
 GRU 22.00 13.33 36.24
@@
 MTGODE 15.61 10.15 25.96
-PAHNN [art_435] 13.58 8.35 22.11
+PAHNN [510] 13.58 8.35 22.11
+PYNet [145] 14.03 9.39 23.84
 RGSL 15.49 9.96 24.80
-RST-LTG [art_406] 13.61 9.05 23.15
+RST-LTG [227] 13.61 9.05 23.15
 SCINet 15.78 9.97 24.60
-SDGL [art_78] 14.93 9.61 24.13
-SNAS4MTF [art_408] 15.07 9.67 24.00
-STAGCN [art_130] 15.36 9.80 24.32
-STCGNN [art_115] 15.55 10.02 24.61
+SDGL [243] 14.93 9.61 24.13
+SNAS4MTF [32] 15.07 9.67 24.00
+STAGCN [112] 15.36 9.80 24.32
+STCGNN [438] 15.55 10.02 24.61
 StemGNN 15.91 10.90 25.44
+STExplainer [369] 14.59 9.80 23.91
 STFGNN 16.64 10.60 26.25
-STIDGCN 14.10 9.15 23.72
+STGNCDE 15.45 9.92 24.81
 STG2Seq 20.17 17.32 30.71
+STGBN [80] 15.35 9.71 24.18
 STGCN 17.50 11.29 27.09
@@
 ST-GTNN 16.51 10.45 25.19
+STIDGCN 14.10 9.15 23.72
 STMetaNet 18.75 14.49 27.50
@@
 STSGCN 17.13 10.96 26.80
-STWave+ [art_453] 13.21 8.63 23.04
+STWave [81] 13.42 8.90 23.40
+STWave+ [82] 13.21 8.63 23.04
 SVR 23.25 14.64 36.16
+TAMP-S2GCNets 13.36 10.15 25.98
 TCN 22.72 14.03 35.79
@@
 VAR 19.19 13.10 29.81
-As stated above, the results consistently indicate that the periodicity aware spatial-temporal adaptive hypergraph neural network (PAHNN) [art_435] proposed by Zhao et al.
+Z-GCNETs [50] 15.76 10.01 25.11
+As stated above, the results almost consistently indicate that the periodicity aware spatial-temporal adaptive hypergraph neural network (PAHNN) [510] proposed by Zhao et al.
 in 2024 is the most accurate model.
@@
 The theme of predictive monitoring is related to sensor-based monitoring of processes, which has become essential in modern industry.
-The evolution of monitoring systems is concomitant with the rapid growth of the Internet of Things (IoT), which aims at a smarter management across various applications [art_9].
+The evolution of monitoring systems is concomitant with the rapid growth of the Internet of Things (IoT), which aims at a smarter management across various applications [221].
 Hence, in this context, multi-sensor systems can be used to detect anomalies in complex scenarios and to monitor the overall status of a system.
@@
 Related data-based approaches include techniques such as anomaly detection, fault diagnosis, and estimation of the equipment’s remaining useful life.
-Despite the specificity of the field, 35 out of 281 papers address the topic of predictive monitoring.
-This includes anomaly detection [art_9, art_10, art_477, art_479, art_481, art_483, art_484, art_489], fault diagnosis [art_53, art_54, art_55, art_56, art_57, art_58, art_192, art_469, art_471, art_474, art_486, art_487, art_491] and RUL estimation [art_152, art_153, art_154, art_155, art_156, art_157, art_158, art_478, art_488].
+Despite the specificity of the field, 40 out of 366 papers address the topic of predictive monitoring.
+This includes anomaly detection [221, 226, 515, 249, 114, 21, 512, 136, 128, 272, 283, 239, 204, 139], fault diagnosis [374, 464, 207, 296, 497, 206, 196, 389, 379, 476, 381] and RUL estimation [400, 46, 382, 291, 441, 195, 408, 488, 194].
 The use of GNNs in these analyses is justified by the necessity of using a model capable of capturing the spatio-temporal correlations among the many sensors present in the system.
@@
 As said, predictive monitoring is a research field which focuses on several techniques, such as anomaly detection, fault diagnosis and remaining useful life estimation.
-Anomaly detection is a technique that aims to identify abnormal data that are not due to random deviations from some regular pattern, but are generated by a different underlying mechanism [ref_anomaly_detection_1].
-It is a prerequisite part of fault diagnosis [ref_anomaly_detection_2].
+Anomaly detection is a technique that aims to identify abnormal data that are not due to random deviations from some regular pattern, but are generated by a different underlying mechanism [131].
+It is a prerequisite part of fault diagnosis [172].
 Fault diagnosis is the process of determining whether a fault has occurred in a machine, including the identification of the time, location, type, and severity of the fault.
@@
 The selected papers implement these three techniques, with anomaly detection being less studied than the other two.
-The main applications developed are industrial (e.g., bearings fault diagnosis [art_54, art_55, art_57, art_155, art_156]) and mechanical (e.g, engines RUL estimation [art_152, art_153, art_154, art_157, art_478, art_488]).
-There are also some papers dealing with energy related tasks, specifically fault diagnosis of photovoltaic systems [art_53] and of energy networks [art_56, art_487].
+The main applications developed are industrial (e.g., bearings fault diagnosis [464, 207, 296, 291, 441]) and mechanical (e.g, engines RUL estimation [400, 46, 382, 195, 488, 194]).
+There are also some papers dealing with energy related tasks, specifically fault diagnosis and degradation analysis of photovoltaic systems [374, 79] and of energy networks [476].
 They are included in this thematic group because of their specific focus on fault classification rather than on energy-related time series forecasting.
@@
 For this reason, we do not make specific reference to forecasting time horizons, unlike in the other subsections, since the goal here is to estimate the RUL rather than to forecast data over future time intervals.
-The majority of the collected papers related to this theme were published in 2024 by authors affiliated with Chinese universities.
-The most popular journal for these topics is IEEE Transactions on Instrumentation and Measurement by IEEE, with 6 out of 35 papers.
+The majority of the collected papers related to this theme were published in 2024, and the most popular journal for these topics is IEEE Transactions on Instrumentation and Measurement by IEEE, with 6 out of 35 papers.
 It can be observed that the majority of the articles are published in IEEE journals.
+As for conference venues, 4 out of 7 conference papers were presented at CIKM.
 5.6.2 Datasets
@@
 Dataset Used by Link
-Air Quality China [art_9] https://www.aqistudy.cn/historydata/
-Air Quality - TPS [art_9] https://www.kaggle.com/amritpal333/tps-july-2021-original-dataset-clean
-C-MAPSS [art_152], [art_153], [art_154], [art_157], [art_478], [art_488] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
-CWRU bearing dataset [art_55], [art_192], [art_491] http://csegroups.case.edu/bearingdatacenter/home
-Herbert river water level [art_483] https://zenodo.org/records/8053359
-Intelligent maintenance system (IMS) [art_475] https://www.nasa.gov/intelligent-systems-division
-JNU bearing dataset [art_55] http://mad-net.org:8765/explore.html?id=9&t=0.9355271549540183
-NASA battery dataset [art_158] http://ti.arc.nasa.gov/project/prognostic-data-repository
-NASA prognostics data [art_58] https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/#bearing
-N-CMAPSS [art_152], [art_154] https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/
-MFPT [art_192] https://api.semanticscholar.org/CorpusID:26870683 (source paper)
-Mars science laboratory rover (MSL) [art_481], [art_489] https://doi.org/10.1145/3219819.3219845 (source paper)
-NREL PV [art_53] https://www.nrel.gov/
-PRONOSTIA [art_155], [art_156] https://hal.science/hal-00719503/document (source paper)
-Pooled server metrics (PSM) [art_479], [art_481] https://doi.org/10.1145/3447548.3467174 (source paper)
-Server Machine Dataset (SMD) [art_479], [art_484] https://github.com/NetManAIOps/OmniAnomaly
-SEU bearing dataset [art_55] https://github.com/cathysiyu/Mechanical-datasets
-Secure water treatment (SWaT) [art_477], [art_479], [art_481], [art_484], [art_489] https://itrust.sutd.edu.sg/itrust-labs_datasets/dataset_info/
-Soil moisture active passive satellite (SMAP) [art_479], [art_481], [art_484], [art_489] https://doi.org/10.1145/3219819.3219845 (source paper)
-Water distribution (WaDi) [art_477], [art_479], [art_484], [art_489] https://itrust.sutd.edu.sg/itrust-labs_datasets/dataset_info/
-Wireless capsule endoscopy images [art_475] https://doi.org/10.1055/s-0043-105488 (source paper)
-XJTUGearbox and XJTUSpurgear [art_192], [art_474] http://dx.doi.org/10.1016/j.ymssp.2021.108653 (source paper)
+Air quality China [221] https://www.aqistudy.cn/historydata/
+Air quality - TPS [221] https://www.kaggle.com/amritpal333/tps-july-2021-original-dataset-clean
+C-MAPSS [400], [46], [382], [195], [488], [194] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
+CWRU bearing dataset [207], [206], [381] http://csegroups.case.edu/bearingdatacenter/home
+HAI security dataset [128] https://github.com/icsdataset/hai
+Herbert river water level [21] https://zenodo.org/records/8053359
+Intelligent maintenance system (IMS) [96] https://www.nasa.gov/intelligent-systems-division
+JNU bearing dataset [207] http://mad-net.org:8765/explore.html?id=9&t=0.9355271549540183
+MFPT [206] https://api.semanticscholar.org/CorpusID:26870683 (source paper)
+Mars science laboratory rover (MSL) [114], [136], [272], [204] https://doi.org/10.1145/3219819.3219845 (source paper)
+NASA battery dataset [408] http://ti.arc.nasa.gov/project/prognostic-data-repository
+NASA prognostics data [497] https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/#bearing
+N-CMAPSS [400], [382] https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/
+NREL PV [374] https://www.nrel.gov/
+PRONOSTIA [291], [441] https://hal.science/hal-00719503/document (source paper)
+Pooled server metrics (PSM) [249], [114] https://doi.org/10.1145/3447548.3467174 (source paper)
+PV degradation [79] https://github.com/Yangxin666/ST-GTrend/tree/main/Datasets
+SEU bearing dataset [207] https://github.com/cathysiyu/Mechanical-datasets
+Soil moisture active passive satellite (SMAP) [249], [114], [512], [136], [204] https://doi.org/10.1145/3219819.3219845 (source paper)
+Server machine dataset (SMD) [249], [512], [272] https://github.com/NetManAIOps/OmniAnomaly
+Secure water treatment (SWaT) [515], [249], [114], [512], [136], [128], [272], [283], [239], [204] https://itrust.sutd.edu.sg/itrust-labs_datasets/dataset_info/
+Water distribution (WaDi) [515], [249], [512], [136], [128], [272], [283], [239], [204] https://itrust.sutd.edu.sg/itrust-labs_datasets/dataset_info/
+Wireless capsule endoscopy images [96] https://doi.org/10.1055/s-0043-105488 (source paper)
+XJTUGearbox and XJTUSpurgear [206], [389] http://dx.doi.org/10.1016/j.ymssp.2021.108653 (source paper)
 The two popular datasets are C-MAPSS and N-CMAPSS, two well-known datasets for RUL prediction provided by NASA.
@@
 Almost no papers use exogenous variables, with the exception of Ref.
-[art_53], which evaluates the impact of meteorological variables on fault diagnosis of photovoltaic systems.
+[374], which evaluates the impact of meteorological variables on fault diagnosis of photovoltaic systems.
+Other popular datasets are SWaT and WaDi.
+The secure water treatment dataset (SWaT) is a cybersecurity dataset that includes data collected from 51 sensors and actuators over a continuous period of 11 days.
+During this time, the system operated under normal conditions for 7 days and was subjected to attack scenarios for 4 days.
+The dataset contains both network traffic and physical process data recorded from all devices.
+The water distribution (WaDi) dataset is an extension of SWaT involving a larger setup of 127 sensors and actuators monitored for 16 consecutive days.
 The time granularity of the studied time series is highly variable, ranging from 1 hour to fractions of seconds.
@@
 5.6.3 Types of models
-More than half of the proposed spatio-temporal GNN models (19 out of 35 papers) fall under the category of pure convolutional GNNs, followed by 10 attentional GNNs, 2 hybrid models, and one recurrent model (Ref.
-[art_470]).
+Almost half of the proposed spatio-temporal GNN models (19 out of 40 papers) fall under the category of pure convolutional GNNs, followed by 15 attentional GNNs, 2 hybrid models, and one recurrent model (Ref.
+[86]).
 One paper (Ref.
-[art_57]) uses both convolutional and attentional GNNs, one paper (Ref.
-[art_475]) uses a variant of the convolutional GNN, and the description of another one (Ref.
-[art_155]) does not clarify which class of the taxonomy it belongs to.
+[296]) uses both convolutional and attentional GNNs, one paper (Ref.
+[96]) uses a variant of the convolutional GNN, and the description of another one (Ref.
+[291]) does not clarify which class of the taxonomy it belongs to.
 Almost all the models have a graph learning module, so that the graph structure is rarely defined a priori by the researchers.
-Three papers [art_55, art_57, art_154] use the visibility graph algorithm proposed by Lacasa et al.
+Three papers [207, 296, 382] use the visibility graph algorithm proposed by Lacasa et al.
 in Ref.
-[visibility_graph] for converting the time series into a graph.
+[201] for converting the time series into a graph.
 The most common loss functions used within this thematic group are the cross entropy function for classification problems, and MSE or RMSE for RUL forecasting problems.
@@
 Model Link
-Convolution-graph attention network (ConvGAT) [art_153] https://github.com/CUG-FDGroup
-Graph Network with Causal Connectivity and Temporal Convolutional Neural Network feature extractor [art_155] https://github.com/mylonasc/gnn-tcnn
-Graph spatiotemporal process and anomaly scorer (GST-Pro) [art_477] https://github.com/huankoh/GST-Pro
-Causality Enhanced Global-Local Graph Neural Network (CEGLo-GNN) [art_482] https://github.com/YueYueXia/CEGLo-GNN
-Graph Deviation Network+ (GDN+) [art_483] https://github.com/KatieBuc/gnnad/releases
+Causality enhanced global-local graph neural network (CEGLo-GNN) [363] https://github.com/YueYueXia/CEGLo-GNN
+Dual time-oriented graph attention networks (DuoGAT) [204] https://github.com/ByeongtaePark/DuoGAT
+Graph deviation network+ (GDN+) [21] https://github.com/KatieBuc/gnnad/releases
+Graph network with causal connectivity and temporal convolutional neural network feature extractor [291] https://github.com/mylonasc/gnn-tcnn
+Spectral multi-view causal graph fusion based anomaly detection (SMV-CGAD) [283] https://github.com/arunvignesh28/SMV-CGAD
+Spatio-temporal graph neural network empowered long-term trend analysis system (ST-GTrend) [79] https://github.com/Yangxin666/ST-GTrend
+Graph attention networks with topological analysis (TopoGDN) [272] https://github.com/ljj-cyber/TopoGDN
 5.6.4 Benchmark models
@@
 Category Model Used by
-Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_483], [art_487]
-Non-GNN machine learning methods Attention-based long-short term memory network [art_153], [art_157], [art_488]
-Feature-attention based bidirectional GRU and CNN model (AGCNN) [AGCNN] [art_153], [art_154], [art_157], [art_478], [art_488]
-Anomalous Rhythm Detection using Adversarially Generated Time Series (BeatGAN) [BeatGAN] [art_481], [art_484]
-Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_152], [art_153], [art_154], [art_157], [art_470], [art_488]
-Convolutional neural network (CNN) [art_9], [art_56], [art_152], [art_153], [art_154], [art_156], [art_157], [art_461], [art_469], [art_471], [art_486], [art_487], [art_488]
-Convolutional neural network with long-short term memory network (CNN-LSTM) [art_154], [art_470]
-Crossformer [Crossformer] [art_152], [art_482]
-Deep autoencoding gaussian model (DAGMM) [DAGMM] [art_481], [art_484]
-Deep belief network (DBN) [DBN] [art_157], [art_488]
-Decision tree (DT) [art_9], [art_56], [art_487]
-Feed-forward neural network (FNN) [art_55], [art_155], [art_469], [art_478]
-Genetic algorithm approach on restricted Boltzmann-based long-short term memory network (GA+RBM+LSTM) [GA+RBM+LSTM] [art_157], [art_488]
-Gradient boosting (GB) [art_56], [art_154], [art_487]
-Gated recurrent unit (GRU) [art_53], [art_56], [art_156], [art_469], [art_487]
-Integrated multi-head dual sparse self-attention network (IMDSSN) [IMDSSN] [art_154], [art_478], [art_488]
-Isolation forest [art_10], [art_481]
-$k$-nearest neighbors (KNN) [art_56], [art_475], [art_487]
-Local outlier factor (LOF) [art_475], [art_481]
-Logistic regression (LR) [art_9], [art_56]
-Long-short term memory network (LSTM) [art_55], [art_56], [art_154], [art_157], [art_158], [art_461], [art_470], [art_487], [art_488]
-Long-short term memory-based variational auto-encoder (LSTM-VAE) [art_10], [art_477], [art_479], [art_484], [art_489]
-OmniAnomaly [OmniAnomaly] [art_477], [art_479], [art_481], [art_484], [art_489]
-Residual network (ResNet) [ResNet] [art_471], [art_491]
-Random forest (RF) [art_56], [art_154], [art_478], [art_487]
-Support vector machine (SVM) [art_9], [art_56], [art_471], [art_478], [art_486], [art_487]
-Deep transformer networks for anomaly detection (TranAD) [TranAD] [art_484], [art_489]
-Transformer [attention] [art_152], [art_154], [art_470], [art_478], [art_479], [art_488], [art_489]
-Unsupervised anomaly detection on multivariate time series (USAD) [5] [art_479], [art_484], [art_489]
-GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_58], [art_482]
-Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_480], [art_482]
-Fused sparse autoencoder and graph network (FuSAGNet) [FuSAGNet] [art_10], [art_479]
-Graph attention network (GAT) [GAT] [art_54], [art_56], [art_192], [art_470], [art_487]
-Graph convolutional network (GCN) [art_54], [art_55], [art_156], [art_192], [art_461], [art_474], [art_486]
-Graph deviation network (GDN) [GDN] [art_10], [art_477], [art_479], [art_481], [art_484]
-Gated graph convolutional network (GGCN) [GGCN] [art_153], [art_478]
-Graph WaveNet (GWN) [GWN] [art_58], [art_480], [art_482]
-Hierarchical attention graph convolutional network (HAGCN) [HAGCN] [art_152], [art_153], [art_478]
-Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_470], [art_483]
-Spatio-temporal fusion attention (STFA) [art_157] [art_153], [art_154], [art_478], [art_488]
-Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_58], [art_152], [art_470], [art_478], [art_480]
+Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [21], [476]
+Non-GNN machine learning methods Auto-encoder (AE) [512], [128], [204], [139]
+Feature-attention based bidirectional GRU and CNN model (AGCNN) [526] [46], [382], [195], [488], [194]
+Attention-based long-short term memory network [46], [195], [194]
+Anomalous rhythm detection using adversarially generated time series (BeatGAN) [518] [114], [512]
+Bidirectional long-short term memory network (BiLSTM) [109] [400], [46], [382], [195], [86], [194]
+Convolutional neural network (CNN) [221], [400], [46], [382], [441], [195], [31], [196], [379], [476], [194]
+Convolutional neural network with long-short term memory network (CNN-LSTM) [382], [86]
+Crossformer [503] [400], [363]
+Deep autoencoding gaussian model (DAGMM) [529] [114], [512], [128], [239], [204]
+Deep belief network (DBN) [474] [195], [194]
+Decision tree (DT) [221], [476]
+Feed-forward neural network (FNN) [207], [291], [196], [488]
+Genetic algorithm approach on restricted Boltzmann-based long-short term memory network (GA+RBM+LSTM) [258] [195], [194]
+Gradient boosting (GB) [382], [476]
+Gated recurrent unit (GRU) [374], [441], [196], [476]
+Integrated multi-head dual sparse self-attention network (IMDSSN) [478] [382], [488], [194]
+Isolation forest [226], [114], [283]
+$k$-nearest neighbors (KNN) [96], [476], [128], [204]
+Local outlier factor (LOF) [19] [96], [114]
+Long-short term memory network (LSTM) [207], [382], [195], [408], [31], [86], [476], [194]
+Long-short term memory-based variational auto-encoder (LSTM-VAE) [226], [515], [249], [512], [136], [128], [272], [139]
+Multivariate anomaly detection with GAN (MAD-GAN) [209] [226], [128], [272], [204]
+OmniAnomaly [358] [515], [249], [114], [512], [136], [272]
+Random forest (RF) [382], [488], [476]
+Support vector machine (SVM) [221], [128], [488], [379], [476], [283], [204], [139]
+Deep transformer networks for anomaly detection (TranAD) [372] [512], [136], [272], [239]
+Transformer [375] [400], [382], [86], [488], [249], [194], [136]
+Unsupervised anomaly detection on multivariate time series (USAD) [4] [249], [512], [136], [128], [204], [139]
+GNN methods Adaptive graph convolutional recurrent network (AGCRN) [8] [497], [363]
+Diffusion convolutional recurrent neural network (DCRNN) [237] [224], [363]
+Fused sparse autoencoder and graph network (FuSAGNet) [128] [226], [249], [239]
+Graph attention network (GAT) [376] [464], [206], [86], [476]
+Graph convolutional network (GCN) [464], [207], [441], [206], [31], [389], [379]
+Graph deviation network (GDN) [66] [226], [515], [249], [114], [512], [128], [272], [239], [204], [139]
+Gated graph convolutional network (GGCN) [386] [46], [488]
+Graph WaveNet (GWN) [425] [497], [224], [363]
+Hierarchical attention graph convolutional network (HAGCN) [230] [400], [46], [488]
+Multivariate time-series anomaly detection via graph attention network (MTAD-GAT) [507] [249], [128], [239], [204]
+Multivariate time series forecasting with graph neural network (MTGNN) [424] [86], [21]
+Spatio-temporal fusion attention (STFA) [195] [46], [382], [488], [194]
+Spatio-temporal graph convolutional network (STGCN) [457] [497], [400], [86], [488], [224]
 In this thematic group there is a high variety of benchmarks used, and before 2024 only few papers share a benchmark model in the GNN category.
@@
 Not surprisingly, almost no statistical or mathematical benchmark models were used, since they are more frequently used in forecasting problems than in classification problems.
-Among the GNN benchmarks, some of them have been specifically developed for remaining useful life prediction or predictive monitoring tasks, such as the hierarchical attention graph convolutional network (HAGCN) [HAGCN] and the spatio-temporal fusion attention (STFA) [art_157].
+Among the GNN benchmarks, some of them have been specifically developed for remaining useful life prediction or predictive monitoring tasks, such as the hierarchical attention graph convolutional network (HAGCN) [230], the spatio-temporal fusion attention (STFA) [195], and the graph deviation network (GCN) [66].
 In the HAGCN model, proposed by Li et al.
@@
 in 2022, it integrates some a priori knowledge about the equipment’s structure with a spatio-temporal deep learning architecture which employs LSTM cells and an attention mechanism.
-Neither of the two original papers includes a link to the source code.
+Finally, the GDN model, proposed by Deng et al.
+in 2021, learns a graph structure that represents the relationships between sensors, and aims to detect deviations from these patterns by using an attentional GNN.
+The source code for the GDN model is available at https://github.com/d-ailin/GDN.
 5.6.5 Results
@@
 For RUL forecasting problems instead, the most commonly used metrics are the RMSE and the score function.
-The only results that can be compared across papers are those for the C-MAPSS dataset.
+The only results that can be compared across many papers are those for the C-MAPSS dataset.
 Tab.
@@
 BLCNN 13.18 302.28 19.09 1557.56 13.75 381.37 20.97 3858.78
-CDSG [art_154] 11.26 188.00 18.13 1740.00 12.03 218.00 19.73 2332.00
+CDSG [382] 11.26 188.00 18.13 1740.00 12.03 218.00 19.73 2332.00
 ChebyNet-EdgePool 15.21 450.22 16.28 1229.22 14.67 421.02 16.26 1159.74
@@
 CNN-Transformer 12.25 198.00 17.08 1575.00 13.39 290.00 19.86 1741.00
-ConvGAT [art_153] 11.34 197.43 14.12 771.61 10.97 235.26 15.51 1231.17
+ConvGAT [46] 11.34 197.43 14.12 771.61 10.97 235.26 15.51 1231.17
 C-Transformer 13.79 475.46 16.11 2214.59 17.10 939.10 19.77 3237.37
@@
 DSAN 13.40 242.00 22.06 2869.00 15.12 497.00 21.03 2677.00
-DS-STFN [art_478] 10.92 161.35 13.77 946.34 10.01 150.42 15.53 1079.91
+DS-STFN [488] 10.92 161.35 13.77 946.34 10.01 150.42 15.53 1079.91
 Earlier CNN 18.45 1287.00 30.29 13570.00 19.82 1596.00 29.16 7886.00
@@
 SMDN 13.40 272.00 - - - - 23.40 4302.00
-STFA [art_157] 11.35 194.44 19.17 2493.09 11.64 224.53 21.41 2760.13
-STP-GNN [art_488] 11.38 192.17 15.81 1094.81 11.02 199.58 17.77 1692.19
+STFA [195] 11.35 194.44 19.17 2493.09 11.64 224.53 21.41 2760.13
+STP-GNN [194] 11.38 192.17 15.81 1094.81 11.02 199.58 17.77 1692.19
 TATFA-Transformer 12.21 261.50 15.07 1359.70 11.23 210.21 18.81 2506.35
 Transformer 12.92 276.27 22.64 1960.25 13.04 284.41 24.81 3892.24
-In general, the best results on the different sub-datasets are achieved by the convolution-graph attention network (ConvGAT) [art_153] and the the dual-stream spatio-temporal fusion network (DS-STFN) [art_478] models.
+In general, the best results on the different sub-datasets are achieved by the convolution-graph attention network (ConvGAT) [46] and the the dual-stream spatio-temporal fusion network (DS-STFN) [488] models.
 The ConvGAT model, proposed by Chen et al.
@@
 The spatial and temporal features are then fused to predict the RUL.
+Other pairs of papers with comparable results are Refs.
+[128] and [204] for SWaT and WaDi datasets and Refs.
+[114] and [272] for the MSL dataset.
+However, due to space limitations, tables comparing just two studies are not included here.
 5.7 Generic
@@
 The reason for this choice is that they are considered to explore broader methodologies or frameworks that are applied to diverse datasets without a declared focus on a particular case.
+For instance, some subsets of papers within the “Generic” group focus on forecasting in presence of missing data (e.g., Refs.
+[458, 286]), or of irregularly sampled time series (e.g., Refs.
+[436, 499, 491, 493, 300]).
+Notably, the majority of these studies have been published in conferences rather than in journals.
+Moreover, some papers in this section also tackle some theoretical issues.
+For example, Ref.
+[271] proves a theorem that provides a theoretical bound on the approximation error of the module that learns the dynamic variable dependence.
+Another example can be found in Ref.
+[505], where the authors theoretically establish that the generalization error bound of their model is related to various components of the GAT model, such as the number of attention heads, the number of neighboring nodes, the Frobenius norm of the weight matrices, and the norm of the input features.
 5.7.1 Overview
-A significant number of papers (44 out of 281) have been categorized under the “Generic” thematic group.
+A significant number of papers (77 out of 366) have been categorized under the “Generic” thematic group.
 The majority of these works commit themselves to address a multivariate time series forecasting problem not necessarily limited to a specific field.
-This “Generic” set includes one paper from 2021, 4 from 2022, 19 from 2023, and 20 from 2024, highlighting the recent interest in applying GNNs in cross-disciplinary contexts.
+This “Generic” set includes 4 papers from 2002, 7 from 2021, 11 from 2022, 28 from 2023, and 27 from 2024, highlighting the recent interest in applying GNNs in cross-disciplinary contexts.
 These papers are scattered across various journals, with the exception of Expert Systems with Applications by Elsevier, IEEE Transactions on Knowledge and Data Engineering by IEEE, and Information Sciences by Elsevier, which published 4 and 3 of the selected papers, respectively.
-The majority of authors are affiliated with institutions in China.
+The three conference venues with the highest number of papers are NeurIPS with 6 papers, CIKM with 5 papers, and KDD with 4 papers.
 5.7.2 Datasets
@@
 Dataset Used by Link
-Beijing Traffic [art_84] https://github.com/BuaaPercy/Traffic-DataSet-of-Beijing-Road
-CCMP wind data [art_73] http://data.remss.com
-CDC US States [art_401] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
-CI earthquakes [art_160] https://doi.org/10.5194/adgeo-43-31-2016 (source paper)
-CityBike NYC [art_464] https://www.citibikenyc.com/system-data
-Chickenpox cases [art_160] https://doi.org/10.48550/arXiv.2102.08100 (source paper)
-C-MAPSS [art_403] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
-CORAv2.0 [art_396] https://mds.nmdis.org.cn/pages/home.html
-CW earthquakes [art_89] https://doi.org/10.5194/adgeo-43-31-2016 (source paper)
-ECG5000 [art_397] https://www.cs.ucr.edu/ẽamonn/time_series_data/
-Electricity consumption [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_88], [art_90], [art_384], [art_405] https://github.com/laiguokun/multivariate-time-series-data
-Electricity Consuming Load (ECL) [art_82], [art_85] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014
-Electricity transformer Temperature (ETT) [art_76], [art_82], [art_85], [art_393], [art_417] https://github.com/zhouhaoyi/ETDataset
-Energy [art_77], [art_388] http://dx.doi.org/10.1016/j.enbuild.2017.01.083 (source paper)
-Enron [art_202] https://doi.org/10.1073/pnas.1800683115 (source paper)
-Eu-Core [art_202] https://doi.org/10.1145/3018661.3018731 (source paper)
-Exchange-rate [art_73], [art_74], [art_76], [art_77], [art_78], [art_79], [art_80], [art_81], [art_85], [art_87], [art_89], [art_90], [art_384], [art_393], [art_405] https://github.com/laiguokun/multivariate-time-series-data
-Facebook [art_202] http://networkrepository.com/socfb
-GEFCOM 2012 [art_89] https://doi.org/10.1016/j.ijforecast.2013.07.001 (source paper)
-Hypertext [art_202] http://networkrepository.com/ia-infect-hyper (source paper)
-IDWR Japan Prefectures [art_401] https://www.niid.go.jp/niid/en/idwr-e.html
-ILI [art_76], [art_85], [art_89] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
-ISRUC-SLEEP Dataset [art_403] https://sleeptight.isr.uc.pt/ISRUC_Sleep/
-Jena Weather [art_76] https://www.bgc-jena.mpg.de/wetter/
-KnowAir [art_400] https://github.com/shuowang-ai/PM2.5-GNN
-LA Traffic [art_404] https://ladot.lacity.gov/residents/traffic-counts
-Loan [art_202] https://ai.ppdai.com/mirror/showCompetitionRisk
-Los-loop [art_383] https://github.com/lehaifeng/T-GCN/tree/master/data
-Large-Scale Traffic and Weather Events (LSTW) [art_404] https://smoosavi.org/datasets/lstw
-METR-LA [art_72], [art_75], [art_79], [art_80], [art_86], [art_160], [art_387], [art_397], [art_401], [art_417], [art_464] https://paperswithcode.com/dataset/metr-la
-Metro [art_400] https://lbs.amap.com/
-MIMIC-III Heart Failure [art_71] https://doi.org/10.1038/sdata.2016.35 (source paper)
-MIMIC-III Infection [art_71] https://doi.org/10.1038/sdata.2016.35 (source paper)
-MOOC network [art_70] http://snap.stanford.edu/jodie/mooc.csv
-Nasdaq [art_77], [art_80] https://doi.org/10.48550/arXiv.1704.02971 (source paper)
-NOAA Weather [art_82], [art_393] https://www.ncei.noaa.gov/data/local-climatological-data/
-NYC-Taxi [art_398] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
-PeMS03 [art_74] https://www.kaggle.com/datasets/elmahy/pems-dataset
-PeMS04 (PeMSD4) [art_73], [art_74], [art_78], [art_79], [art_88], [art_384], [art_396], [art_397], [art_400], [art_405] https://paperswithcode.com/dataset/pems04
-PeMS08 (PeMSD8) [art_73], [art_74], [art_78], [art_79], [art_88], [art_383], [art_384], [art_396], [art_397], [art_400], [art_405] https://paperswithcode.com/dataset/pems08
-PEMS-BAY [art_72], [art_75], [art_160], [art_401] https://paperswithcode.com/dataset/pems-bay
-PhysioNet [art_71] https://archive.physionet.org/challenge/2012/papers/
-Reddit network [art_70] http://snap.stanford.edu/jodie/reddit.csv
-Seattle-loop [art_401] https://github.com/zhiyongc/Seattle-Loop-Data
-Sleep-EDF-20 [art_395] https://gist.github.com/emadeldeen24/a22691e36759934e53984289a94cb09b
-Solar-Energy [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_88], [art_90], [art_384], [art_387], [art_405] https://github.com/laiguokun/multivariate-time-series-data
-Secure Water Treatment (SWaT) [art_387] https://itrust.sutd.edu.sg/itrust-labs_datasets/
-Stocktwits [art_398] https://stocktwits.com/
-TAIEX [art_84] https://finance.yahoo.com/
-TaoBao [art_202] https://tianchi.aliyun.com/competition/entrance/
-Taiwan PM2.5 [art_464] https://taqm.epa.gov.tw/taqm/en/default.aspx
-T-drive [art_383] https://www.kaggle.com/datasets/arashnic/tdriver
-Tiingo [art_398] https://www.tiingo.com/
-Traffic [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_85], [art_87], [art_90], [art_388] https://github.com/laiguokun/multivariate-time-series-data
-UCI Human Activity Recognition [art_403] https://archive.ics.uci.edu/dataset/240/human+activity+recognition+using+smartphones
-University of East Anglia (UEA) datasets [art_385], [art_386], [art_389], [art_407] https://timeseriesclassification.com/
-US stock market price [art_89] https://doi.org/10.48550/arXiv.1810.09936 (source paper)
-Water Distribution (WaDi) [art_387] https://itrust.sutd.edu.sg/itrust-labs_datasets/
-WESAD dataset [art_404] https://ubi29.informatik.uni-siegen.de/usi/data_wesad.html
-Wikipedia network [art_70] http://snap.stanford.edu/jodie/wikipedia.csv
-Wind-Speed [art_76] https://www.kaggle.com/datasets/fedesoriano/wind-speed-prediction-dataset
-Wind Generation [art_405] https://www.kaggle.com/datasets/sohier/30-years-of-european-wind-generation
-WISDM [art_395] https://archive.ics.uci.edu/dataset/507/wisdm+smartphone+and+smartwatch+activity+and+biometrics+dataset)
+Beijing Traffic [47] https://github.com/BuaaPercy/Traffic-DataSet-of-Beijing-Road
+Bytom [50], [51], [48] https://etherscan.io/
+CCMP wind data [241] http://data.remss.com
+CDC US States [360] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
+CER-E [59] https://www.ucd.ie/issda/data/commissionforenergyregulationcer/
+Chickenpox [15], [327] https://www.kaggle.com/datasets/die9origephit/chickenpox-cases-hungary
+China AQI [458] https://quotsoft.net/air/
+CI earthquakes [15] https://doi.org/10.5194/adgeo-43-31-2016 (source paper)
+C-MAPSS [401] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
+CORAv2.0 [137] https://mds.nmdis.org.cn/pages/home.html
+COVID-19 CA [51], [48], [451], [49] https://www.dropbox.com/sh/n0ajd5l0tdeyb80/AABGn-ejfV1YtRwjf_L0AOsNa?dl=0
+COVID-19 England [327] https://pytorch-geometric-temporal.readthedocs.io/en/latest/notes/introduction.html#benchmark-datasets
+COVID-19 TX [51], [48] https://www.dropbox.com/sh/n0ajd5l0tdeyb80/AABGn-ejfV1YtRwjf_L0AOsNa?dl=0
+CSSE COVID-19 [432], [451] https://github.com/CSSEGISandData/COVID-19
+CW earthquakes [57] https://doi.org/10.5194/adgeo-43-31-2016 (source paper)
+Decentraland [50], [51] https://etherscan.io/
+ECG5000 [152], [432], [26], [451] https://www.cs.ucr.edu/~eamonn/time_series_data/
+Electricity consumption [241], [99], [177], [243], [118], [40], [287], [355], [473], [138], [240], [83], [448], [279], [424], [61], [26], [271], [153], [451], [268] https://github.com/laiguokun/multivariate-time-series-data
+Electricity consuming load (ECL) [395], [186], [356] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014
+Electricity transformer temperature (ETT) [504], [395], [186], [231], [443], [153], [356] https://github.com/zhouhaoyi/ETDataset
+Energy [76], [125] http://dx.doi.org/10.1016/j.enbuild.2017.01.083 (source paper)
+ENG-RAD [286] https://open-meteo.com/
+Enron [155] https://doi.org/10.1073/pnas.1800683115 (source paper)
+Eu-Core [155] https://doi.org/10.1145/3018661.3018731 (source paper)
+European wind generation [448] https://www.kaggle.com/datasets/sohier/30-years-of-european-wind-generation
+Exchange-rate [241], [99], [504], [76], [243], [118], [40], [287], [186], [397], [57], [138], [231], [240], [448], [424], [61], [271], [153], [268] https://github.com/laiguokun/multivariate-time-series-data
+Facebook [155] http://networkrepository.com/socfb
+GEFCOM 2012 [57] https://doi.org/10.1016/j.ijforecast.2013.07.001 (source paper)
+Golem [51] https://www.dropbox.com/sh/n0ajd5l0tdeyb80/AABGn-ejfV1YtRwjf_L0AOsNa?dl=0
+GraphMSO [286] https://github.com/marshka/hdtts/tree/main/config/dataset
+Hypertext [155] http://networkrepository.com/ia-infect-hyper (source paper)
+HZMetro [279] https://github.com/HCPLab-SYSU/PVCGN/tree/master/data
+IDWR Japan prefectures [360] https://www.niid.go.jp/niid/en/idwr-e.html
+ILI [504], [186], [57] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
+ISRUC-SLEEP [401] https://sleeptight.isr.uc.pt/ISRUC_Sleep/
+Jena weather [504], [153] https://www.bgc-jena.mpg.de/wetter/
+KnowAir [390] https://github.com/shuowang-ai/PM2.5-GNN
+LA Traffic [337] https://ladot.lacity.gov/residents/traffic-counts
+Loan [155] https://ai.ppdai.com/mirror/showCompetitionRisk
+Los-loop [244] https://github.com/lehaifeng/T-GCN/tree/master/data
+Large-scale traffic and weather events (LSTW) [337] https://smoosavi.org/datasets/lstw
+METR-LA [60], [177], [118], [40], [52], [15], [396], [152], [360], [443], [338], [458], [83], [277], [424], [26], [451], [59], [300] https://www.kaggle.com/datasets/annnnguyen/metr-la-dataset
+Metro [390] https://lbs.amap.com/
+MIMIC [129], [436], [491] https://doi.org/10.1038/sdata.2016.35 (source paper)
+Montevideo buses [327] https://pytorch-geometric-temporal.readthedocs.io/en/latest/notes/introduction.html#benchmark-datasets
+MOOC network [232] http://snap.stanford.edu/jodie/mooc.csv
+MTM-1 hand motions [327] https://pytorch-geometric-temporal.readthedocs.io/en/latest/notes/introduction.html#benchmark-datasets
+Nasdaq [76], [40] https://dl.acm.org/doi/10.5555/3172077.3172254 (source paper)
+Neonatal EEG [83] https://zenodo.org/records/2547147
+NOAA weather [395], [231] https://www.ncei.noaa.gov/data/local-climatological-data/
+NREL [432], [286], [451] https://www.nrel.gov/
+NYC-Bike [448], [279], [277], [49] https://citibikenyc.com/system-data
+NYC-Taxi [453], [450], [448], [279], [49] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
+PAMAP2 [289], [499], [456] https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
+Pedal Me deliveries [327] https://pytorch-geometric-temporal.readthedocs.io/en/latest/notes/introduction.html#benchmark-datasets
+PeMS03 (PeMSD3) [99], [51], [26], [143] https://www.kaggle.com/datasets/elmahy/pems-dataset
+PeMS04 (PeMSD4) [241], [99], [243], [118], [473], [137], [152], [390], [240], [458], [50], [51], [26], [48], [369], [368] https://www.kaggle.com/datasets/elmahy/pems-dataset
+PeMS07 (PeMSD7) [26], [271], [369], [368] https://www.kaggle.com/datasets/elmahy/pems-dataset
+PeMS08 (PeMSD8) [241], [99], [243], [118], [473], [244], [137], [152], [390], [240], [458], [50], [51], [26], [48], [409], [369], [368] https://www.kaggle.com/datasets/elmahy/pems-dataset
+PEMS-BAY [60], [177], [15], [360], [338], [458], [83], [424], [26], [271], [59], [143], [300] https://zenodo.org/records/4263971
+PhysioNet [129], [436], [499], [491] https://archive.physionet.org/challenge/2012/papers/
+Reddit network [232] http://snap.stanford.edu/jodie/reddit.csv
+Seattle-loop [360] https://github.com/zhiyongc/Seattle-Loop-Data
+Sleep-EDF-20 [100] https://gist.github.com/emadeldeen24/a22691e36759934e53984289a94cb09b
+Solar-Energy [241], [99], [177], [243], [118], [40], [287], [355], [397], [473], [138], [396], [240], [83], [448], [424], [61], [26], [271], [268] https://github.com/laiguokun/multivariate-time-series-data
+Secure water treatment (SWaT) [396] https://itrust.sutd.edu.sg/itrust-labs_datasets/
+SHMetro [279] https://github.com/HCPLab-SYSU/PVCGN/tree/master/data
+Stock market data [505] https://www.kaggle.com/datasets/paultimothymooney/stock-market-data
+Stocktwits [453] https://stocktwits.com/
+Synthetic periodic data synthetic [300] https://github.com/joeloskarsson/tgnn4i
+TAIEX [47] https://finance.yahoo.com/
+Taiwan PM2.5 [277] https://taqm.epa.gov.tw/taqm/en/default.aspx
+TaoBao [155] https://tianchi.aliyun.com/competition/entrance/
+T-drive [244] https://www.kaggle.com/datasets/arashnic/tdriver
+Tiingo [453] https://www.tiingo.com/
+Traffic [177], [243], [118], [40], [287], [355], [186], [397], [138], [125], [424], [61], [271], [153], [451], [268] https://github.com/laiguokun/multivariate-time-series-data
+UCI Beijing PM2.5 [505] https://archive.ics.uci.edu/ml/datasets/Beijing+PM2.5+Data
+UCI human activity recognition [401], [456] https://archive.ics.uci.edu/dataset/240/human+activity+recognition+using+smartphones
+UCR datasets [472] https://www.cs.ucr.edu/~eamonn/time_series_data_2018/
+University of East Anglia (UEA) datasets [74], [113], [261], [140], [456] https://timeseriesclassification.com/
+US historical climatology network (USHCN) [436], [491], [300] https://www.osti.gov/biblio/1389426
+US stock market price [57] http://dx.doi.org/10.24963/ijcai.2019/810 (source paper)
+Water distribution (WaDi) [396] https://itrust.sutd.edu.sg/itrust-labs_datasets/
+WESAD dataset [337] https://ubi29.informatik.uni-siegen.de/usi/data_wesad.html
+Wiki [451] https://www.kaggle.com/c/web-traffic-time-series-forecasting/data
+Wikipedia network [232] http://snap.stanford.edu/jodie/wikipedia.csv
+Windmill [327] https://pytorch-geometric-temporal.readthedocs.io/en/latest/notes/introduction.html#benchmark-datasets
+Wind-speed [504] https://www.kaggle.com/datasets/fedesoriano/wind-speed-prediction-dataset
+Wind generation [240] https://www.kaggle.com/datasets/sohier/30-years-of-european-wind-generation
+WISDM [100], [289] https://archive.ics.uci.edu/dataset/507/wisdm+smartphone+and+smartwatch+activity+and+biometrics+dataset
 It should be noted that multiple recurring datasets can be considered as benchmarks for this community.
@@
 Interested readers are referred to Refs.
-[art_385, art_386, art_389, art_407].
+[74, 113, 261, 140, 456].
 In general, the data granularity varies significantly, ranging from fractions of seconds to one month, indicating that there is no particular focus on a specific granularity.
@@
 5.7.3 Types of models
-Among the 43 papers in the “Generic” thematic group, 33 use a purely convolutional GNN approach, and four employ a purely attentional model.
+Among the 77 papers in the “Generic” thematic group, 56 use a purely convolutional GNN approach, and 10 employ a purely attentional model.
 Five paper lack sufficient detail to be classified within the taxonomy chosen (Refs.
-[art_85, art_385, art_386, art_389, art_407]), and another one declares that it uses an aggregating approach that is simpler than either convolutional or attentional approaches [art_81].
-Another paper proposes a so-called multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) model, where the continuous dynamics of simplified graph propagation is described by an ordinary differential equation (ODE) [art_75].
+[186, 74, 113, 261, 140]), and another one declares that it uses an aggregating approach that is simpler than either convolutional or attentional approaches [287].
+Moreover, in Ref.
+[59] the authors use an approach that can be proved to be equivalent to a convolutional model; Ref.
+[59] perform a general analysis that extends beyond a single model; Ref.
+[327] explores various models, all based on convolutional architectures; Ref.
+[300] proposes an architecture that can adopt either a convolutional or an attentional aggregation.
+Another paper proposes a so-called multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) model, where the continuous dynamics of simplified graph propagation is described by an ordinary differential equation (ODE) [177].
 This is not the only paper in this section using differential equations.
 In Ref.
-[art_401], two domain specific ODEs (one for traffic speed evolution, and the other for disease propagation) are integrated into the convolutional GNN model.
-The authors argue that this dynamics-informed model approach enhances robustness, allow easier generalization when the equation of the dynamics is known, and reduces the number of model parameters.Additionally, during inference, the model only requires a single data point, as the ODEs explicitly incorporate the immediate dynamics.
+[360], two domain specific ODEs (one for traffic speed evolution, and the other for disease propagation) are integrated into the convolutional GNN model.
+The authors argue that this dynamics-informed model approach enhances robustness, allow easier generalization when the equation of the dynamics is known, and reduces the number of model parameters.
+Additionally, during inference, the model only requires a single data point, as the ODEs explicitly incorporate the immediate dynamics.
 As for the definition of the graph structure, the majority of these models use self-learning techniques.
-This is a straightforward choice, since all the papers in this group want to ensure the applicability of the model in different contexts, and the use of pre-defined rules for the definition of the graph limits the model’s universal applicability.
+This is a straightforward choice, since all the papers in this group want to ensure the applicability of the model in different contexts, and the use of pre-defined rules for the definition of the graph limits the model’s universal applicability and negatively impacts knowledge transfer.
 More than half of the papers employ a dynamic graph structure, which adapts to the varying temporal patterns observed in the data.
+In Ref.
+[451] the authors introduce a novel data structure called hypervariate graph, where each time serie values is represented as a graph node, and slicing windows are modeled as fully connected space-time graphs.
+Unlike traditional approaches that treat spatial and temporal aspects separately, according to the authors, this approach captures spatiotemporal correlations in an integrated manner.
 The two most commonly utilized loss functions are the MSE and the MAE.
 Since the graph structure is typically learned by the models themselves, many papers include a regularization term in the loss function for the optimization of the graph structure (e.g., Refs.
-[art_73, art_78, art_79, art_86]).
-The majority of the papers specify the language and libraries used for the code, namely Python with PyTorch, PyTorch Geometric [PyTorch_Geometric], and Torch Spatiotemporal [Cini_Torch_Spatiotemporal_2022].
-However, only slightly less than half of the papers provide a link to the source code.
+[241, 243, 118, 52, 338]).
+The majority of the papers specify the language and libraries used for the code, namely Python with PyTorch, PyTorch Geometric [91], and Torch Spatiotemporal [58].
+However, only slightly more than half of the papers provide a link to the source code.
 The links to the code repositories are listed in Tab.
 LABEL:t:link_codici_generic.
+Ref.
+[59] is a study that does not introduce a new model but rather investigates the relationship between global and local patterns in graph-based spatiotemporal forecasting.
+The authors provide access to their source code at https://github.com/Graph-Machine-Learning-Group/taming-local-effects-stgnns.
+Additionally, this group includes the work that introduced the PyTorch Geometric Temporal library [327], a framework for dynamic graph learning.
+The corresponding source code is publicly available at https://github.com/benedekrozemberczki/pytorch_geometric_temporal.
+In addition, the authors compare various open-source libraries for geometric deep learning.
 Table 34: List of source codes of the “Generic” models in the review.
 Model Link
-Adaptive dependency learning neural network (ADLNN) [art_83] https://github.com/AbishekSriramulu/ADLGNN.git
-Attention-adjusted Graph Spatio-Temporal Network (AGSTN) [art_464] https://github.com/l852888/AGSTN
-Cross-attention temporal dynamic graph neural network (CATodyNet) [art_386] https://github.com/HQYWY/CATodyNet
-Graph neural network with neural Granger causality (CauGNN) [art_77] https://github.com/RRRussell/CauGNN
-Dynamic diffusion-variational GNN (DVGNN) [art_383] https://github.com/gorgen2020/DVGNN
-Gformer [art_82] https://github.com/wxh453751461/Gformer
-GraphSensor [art_395] https://github.com/jldx-gjq/Graphnet/tree/main/graphsensor-code-main
-Multi-scale adaptive graph neural network (MAGNN) [art_80] https://github.com/shangzongjiang/MAGNN
-Multi-task Dynamic Graph generation framework for multivariate time series forecasting (MDG) [art_397] https://github.com/gortwwh/MDG
-Multivariate time series deep spatiotemporal forecasting model with a graph neural network (MDST-GNN) [art_90] https://github.com/yiminghzc/MDST-GNN
-Multivariate Time Series Classification based on Fusion Features (MSTC_FF) [art_385] https://github.com/dumingsen/MTSC_FF
-Multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) [art_75] https://github.com/GRAND-Lab/MTGODE
-Static and dynamic graph learning network (SDGL) [art_78] https://github.com/ZhuoLinLi-shu/SDGL
-Sparse graph learning from spatiotemporal time series [art_72] https://github.com/andreacini/sparse-graph-learning
-Temporal decomposition enhanced graph neural network for multivariate time series forecasting (TDG4MSF) [art_81] https://github.com/TYUT-Theta/MHZN.git
-Temporal graph convolution and attention (T-GAN) [art_202] https://github.com/malei666666/T_GAN
-Temporal Dynamic Graph Neural Network (TodyNet) [art_389] https://github.com/liuxz1011/TodyNet
-Two GNN models with different graph structures [art_160] https://github.com/StefanBloemheuvel/graph_comparison
+Adaptive dependency learning neural network (ADLNN) [355] https://github.com/AbishekSriramulu/ADLGNN.git
+Attention-adjusted graph spatio-temporal network (AGSTN) [277] https://github.com/l852888/AGSTN
+Cross-attention temporal dynamic graph neural network (CATodyNet) [113] https://github.com/HQYWY/CATodyNet
+Graph neural network with neural Granger causality (CauGNN) [76] https://github.com/RRRussell/CauGNN
+Spatio-temporal contrastive learning (CL4ST) [368] https://github.com/HKUDS/CL4ST
+CrossGNN [153] https://github.com/hqh0728/CrossGNN
+Dynamic diffusion-variational GNN (DVGNN) [244] https://github.com/gorgen2020/DVGNN
+Evolving multi-scale graph neural network (ESG) [448] https://github.com/LiuZH-19/ESG
+Fourier graph neural network (FourierGNN) [451] https://github.com/aikunyi/FourierGNN
+Graph interpolation attention recursive network (GinAR) [458] https://github.com/ChengqingYu/GinAR
+Gformer [395] https://github.com/wxh453751461/Gformer
+Graphs for forecasting irregularly sampled time series with missing values (GraFITi) [436] https://github.com/yalavarthivk/GraFITi
+GraphSensor [100] https://github.com/jldx-gjq/Graphnet/tree/main/graphsensor-code-main
+Graph for time series (GTS) [338] https://github.com/chaoshangcs/GTS
+HAR-Sensor [289] https://github.com/riktimmondal/HAR-Sensor
+Hierarchical downsampling time-then-space (HD-TTS) [286] https://github.com/marshka/hdtts
+HyperTime [456] https://gitfront.io/r/user-9180183/WXPgc7aQjSKE/HyperTime/
+Multi-scale adaptive graph neural network (MAGNN) [40] https://github.com/shangzongjiang/MAGNN
+Multi-task dynamic graph generation framework for multivariate time series forecasting (MDG) [152] https://github.com/gortwwh/MDG
+Multivariate time series deep spatiotemporal forecasting model with a graph neural network (MDST-GNN) [138] https://github.com/yiminghzc/MDST-GNN
+Multivariate time series classification based on fusion features (MSTC_FF) [74] https://github.com/dumingsen/MTSC_FF
+Multivariate time series forecasting with graph neural networks (MTGNN) [424] https://github.com/nnzhan/MTGNN
+Multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) [177] https://github.com/GRAND-Lab/MTGODE
+RAINDROP [499] https://github.com/mims-harvard/Raindrop
+Static and dynamic graph learning network (SDGL) [243] https://github.com/ZhuoLinLi-shu/SDGL
+Similarity-aware time-series classification (SimTSC) [472] https://github.com/daochenzha/SimTSC
+Sparse graph learning from spatiotemporal time series [60] https://github.com/andreacini/sparse-graph-learning
+Spatial relation decomposition (SRD) [83] https://github.com/Arthur-Null/SRD/tree/main
+Spatio-temporal graph neural networks (STExplainer) [369] https://github.com/HKUDS/STExplainer
+Time-aware multipersistence spatio-supra graph convolutional network (TAMP-S2GCNet) [51] https://www.dropbox.com/sh/n0ajd5l0tdeyb80/AABGn-ejfV1YtRwjf_L0AOsNa?dl=0
+Temporal decomposition enhanced graph neural network for multivariate time series forecasting (TDG4MSF) [287] https://github.com/TYUT-Theta/MHZN.git
+Temporal graph convolution and attention (T-GAN) [155] https://github.com/malei666666/T_GAN
+Temporal graph neural networks for irregular data (TGNN4I) [300] https://github.com/joeloskarsson/tgnn4i
+Temporal dynamic graph neural network (TodyNet) [261] https://github.com/liuxz1011/TodyNet
+Transformable patching graph neural network (T-PATCHGNN) [491] https://github.com/usail-hkust/t-PatchGNN
+Temporal polynomial graph neural network (TPGNN) [271] https://github.com/zyplanet/TPGNN
+Two GNN models with different graph structures [15] https://github.com/StefanBloemheuvel/graph_comparison
+Unified Spatio-Temporal Diffusion Models (USTD) [143] https://github.com/hjf1997/USTD
+Weight-bounded-GAT [505] https://github.com/zzh237/Weight-bounded-GAT
+Zigzag filtration curve based supra-hodge convolution networks (ZFC-SHCN) [48] https://github.com/zfcshcn/ZFC-SHCN
+Time zigzags at graph convolutional network (Z-GCNET) [50] https://github.com/Z-GCNETs/Z-GCNETs
 5.7.4 Benchmark models
@@
 Category Model Used by
-Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_73], [art_74], [art_78], [art_79], [art_80], [art_81], [art_83], [art_84], [art_88], [art_89], [art_384], [art_393], [art_397], [art_400], [art_405], [art_464]
-Gaussian process (GP) [art_73], [art_74], [art_78], [art_80], [art_81], [art_83]
-Historical average (HA) [art_78], [art_88], [art_400]
-Vector autoregression (VAR) [art_73], [art_74], [art_77], [art_78], [art_79], [art_81], [art_84], [art_88], [art_384], [art_400], [art_405]
-Non-GNN machine learning methods Autoformer [Autoformer] [art_74], [art_76], [art_85]
-DeepAR [DeepAR] [art_69], [art_393]
-DeepSleepNet [DeepSleepNet] [art_395], [art_403]
-Dual self-attention network for multivariate time series forecasting (DSANet) [DSANet] [art_73], [art_78], [art_405]
-Feed-forward neural network (FNN) [art_84], [art_89]
-Gated recurrent unit (GRU) [art_73], [art_78], [art_82], [art_88], [art_90], [art_384], [art_398], [art_404], [art_405]
-Hybrid framework based on fully Dilated CNN (HyDCNN) [HyDCNN] [art_75], [art_83]
-Informer [Informer] [art_74], [art_76], [art_82], [art_85], [art_397]
-$k$-nearest neighbors (KNN) [art_84], [art_387]
-LogTrans [LogTrans] [art_82], [art_85]
-Long-short term memory network (LSTM) [art_71], [art_74], [art_79], [art_82], [art_84], [art_85], [art_86], [art_88], [art_89], [art_396], [art_397], [art_398], [art_400], [art_464]
-Long-short term time series network (LSTNet) [LSTNet] [art_73], [art_74], [art_75], [art_77], [art_78], [art_80], [art_81], [art_82], [art_83], [art_85], [art_87], [art_89], [art_90], [art_387], [art_393]
-Multi-Level Construal Neural Network (MLCNN) [MLCNN] [art_77], [art_87]
-Neural basis expansion analysis for interpretable time series forecasting (N-BEATS) [N-BEATS] [art_69], [art_85]
-Recurrent neural network with fully connected gated recurrent units (RNN-GRU) [art_73], [art_77], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_405]
-Reformer [Reformer] [art_76], [art_82], [art_85]
-SCINet [SCINet] [art_74], [art_90]
-Support vector machine (SVM) [art_84], [art_396], [art_464]
-Temporal convolutional network (TCN) [TCN] [art_88], [art_387], [art_396], [art_397]
-Temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM] [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_90], [art_405]
-Transformer [attention] [art_71], [art_77], [art_393]
-Hybrid of the multilayer perception and autoregressive model (VAR-MLP) [VAR-MLP] [art_73], [art_74], [art_78], [art_79], [art_80], [art_81], [art_83], [art_87], [art_90], [art_405]
-GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_73], [art_74], [art_78], [art_79], [art_80], [art_88], [art_384], [art_396], [art_400], [art_464]
-Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_78], [art_79], [art_383], [art_384], [art_396], [art_397], [art_400], [art_464]
-Attention based Spatial-Temporal Graph Neural Network (ASTGNN) [ASTGNN] [art_74], [art_396]
-Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_73], [art_74], [art_75], [art_78], [art_79], [art_84], [art_86], [art_87], [art_88], [art_383], [art_384], [art_396], [art_397], [art_400], [art_405]
-Dynamic Graph Convolutional Recurrent Network (DGCRN) [art_100] [art_397], [art_398]
-Dynamic and Multi-faceted SpatioTemporal Graph Convolution Network (DMSTGCN) [DMSTGCN] [art_396], [art_417]
-Dynamic Spatial-Temporal Aware Graph Neural Network (DSTAGNN) [DSTAGNN] [art_383], [art_384]
-Graph auto encoder (GAE) [GAE] [art_70], [art_202]
-Graph multi-attention network (GMAN) [GMAN] [art_73], [art_75], [art_86], [art_88]
-Discrete graph structure learning for time series (GTS) [GTS] [art_72], [art_78], [art_384], [art_397], [art_401]
-Graph WaveNet (GWN) [GWN] [art_74], [art_75], [art_79], [art_80], [art_86], [art_87], [art_396], [art_397], [art_417]
-Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_72], [art_73], [art_74], [art_75], [art_78], [art_79], [art_80], [art_81], [art_83], [art_86], [art_87], [art_88], [art_90], [art_393], [art_396], [art_397], [art_401], [art_405], [art_417]
-Multivariate Time series with dynamic Graph neural Ordinary Differential Equations (MTGODE) [art_75] [art_88], [art_401], [art_417]
-Node2vec [node2vec] [art_70], [art_202]
-Spectral temporal graph neural network for multivariate time-series forecasting (StemGNN) [StemGNN] [art_397], [art_405]
-Spatial-temporal fusion graph neural network (STFGNN) [STFGNN] [art_73], [art_86], [art_405]
-Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_73], [art_75], [art_78], [art_79], [art_84], [art_86], [art_88], [art_384], [art_397], [art_400], [art_401]
-Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_75], [art_88], [art_383]
-Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_73], [art_74], [art_78], [art_387], [art_397], [art_400], [art_405]
-Temporal graph convolutional network (T-GCN) [T-GCN] [art_383], [art_384], [art_393]
-Time zigzags at graph convolutional networks for time series forecasting (Z-GCNETs) [Z-GCNETs] [art_400], [art_405]
+Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [241], [99], [243], [118], [40], [287], [355], [47], [473], [57], [231], [152], [390], [240], [450], [338], [83], [448], [277], [424], [61], [271], [49]
+Gaussian process (GP) [241], [99], [243], [40], [287], [355], [83], [448], [424], [61], [271]
+Historical average (HA) [243], [473], [390], [338], [279], [61], [50], [369], [368]
+Vector autoregression (VAR) [241], [99], [76], [243], [118], [287], [47], [473], [390], [240], [432], [338], [50], [451], [356], [369], [368]
+Non-GNN machine learning methods Autoformer [414] [99], [504], [186], [432], [83], [153], [451]
+Crossformer [503] [99], [491], [153]
+DeepAR [329] [36], [231], [409]
+Attentive hierarchical recurrent networks for crime prediction (DeepCrime) [149] [369], [368]
+DeepGLO [335] [432], [51], [26], [48], [451], [49]
+DeepSleepNet [364] [100], [401]
+DeepState [321] [51], [26], [48], [49]
+DLinear [469] [504], [436], [491], [153]
+Deep multi-view spatial-temporal network (DMVST-Net) [447] [450], [49]
+Dual self-attention network for multivariate time series forecasting (DSANet) [157] [241], [243], [240], [50], [368]
+Exponential smoothing transformers for time-series forecasting (ETSformer) [412] [504], [153]
+Frequency enhanced decomposed transformer (FEDformer) [522] [99], [432], [436], [153]
+Feed-forward neural network (FNN) [47], [57], [450], [338], [472]
+Gated recurrent unit (GRU) [241], [243], [395], [473], [138], [453], [337], [240], [286], [499], [491], [50], [436], [356], [300]
+Hybrid framework based on fully Dilated CNN (HyDCNN) [233] [177], [355]
+InceptionTime [162] [83], [472]
+Informer [520] [99], [504], [395], [186], [152], [432], [436], [279], [61], [271], [451], [356]
+$k$-nearest neighbors (KNN) [47], [396], [456], [472]
+LogTrans [294] [395], [186]
+Long-short term memory network (LSTM) [129], [99], [118], [395], [47], [186], [52], [473], [57], [137], [152], [453], [390], [338], [436], [279], [277], [50], [51], [26], [271], [48], [356], [49]
+Long-short term time series network (LSTNet) [202] [241], [99], [177], [76], [243], [40], [287], [395], [355], [186], [397], [57], [138], [396], [231], [432], [83], [448], [424], [61], [51], [26], [271], [48], [451], [356], [49], [268]
+Multi-level construal neural network (MLCNN) [54] [76], [397]
+Multi-time attention networks (mTANs) [349] [436], [499], [491]
+Neural basis expansion analysis for interpretable time series forecasting (N-BEATS) [299] [36], [186], [51], [26], [48], [49]
+Channel-independent patch time series Transformer (PatchTST) [295] [491], [153]
+Reformer [191] [504], [395], [186], [432], [451]
+Recurrent neural network (RNN) [57], [59]
+Recurrent neural network with fully connected gated recurrent units (RNN-GRU) [241], [76], [243], [118], [40], [287], [355], [397], [240], [448], [424], [271]
+Rocket [64] [140], [83]
+SCINet [264] [99], [138]
+Set functions for time series (SeFT) [141] [499], [491]
+State frequency memory (SFM) [484] [432], [51], [26], [48], [451]
+Shapelet-neural network (ShapeNet) [215] [113], [261], [456]
+Spatial-temporal dynamic network (STDN) [446] [450], [369], [49], [368]
+Spatial-temporal identity (STID) [340] [99], [458]
+Spatial-temporal transformer network (STtrans) [418] [369], [368]
+Spatio-temporal residual networks (ST-ResNet) [480] [369], [49], [368]
+Support vector machine (SVM) [47], [137], [338], [277], [49], [368]
+TapNet [495] [74], [261], [456], [472]
+Temporal convolutional network (TCN) [9] [473], [396], [137], [152], [432], [51], [26], [48], [451], [49]
+Temporal 2D-variation modeling for general time series analysis (TimesNet) [413] [504], [491], [456], [153]
+Temporal pattern attention long-short term memory network (TPA-LSTM) [347] [241], [99], [177], [243], [118], [40], [287], [355], [138], [240], [448], [424], [61], [271], [268]
+TimeGrad [322] [143], [409]
+Transformer [375] [129], [76], [231], [499], [300]
+Hybrid of the multilayer perception and autoregressive model (VAR-MLP) [485] [241], [99], [243], [118], [40], [287], [355], [397], [138], [240], [448], [424], [61], [271], [268]
+WEASEL + multivariate unsupervised symbols and derivatives (WEASEL+MUSE) [333] [74], [113], [261], [456]
+Extreme gradient boosting (XGBoost) [450], [448], [279]
+GNN methods Adaptive graph convolutional recurrent network (AGCRN) [8] [241], [99], [243], [118], [40], [473], [137], [390], [432], [286], [279], [50], [51], [48], [451], [59], [369], [49], [327], [368]
+Attention-based spatial-temporal graph convolutional network (ASTGCN) [116] [243], [118], [244], [137], [152], [390], [277], [50], [369], [505], [368]
+Attention based spatial-temporal graph neural network (ASTGNN) [117] [99], [137]
+Coupled layer-wise convolutional recurrent neural network (CCRNN) [449] [448], [279]
+CrossGNN [153] [491], [153]
+Diffusion convolutional recurrent neural network (DCRNN) [237] [241], [99], [177], [243], [118], [47], [52], [397], [473], [244], [137], [152], [390], [240], [338], [286], [83], [448], [279], [277], [424], [50], [51], [271], [48], [59], [143], [409], [369], [49], [327], [368]
+Dynamic graph convolutional recurrent network (DGCRN) [212] [152], [453]
+Dynamic and multi-faceted spatiotemporal graph convolution network (DMSTGCN) [127] [137], [443], [83], [369], [368]
+Dynamic spatial-temporal aware graph neural network (DSTAGNN) [203] [244], [369]
+Evolving multi-scale graph neural network (ESG) [448] [99], [279]
+Graph auto encoder (GAE) [189] [232], [155]
+Graph deviation network (GDN) [66] [396], [505]
+Graph multi-attention network (GMAN) [513] [241], [177], [52], [473], [424], [369], [505], [368]
+Graph-based multi-step dependency relation (GMSDR) [259] [143], [409], [369], [368]
+Discrete graph structure learning for time series (GTS) [338] [60], [243], [152], [360], [448], [279]
+Graph WaveNet (GWN) [425] [99], [177], [118], [40], [52], [397], [137], [152], [443], [432], [286], [491], [279], [424], [50], [51], [26], [271], [48], [451], [59], [143], [369], [49], [368]
+Hypergraph convolutional recurrent neural network (HGC-RNN) [450] [390], [49]
+Long short-term graph convolutional networks (LSGCN) [154] [50], [368]
+Multi-range attentive bicomponent graph convolutional network (MRA-BGCN) [44] [52], [424]
+Multivariate time series forecasting with graph neural network (MTGNN) [424] [60], [241], [99], [177], [243], [118], [40], [287], [355], [52], [397], [473], [138], [231], [137], [152], [360], [240], [443], [432], [491], [83], [448], [61], [271], [153], [451], [268], [505]
+Multivariate time series with dynamic graph neural ordinary differential equations (MTGODE) [177] [473], [360], [443]
+Node2vec [110] [232], [155]
+Spectral temporal graph neural network for multivariate time-series forecasting (StemGNN) [26] [152], [240], [432], [491], [448], [51], [26], [271], [48], [451], [369], [49], [368]
+Spatial-temporal fusion graph neural network (STFGNN) [228] [241], [52], [240], [369], [368]
+Spatio-temporal graph convolutional network (STGCN) [457] [241], [177], [243], [118], [47], [52], [473], [152], [390], [360], [448], [424], [50], [51], [26], [271], [48], [143], [409], [369], [49], [505], [368]
+Spatio-temporal graph neural controlled differential equation (STG-NCDE) [56] [177], [143], [409], [369], [368]
+Spatial-temporal graph ODE networks (STGODE) [84] [177], [473], [244], [143], [369], [368]
+Deep-meta-learning based model (ST-MetaNet) [306] [52], [424], [369], [49], [368]
+Spatiotemporal multi-graph convolution network (ST-MGCN) [101] [277], [49]
+Spatio-temporal synchronous graph convolutional network (STSGCN) [353] [241], [99], [243], [396], [152], [390], [240], [448], [50], [143], [369], [368]
+Spatial-temporal sequential hypergraph network (ST-SHN) [427] [369], [368]
+Time-aware multipersistence spatio-supra graph convolutional network (TAMP-S2GCNet) [51] [369], [368]
+Temporal graph convolutional network (T-GCN) [508] [244], [231], [327]
+Temporal dynamic graph neural network (TodyNet) [261] [113], [140], [456]
+Time zigzags at graph convolutional networks for time series forecasting (Z-GCNETs) [50] [390], [240], [51], [26], [48], [369], [49], [368]
 This “Generic” thematic group includes a wide variety of benchmark models, also due to the large number of papers within this group.
 As for mathematical and statistical models, autoregressions and ARIMA models are very popular for their simplicity and interpretability.
-Among non-GNN machine learning models, the most widely used models are LSTM networks, the so-called temporal pattern attention long-short term memory network (TPA-LSTM) [TPA-LSTM], and the long-short term time series network (LSTNet) [LSTNet].
+Among non-GNN machine learning models, the most widely used models are LSTM networks, the so-called temporal pattern attention long-short term memory network (TPA-LSTM) [347], and the long-short term time series network (LSTNet) [202].
 The TPA-LSTM model, introduced by Shih et al.
@@
 The source code for the model is accessible at https://github.com/laiguokun/LSTNet.
-Among GNN models, the most common ones are DCRNN [DCRNN] and MTGNN [MTGNN] models, already discussed in the previous subsections.
+Among GNN models, the most common ones are DCRNN [237], MTGNN [424], and GWN [425] models, already discussed in the previous subsections.
 5.7.5 Results
@@
 This subsection about results compares the results obtained for the Electricity consumption, Exchange rate, Solar energy and Traffic datasets.
-The results of the proposed models and benchmarks on METR-LA and PEMS datasets were already reported with a in Subsec.
+The results of the proposed models and benchmarks on the “Mobility” datasets were already reported with a in Subsec.
 5.5.5.
@@
 Metrics RRSE CORR RRSE CORR RRSE CORR RRSE CORR
-ADLGNN [art_83] 0.0719 0.9506 0.0809 0.9386 0.0887 0.9312 0.0930 0.9294
+ADLGNN [355] 0.0719 0.9506 0.0809 0.9386 0.0887 0.9312 0.0930 0.9294
 AGCRN 0.0766 0.9408 0.0894 0.9309 0.0921 0.9222 0.0967 0.9183
-AGLG-GRU [art_79] 0.0738 0.9434 0.0864 0.9302 0.0912 0.9283 0.0947 0.9274
+AGLG-GRU [118] 0.0738 0.9434 0.0864 0.9302 0.0912 0.9283 0.0947 0.9274
 AR 0.0995 0.8845 0.1035 0.8632 0.1050 0.8591 0.1054 0.8595
 ARIMA 0.0917 0.8902 0.1002 0.8834 0.1028 0.8604 0.1042 0.8487
-ASTCRF [art_384] 0.0742 0.9593 0.0872 0.9511 0.0913 0.9388 - -
 Autoformer 0.1258 0.9147 0.1344 0.9001 0.1357 0.8921 0.1554 0.8704
 Crossformer 0.0742 0.9452 0.0855 0.9351 0.0901 0.9280 0.0967 0.9205
-DPGNN [art_405] 0.0720 0.9527 0.0810 0.9422 0.0893 0.9357 0.0935 0.9291
-DSTGN [art_73] 0.0713 0.9518 0.0821 0.9424 0.0887 0.9357 - -
-DSTIGNN [art_74] 0.0733 0.9515 0.0820 0.9412 0.0899 0.9338 0.0954 0.9289
-ESG 0.0718 0.9494 0.0844 0.9372 0.0898 0.9321 0.0962 0.9279
+DPGNN [240] 0.0720 0.9527 0.0810 0.9422 0.0893 0.9357 0.0935 0.9291
+DSTGN [241] 0.0713 0.9518 0.0821 0.9424 0.0887 0.9357 - -
+DSTIGNN [99] 0.0733 0.9515 0.0820 0.9412 0.0899 0.9338 0.0954 0.9289
+ESG [448] 0.0718 0.9494 0.0844 0.9372 0.0898 0.9321 0.0962 0.9279
 FEDformer 0.0889 0.9321 0.1006 0.9191 0.1154 0.9080 0.1202 0.9012
@@
 LSTNet 0.0864 0.9283 0.0931 0.9135 0.1007 0.9077 0.1007 0.9119
-MAGNN [art_80] 0.0745 0.9476 0.0876 0.9323 0.0908 0.9282 0.0963 0.9217
-MDST-GNN [art_90] 0.0738 0.9454 0.0833 0.9346 0.0884 0.9264 0.0922 0.9222
-MTGNN 0.0745 0.9474 0.0878 0.9316 0.0916 0.9278 0.0953 0.9234
-MTGODE [art_75] 0.0736 0.9430 0.0809 0.9340 0.0891 0.9279 - -
+MAGL [268] 0.0728 0.9476 0.0824 0.9377 0.0893 0.9299 0.0951 0.9240
+MAGNN [40] 0.0745 0.9476 0.0876 0.9323 0.0908 0.9282 0.0963 0.9217
+MDST-GNN [138] 0.0738 0.9454 0.0833 0.9346 0.0884 0.9264 0.0922 0.9222
+MTGNN [424] 0.0745 0.9474 0.0878 0.9316 0.0916 0.9278 0.0953 0.9234
+MTGODE [177] 0.0736 0.9430 0.0809 0.9340 0.0891 0.9279 - -
 MTHetGNN 0.0749 0.9456 0.0892 0.9307 0.0959 0.8783 0.0969 0.8782
@@
 SARIMA 0.0906 0.9055 0.0999 0.8829 0.1026 0.8674 0.1036 0.8621
-SCINet 0.0740 0.9494 0.0845 0.9387 0.0929 0.9305 0.0967 0.927
-SDGL [art_78] 0.0698 0.9534 0.0805 0.9445 0.0889 0.9351 0.0935 0.9301
+SCINet 0.0740 0.9494 0.0845 0.9387 0.0929 0.9305 0.0967 0.9270
+SDGL [243] 0.0698 0.9534 0.0805 0.9445 0.0889 0.9351 0.0935 0.9301
 SDLGNN 0.0726 0.9502 0.0820 0.9384 0.0896 0.9304 0.0947 0.9257
 SDLGNN-Corr 0.0737 0.9475 0.0841 0.9346 0.0923 0.9263 0.0971 0.9227
-STG-NCDE 0.6152 0.8739 0.6584 0.8663 0.7302 0.8728 - -
 StemGNN 0.0799 0.9490 0.0909 0.9397 0.0989 0.9342 0.1019 0.9209
-TDG4-MSF [art_81] 0.0731 0.9499 0.0828 0.9371 0.0894 0.9306 0.0969 0.9246
+TDG4-MSF [287] 0.0731 0.9499 0.0828 0.9371 0.0894 0.9306 0.0969 0.9246
 Theta 0.0975 0.8906 0.1029 0.8723 0.1040 0.8576 0.1041 0.8595
+TPGNN [271] 0.0627 0.9417 0.0685 0.9362 0.0699 0.9285 0.0936 0.9293
 TPA-LSTM 0.0823 0.9439 0.0916 0.9337 0.0964 0.9250 0.1006 0.9133
@@
 VAR-MLP 0.1393 0.8708 0.1620 0.8389 0.1557 0.8192 0.1274 0.8679
-The accuracy of the various models varies depending on the studied forecasting horizon but, overall, one can notice that the GNN models perform better than the others, with the static and dynamic graph learning network (SDGL) [art_78] being the most accurate.
+The accuracy of the various models varies depending on the studied forecasting horizon but, overall, one can notice that the GNN models perform better than the others, with the static and dynamic graph learning network (SDGL) [243] and the temporal polynomial graph neural network (TPGNN) [271] being the most accurate.
 The SDGL model, proposed by Li et al.
@@
 The graph input then passes through gated temporal convolutional networks and graph convolution modules, which are equipped with residual and skip connections that deliver the information to the output layer.
+The TPGNN, proposed by Liu et al.
+in 2022, builds a series of matrix polynomials that share a common matrix basis.
+The graph learning module first captures the overall correlations using a static matrix basis.
+Then, a set of time-varying coefficients and the matrix basis are used to create a dynamic graph.
+Overall, the TPGNN model uses an encoder-decoder architecture and makes predictions in an auto-regressive manner.
 Tab.
@@
 AGCRN 0.0269 0.9717 0.0331 0.9615 0.0374 0.9531 0.0476 0.9334
-AGLG-GRU [art_79] 0.0191 0.9792 0.0233 0.9701 0.0328 0.9548 0.0449 0.9372
+AGLG-GRU [118] 0.0191 0.9792 0.0233 0.9701 0.0328 0.9548 0.0449 0.9372
 AR 0.0228 0.9734 0.0279 0.9656 0.0353 0.9526 0.0445 0.9357
 ARIMA 0.0198 0.9754 0.0261 0.9721 0.0344 0.9548 0.0445 0.9301
-ASTCRF [art_384] 0.0165 0.9792 0.0235 0.9709 0.0344 0.9572 - -
 Autoformer 0.0291 0.9631 0.0379 0.9397 0.0441 0.9201 0.0501 0.9012
 Crossformer 0.0226 0.9759 0.0269 0.9656 0.0358 0.9436 0.0489 0.9291
-DPGNN [art_405] 0.0184 0.9797 0.0246 0.9715 0.0346 0.9564 0.0456 0.9386
-DSTGN [art_73] 0.0179 0.9782 0.0250 0.9715 0.0346 0.9562 - -
-DSTIGNN [art_74] 0.0173 0.9818 0.0244 0.9723 0.0333 0.9571 0.0430 0.9407
-ESG 0.0181 0.9792 0.0246 0.9717 0.0345 0.9564 0.0468 0.9392
+DPGNN [240] 0.0184 0.9797 0.0246 0.9715 0.0346 0.9564 0.0456 0.9386
+DSTGN [241] 0.0179 0.9782 0.0250 0.9715 0.0346 0.9562 - -
+DSTIGNN [99] 0.0173 0.9818 0.0244 0.9723 0.0333 0.9571 0.0430 0.9407
+ESG [448] 0.0181 0.9792 0.0246 0.9717 0.0345 0.9564 0.0468 0.9392
 FEDformer 0.0256 0.9701 0.0287 0.9555 0.0379 0.9385 0.0487 0.9190
@@
 LSTNet 0.0226 0.9735 0.0280 0.9658 0.0356 0.9511 0.0449 0.9354
-MAGNN [art_80] 0.0183 0.9778 0.0246 0.9712 0.0343 0.9557 0.0474 0.9339
-MDST-GNN [art_90] 0.0172 0.9811 0.0245 0.9727 0.0337 0.9578 0.0431 0.9392
-MTGNN 0.0194 0.9786 0.0259 0.9708 0.0349 0.9551 0.0456 0.9372
+MAGL [268] 0.0178 0.9797 0.0246 0.9716 0.0337 0.9567 0.0444 0.9361
+MAGNN [40] 0.0183 0.9778 0.0246 0.9712 0.0343 0.9557 0.0474 0.9339
+MDST-GNN [138] 0.0172 0.9811 0.0245 0.9727 0.0337 0.9578 0.0431 0.9392
+MTGNN [424] 0.0194 0.9786 0.0259 0.9708 0.0349 0.9551 0.0456 0.9372
 MTHetGNN 0.0198 0.9769 0.0259 0.9701 0.0345 0.9539 0.0451 0.9360
@@
 SCINet 0.0171 0.9787 0.0240 0.9704 0.0331 0.9553 0.0436 0.9396
-SDGL [art_78] 0.0180 0.9808 0.0249 0.9730 0.0342 0.9583 0.0455 0.9402
+SDGL [243] 0.0180 0.9808 0.0249 0.9730 0.0342 0.9583 0.0455 0.9402
 StemGNN 0.0506 0.8871 0.0674 0.8703 0.0676 0.8499 0.0685 0.8738
-TDG4-MSF [art_81] 0.0172 0.9825 0.0244 0.9718 0.0330 0.9569 0.0433 0.9386
+TDG4-MSF [287] 0.0172 0.9825 0.0244 0.9718 0.0330 0.9569 0.0433 0.9386
 Theta 0.0497 0.9738 0.0257 0.9656 0.0342 0.9510 0.0441 0.9323
 TPA-LSTM 0.0174 0.9790 0.0241 0.9709 0.0341 0.9564 0.0444 0.9381
+TPGNN [271] 0.0174 0.9792 0.0250 0.9687 0.0350 0.9509 0.0458 0.9306
 TRMF 0.0351 0.9142 0.0875 0.8123 0.0494 0.8993 0.0563 0.8678
@@
 in Ref.
-[GTS] and used as a benchmark in Ref.
-[art_78], demonstrates itself to have a superior performance.
-It is therefore a suitable benchmark for comparison.
-The GTS model was introduced by Shang et al.
-in 2021 for traffic and sensor network forecasting.
+[338] for traffic and sensor network forecasting, and used as a benchmark in Ref.
+[243], demonstrates itself to have a superior performance.
 It includes a probabilistic graph learning module that optimizes the mean performance over the graph distribution, and a recurrent graph convolutional module to generate the forecasts.
@@
 Metrics RRSE CORR RRSE CORR RRSE CORR RRSE CORR
-ADLGNN [art_83] 0.1708 0.9866 0.2188 0.9768 0.2897 0.9551 0.4128 0.9060
+ADLGNN [355] 0.1708 0.9866 0.2188 0.9768 0.2897 0.9551 0.4128 0.9060
 AGCRN 0.1840 0.9841 0.2432 0.9708 0.3185 0.9487 0.4141 0.9087
-AGLG-GRU [art_79] 0.1762 0.9842 0.2302 0.9682 0.3021 0.9532 0.4130 0.9084
+AGLG-GRU [118] 0.1762 0.9842 0.2302 0.9682 0.3021 0.9532 0.4130 0.9084
 AR 0.2435 0.9710 0.3790 0.9263 0.5911 0.8107 0.8699 0.5314
 ARIMA 0.2328 0.9739 0.3413 0.9402 0.4531 0.8886 0.5810 0.8133
-ASTCRF [art_384] 0.1776 0.9855 0.2347 0.9743 0.3107 0.9503 - -
-Autoformer 0.1935 0.9784 0.2604 0.9651 0.3959 0.9111 0.6064 0.818
+Autoformer 0.1935 0.9784 0.2604 0.9651 0.3959 0.9111 0.6064 0.8180
 Crossformer 0.1758 0.9801 0.2304 0.9679 0.3101 0.9437 0.4115 0.9001
-DPGNN [art_405] 0.1794 0.9850 0.2364 0.9724 0.3003 0.9543 0.4091 0.9117
-DSTGN [art_73] 0.1787 0.9867 0.2358 0.9721 0.3079 0.9513 - -
-DSTIGNN [art_74] 0.1684 0.9869 0.2165 0.9773 0.2863 0.9584 0.4064 0.9118
-ESG 0.1708 0.9865 0.2278 0.9743 0.3073 0.9519 0.4101 0.9100
+DPGNN [240] 0.1794 0.9850 0.2364 0.9724 0.3003 0.9543 0.4091 0.9117
+DSTGN [241] 0.1787 0.9867 0.2358 0.9721 0.3079 0.9513 - -
+DSTIGNN [99] 0.1684 0.9869 0.2165 0.9773 0.2863 0.9584 0.4064 0.9118
+ESG [448] 0.1708 0.9865 0.2278 0.9743 0.3073 0.9519 0.4101 0.9100
 FEDformer 0.1901 0.9827 0.2444 0.9606 0.3502 0.9275 0.4851 0.8876
@@
 LSTNet 0.1843 0.9843 0.2559 0.9690 0.3254 0.9467 0.4643 0.887
-MAGNN [art_80] 0.1771 0.9853 0.2361 0.9724 0.3015 0.9539 0.4108 0.9097
-MDST-GNN [art_90] 0.1764 0.9855 0.2321 0.9735 0.3082 0.9519 0.4119 0.9103
-MTGNN 0.1778 0.9852 0.2348 0.9726 0.3109 0.9509 0.4270 0.9031
-MTGODE [art_75] 0.1693 0.9868 0.2171 0.9771 0.2901 0.9577 - -
+MAGL [268] 0.1678 0.9869 0.2220 0.9759 0.2998 0.9550 0.4188 0.9057
+MAGNN [40] 0.1771 0.9853 0.2361 0.9724 0.3015 0.9539 0.4108 0.9097
+MDST-GNN [138] 0.1764 0.9855 0.2321 0.9735 0.3082 0.9519 0.4119 0.9103
+MTGNN [424] 0.1778 0.9852 0.2348 0.9726 0.3109 0.9509 0.4270 0.9031
+MTGODE [177] 0.1693 0.9868 0.2171 0.9771 0.2901 0.9577 - -
 MTHetGNN 0.1838 0.9845 0.2600 0.9681 0.3169 0.9486 0.4231 0.9031
@@
 SCINet 0.1775 0.9853 0.2301 0.9739 0.2997 0.9550 0.4081 0.9112
-SDGL [art_78] 0.1699 0.9866 0.2222 0.9762 0.2924 0.9565 0.4047 0.9119
+SDGL [243] 0.1699 0.9866 0.2222 0.9762 0.2924 0.9565 0.4047 0.9119
 SDLGNN 0.1720 0.9864 0.2249 0.9757 0.3024 0.9547 0.4184 0.9051
@@
 STG-NCDE 0.2346 0.9748 0.2908 0.9605 0.5149 0.8639 - -
-TDG4-MSF [art_81] 0.1746 0.9858 0.2348 0.9727 0.3082 0.9527 0.4031 0.9143
+TDG4-MSF [287] 0.1746 0.9858 0.2348 0.9727 0.3082 0.9527 0.4031 0.9143
 Theta 0.2442 0.9685 0.3327 0.9445 0.4488 0.8979 0.6092 0.8139
 TPA-LSTM 0.1803 0.9850 0.2347 0.9742 0.3234 0.9487 0.4389 0.9081
-TRMF 0.2473 0.9703 0.3470 0.9418 0.5597 0.8475 0.9005 0.5598
+TPGNN [271] 0.1850 0.9840 0.2412 0.9716 0.3059 0.9529 0.3498 0.9710
 VAR-MLP 0.1922 0.9829 0.2679 0.9655 0.4244 0.9058 0.6841 0.7149
-The results indicate that the dynamic spatiotemporal interactive graph neural network (DSTIGNN) model [art_74] and the static and dynamic graph learning network (SDGL) model [art_78] are the most accurate on the solar energy dataset, for shorter and longer time horizons, respectively.
+The results indicate that the dynamic spatiotemporal interactive graph neural network (DSTIGNN) model [99] is the most accurate on the solar energy dataset.
 The DSTIGNN model, introduced by Gao et al.
@@
 Once the graph structure is determined, a stack of graph convolutional layers, enhanced with residual connections, along with a convolutional output module, generate the final forecast.
-The description of the SDGL instead has already been provided in this subsection, in the discussion of the results on the Electricity consumption dataset.
 Finally, Tab.
@@
 Metrics RRSE CORR RRSE CORR RRSE CORR RRSE CORR
-ADLGNN [art_83] 0.4047 0.9028 0.4201 0.8928 0.4299 0.8876 0.4416 0.8818
+ADLGNN [355] 0.4047 0.9028 0.4201 0.8928 0.4299 0.8876 0.4416 0.8818
 AGCRN 0.4379 0.8850 0.4635 0.8670 0.4694 0.8679 0.4707 0.8664
-AGLG-GRU [art_79] 0.4173 0.8958 0.4722 0.8541 0.4427 0.8755 0.4526 0.8842
+AGLG-GRU [118] 0.4173 0.8958 0.4722 0.8541 0.4427 0.8755 0.4526 0.8842
 AR 0.5991 0.7752 0.6218 0.7568 0.6252 0.7544 0.6300 0.7519
@@
 LSNet 0.4777 0.8721 0.4893 0.8690 0.4950 0.8614 0.4973 0.8588
-MAGNN [art_80] 0.4097 0.8992 0.4555 0.8753 0.4423 0.8815 0.4434 0.8813
-MDST-GNN [art_90] 0.4162 0.8958 0.4461 0.8803 0.4377 0.8841 0.4452 0.8792
-MTGNN 0.4162 0.8963 0.4754 0.8667 0.4461 0.8794 0.4535 0.8810
-MTGODE [art_75] 0.4127 0.900 0.4259 0.8945 0.4329 0.8899 - -
+MAGL [268] 0.4058 0.9013 0.4189 0.8935 0.4308 0.8871 0.4417 0.8810
+MAGNN [40] 0.4097 0.8992 0.4555 0.8753 0.4423 0.8815 0.4434 0.8813
+MDST-GNN [138] 0.4162 0.8958 0.4461 0.8803 0.4377 0.8841 0.4452 0.8792
+MTGNN [424] 0.4162 0.8963 0.4754 0.8667 0.4461 0.8794 0.4535 0.8810
+MTGODE [177] 0.4127 0.9000 0.4259 0.8945 0.4329 0.8899 - -
 MTHetGNN 0.4826 0.8643 0.5198 0.8452 0.5147 0.8744 0.5250 0.8418
@@
 SCINet 0.4216 0.8920 0.4414 0.8809 0.4495 0.8772 0.4453 0.8825
-SDGL [art_78] 0.4142 0.9010 0.4475 0.8825 0.4584 0.8760 0.4571 0.8766
+SDGL [243] 0.4142 0.9010 0.4475 0.8825 0.4584 0.8760 0.4571 0.8766
 SDLGNN 0.4053 0.9017 0.4209 0.8925 0.4313 0.8868 0.4444 0.8801
 SDLGNN-Corr 0.4227 0.8937 0.4378 0.8846 0.4576 0.8746 0.4579 0.8784
-TDG4-MSF [art_81] 0.4029 0.9014 0.4196 0.8925 0.4294 0.8864 0.4366 0.8834
+TDG4-MSF [287] 0.4029 0.9014 0.4196 0.8925 0.4294 0.8864 0.4366 0.8834
 Theta 0.6071 0.7768 0.6241 0.7606 0.6271 0.7591 0.6012 0.7830
 TPA-LSTM 0.4487 0.8812 0.4658 0.8717 0.4641 0.8717 0.4765 0.8629
+TPGNN [271] 0.3989 0.9232 0.4715 0.8945 0.4476 0.9028 0.4696 0.8858
 TRMF 0.6708 0.6964 0.6261 0.7430 0.5956 0.7748 0.6442 0.7278
@@
 The accuracy of the models varies by horizon and metric.
-However, the best one in terms of RRSE over all horizons is the temporal decomposition enhanced graph neural network for multivariate time series forecasting (TDG4MSF) [art_81], proposed by Miao et al.
-in 2021.
-The TDG4MSF model is composed of four components: temporal decomposition enhanced representation learning for extracting short and long term periodic patterns, graph structure learning, gated graph neural network depending on linear transformation for representation learning, and MLP-based output layer.
+Overall, the best models are the temporal polynomial graph neural network (TPGNN) [271] already discussed for the Electricity consumption dataset, and the temporal decomposition enhanced graph neural network for multivariate time series forecasting (TDG4-MSF) [287].
+The TDG4-MSF model, proposed by Miao et al.
+in 2021, is composed of four components: temporal decomposition enhanced representation learning for extracting short and long term periodic patterns, graph structure learning, gated graph neural network depending on linear transformation for representation learning, and MLP-based output layer.
+There are also other pairs of papers whose results can be compared.
+However, due to space limitations, we do not include tables comparing the results of just two papers in this “Generic” subsection.
+Instead, the reader is referred to the studies Refs.
+[432] and [451] for the CSSE COVID-19 and ECG 5000 datasets; Refs.
+[50] and [51] for the Bytom dataset; Refs.
+[504] and [153] for the Jena weather dataset.
 5.8 Other topics
-This final subsection includes 29 papers that fall outside the previously discussed groups.
-The case studies include cellular traffic prediction, recommendation systems for user preferences, and other forecasting or classification problems.
+This final subsection includes 34 papers that fall outside the previously discussed groups.
+The case studies include cellular traffic prediction, modulation classification, and other forecasting or classification problems.
 Given the heterogeneity of the papers, only a general overview is provided here, and no comparisons of the approaches or discussions of benchmarks are presented.
@@
 However, since only a few papers would belong to each potential additional theme, they are grouped under the label “Other topics”.
-A frequently discussed topic in this “Other topics” category is cellular traffic prediction [art_138, art_144, art_148, art_468, art_457], in terms of SMS messages, calls, internet connections, phone’s traffic statistics or locations, which are essential for the cellular network resource management system.
-Another topic included is modulation classification [art_12, art_13, art_303, art_459, art_463], which is the process of determining the modulation used at the transmitter based on observations of the received signal [modulation_classification_definition].
-Two further areas of investigation to be added are the study of user preferences [art_147, art_151] to enable personalized experiences and recommendations, and anomalous IoT intrusion detection [4, art_460].
-Other applications discussed in the papers of this subsection are encrypted traffic classification [art_118], urban spatio-temporal event prediction [art_136], prediction of the quality of operational processes [art_137], seismic intensity forecasting [art_139] (with source code available at https://github.com/StefanBloemheuvel/GCNTimeseriesRegression), prediction of spatio-temporal dynamics with complex structures [art_140], tailing dam monitoring [art_141], forecasting of educational video engagement [art_142], subsurface production forecasting [art_143], time series prediction for data centers maintenance [art_145], hydraulic runoff [art_146], water demand forecasting [art_149], human activity recognition [art_159] (with source code available at https://github.com/riktimmondal/HAR-Sensor ), crime prediction [art_465] (with source code available at https://github.com/ZJUDataIntelligence/HDM-GNN), polypropylene production data [art_455], and tungsten flotation process [art_458] (with source code available at https://github.com/ywmcsu/SASGNN).
+A frequently discussed topic in this “Other topics” category is cellular traffic prediction [404, 34, 459, 170, 254, 108, 252], in terms of SMS messages, calls, internet connections, phone’s traffic statistics or locations, which are essential for the cellular network resource management system.
+Another topic included is modulation classification [380, 435, 6, 487, 454], which is the process of determining the modulation used at the transmitter based on observations of the received signal [124].
+Four further areas of investigation to be added are the study of product demand and user preferences [92, 501] to enable personalized experiences and recommendations, anomalous IoT intrusion detection [3, 394], encrypted traffic classification [70, 475].
+Other applications discussed in the papers of this subsection are urban spatio-temporal event prediction [175], prediction of the quality of operational processes [492], seismic intensity forecasting [16] (with source code available at https://github.com/StefanBloemheuvel/GCNTimeseriesRegression), prediction of spatio-temporal dynamics with complex structures [410], tailing dam monitoring [328], subsurface production forecasting [266], time series prediction for data centers maintenance [344], hydraulic runoff [440], water demand forecasting [465], crime prediction [519] (with source code available at https://github.com/ZJUDataIntelligence/HDM-GNN), polypropylene production data [130], tungsten flotation process [399] (with source code available at https://github.com/ywmcsu/SASGNN), talent and demand supply for labor market [121] (with source code available at https://github.com/gzn00417/DH-GEM), anomaly detection in microservice systems [150] (with source code available at https://github.com/ant-research/microservice_system_twin_graph_based_anomaly_detection ), and emergency supply forecasting for logistics systems [255].
 5.8.2 Datasets
@@
 The two datasets for modulation classification used in Refs.
-[art_12, art_13, art_303, art_463] are RML2016.10a (https://www.kaggle.com/datasets/raindrops12/rml201610a, source paper: http://dx.doi.org/10.1007/978-3-319-44188-7_16) and RML2016.10b (https://www.kaggle.com/datasets/marwanabudeeb/rml201610b, source paper: https://pubs.gnuradio.org/index.php/grcon/article/view/11).
+[380, 435, 6, 454] are RML2016.10a (https://www.kaggle.com/datasets/raindrops12/rml201610a, source paper: http://dx.doi.org/10.1007/978-3-319-44188-7_16) and RML2016.10b (https://www.kaggle.com/datasets/marwanabudeeb/rml201610b, source paper: https://pubs.gnuradio.org/index.php/grcon/article/view/11).
 They include 220,000 and 1.2 million samples respectively, in many modulation types.
@@
 As for the cellular traffic prediction dataset used in Refs.
-[art_138, art_144, art_468], it was proposed by Telecom Italia and MIT Media Lab during the “Telecom Italia Big Data Challenge”, and it is available at https://dataverse.harvard.edu/dataverse/bigdatachallenge.
+[303, 404, 34, 170, 108], it was proposed by Telecom Italia and MIT Media Lab during the “Telecom Italia Big Data Challenge”, and it is available at https://dataverse.harvard.edu/dataverse/bigdatachallenge.
 The investigated geographical area is divided in a grid, and each cell records the number of SMS messages, calls, and wireless network traffic data in a 10 minutes interval.
 Finally, the Bot-IoT dataset used in Refs.
-[4, art_460] and available at https://www.unsw.adfa.edu.au/unsw-canberra-cyber/cybersecurity/ADFA-NB15-Datasets/bot_iot.php, was created by designing a realistic network environment in the Cyber Range Lab of UNSW Canberra, and incorporates a combination of normal and botnet traffic.
+[3, 394] and available at https://www.unsw.adfa.edu.au/unsw-canberra-cyber/cybersecurity/ADFA-NB15-Datasets/bot_iot.php, was created by designing a realistic network environment in the Cyber Range Lab of UNSW Canberra, and incorporates a combination of normal and botnet traffic.
 6 Discussion
-As stated in the Introduction, the aim of this SLR is to provide a comprehensive and detailed overview of spatio-temporal GNN models for time series classification and forecasting in different fields, and to assess their performance.
-Existing literature was collected and synthesized to provide an overview of datasets, models, and tables of results that will hopefully assist researchers in their future work.
-In the Introduction two sets of research questions were stated: a generic set and a specific set.
-This section will collect the answers to those research questions.
+As stated in the Introduction, the aim of this SLR is to provide a comprehensive but also detailed overview of spatio-temporal GNN models for time series classification and forecasting used across different fields, assessing their performance and highlighting patterns that could guide future research.
+Existing literature was collected and synthesized to provide both an overview and a distillation of datasets, models, and tables of results, addressing the two sets of research questions posed in the Introduction.
+Although our intention was to provide a broad and general perspective on widely used spatio-temporal GNN models and identify patterns that could be generalized across different domains and datasets, our analysis revealed that such generalizability is highly elusive.
+This characteristics stems from the primary strength of spatio-temporal GNNs: the possibility of explicitly modeling relationships between different variables, which is inherently dataset-dependent and highly heterogeneous across different applications.
+For instance, models for human mobility (e.g., traffic flow, epidemic spread) share intrinsic similarities but differ in temporal scales, spatial resolution, and interaction types.
+Similarly, environmental applications such as weather forecasting and energy demand prediction are both influenced by climatic factors but rely on different variables and relationships.
+We believe that a comprehensive perspective can be drawn after such an extensive examination of the available literature.
+Hence, we attempt now to discuss common patterns and insights that are not immediately apparent across different domains.
+In this section, we first address the research questions posed in the Introduction, then summarize prevailing research trends, and finally present a meta-analysis aimed at uncovering common insights.
+6.1 Research questions
 Most of the general overview questions (GQs) were addressed in the overview presented in Sec.
@@
 However, there is a recent growing interest in models that learn the graph structure and the edge weights themselves.
-This trend is expected to become more prevalent in the future.
+While this trend toward adaptive graph learning is expected to gain momentum, the literature remains divided on whether learned or pre-defined adjacency matrices are preferable.
+The choice between these approaches should be based on the availability of domain knowledge and model performance requirements, requiring further research to establish clearer guidelines.
 SQ3 (Taxonomy).
 As observed in Ref.
-[survey_spatiotemporal], there are two main approaches for the design of spatio-temporal GNNs: one that treats the spatial and temporal components in separate modules, and another that integrates and processes them together.
+[467], there are two main approaches for the design of spatio-temporal GNNs: one that treats the spatial and temporal components in separate modules, and another that integrates and processes them together.
 The analysis of the collected papers reveals that the most common approach in literature is that of separate modules.
@@
 Hence, the proposed taxonomy of GNN models takes its main perspective from the point of view of the spatial component.
-The review indicates that the convolutional and the attentional approaches are the prevalent ones.
-Approximately 62% of the models are purely convolutional, 21% are purely attentional, and 9% are hybrid convolutional-attentional.
+The review indicates that convolutional approaches are the most prevalent, followed by attentional methods.
+Approximately 60% of the models are purely convolutional, 22% are purely attentional, and 10% are hybrid convolutional-attentional.
+Additional statistics related to taxonomy distributions by year and research field are available in the GitHub repository associated with this review.
 With regard to the temporal component, the recurrent structures of GRUs and the attention mechanism are widely used.
@@
 Notably, in “Mobility”, “Environment”, and “Generic” groups (which are also the most investigated fields), there are many reference GNN benchmark models.
-Among the most prevalent benchmarks there are: the graph convolutional neural network with long-short term memory (GC-LSTM) [GC-LSTM], spatio-temporal graph convolutional network (STGCN) [STGCN], attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN], diffusion convolutional recurrent neural network (DCRNN) [DCRNN], spatio-temporal graph convolutional network (STGCN) [STGCN], and multivariate time series forecasting with graph neural network (MTGNN) [MTGNN].
+The most widely used GNN benchmarks, spanning all thematic groups, include: the graph convolutional neural network with long-short term memory (GC-LSTM) [316], spatio-temporal graph convolutional network (STGCN) [457], attention-based spatial-temporal graph convolutional network (ASTGCN) [116], diffusion convolutional recurrent neural network (DCRNN) [237], and multivariate time series forecasting with graph neural network (MTGNN) [424].
+These GNN benchmarks, mostly published between 2018 and 2020, represent some of the earliest and most cited works in spatio-temporal GNN literature.
+Almost all of them employ graph convolutional mechanisms, often Chebyshev-based or similar variants, reflecting the state-of-the-art at the time.
+In most cases, the adjacency matrix is pre-defined, often based on distances between measurement sites, whereas just one approach incorporates a learned adjacency matrix (i.e., MTGNN, which is one of the first models to use self-learned graph structures).
+In terms of application domains, four out of five of these benchmarks were originally developed to focus on “Mobility” or “Generic” tasks, with only GC-LSTM originally addressing environmental applications.
+Among these benchmarks, ASTGCN also explores attention mechanisms alongside graph convolutions and constitutes one of the earliest hybrid architectures.
+Overall, these benchmarks illustrate the early development of GNNs: graph convolution dominates, adjacency matrices are usually static, and mobility-related tasks are the most commonly studied.
 SQ5 (Benchmark datasets).
@@
 Hence, it would be beneficial for the community to agree on, and start adopting, some of the most commonly used datasets, with the goal of developing a shared benchmark dataset.
-A possible common benchmark could be the traffic one.
+Some possible common benchmark could be the ones presented in the “Mobility” and the “Generic” thematic group.
 Hopefully, this review could serve this purpose by providing comprehensive tables of the results of the models on different datasets to facilitate comparison between spatio-temporal GNN models.
@@
 However, the most common ones are mean absolute error (MAE), mean squared error (MSE), and mean absolute percentage error (MAPE) for forecasting problems, and accuracy for classification problems.
-A meta-analysis of papers from different areas indicates that some thematic groups have seen more intense research on GNN models than others.
-For instance, the papers in the “Mobility” and the “Generic” groups present more complex models than others, this being likely due to the more extensive research conducted in these fields, as also evidenced by the larger number of published papers.
-In contrast, the “Finance” and “Health” thematic groups have less papers, and often have only a limited number of benchmarks available for comparison.
+6.2 Observed trends
+Our analysis of papers from different domains indicates that some thematic groups have seen more intense research on GNN models than others.
+For instance, the papers in the “Mobility” and “Generic” groups generally present more complex models, likely reflecting the greater research effort in these fields, as evidenced by the larger number of published papers.
+In contrast, the “Finance” and “Health” thematic groups have fewer papers and often offer only a limited number of benchmarks for comparison.
 Another notable example is the “Energy” group, which has a relatively large number of papers, particularly in 2024.
-However, the energy community tends to align more closely with traditional econometrics or non-GNN machine learning models, and the proposed GNN models in this field are generally less complex than those found in other groups.
-This suggests that some fields are currently more prone to the development of innovative and complex models, whereas others require a greater focus on adapting existing models to real-world problems.
-Furthermore, papers in the “Energy” group often test their models on proprietary dataset, which further complicates the development of models across the entire group.
+However, the related community tends to rely more on traditional econometrics or non-GNN machine learning approaches, and the GNN models proposed are generally less complex than those in other domains.
+Moreover, many papers in the “Energy” thematic group test their models on proprietary datasets, which further complicates model development across the group.
+Overall, this analysis suggests that some fields are more prone to the development of innovative and complex models, whereas others focus on adapting existing models to specific real-world problems.
 Consequently, we expect some form of cross-pollination between the communities, which is likely to reduce these differences over time.
-A final aspect to highlight is the importance of reproducibility and comparability of these studies, which is essential for the advancement of scientific knowledge.
-In some fields, a greater attention to reproducibility can be observed, also thanks to the availability of datasets, a larger number of benchmarks, and more detailed descriptions of the models.
+For readers interested in a more detailed view of how GNN taxonomies evolve across fields and over time, interactive charts are available in the GitHub repository accompanying this paper.
+Another important aspect related to research trends is the focus on reproducibility and comparability of studies, which, more in general, is essential for advancing scientific knowledge.
+Some fields pay greater attention to reproducibility, facilitated by the availability of datasets, a larger number of benchmarks, and more detailed model descriptions.
 Tab.
@@
 Group Datasets (avg.) Benchmarks (avg.) Code availability
-Energy 1.6 5.8 3%
-Environment 1.3 6.5 26%
-Finance 1.8 5.6 12%
-Generic 4.2 10.0 41%
-Health 1.7 7.7 53%
-Mobility 2.6 8.4 27%
-Other 2.1 5.6 14%
-Predictive monitoring 2.1 8.1 14%
+Energy 1.6 5.7 3%
+Environment 1.3 6.6 23%
+Finance 1.5 6.3 7%
+Generic 4.5 10.6 56%
+Health 1.7 8.3 48%
+Mobility 2.8 9.4 35%
+Other 1.9 6.3 15%
+Predictive monitoring 2.2 7.9 22%
+6.3 Meta-analysis across groups
+Here, we conduct a meta-analysis of the reviewed spatio-temporal GNNs models across multiple application domains.
+By definition, all models incorporate both spatial and temporal components, either arranged sequentially (space-then-time) or jointly (space-and-time).
+Beyond this common feature, designs vary widely, and our meta-analysis focuses on the key aspects that characterize these models, namely graph construction and spatial aggregation, aiming to reveal patterns where possible.
+However, the highly fragmented research landscape makes identifying these trends challenging.
+We observed that many models are tailored to specific case studies, and differences in domain focus, limited reproducibility, and inconsistent experimental setups further complicate cross-domain comparisons and the extraction of generalizable insights.
+For this reason, rather than attempting a meta-analysis across the entire fragmented landscape of reviewed papers, we focused on the best-performing models within each domain to identify shared characteristics and emerging trends.
+In particular, we selected those models that, in the results tables, ranked among the top performers for the associated datasets and that were briefly discussed below each table.
+These models were gathered across all groups and are available in the GitHub repository associated with this work.
+In this selected subset, models from the “Generic” and “Mobility” thematic groups are more numerous.
+This is not because they are inherently more important, but because these models offer greater comparability, having been evaluated across multiple common datasets in their original studies.
+For this reason, they are particularly suitable for this broader meta-analysis, which does not focus on highly specific use cases.
+As for the determination of the graph structure, most models in the selected subset have a self-learning module.
+Pre-defined adjacency matrices based on physical distances, correlations, cosine similarities, or similar techniques remain more common in studies focused on specific tasks.
+From a historical perspective, early models relied more heavily on pre-defined adjacency matrices, whereas more recent works increasingly favor learned or hybrid adjacency matrices, often combining learned and pre-defined structures to enhance adaptivity and robustness.
+This shift toward learned or hybrid adjacency matrices, often implemented via self-learning modules, supports broader applicability across diverse domains.
+Among self-learning modules in spatio-temporal GNNs, two predominant approaches have emerged with comparable prevalence in the selected subset of papers: the method of the GWN model [425] and that of the MTGNN model [424], both developed by Wu et al, and discussed in Subsec.
+4.4.
+As for the spatial aggregation module, the subset reveals that graph convolution remains the dominant spatial aggregation mechanism in high-performing spatio-temporal GNNs, while attention-based spatial aggregation is less common.
+This suggests that the practical advantages of convolutional aggregation, in terms of efficiency and training stability, continue to make it preferable to attention mechanisms.
+The aggregation mechanism of recurrent GNNs, which defined the original GNN formulation, is instead almost entirely unused.
 7 Limits, challenges, and future research directions
@@
 As highlighted in Ref.
-[cit_rev2], the evaluation of GNN models has improved thanks to the introduction of the Open Graph Benchmark (OGB) [NEURIPS2020_fb60d411], which provides a standardized evaluation framework and a variety of benchmark graph datasets.
+[273], the evaluation of GNN models has improved thanks to the introduction of the Open Graph Benchmark (OGB) [146], which provides a standardized evaluation framework and a variety of benchmark graph datasets.
 However, there is currently no standardized benchmark for spatio-temporal GNNs.
@@
 Some models are claimed to be explainable because of the integration of non-linear physical laws (e.g., Ref.
-[art_401]), while other papers claim explainability on the basis of the possibility to visualize a graph (e.g., Ref.
-[art_360]) or because they present a feature aggregation form that can be intuitively understood.
-In addition, other approaches focus on determining feature importance through tools such as Captum [Captum], which enables the interpretation and understanding of PyTorch model predictions by exploring the features that contribute to a specific outcome (e.g., Ref.
-[art_361]).
+[360]), while other papers claim explainability on the basis of the possibility to visualize a graph (e.g., Ref.
+[106]) or because they present a feature aggregation form that can be intuitively understood.
+In addition, other approaches focus on determining feature importance through tools such as Captum [192], which enables the interpretation and understanding of PyTorch model predictions by exploring the features that contribute to a specific outcome (e.g., Ref.
+[325]).
 This highlights a research gap in this area.
-However, recent literature is beginning to investigate this important topic.
-For instance, few papers studied perturbation-based explanation methods, which however have proven to be somewhat ineffective, as reported in Ref.
-[interpret_3].
-Other approaches, such as GNNExplainer [interpret_2], aim to identify a small subset of node features that have a crucial role in GNN’s predictions, as in Ref.
-[art_349, art_357].
+However, recent literature is beginning to investigate this important topic, as also discussed in [369].
+For instance, few papers studied perturbation-based explanation methods.
+Other approaches, such as GNNExplainer [455], aim to identify a small subset of node features that have a crucial role in GNN’s predictions, as in Ref.
+[377, 184].
 Recently, some researchers have started to explore graph counterfactuals as a means of generating explanations, as reviewed in Ref.
-[interpret_1].
+[314].
+For a recent survey on GNNs for explainable AI see Ref.
+[292].
 Quasi-equilibrium learning.
@@
 4.2, GNNs were originally introduced in Ref.
-[Scarselli2009] by promoting an extension of the backpropagation algorithm which exploits a fixed-point principle.
+[332] by promoting an extension of the backpropagation algorithm which exploits a fixed-point principle.
 The subsequent introduction of the notion of convolutional graph has fundamentally changed such computational scheme.
@@
 This is a case where the original computational approach proposed in Ref.
-[Scarselli2009] seems to be more adequate, especially when considering that the rate of weight updates is generally much slower than the rate of external temporal signals.
+[332] seems to be more adequate, especially when considering that the rate of weight updates is generally much slower than the rate of external temporal signals.
 This suggests the necessity of revisiting the process of convergence to the fixed point, and of exploring computational schemes which take place at quasi-equilibrium points.
 We expect that this approach can significantly alleviate the computational burden that is typical of the convergence to fixed-points required for spatio-temporal graph collections.
+Broader Directions.
+The field of time series modeling by means of GNNs is expected to expand in a set of further directions.
+While most of the reviewed studies focus on point forecasting with deterministic models, relatively few address probabilistic forecasting (e.g., Refs.
+[409, 143]), which is crucial for assessing uncertainty in predictions.
+Another potential direction could be the integration of spatio-temporal GNNs with transformer-based architectures or large language models.
+Although studies are still limited, such models could help capture long term dependencies and complex patterns in time series data, yet not forgetting all possible limitations of LLMs in time series applications [367].
+Exploring these directions may help overcome current limitations in modeling uncertainty and complex temporal relationships, providing more informative and reliable results.
 8 Conclusions
@@
 Another key point which results from the presented overview is that the current literature on spatio-temporal GNNs is very fragmented, as it was pictorially reported in Fig.
-3.
+5.
 This can be attributed to the fact that the involved researchers come from different communities which focus on distinct application domains.
@@
 The objective of this review was also to collect information on datasets, proposed models, links to source codes, benchmarks, and results, in order to provide common background information for future studies.
+To support this goal, the GitHub repository associated with this review offers interactive tools for sorting, filtering, and visualizing the results achieved by the different models, making it easier for researchers to explore and compare existing work.
 A further major key point resulting from our review, which we believe should be emphasized in these conclusions, is the need for the spatio-temporal GNN research community to work on common datasets and develop comparable and reproducible models.
@@
 Supplementary material
-Link to the GitHub repository with the links to the 281 reviewed papers, organized by groups: https://github.com/FlaGer99/SLR-Spatio-Temporal-GNN.git.
+Link to the GitHub repository containing information on the 366 reviewed papers organized by groups, along with interactive tables and charts: https://github.com/FlaGer99/SLR-Spatio-Temporal-GNN.git.
 CRediT author statement
@@
 Carlo Lucheroni: Writing - Review & Editing, Supervision.
-Marco Piangerelli: Conceptualization, Methodology, Writing - Review & Editing, Supervision.
-Martina Zannotti: Conceptualization, Methodology, Validation, Investigation, Writing - Original Draft, Writing - Review & Editing, Visualization.
+Marco Piangerelli: Conceptualization, Methodology, Writing - Original Draft, Writing - Review & Editing, Supervision.
+Martina Zannotti: Conceptualization, Methodology, Validation, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization.
+Author names appear in alphabetical order and do not reflect relative contributions.
 Acknowledgments
 This work has been funded by the European Union - NextGenerationEU, Mission 4, Component 2, under the Italian Ministry of University and Research (MUR) National Innovation Ecosystem grant ECS00000041 - VITALITY - CUPJ13C22000430001; and the NextGenerationEU - National Recovery and Resilience Plan (PNRR) grant CUP J11J23001450006, with the contribution of Syeew srl.
-Appendix A List of selected journal papers
-Tab.
-LABEL:t:selected_papers lists all the journal papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (classification or forecasting).
-Table 41: List of the 281 selected journal papers, with year, group they belong to, case study, and nature of the task.
-Ref. Year Group Case study Task
-[art_15] 2023 Energy Heat load Forecasting
-[art_16] 2023 Energy Electricity net load Forecasting
-[art_17] 2023 Energy Residential load Forecasting
-[art_18] 2023 Energy Electricity load Forecasting
-[art_19] 2023 Energy Photovoltaic power Forecasting
-[art_20] 2022 Energy Wind power Forecasting
-[art_21] 2022 Energy Photovoltaic power Forecasting
-[art_22] 2022 Energy Photovoltaic power Forecasting
-[art_23] 2022 Energy Power system transient dynamics Forecasting
-[art_24] 2020 Energy Wind power Forecasting
-[art_25] 2022 Energy Urban energy consumption Forecasting
-[art_26] 2023 Energy Solar power Forecasting
-[art_27] 2022 Energy Wind power Forecasting
-[art_332] 2024 Energy Wind power Forecasting
-[art_333] 2024 Energy Wind power Forecasting
-[art_334] 2024 Energy Air conditioning control Forecasting
-[art_335] 2024 Energy Air conditioning control Forecasting
-[art_337] 2024 Energy Electrical demand and prices Forecasting
-[art_338] 2024 Energy Electric load Forecasting
-[art_339] 2024 Energy Electric load Forecasting
-[art_340] 2024 Energy Chiller energy consumption Forecasting
-[art_341] 2024 Energy Electricity demand Forecasting
-[art_342] 2024 Energy Electric power anomaly detection Classification
-[art_343] 2024 Energy Electricity demand Forecasting
-[art_344] 2024 Energy Electricity consumption Forecasting
-[art_345] 2024 Energy District heating networks Forecasting
-[art_347] 2024 Energy Power quality disturbances Classification
-[art_348] 2024 Energy Residential load Forecasting
-[art_349] 2024 Energy Photovoltaic energy Forecasting
-[art_350] 2024 Energy Electric vehicle charging stations Forecasting
-[art_351] 2024 Energy Natural gas consumption Forecasting
-[art_353] 2024 Energy Networked microgrids Forecasting
-[art_390] 2024 Energy Multivariate time series Forecasting
-[art_1] 2023 Environment Air quality index Forecasting
-[art_2] 2023 Environment PM 2.5 concentration Forecasting
-[art_3] 2023 Environment PM 2.5 concentration Forecasting
-[art_4] 2023 Environment PM 2.5 concentration Forecasting
-[art_5] 2023 Environment Air quality index Forecasting
-[art_6] 2023 Environment PM 2.5 concentration Forecasting
-[art_7] 2023 Environment PM 2.5 concentration Forecasting
-[art_8] 2022 Environment PM 2.5 concentration Forecasting
-[art_28] 2023 Environment Sea temperature Forecasting
-[art_29] 2023 Environment Wind speed Forecasting
-[art_30] 2023 Environment Water quality Forecasting
-[art_31] 2023 Environment Wind speed Forecasting
-[art_32] 2023 Environment Significant wave height Forecasting
-[art_33] 2023 Environment Sea temperature Forecasting
-[art_34] 2023 Environment Wind speed Forecasting
-[art_35] 2023 Environment Sea temperature Forecasting
-[art_36] 2023 Environment Sea temperature Forecasting
-[art_38] 2022 Environment Frost Both
-[art_39] 2022 Environment Sea temperature Forecasting
-[art_40] 2021 Environment Wind speed Forecasting
-[art_41] 2021 Environment Sea temperature Forecasting
-[art_42] 2022 Environment Water quality Forecasting
-[art_43] 2023 Environment Groundwater level Forecasting
-[art_44] 2023 Environment Wind speed Forecasting
-[art_45] 2023 Environment Rainfall Forecasting
-[art_46] 2023 Environment Wind speed Forecasting
-[art_47] 2023 Environment Sea temperature Forecasting
-[art_354] 2024 Environment PM2.5 prediction Forecasting
-[art_355] 2024 Environment Sea clutter Forecasting
-[art_356] 2024 Environment Earthquakes Forecasting
-[art_357] 2024 Environment Plankton dynamics Forecasting
-[art_359] 2024 Environment Groundwater data Forecasting
-[art_360] 2024 Environment Carbon emission Forecasting
-[art_361] 2024 Environment PM2.5 prediction Forecasting
-[art_363] 2024 Environment Water demand Forecasting
-[art_364] 2024 Environment Water quality and flow rate Forecasting
-[art_365] 2024 Environment Weather Forecasting
-[art_366] 2024 Environment Wind speed scenario Forecasting
-[art_367] 2024 Environment Carbon emission Forecasting
-[art_368] 2024 Environment Air quality Forecasting
-[art_370] 2024 Environment Temperature Forecasting
-[art_371] 2024 Environment Air quality Forecasting
-[art_373] 2024 Environment Greenhouse gas concentrations Forecasting
-[art_59] 2022 Finance Stock prediction Classification
-[art_60] 2023 Finance Stock prediction Classification
-[art_62] 2023 Finance Stock prediction Classification
-[art_63] 2023 Finance Stock prediction Classification
-[art_64] 2022 Finance Investment prediction Forecasting
-[art_65] 2023 Finance Stock prediction Classification
-[art_66] 2022 Finance Financial prediction Classification
-[art_67] 2023 Finance Trading Classification
-[art_68] 2023 Finance Sale prediction Forecasting
-[art_374] 2024 Finance Oil and metal market Classification
-[art_375] 2024 Finance Supply chain demand Forecasting
-[art_377] 2024 Finance Stock price Classification
-[art_378] 2024 Finance Stock market volatility Forecasting
-[art_379] 2024 Finance Agricultural price Forecasting
-[art_381] 2024 Finance Cryptocurrencies’ price Forecasting
-[art_382] 2024 Finance Stock trend Classification
-[art_69] 2023 Generic Batch workloads Forecasting
-[art_70] 2023 Generic Node classification for websites Classification
-[art_71] 2023 Generic Clinical risk classification Classification
-[art_72] 2023 Generic Sensor network Forecasting
-[art_73] 2023 Generic Multivariate time series Forecasting
-[art_74] 2023 Generic Multivariate time series Forecasting
-[art_75] 2023 Generic Multivariate time series Forecasting
-[art_76] 2023 Generic Long term series forecasting Forecasting
-[art_77] 2023 Generic Multivariate time series Forecasting
-[art_78] 2023 Generic Multivariate time series Forecasting
-[art_79] 2023 Generic Multivariate time series Forecasting
-[art_80] 2023 Generic Multivariate time series Forecasting
-[art_81] 2023 Generic Multivariate time series Forecasting
-[art_82] 2023 Generic Multivariate long sequences Forecasting
-[art_83] 2023 Generic Multivariate time series Forecasting
-[art_84] 2022 Generic Multivariate time series Both
-[art_85] 2023 Generic Multivariate time series Forecasting
-[art_86] 2022 Generic Multivariate time series Forecasting
-[art_87] 2022 Generic Multivariate time series Forecasting
-[art_88] 2023 Generic Multivariate time series Forecasting
-[art_89] 2023 Generic Multivariate time series Forecasting
-[art_90] 2022 Generic Multivariate time series Forecasting
-[art_160] 2023 Generic Impact of graph construction Forecasting
-[art_202] 2021 Generic Generic Both
-[art_383] 2024 Generic Multivariate time series Forecasting
-[art_384] 2024 Generic Multivariate time series Forecasting
-[art_385] 2024 Generic Multivariate time series Classification
-[art_386] 2024 Generic Multivariate time series Classification
-[art_387] 2024 Generic Multivariate time series Both
-[art_388] 2024 Generic Multivariate time series Forecasting
-[art_389] 2024 Generic Multivariate time series Classification
-[art_393] 2024 Generic Multivariate time series Forecasting
-[art_395] 2024 Generic Multivariate time series Classification
-[art_396] 2024 Generic Multivariate time series Forecasting
-[art_397] 2024 Generic Multivariate time series Forecasting
-[art_398] 2024 Generic Multivariate time series Forecasting
-[art_400] 2024 Generic Multivariate time series Forecasting
-[art_401] 2024 Generic Multivariate time series Forecasting
-[art_403] 2024 Generic Multivariate time series Both
-[art_404] 2024 Generic Multivariate time series Forecasting
-[art_405] 2024 Generic Multivariate time series Forecasting
-[art_407] 2024 Generic Multivariate time series Classification
-[art_417] 2024 Generic Multivariate time series Forecasting
-[art_464] 2024 Generic Multivariate time series Forecasting
-[art_14] 2023 Health Epilepsy diagnosis Classification
-[art_48] 2023 Health Epidemic prediction Forecasting
-[art_49] 2023 Health Epidemic prediction Forecasting
-[art_50] 2022 Health Epidemic prediction Forecasting
-[art_51] 2021 Health Epidemic prediction Forecasting
-[art_52] 2022 Health Epidemic prediction Forecasting
-[art_92] 2023 Health Disease diagnosis Classification
-[art_189] 2023 Health Clinical risk classification Classification
-[art_410] 2024 Health Infectious disease Forecasting
-[art_411] 2024 Health EEG emotion recognition Classification
-[art_412] 2024 Health EEG emotion recognition Classification
-[art_413] 2024 Health Epidemic Forecasting
-[art_414] 2024 Health Epidemic Forecasting
-[art_415] 2024 Health Epidemic Forecasting
-[art_416] 2024 Health Epidemic Forecasting
-[art_93] 2022 Mobility Urban traffic Forecasting
-[art_94] 2021 Mobility Crowd flow Forecasting
-[art_96] 2022 Mobility Crowd flow Forecasting
-[art_97] 2023 Mobility Urban traffic Forecasting
-[art_98] 2022 Mobility Urban traffic Forecasting
-[art_99] 2023 Mobility Urban traffic Forecasting
-[art_100] 2023 Mobility Urban traffic Forecasting
-[art_102] 2021 Mobility Passenger demand Forecasting
-[art_103] 2023 Mobility Urban traffic Forecasting
-[art_104] 2022 Mobility Internet of vehicles traffic Forecasting
-[art_105] 2020 Mobility Urban traffic Forecasting
-[art_106] 2023 Mobility Urban traffic Forecasting
-[art_108] 2023 Mobility Urban traffic Forecasting
-[art_109] 2023 Mobility Urban traffic Forecasting
-[art_110] 2023 Mobility Urban traffic Forecasting
-[art_113] 2023 Mobility Urban traffic Forecasting
-[art_114] 2023 Mobility Bike station demand Forecasting
-[art_115] 2023 Mobility Urban traffic Forecasting
-[art_116] 2023 Mobility Urban traffic Forecasting
-[art_117] 2023 Mobility Urban traffic Forecasting
-[art_119] 2022 Mobility Taxi demand Forecasting
-[art_120] 2022 Mobility Urban traffic Forecasting
-[art_121] 2022 Mobility Metro passengers Forecasting
-[art_122] 2022 Mobility Urban traffic Forecasting
-[art_123] 2021 Mobility Urban traffic Forecasting
-[art_124] 2022 Mobility Urban traffic Forecasting
-[art_125] 2023 Mobility Urban traffic Forecasting
-[art_126] 2020 Mobility Urban traffic Forecasting
+Appendix A List of selected papers
+Tab. LABEL:t:selected_papers lists all the papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (classification or forecasting).
+Table 41: List of the selected papers, with year, group they belong to, case study, and nature of the task.
+Reference Year Group Case study Task
+[460] 2020 Energy Wind power Forecasting
+[256] 2021 Energy Electric load Forecasting
+[132] 2022 Energy Wind power Forecasting
+[350] 2022 Energy Photovoltaic power Forecasting
+[351] 2022 Energy Photovoltaic power Forecasting
+[509] 2022 Energy Power system transient dynamics Forecasting
+[147] 2022 Energy Urban energy consumption Forecasting
+[216] 2022 Energy Wind power Forecasting
+[405] 2023 Energy Heat load Forecasting
+[486] 2023 Energy Electricity load Forecasting
+[398] 2023 Energy Residential load Forecasting
+[93] 2023 Energy Electricity load Forecasting
+[317] 2023 Energy Solar power Forecasting
+[439] 2024 Energy Wind power Forecasting
+[452] 2024 Energy Wind power Forecasting
+[531] 2024 Energy Air conditioning control Forecasting
+[148] 2024 Energy Air conditioning control Forecasting
+[323] 2024 Energy Electrical demand and prices Forecasting
+[104] 2024 Energy Electric load Forecasting
+[477] 2024 Energy Electric load Forecasting
+[265] 2024 Energy Chiller energy consumption Forecasting
+[288] 2024 Energy Electricity demand Forecasting
+[161] 2024 Energy Electric power anomalies Classification
+[324] 2024 Energy Electricity demand Forecasting
+[156] 2024 Energy Electricity consumption Forecasting
+[18] 2024 Energy District heating networks Forecasting
+[24] 2024 Energy Power quality disturbances Classification
+[87] 2024 Energy Residential load Forecasting
+[377] 2024 Energy Photovoltaic energy Forecasting
+[500] 2024 Energy Electric vehicle charging stations Forecasting
+[315] 2024 Energy Natural gas consumption Forecasting
+[463] 2024 Energy Networked microgrids Forecasting
+[319] 2024 Energy Multivariate time series Forecasting
+[102] 2021 Environment Wind speed Forecasting
+[361] 2021 Environment Sea temperature Forecasting
+[42] 2022 Environment Water temperature, stream flow Forecasting
+[310] 2022 Environment PM 2.5 concentration Forecasting
+[257] 2022 Environment Frost Both
+[392] 2022 Environment Sea temperature Forecasting
+[5] 2023 Environment Soil Moisture Forecasting
+[41] 2023 Environment Air quality index Forecasting
+[267] 2023 Environment PM 2.5 concentration Forecasting
+[357] 2023 Environment PM 2.5 concentration Forecasting
+[223] 2023 Environment PM 2.5 concentration Forecasting
+[72] 2023 Environment Air quality index Forecasting
+[293] 2023 Environment PM 2.5 concentration Forecasting
+[183] 2023 Environment PM 2.5 concentration Forecasting
+[437] 2023 Environment Sea temperature Forecasting
+[11] 2023 Environment Wind speed Forecasting
+[210] 2023 Environment Water quality Forecasting
+[326] 2023 Environment Wind speed Forecasting
+[502] 2023 Environment Significant wave height Forecasting
+[185] 2023 Environment Sea temperature Forecasting
+[12] 2023 Environment Wind speed Forecasting
+[304] 2023 Environment Sea temperature Forecasting
+[98] 2023 Environment Sea temperature Forecasting
+[10] 2023 Environment Groundwater level Forecasting
+[97] 2023 Environment Wind speed Forecasting
+[311] 2023 Environment Rainfall Forecasting
+[73] 2023 Environment Wind speed Forecasting
+[345] 2023 Environment Sea temperature Forecasting
+[133] 2024 Environment Weather Forecasting
+[248] 2024 Environment PM2.5 prediction Forecasting
+[229] 2024 Environment Sea clutter Forecasting
+[184] 2024 Environment Plankton dynamics Forecasting
+[366] 2024 Environment Groundwater data Forecasting
+[106] 2024 Environment Carbon emission Forecasting
+[325] 2024 Environment PM2.5 prediction Forecasting
+[416] 2024 Environment Water demand Forecasting
+[417] 2024 Environment Water quality and flow rate Forecasting
+[336] 2024 Environment Weather Forecasting
+[269] 2024 Environment Wind speed scenario Forecasting
+[45] 2024 Environment Air quality Forecasting
+[403] 2024 Environment Temperature Forecasting
+[426] 2024 Environment Air quality Forecasting
+[17] 2024 Environment Greenhouse gas concentrations Forecasting
+[343] 2022 Finance Energy markets Forecasting
+[429] 2022 Finance Stock prices Classification
+[151] 2022 Finance Stock prediction Classification
+[225] 2022 Finance Investment prediction Forecasting
+[53] 2022 Finance Financial prediction Classification
+[307] 2023 Finance Stock prediction Classification
+[371] 2023 Finance Stock prediction Classification
+[391] 2023 Finance Stock prediction Classification
+[362] 2023 Finance Stock prediction Classification
+[158] 2023 Finance Trading Classification
+[444] 2023 Finance Sale prediction Forecasting
+[297] 2024 Finance Supply chain demand Forecasting
+[205] 2024 Finance Stock market volatility Forecasting
+[37] 2024 Finance Stock trend Classification
+[68] 2020 Health Epidemics Forecasting
+[334] 2021 Health Covid 19 severity Forecasting
+[514] 2021 Health Epidemics Forecasting
+[290] 2021 Health Epidemic prediction Forecasting
+[38] 2022 Health Epilepsy diagnosis Classification
+[160] 2022 Health Epidemics Forecasting
+[282] 2022 Health Epidemics Forecasting
+[235] 2022 Health Epidemic prediction Forecasting
+[298] 2022 Health Epidemic prediction Forecasting
+[23] 2023 Health Epilepsy diagnosis Classification
+[384] 2023 Health Sleep stage, emotion recognition Classification
+[383] 2023 Health Epilepsy diagnosis Classification
+[483] 2023 Health Epidemic prediction Forecasting
+[462] 2023 Health Epidemic prediction Forecasting
+[498] 2023 Health Disease diagnosis Classification
+[434] 2023 Health Clinical risk classification Classification
+[270] 2024 Health Infectious disease Forecasting
+[527] 2024 Health EEG emotion recognition Classification
+[119] 2024 Health EEG emotion recognition Classification
+[445] 2024 Health Epidemic Forecasting
+[528] 2024 Health Epidemic Forecasting
+[285] 2024 Health Epidemic Forecasting
+[187] 2024 Health Epidemic Forecasting
+[134] 2020 Mobility E-Scooter distribution Forecasting
+[135] 2020 Mobility Bike sharing Forecasting
+[393] 2020 Mobility Urban traffic Forecasting
+[8] 2020 Mobility Urban traffic Forecasting
+[33] 2020 Mobility Urban traffic Forecasting
+[496] 2020 Mobility Urban traffic Forecasting
+[275] 2020 Mobility Urban traffic Forecasting
+[308] 2020 Mobility Urban traffic Forecasting
+[301] 2020 Mobility Metro flow Forecasting
+[431] 2020 Mobility Urban traffic Forecasting
+[39] 2020 Mobility Urban traffic Forecasting
 [127] 2021 Mobility Urban traffic Forecasting
-[art_128] 2022 Mobility Urban traffic Forecasting
-[art_129] 2022 Mobility Flight delay Forecasting
-[art_130] 2022 Mobility Urban traffic Forecasting
-[art_131] 2023 Mobility Urban traffic Forecasting
-[art_132] 2022 Mobility Urban traffic Forecasting
-[art_133] 2022 Mobility Bicycle demand Forecasting
-[art_135] 2022 Mobility Urban traffic Forecasting
-[art_406] 2024 Mobility Multivariate time series Forecasting
-[art_408] 2024 Mobility Multivariate time series Forecasting
-[art_419] 2024 Mobility Urban traffic Forecasting
-[art_420] 2024 Mobility Urban traffic Forecasting
-[art_421] 2024 Mobility Metro flow traffic Forecasting
-[art_422] 2024 Mobility Urban traffic Forecasting
-[art_424] 2024 Mobility Urban traffic Forecasting
-[art_425] 2024 Mobility Network traffic prediction Forecasting
-[art_427] 2024 Mobility Urban traffic Forecasting
-[art_430] 2024 Mobility Urban traffic Forecasting
-[art_431] 2024 Mobility Urban traffic Forecasting
-[art_432] 2024 Mobility Urban traffic Forecasting
-[art_433] 2024 Mobility Urban traffic Forecasting
-[art_434] 2024 Mobility Airport performances Forecasting
-[art_435] 2024 Mobility Urban traffic Forecasting
-[art_438] 2024 Mobility Urban traffic Forecasting
-[art_439] 2024 Mobility Urban traffic Forecasting
-[art_441] 2024 Mobility Urban traffic Forecasting
-[art_442] 2024 Mobility Urban traffic Forecasting
-[art_443] 2024 Mobility Flight delay prediction Forecasting
-[art_444] 2024 Mobility Urban traffic Forecasting
-[art_445] 2024 Mobility Urban traffic Forecasting
-[art_446] 2024 Mobility Bike sharing demand Forecasting
-[art_447] 2024 Mobility Urban traffic Forecasting
-[art_448] 2024 Mobility Urban traffic Forecasting
-[art_449] 2024 Mobility Human mobility Forecasting
-[art_451] 2024 Mobility Urban traffic Forecasting
-[art_452] 2024 Mobility Bus station Profiling
-[art_453] 2024 Mobility Urban traffic Forecasting
-[art_454] 2024 Mobility Bus stop mobility Forecasting
-[art_12] 2023 Other topics Modulation classification Classification
-[art_13] 2022 Other topics Modulation classification Classification
-[art_118] 2023 Other topics Encrypted traffic classification Classification
-[art_136] 2022 Other topics Spatio-temporal events Forecasting
-[art_137] 2023 Other topics Quality indicators prediction Forecasting
-[art_138] 2023 Other topics Cellular traffic prediction Forecasting
-[art_139] 2023 Other topics Seismic data Forecasting
-[art_140] 2023 Other topics Structural dynamics Forecasting
-[art_141] 2023 Other topics Tailing dam monitoring Forecasting
-[art_142] 2023 Other topics Educational videos engagement Forecasting
-[art_143] 2023 Other topics Subsurface production Forecasting
-[art_144] 2023 Other topics Cellular Network Traffic Forecasting
-[art_145] 2023 Other topics Data centers maintenance Forecasting
-[art_146] 2023 Other topics Hydraulic runoff Forecasting
-[art_147] 2022 Other topics App popularity Forecasting
-[art_148] 2021 Other topics Cellular Network Traffic Forecasting
-[art_149] 2022 Other topics Water demand Forecasting
-[art_151] 2023 Other topics User preference Edge prediction
-[art_159] 2021 Other topics Human activity recognition Classification
-[art_303] 2023 Other topics Modulation classification Classification
-[art_455] 2024 Other topics Polypropylene production data Forecasting
-[4] 2024 Other topics Botnet detection Classification
-[art_457] 2024 Other topics Cellular KPI prediction Forecasting
-[art_458] 2024 Other topics Tungsten floatation plant Forecasting
-[art_459] 2024 Other topics Modulation recognition Classification
-[art_460] 2024 Other topics IoT intrusion detection Classification
-[art_463] 2024 Other topics Modulation recognition Classification
-[art_465] 2024 Other topics Crime prediction Forecasting
-[art_468] 2024 Other topics Mobile traffic prediction Forecasting
-[art_9] 2022 Predictive monitoring Anomaly detection Classification
-[art_10] 2023 Predictive monitoring Anomaly detection Forecasting
-[art_53] 2023 Predictive monitoring Fault diagnosis Classification
-[art_54] 2023 Predictive monitoring Fault diagnosis Classification
-[art_55] 2021 Predictive monitoring Fault diagnosis Classification
-[art_56] 2023 Predictive monitoring Fault diagnosis Classification
-[art_57] 2023 Predictive monitoring Fault diagnosis Classification
-[art_58] 2023 Predictive monitoring Fault prediction Forecasting
-[art_152] 2023 Predictive monitoring Remaining useful life Forecasting
-[art_153] 2023 Predictive monitoring Remaining useful life Forecasting
-[art_154] 2023 Predictive monitoring Remaining useful life Forecasting
-[art_155] 2021 Predictive monitoring Remaining useful life Forecasting
-[art_156] 2022 Predictive monitoring Remaining useful life Forecasting
-[art_157] 2022 Predictive monitoring Remaining useful life Forecasting
-[art_158] 2023 Predictive monitoring Remaining useful life Forecasting
-[art_192] 2023 Predictive monitoring Fault diagnosis Classification
-[art_461] 2024 Predictive monitoring Motor temperature Forecasting
-[art_469] 2024 Predictive monitoring Fault diagnosis Classification
-[art_470] 2024 Predictive monitoring Root cause localization Forecasting
-[art_471] 2024 Predictive monitoring Mechanical fault diagnosis Classification
-[art_474] 2024 Predictive monitoring Mechanical fault diagnosis Classification
-[art_475] 2024 Predictive monitoring Outlier detection Classification
-[art_477] 2024 Predictive monitoring Anomaly detection Forecasting
-[art_478] 2024 Predictive monitoring Remaining useful life Forecasting
-[art_479] 2024 Predictive monitoring Anomaly detection Classification
-[art_480] 2024 Predictive monitoring Sensor data Forecasting
-[art_481] 2024 Predictive monitoring Anomaly detection Classification
-[art_482] 2024 Predictive monitoring Process factors Forecasting
-[art_483] 2024 Predictive monitoring Anomaly detection Classification
-[art_484] 2024 Predictive monitoring Anomaly detection Classification
-[art_486] 2024 Predictive monitoring Fault diagnosis Classification
-[art_487] 2024 Predictive monitoring Fault diagnosis Classification
-[art_488] 2024 Predictive monitoring Remaining useful life Forecasting
-[art_489] 2024 Predictive monitoring Anomaly detection Classification
-[art_491] 2024 Predictive monitoring Fault diagnosis Classification
+[84] 2021 Mobility Urban traffic Forecasting
+[222] 2021 Mobility Urban traffic Forecasting
+[219] 2021 Mobility Urban traffic Forecasting
+[428] 2021 Mobility Crowd flow Forecasting
+[402] 2021 Mobility Passenger demand Forecasting
+[341] 2022 Mobility Urban traffic Forecasting
+[276] 2022 Mobility Urban traffic Forecasting
+[342] 2022 Mobility Urban traffic Forecasting
+[203] 2022 Mobility Urban traffic Forecasting
+[259] 2022 Mobility Urban traffic Forecasting
+[85] 2022 Mobility Urban traffic Forecasting
+[213] 2022 Mobility Urban traffic Forecasting
+[370] 2022 Mobility Urban traffic Forecasting
+[320] 2022 Mobility Urban traffic Forecasting
+[274] 2022 Mobility Urban traffic Forecasting
+[211] 2022 Mobility Crowd flow Forecasting
+[220] 2022 Mobility Urban traffic Forecasting
+[489] 2022 Mobility Connected internet of vehicles Forecasting
+[251] 2022 Mobility Taxi demand Forecasting
+[490] 2022 Mobility Urban traffic Forecasting
+[506] 2022 Mobility Metro passengers Forecasting
+[181] 2022 Mobility Urban traffic Forecasting
+[278] 2022 Mobility Urban traffic Forecasting
+[22] 2022 Mobility Urban traffic Forecasting
+[25] 2022 Mobility Flight delay Forecasting
+[112] 2022 Mobility Urban traffic Forecasting
+[27] 2022 Mobility Urban traffic Forecasting
+[71] 2022 Mobility Bicycle demand Forecasting
+[169] 2022 Mobility Urban traffic Forecasting
+[178] 2023 Mobility Urban traffic Forecasting
+[81] 2023 Mobility Urban traffic Forecasting
+[524] 2023 Mobility Urban traffic Forecasting
+[122] 2023 Mobility Urban traffic Forecasting
+[80] 2023 Mobility Urban traffic Forecasting
+[422] 2023 Mobility Urban traffic Forecasting
+[387] 2023 Mobility Urban traffic congestion Forecasting
+[123] 2023 Mobility Visitation patterns Forecasting
+[214] 2023 Mobility Urban traffic congestion Classification
+[281] 2023 Mobility Urban traffic Forecasting
+[145] 2023 Mobility Urban traffic Forecasting
+[411] 2023 Mobility Urban traffic Forecasting
+[62] 2023 Mobility Urban traffic Forecasting
+[218] 2023 Mobility Urban traffic Forecasting
+[212] 2023 Mobility Urban traffic Forecasting
+[263] 2023 Mobility Urban traffic Forecasting
+[262] 2023 Mobility Urban traffic Forecasting
+[173] 2023 Mobility Urban traffic Forecasting
+[217] 2023 Mobility Urban traffic Forecasting
+[200] 2023 Mobility Urban traffic Forecasting
+[198] 2023 Mobility Urban traffic Forecasting
+[245] 2023 Mobility Bike sharing demand Forecasting
+[438] 2023 Mobility Urban traffic Forecasting
+[120] 2023 Mobility Urban traffic Forecasting
+[111] 2023 Mobility Urban traffic Forecasting
+[511] 2023 Mobility Urban traffic Forecasting
+[69] 2023 Mobility Urban traffic Forecasting
+[171] 2024 Mobility Urban traffic Forecasting
+[493] 2024 Mobility Urban traffic Forecasting
+[144] 2024 Mobility Urban traffic Forecasting
+[95] 2024 Mobility Urban traffic Forecasting
+[530] 2024 Mobility Urban traffic Forecasting
+[63] 2024 Mobility Urban traffic Forecasting
+[523] 2024 Mobility Urban traffic Forecasting
+[461] 2024 Mobility Urban traffic Forecasting
+[227] 2024 Mobility Urban traffic Forecasting
+[32] 2024 Mobility Urban traffic Forecasting
+[163] 2024 Mobility Urban traffic Forecasting
+[159] 2024 Mobility Urban traffic Forecasting
+[385] 2024 Mobility Metro flow traffic Forecasting
+[466] 2024 Mobility Urban traffic Forecasting
+[407] 2024 Mobility Urban traffic Forecasting
+[312] 2024 Mobility Network traffic volume Forecasting
+[13] 2024 Mobility Traffic speed Forecasting
+[199] 2024 Mobility Traffic speed Forecasting
+[330] 2024 Mobility Urban traffic Forecasting
+[352] 2024 Mobility Urban traffic Forecasting
+[242] 2024 Mobility Urban traffic Forecasting
+[420] 2024 Mobility Airport performances Forecasting
+[510] 2024 Mobility Urban traffic Forecasting
+[406] 2024 Mobility Urban traffic Forecasting
+[348] 2024 Mobility Urban traffic Forecasting
+[378] 2024 Mobility Urban traffic Forecasting
+[419] 2024 Mobility Flight delay prediction Forecasting
+[208] 2024 Mobility Traffic speed Forecasting
+[250] 2024 Mobility Urban traffic Forecasting
+[89] 2024 Mobility Bike sharing demand Forecasting
+[126] 2024 Mobility Urban traffic Forecasting
+[388] 2024 Mobility Urban traffic Forecasting
+[82] 2024 Mobility Urban traffic Forecasting
+[193] 2024 Mobility Bus stop mobility pattern Forecasting
+[207] 2021 Predictive monitoring Fault diagnosis Classification
+[291] 2021 Predictive monitoring Remaining useful life Forecasting
+[128] 2022 Predictive monitoring Anomaly detections Forecasting
+[221] 2022 Predictive monitoring Anomaly detection Classification
+[441] 2022 Predictive monitoring Remaining useful life) Forecasting
+[195] 2022 Predictive monitoring Remaining useful life Forecasting
+[204] 2023 Predictive monitoring Anomaly detection Classification
+[139] 2023 Predictive monitoring Anomaly detection Classification
+[226] 2023 Predictive monitoring Anomaly detection Forecasting
+[374] 2023 Predictive monitoring Fault diagnosis Classification
+[464] 2023 Predictive monitoring Fault diagnosis Classification
+[296] 2023 Predictive monitoring Fault diagnosis Classification
+[497] 2023 Predictive monitoring Fault prediction Forecasting
+[400] 2023 Predictive monitoring Remaining useful life Forecasting
+[46] 2023 Predictive monitoring Remaining useful life Forecasting
+[382] 2023 Predictive monitoring Remaining useful life Forecasting
+[408] 2023 Predictive monitoring Remaining useful life Forecasting
+[206] 2023 Predictive monitoring Fault diagnosis Classification
+[79] 2024 Predictive monitoring Photovoltaic degradation Forecasting
+[272] 2024 Predictive monitoring Anomaly detection Classification
+[283] 2024 Predictive monitoring Anomaly detection Classification
+[239] 2024 Predictive monitoring Anomaly detection Classification
+[31] 2024 Predictive monitoring Motor temperature Forecasting
+[196] 2024 Predictive monitoring Fault diagnosis and detection Classification
+[86] 2024 Predictive monitoring Root cause localization Forecasting
+[389] 2024 Predictive monitoring Fault diagnosis Classification
+[96] 2024 Predictive monitoring Outlier detection Classification
+[515] 2024 Predictive monitoring Anomaly detection Classification
+[488] 2024 Predictive monitoring Remaining useful life Forecasting
+[249] 2024 Predictive monitoring Anomaly detection Classification
+[224] 2024 Predictive monitoring Sensor data Forecasting
+[114] 2024 Predictive monitoring Anomaly detection Classification
+[363] 2024 Predictive monitoring Process factors Forecasting
+[21] 2024 Predictive monitoring Anomaly detection Classification
+[512] 2024 Predictive monitoring Anomaly detection Classification
+[379] 2024 Predictive monitoring Fault diagnosis Classification
+[476] 2024 Predictive monitoring Fault diagnosis Classification
+[194] 2024 Predictive monitoring Remaining useful life Forecasting
+[136] 2024 Predictive monitoring Anomaly detection Classification
+[381] 2024 Predictive monitoring Fault diagnosis Classification
+[252] 2021 Other Cellular traffic Forecasting
+[92] 2021 Other Product demand Forecasting
+[303] 2021 Other Cellular traffic Forecasting
+[459] 2021 Other Cellular Network Traffic Forecasting
+[121] 2022 Other Talent demand and supply Forecasting
+[435] 2022 Other Modulation classification Classification
+[175] 2022 Other Spatial temporal event prediction Forecasting
+[501] 2022 Other App popularity Forecasting
+[465] 2022 Other Water demand Forecasting
+[475] 2023 Other Encrypted traffic Classification
+[108] 2023 Other Cellular traffic Forecasting
+[380] 2023 Other Modulation classification Classification
+[70] 2023 Other Encrypted traffic classification Classification
+[492] 2023 Other Quality indicators prediction Forecasting
+[404] 2023 Other Cellular traffic prediction Forecasting
+[16] 2023 Other Seismic data Forecasting
+[410] 2023 Other Structural dynamics Forecasting
+[328] 2023 Other Tailing dam monitoring Forecasting
+[266] 2023 Other Subsurface production Forecasting
+[34] 2023 Other Cellular Network Traffic Forecasting
+[344] 2023 Other Data centers maintenance Forecasting
+[440] 2023 Other Hydraulic runoff Forecasting
+[6] 2023 Other Modulation classification Classification
+[150] 2024 Other Anomaly in microservice system Classification
+[255] 2024 Other Emergency supply Forecasting
+[130] 2024 Other Polypropylene production Forecasting
+[3] 2024 Other Botnet detection Classification
+[254] 2024 Other Cellular KPI prediction Forecasting
+[399] 2024 Other Tungsten floatation plant Forecasting
+[487] 2024 Other Modulation recognition Classification
+[394] 2024 Other IoT intrusion detection Classification
+[454] 2024 Other Modulation recognition Classification
+[519] 2024 Other Crime prediction Forecasting
+[170] 2024 Other Mobile traffic prediction Forecasting
+[450] 2020 Generic Multivariate time series Forecasting
+[277] 2020 Generic Multivariate time series Forecasting
+[424] 2020 Generic Multivariate time series Forecasting
+[26] 2020 Generic Multivariate time series Forecasting
+[338] 2021 Generic Multivariate time series Forecasting
+[61] 2021 Generic Multivariate time series Forecasting
+[50] 2021 Generic Multivariate time series Forecasting
+[356] 2021 Generic Multivariate time series Forecasting
+[327] 2021 Generic Multivariate time series Forecasting
+[289] 2021 Generic Multivariate time series Classification
+[155] 2021 Generic Generic Both
+[499] 2022 Generic Multivariate time series Forecasting
+[448] 2022 Generic Multivariate time series Forecasting
+[51] 2022 Generic Time series Forecasting
+[271] 2022 Generic Multivariate time series Forecasting
+[48] 2022 Generic Multivariate time series Forecasting
+[268] 2022 Generic Multivariate time series Forecasting
+[472] 2022 Generic Multivariate time series Classification
+[47] 2022 Generic Multivariate time series Forecasting
+[52] 2022 Generic Multivariate time series Forecasting
+[397] 2022 Generic Multivariate time series Forecasting
+[138] 2022 Generic Multivariate time series Forecasting
+[83] 2023 Generic Multivariate time series Both
+[153] 2023 Generic Multivariate time series Forecasting
+[451] 2023 Generic Multivariate time series Forecasting
+[59] 2023 Generic Multivariate time series Forecasting
+[409] 2023 Generic Multivariate time series Forecasting
+[369] 2023 Generic Multivariate time series Forecasting
+[49] 2023 Generic Multivariate time series Forecasting
+[368] 2023 Generic Multivariate time series Forecasting
+[300] 2023 Generic Multivariate time series Forecasting
+[36] 2023 Generic Batch workloads Forecasting
+[232] 2023 Generic Node classification for websites Classification
+[129] 2023 Generic Clinical risk classification Classification
+[60] 2023 Generic Sensor network Forecasting
+[241] 2023 Generic Multivariate time series Forecasting
+[99] 2023 Generic Multivariate time series Forecasting
+[177] 2023 Generic Multivariate time series Forecasting
+[504] 2023 Generic Long term series forecasting Forecasting
+[76] 2023 Generic Multivariate time series Forecasting
+[243] 2023 Generic Multivariate time series Forecasting
+[118] 2023 Generic Multivariate time series Forecasting
+[40] 2023 Generic Multivariate time series Forecasting
+[287] 2023 Generic Multivariate time series Forecasting
+[395] 2023 Generic Multivariate time series Forecasting
+[355] 2023 Generic Multivariate time series Forecasting
+[186] 2023 Generic Multivariate time series Forecasting
+[473] 2023 Generic Multivariate time series Forecasting
+[57] 2023 Generic Multivariate time series Forecasting
+[15] 2023 Generic Impact of graph construction Forecasting
+[432] 2024 Generic Multivariate time series Forecasting
+[458] 2024 Generic Multivariate time series Forecasting
+[436] 2024 Generic Multivariate time series Forecasting
+[286] 2024 Generic Multivariate time series Forecasting
+[491] 2024 Generic Multivariate time series Forecasting
+[279] 2024 Generic Multivariate time series Forecasting
+[456] 2024 Generic Multivariate time series Classification
+[143] 2024 Generic Multivariate time series Forecasting
+[505] 2024 Generic Multivariate time series Forecasting
+[244] 2024 Generic Multivariate time series Forecasting
+[74] 2024 Generic Multivariate time series Classification
+[113] 2024 Generic Multivariate time series Classification
+[396] 2024 Generic Multivariate time series Both
+[125] 2024 Generic Multivariate time series Forecasting
+[261] 2024 Generic Multivariate time series Classification
+[231] 2024 Generic Multivariate time series Forecasting
+[100] 2024 Generic Multivariate time series Classification
+[137] 2024 Generic Multivariate time series Forecasting
+[152] 2024 Generic Multivariate time series Forecasting
+[453] 2024 Generic Multivariate time series Forecasting
+[390] 2024 Generic Multivariate time series Forecasting
+[360] 2024 Generic Multivariate time series Forecasting
+[401] 2024 Generic Multivariate time series Both
+[337] 2024 Generic Multivariate time series Forecasting
+[240] 2024 Generic Multivariate time series Forecasting
+[140] 2024 Generic Multivariate time series Classification
+[443] 2024 Generic Multivariate time series Forecasting